A Bibliography of Publications in The International Journal of Quantum Chemistry:

1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.15

Title word cross-reference

(1s)^{2}(2s)(2p) [Tat73]. (1s2s,^{3} S) [RM79]. (ns) [CH74]. + [FM78a, JŻ79a, Koc74, PK71, RM74, RW75, RM79]. 1/z [Lin70]. 12 [HLMV77]. $18.10 [Cal77b]. 1: 1 [SP71]. 1s [Lec76a, Lec76b]. 1s2p3s [SA77]. 1s2s^{3} [LAM71]. 1s3d [AC70]. 1s3p [AC70]. 1s3s^{3} [LAM71]. 1s^{2}(2p)^{2} [Lar78b]. 1s^{2}2s2p^{n} [HT75]. 1sσ [AM77a]. 1sσ_{g} [AM77b, TS72]. 2 [Han70a]. $22.50 [Gos78a]. $24.95 [Mår78]. $29.95 [Sab75a]. 2d [Cha77]. 2l + 2 [Mos73]. 2p [Tat73]. 2pπ [SHI73]. 2pσ [AM77a]. 2pσ_{u} [TS72]. 2s [GB72]. 2sσ [AM77a]. $30.00 [Shu77a]. $38.50 [Ano78c]. 3_{10} [Meh79]. 3pσ [AM77a]. 3sσ [AM77a]. $43.00 [Cal76b]. $48.50 [Dun78]. $49.50 [Shu77b]. $56.00 [Sti79]. 5 · 50 [Lin73]. $69.50 [Lar78c]. $74 [Ahl77a]. < 2^{N/2−S}, 1^{2S} >↓< 2^{N_{1}/2−S_{1}}, 1^{2S_{1}} > ⊗ < 2^{N_{2}/2} − S_{2} >, 1^{2S_{2}} > [SD79b]. ^{+} [BB70, BS70, Doo78, HL79, HG75, HT75, KP71a, Koł76, KJ78, LH74, Lar75, Lar78a, Lar78b, LV71c, PK71, Pul74, PPB78, RS79, SD78, TS77]. ^{++} [AM77a, JP77, RS79, SW78]. ^{+2} [KD79b]. _{2}^{+} [Gub79, Hop78, SW77]. _{5}^{+} [MHA70]. ^{−} [Goo76, HPK72, KSP78, KPD78, KJ78, LCPP79, RS79, TTS^{+}79]. ^{−1} [SHR76, You75]. ^{−3} [GML77]. ^{1} [BS70, DRCT78, GL74, JE79, JŻ79a, MSM70, RKOU77, SHR76, TTS71, Tat73]. ^{1,3} [AC70]. ^{13} [NN72]. ^{15} [RKOU77]. ^{1}S [Lar78b]. ^{1}Σ [JŻ79a, Koł75, Lec76a, Lec76b]. ^{2} [GH72b, LL75, LG75, MLB76, RS79, Sim78c]. ^{2+} [BMD70, DS74, JMP79, KP71b, PPB78]. ^{2−} [BBL77, MMR74, RC71]. ^{2}Σ^{−} [HE77]. ^{3} [BK74, DRCT78, DD71, GL74, MBM71, TTS71, Tat73]. ^{3+} [DS74]. 3− [RC71]. ^{3}Π_{u} [Tru73a]. ^{3}Σ_{u} [TTS^{+}79]. ^{4} [AS78]. ^{4}P^{θ} [SA77]. ^{e} [AS78]. ^{II} [KPD78]. ^{N} [JSK75b]. ^{o} [AS78]. ′Π_{u} [Rit76]. ′′Σ [Rit76]. _{1} [DRCT78, GBT79, JŻ79a]. _{1/2} [Sim78c]. _{1s} [FWFM77]. _{2} [ÅSW77, Ada78, Ada73, Ada77, AAC76, BT78a, BB74, BP75a, BL70, BS76b, BDT77, BMD70, Bur77, BWRD79, Cur77, Cur78, EH73, FM78a, FMZ70, GG79a, GBT79, GV77, GS70a, GGJ77, GH72b, GBK77b, Han79, Har70a, Har77a, HLGC73, HNT78, HLYC77, JTR77, JŻ79a, JR74, KP71a, Koc74, KPD78, KPÖ76, LL72, LGP75, LGM77, Lec76a, Lec76b, LG77, LE72, LGR73, LK78, MM72a, Mes73a, Min79, MMR74, MG77, OSL77, PK71, PK76b, PD79, RR77, RW75, SR77, Spe72a, SBD^{+}70, TS77, TBB70, TTS^{+}79, TG76, Tru72, Tru73b, Tru73a, VG78, WK73, WB73, ZL78]. _{2}^{+} [BC79, CJ73b, LGR73]. _{2}^{++} [LGR73]. _{2}^{−} [HR79]. _{2α} [KD79b]. _{2v} [Ros78a]. _{3} [Bai74, BMD70, Bur77, BW76c, CE73, Cur77, EL73, FM78a, FMZ70, FJP70, GG79a, GBT79, GH72b, HNT78, HPK72, HLYC77, JR74, KC72, LL72, Liu71a, MS71b, RMB78, RC70, RC71, ST72, TG76, WN71]. _{3}^{+} [HPK79, LGR73]. _{3h} [TP77b, Wei76]. _{4} [And79, EP71, FMZ70, GG79a, Har70a, HP77b, JŻ79a, LB79a, Mes73a, ST72]. _{4}^{+} [Bai74, KR71]. _{5} [BDT77, LGP75, LGM77, MHA70, RR77]. _{5}^{+} [KR71]. _{6} [ACK^{+}75, BG78, BBL77, Cur78, GG79a, GML77, HP77a, Spe72a, TDB77]. _{6}^{3−} [KL78]. _{8} [MMR74]. _{>} [WRS70]. _{α} [BFH74, Sla73a]. _{C13−C13} [Mak71]. C13−H1 [Mak71]. _{h} [Ros78a]. _{II}, III [BP75a]. _{ij} [MM74]. _{∞h} [Ros78a]. _{∞v} [Ros78a]. _{n} [KC77, TG76]. _{n+1} [TG76]. _{x} [KKA78]. A [LS72]. α [AD77, BB75, Con72, Dan72, DC73, DCS77, GML77, HC72, JW72, Joh73, JM76a, Joh77, KC79, Kau73, Li76, Meh79, MS76a, Min79, MMR74, PKK75, PK77c, RK77, RK79, Sam75, SF76, Sha76, SS71a, Sla75, WC75b, WK73, YJ76]. α_{R} [RK77]. Be [OJB77]. β [BKW75, BCL79, Boy77, Dav77b, EB74, KVSG79, RRS75, RU76, VU79, ZG79]. · [DMS^{+}78, LL72, PPP79]. CH^{+} [OJB77]. d [BSZ78, CW77, FGV77, KC77, KP72, MGM78, MMS70, RS78a]. ϵ_{2} [Lin70]. f [KC77, MMS70]. f^{6} [MS71a, SM71a]. f^{7} [SM71b, SM71c]. g [Kij72]. G(a) [LA76]. G^{1}Π ∼ I^{1}Π [BEF78]. γ [Kot76]. j [BW76b, But76, BW76c]. J(0) [Kai77]. j_{L}(kr)Y_{L}^{M}(θ, φ) [SF75]. jm [BW76b, But76, BW76c]. K_{2s2p} [Tat73]. L^{2} [FB76]. l^{n}l [JSK75a]. m [Goł79]. N [Boy77, Col70, Col71, Col78, Gal72a, Jun78, KK76b, Kum77, LD71, Mat74a, Per78, TPC77, Gal74, Lét74]. n ≥ 3 [Lét74]. ns [CH75]. OAO [CK74]. ω [MSKM77]. p [KC77, Mat79b, RS78a]. Φ [SPČ72]. π [Bra75, Bra79a, ČPP75, HTG77a, JPTM78, KTO^{+}78, KJ75, KT76, KL79, Loc77d, MCST72, MK75a, Mic76b, MK72, NT79, PN76, San72, SPČ79, SIM75]. π → π [DRCT78]. r^{N}Y_{L}^{M}(θ, φ) [SF75, ST70]. r_{12}^{−1} [SR72a, Sal73]. r_{<} [WRS70]. → [FM78a, JŻ79a, PK71]. s^{2}p^{4}d [YW73]. Σ [SIM75, BB72c, GBS70, HTG77a]. SO(4, 2) [Oda73]. sp [DF76]. sp^{6} [YW73]. SU(2) ⊃ D′_{∞} ⊃ D′_{4} ⊃ D′_{2} [KG77]. ⊃ T ⊃ C_{3} ⊃ C_{1} [BW76c]. U(n) ↓ U(n_{1}) ⊗ U(n_{2}) [SD79b]. x [RB75a]. Z [MW75c]. ζ [Sel75].

* [DRCT78, NT79].

-adrenergic [Dav77b, KVSG79]. -adrenomimetics [ZG79]. -alkylnitrosamines [TPC77]. -band [MGM78]. -bends [BCL79]. -Berechnungen [BKW75]. -bonds [Loc77d]. -completeness [KK76b]. -Electron [Bra75, Mic76b, Bra79a, Gal74, JPTM78, MK75a, PN76, SHI73, San72, SPČ79, SIM75, Gal72a, Jun78]. -electronic [ČPP75, KJ75, MK72, SPČ72]. -electrons [SIM75]. -expansion [Lin70, MW75c]. -fermion [Per78]. -fold [Lét74]. -functions [Sha76]. -glycine [RRS75]. -helical [VU79]. -hybrid [DF76]. -ionized [KTO^{+}78]. -Lactam [Boy77]. -line [Kot76]. -matrices [Han70a]. -matrix [Kij72]. -method [DC73]. -molecule [GS70a]. -monosubstituierten [FJP70]. -monosubstituted [FJP70]. -orbitals [FGV77, RS78a, Tat73]. -particle [Col78, Mat74a]. -Picolins [BKW75]. -radicals [KT76]. -Representability [Col71, Col70, Kum77, LD71]. -scattered [BB75]. -state [Mat79b]. -structures [BCL79]. -Substituted [Boy77]. -systems [BB72c, GBS70]. -technique [MSKM77]. -type [LS72, LGR73].

/2 [ZR78]. /3 [ÅFL78, LBÅF78]. /VRDDO [KPPH78, PK77a].

1-fluoroacetic [BL75]. 10-epoxide [LPP78]. 10.50 [Com77a]. 11- [WMB75]. 12.00 [Com77b]. 128K [Spa78]. 134 [Cal77b]. 14th [Pul79b]. 17.50/$29.00 [Cal79b]. 1935-1974 [Cus75]. 1969 [BP70]. 1970 [BP71]. 1971 [BP72, Mar73b, Sla71b, Cal76b]. 1972 [BP73, Cal73b]. 1973 [Sla73b, Sab75a]. 1974 [Cal75c]. 1975 [Lar76]. 1978 [Pul78b, Sti79]. 1s [ÅSW77, DY73, GMZ73]. 1st [Tho76]. 1st-row [Tho76].

2 [DMS^{+}78, DD71, KPKC73, SL73, Wel71, ZBH^{+}78]. 2-benzyne [MPS70]. 2-matrix [MSL73]. 2-particle [LD71]. 20-parameter [BS71]. 225 [Cal76a]. 240 [Cal73a]. 2nd [Cal79b]. 2s [DY73].

358 [Sab75a]. 378 [Lin73]. 3G [ZSR79].

5-fluorouracil [GJP79]. 5-hydroxyindole [WO77]. 5-trifluoromethyluracil [PLSL78].

8-dihydrodiol-9 [LPP78]. 8-tetracyanoquinodimethane [DT75].

a-growing [Lak76]. Ab-initio [KPP77, Meh79, PK73, ZSR79, HKP79, KPKC73, Kau79b, RL79]. ABH [TG76]. abnormal [KHI79, KKP77]. abortifacient [KD79b]. above [Enz73]. absolute [ČZ79]. absorber [Sch71b]. absorbing [BBT71]. Absorption [HMV77, HMV78, AeAs78, BMD70, Bow71, BHLM78, GPC^{+}71, LGM77, NT79, Sch76b, Sch71c]. Abstract [Eig78]. Academic [Com77a, Com77b]. Academy [Cal73b, Pul79b]. acceptor [DMS^{+}78, ZV79]. acceptors [SRE^{+}79]. account [MCF^{+}70]. accuracy [JTR77, MR73]. Accurate [DS74, MH72, PÖ77, RS79, YG77, Bar70, CA72, JM73, HS79]. acetyl [RRNS75]. acetylene [TH78]. acid [BL75, CS74, DS79, FB72, Kla78, KD77, KL79, LL72, LDP79, LP79c, LSB^{+}79, NKKY79, NJF^{+}77, PL77, Pol79, PGPP79, SRE^{+}79, WTB^{+}76, PDEB78]. acid-base [LL72]. Acid-catalyzed [PDEB78]. acide [LL72]. acide-base [LL72]. Acidität [BKW75]. acidity [BKW75]. acids [AMJS79, BC75, IZZ^{+}79, KPP78b, SD77, SMB^{+}79, Sun74a, Sun77, SWF79, UP71, VBC78]. across [Hos78, Mar79]. Actin [Lak75]. acting [WOEG78]. actinomycin [PLYM78]. action [BCR^{+}78, DBBJ76, DMS^{+}78, KKK74, KK75, KD79b, LP79b]. activation [NKKY79, Tal77, LSB^{+}79]. active [AMSS79, BBG79, BWRD79, Cha71, DSMM79, JG74, Kla78, LHK77, LK78, PD79, SS79, Sun74b, Sun75b, ST78, ST79b]. activities [Jan77]. activity [BC75, BCL77, INY^{+}71, Kau79a, PS79a, Pur75, Ran78, SJ79, SH74, SH75b, WCK^{+}76, ZG79]. adaptation [BF71, KC74]. adapted [Amo71, Dav74, Ell73, Kle70a, Kle74a, Kut78, Lim76a, VF76]. additive [GCP79]. Additivity [Sha77a, Sha78a, CW78]. address [Ede70, Lon74, Lon73]. adenine [GJP79, KD77, ZBH^{+}78]. Adenosine [KP79b]. adequacy [CSDV78]. Adiabatic [Dmi75a, Dmi75b, HL79, SG77a]. adjoint [Sco73]. adjustments [Vla70]. admixture [DH73]. adrenergic [Dav77b, KVSG79]. adrenomimetics [ZG79]. adsorbates [Wen79]. adsorbed [Sim79]. adsorption [BFH74, KGB76]. Advances [Cal75c]. Advantages [SW77]. affecting [AF73, CT79, Kar78]. affinities [BB78, CS74, ED79, PK77c, SMH77, SG77a]. AG [Cal76b]. agents [KK72, KK75, Pur75]. agonist [Kan74]. agreement [Sin73a]. AH [KC77]. aid [Nee75]. Akademie [Cal73a]. Akademie-Verlag [Cal73a]. Al [BMD70, Cur78, Pos70]. Ala [RSK^{+}78, VU79]. Ala-Ala-GIy [VU79]. Ala-Pro-Gly-Ala-Pro-Gly [RSK^{+}78]. alanine [RRS75]. Albert [Lak76]. alcalins [JM76b, Mar73a]. AlCN [Tho76]. alcohols [Cur77]. aldehydes [LDP79]. algebra [WG74]. algebraic [ČP77, Kry78, Ell73, WP79]. algebras [APČ77, PAČ77, Wyb73a]. algorithm [Hac74, Sch77a]. Algorithms [Pau77]. alkali [CKD73, CS74, KW76, Mar73a, Moh77, Moh78, Moh79, PBG76, Sim78c, Tal77, YD75, JM76b]. alkali-like [YD75]. alkaline [PBG76, LB73a]. alkaline-earth [PBG76, LB73a]. alkanediols [VH75]. alkanes [BB72c]. alkanols [VH75]. alkenes [Par72]. Alkyl [BB72c]. alkylnitrosamines [TPC77]. all- [WMB75]. All-electron [MMB70]. All-electrons [AL71]. All-pair [NS75]. all-reactive [Hil79]. All-valence-electron [MW74a]. allene [TS79]. allergy [Kau74a]. allosteric [KN79]. allowed [LP74b]. allowing [HHR70]. alloys [CM75b, Fau71, Fau76, DV79, MM77b, Sto71]. allyl [TS78]. Almon [Cal76a]. AlNC [Tho76]. alpha [De 76, Duf71, JW78]. alternancy [MW77]. Alternant [DT75, BK71, ČPH74, ČPP75, JPTM78, KBK72, KU70, PČH74, Tyu75, Tyu78]. Alternate [HC72, But76, HS71, SD79a]. alternations [TTN76]. Alternative [BÖ78, BS74b, Han70b, Feu77, Kib75a, Sac72]. Aluminum [Sco71, SC79]. am [BKW75]. Amaldi [Csa79a, Csa79b]. Ambiguities [Sab73]. amide [PS79b, RRNS75]. amine [LSB^{+}79, VH75]. amines [HL78b, KM72, Nee75, PL79]. amino [Kap73, LP79c, WTB^{+}76, Kla78]. amino-substituted [Kap73]. ammonia [JŻ79b]. ammoniation [PA74]. ammonium [PA74, PBG76]. AMO [CS73, CODM77, CD78, DT75, LV71a, SC73, Spe73]. among [CP70, HYL78, KM72]. AMOS [DD77b, RD79]. amphetamine [DH78, WN78]. amplitude [Hun74]. amplitudes [Hun75a, Hun75b, KPÖ76, Nut78]. Amsterdam [Ahl77a, Ano78c, Lar78c]. anaesthesia [DBJ74, DBBJ76]. analogs [Han79, RSK^{+}78, WN78, ST79b]. analogues [Tho76]. analogy [Sch71c, SE79]. analyses [FTN77, KPKC73]. Analysis [BS76b, LS71, SDB76, Spe71a, Spe71b, Spe72a, Ada78, AAC76, AJ79, BB75, Ber76, BC79, BK75a, CK75, CS76, CL73, CA72, DSC^{+}77, GK78, Hal75a, HPW79, Hun74, Kau70b, KP79a, LS70, Lar78a, LTR74, MG77, PD77, Ran78, RRS75, Rou73, Rou74, TC70, Wei75, WTB^{+}76, ZL78, AF73, RS78a]. Analytic [BO79, Cal72, GGJ77, JS77, KD79a, BSZ78, Doo78, ES79, GGJ75, Jor75, SK72, Wei74]. Analytical [ST70, DS74]. ancient [Lak75]. anesthetics [DMS^{+}78, KK75]. Angular [MP71b, RS78a, RS78b, RS78c, Tat76, BBT71, But76, Cro77, Kib75b, Sch71a, Sur71]. Angularly [LL75]. anhydrase [PD79, SS79, SS74]. aniline [JL70]. anion [Hin77, LCPP79, MMR74, NI70a, PGPP79, PPB78]. anionic [KL79]. anions [KPP78b, Sha77a, Sha78a, Sim78c]. Anisotropic [WMV77]. anisotropies [KW76]. anisotropy [Fog71]. Annales [Cal75b]. annelation [NT79]. Announcement [Ano71a, Ano73a, Ano74a, Ano74b, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Hud72]. Announcements [Ano70a, Ano70b, Ano70c, Ano71b, Ano72a, Ano72b, Ano73b, Ano73c, Ano75a, Ano76a, Ano76b, Ano76c, Ano76d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k]. ansätze [GbK77a]. antagonists [Dav77b, KKK74, KK75]. anthracene [BOP76, Chr73]. anthracenes [SS78b]. anti [BP71, Sla70b]. anti-aromaticity [BP71]. anti-symmetric [Sla70b]. antibinding [ZL78]. antibiotics [SMP77, SMP77]. anticonvulsant [AJ79, Wei77]. anticrossing [BK77]. antiferromagnets [Hub71]. antimony [Kot76]. antiparallel [BCL79]. antipsychotic [KK72, KM72]. antisymmetric [Sim72]. antisymmetrization [Lim77a]. antisymmetrized [Lya78]. antitumor [Gab76]. Anwendung [PTUK69, Cal73a]. any [KD79a]. applicability [DCS77, KP71a, KSP78, SM76b, SM76a]. Application [APČ77, CD74, CSY^{+}74, Kla78, KR78, LB73b, MMR74, Ner75, Nir75, OJB77, Pai71, PAČ77, Ros78b, Tru73b, ANR75, BBA73, CE73, ES73a, Ell73, Gra77, Hal75b, DV79, JM76b, JMP78, KL75, Lar79, NI70a, PLYM78, PGQ77, PKSB78, Rei76, RC70, RC71, RC72, RPK78, SW77, VK74, Bur72, Bur77, BW76c, Goo76, GL74, LH74, LG75, LGR73, Meh78, NI70b, PNJB78, PTUK69, RKE79, SPČ79]. Applications [ERW75, Jor75, Lan77, Røe73, Sun75a, WO76, BO73, Chr79c, DL75, HS70b, Jan76, Joh73, JM76a, KKW73, MS76a, Nal78, Pea74b, WO75, BR78, Csi77, JL70, LG77, MK75a, SDM74, Lar78c]. applied [BEF78, CS73, Dav77b, EB74, Lar78b, LV71b, LV72, SK71, SC73, Spe73]. Approach [Ave79b, AB79, Ave79a, Har71b, OJB77, AC76, AM72b, Ave79c, Boy72b, ČPP75, ČP77, Csa77b, Del73, DC75, DDD70, ES79, FU76, FBL79, Fro74, Gab76, Gin78, Har70b, Joh70, Kau76, KKP77, KE79, Lat76, LB79b, LÖ77, Lin78, Nir75, Pal75, PW79, Pop79, Rou79, SPČ79, SP75a, Sha77b, Sha78b, TB78, TOED79, WP79, ZF76, ZF77, Ano76e]. approaches [APČ77, CL73, Kau77, PAČ77, Rat78]. approximant [ČPP75, CM78]. Approximants [GB71, GBL^{+}73, Røe73, Røe75]. Approximate [Csa79a, KP78, NN74, Sco74, SC78a, Abs72, BDP73, CA78, Cas79, CZ75, DY76, JTR77, JPTM78, KP71a, MDC79, NN77, Nor75, PN76, Sam75, SWP77, GZ72a]. approximately [KPP77]. approximating [GPP79]. Approximation [LP74c, ÖL78, Aro78, BA78, BBP71, BBL72, Bow79, Bre73, CM75a, CL76, CESW70, Dav72, DF76, DLG79, Ess77, ES70, FFCB79, Fro70, GJP79, GLS79, HLGC73, HNT78, HTG77a, HTG77b, KT77, KM76, Lar75, LV71b, LV72, Laz79, Lin71, Meh78, MLM75, Nel79, NMS77, ÖL79a, PNC^{+}78, PW74a, RL77, Rat71, SM77, Sco71, Spe71b, Spe73, SS77b, SO77, Tru73b, Van70, WO76, Wil74]. Approximations [Kas76, Spe72a, DR75, Ive74, JW76, Lan74, PK75]. April [BP70, BP71, BP72, BP73, Pul78b, Sti79]. APW [BK75c, Koe74]. aqueous [Tal77]. arbitrary [BF70, BW76b, But76, BW76c, CM72a, KK77b, SP75b]. ARCANA [AMC73, CAMC71, KPKC73]. ArF [SD78]. arginine [DAP79]. arising [HT75, Tat73]. Aromatic [LSB^{+}79, BOP76, DEE^{+}75, Her74, KPP78b, KJ75, LP79a, Mem79, Nee75, PD77, Pul79a, SG77a, Sha78a, UP71]. Aromaticity [BP71]. aroms [Suz77]. arrangements [JM76b]. Array [BL77b]. Aspects [KPP78c, SLM70, BS75a, BP70, DEC78, ED79, GSE78, HC78, KC73, LP79a, SW78, Vol79]. ass [MKT77]. Assessment [Mag79]. associated [EL73, Pop71]. astronomical [Lay77]. Asymptotic [Kut78, Ra76a, BS71, CS78, DMM75, WB76]. Äthylenen [FJP70]. Atkins [Gos78b]. Atom [Nes75, OJB77, AM74, BK75a, CM72a, CSH74, CH74, CH75, DCS77, Goo76, JG72, KP76, Kle74b, KL77b, LA70, LKH77, MMS70, PS73, Pol79, RB77, Rei72a, Spe70, WC75a, WK72, WJC70, WAL71, vDR79]. atom-diatom [KL77b, LKH77]. Atom-molecule [Nes75]. atom-pair [KP76]. atomes [Mar73a]. Atomic [Fla79, GJP79, Lud77, SDS73, Thu75, WM79, Ada73, Ahl77a, BS74a, BS76a, BN74, BT71, BP75b, BEH77a, BEH77b, BL77d, BB78, CL76, ČP71, CM78, CA72, CA73, FKS76, Fri72, Fro74, GF76, GBV77, GK76, GL74, GRM75, HKM71, Hea71, Höj79a, Höj79b, IS75, JMM71, JS70, JNŽ72, JL74, Kar70a, KZ73, Lan77, Lar79, LG75, LAM71, MMM71, NN72, NN74, PK72, PD78b, Raf75, Rou79, RGM75, SK72, SPČ79, Sch72, SDD72, ST70, Sim72, SH75a, SHR76, Sla71a, SJH79, Sol71, SR72c, SW77, SW78, TGG79, TS71, TNB79, YJ76, JBG70]. atomic-integral [Rou79]. Atoms [Nes71, AZ72, Ave72, BM74, BK76, BW73, BB73, Cal72, Cal73c, Cal74, CP70, CW77, CL78, Csa75, Csa77b, Csa79b, Dah78, Dea72, DY73, DCS77, DDD70, GGJ75, GB75, GB72, Jak76d, Jam70, JBP76, JP73, Jun78, KM74a, KP72, Koc74, Koł74, Koł75, Lin73, Mar73a, Mil71, MS76b, Mor78, MI78, MDS75, MSM70, MBM71, NS72, Nes70, Nes77, NY76, OK75, PF74, Raf75, Rob71, Rob72b, RS79, RK76, SHO71, ST72, SDJ79, Sin73b, SJH79, Spe71b, Wel77a, Wel77b, Wen79, WRS70, WS79, YD75]. attachment [Nes77, SL79]. attack [HPK79, PLSL78]. attempt [BCR^{+}78]. attention [WV74]. attenuation [PP71]. attraction [LR77]. Attractive [GBTB78]. Attractive-dominant [GBTB78]. aufbau [Nov73]. Auger [TNB79]. augmented [ML76, WFS77]. augmented-plane-wave [WFS77]. auramine [BZSD79]. Author [Cal76a, Cal77b, Lar76, Lin73]. Authors [Cal73a, Sab75a]. auto [TGSL70]. auto-cohérente [TGSL70]. autocoherent [Gou73]. autodetaching [Sim78c]. autoionization [Naz72]. autoionizing [MCW78]. Automatic [BSZ78]. autonomiterque [BP79]. auxiliary [Gru70, LA76]. Auxin [BCL77]. average [KZ73]. Avery [Lin73]. Axial [Han79, BM71b, RS78b]. axis [Lét74]. aza [Sha78a]. aza-aromatic [Sha78a]. azulene [JL70].

B [Cal73b, Dun78, Sti79, CC79, HE77, HT75, Jak76b, JPTM78, Koł75, LH74, Lar75, Lar78a, Lar78b, MDS75, PPP79, PZP79, RR77, Rit76, SBD^{+}70]. B-DNA [PPP79, PZP79]. Back [LB79a]. backbone [SY77]. bacterial [KKKP79]. balance [KKP77]. Band [FRM74, PN71, AL71, ABD^{+}79, CD78, Con70, KKA78, Lad74a, Lad74b, Lad75a, Lad76, LSOC77, MGM78, Ove73, Pai73, PNJB78, Pet78, Sco71, SL73, Swi71, ZF76]. band-structure [Ove73]. bandmatrices [Wel71]. Bands [HP77a, Bor70, CZNL70, CD74, CD78, Csa78a, Dav72, DF76, Euw71, FN73, Lad70, Lip72]. barbiturates [AJ79]. barrier [CL77, MEW70, WC75b]. base [ENRM78, FB72, GRE^{+}77, KKK79, LL72, MW74a, Mar73a, PZP79, RCE75, Sun77]. base-pair [ENRM78]. based [AM74, DMS^{+}78, FRM74, HSBD78, KT77, Kau74b, Pau78, Sam75, SRP76, SRP77, TC70]. Basel [Cal76b]. bases [ANR75, Bri73, CLP79b, GJP79, KD77, KS79, KL79, LP79c, LSB^{+}79, MHPS70, Mos73, NKKY79, NY76, PH77, Raf75, Sey76, SRE^{+}79, SM79, WTB^{+}76]. Basic [Ano76e, BCL79, Spe71a, Hos72]. Basis [BL72, Kas76, Tan74, ÅSW77, Abs78, Amo71, BS75b, CA78, CHS77, Cur78, Dah78, DRCT78, DPGC79, ES79, Euw71, Fae79, FGV77, Fin73, HS70a, Höj79a, Höj79b, Hur77, KB77, KMO77, KPÖ76, LSOC77, Lec76a, LCK76, Lev78a, LGR73, LPT^{+}74, Mal71a, Mar73a, Mcd79, MYTC77, MDC79, NI74, OKM74, Pel77, PAM75, RPK78, SWN78, SBM72, SJ79, SC78b, SSH78, TB74, Van74, Wel76, Wel77b, Bur77, DLG79]. bath [HMV78]. bathochromic [MP71a]. be [SH75a, Bun77, GF76, Jak76b, RS79, SP75d, SBD^{+}70]. beam [SMMA73]. BeCH [RMB78]. BeF [SLJ78]. BeH [GG79a, Mul71a]. Behandlung [FJP70]. behavior [BG76, DMM75, Kut78, Tal77]. beim [KBA76]. bends [BCL79]. beneath [VH75]. benzaldehyde [LM72]. Benzene [SPČ72, AB73, HG75, KP76, NI70a, NI70b, NT79, PLMB73, San71]. benzene/H [HG75]. benzenes [MR72]. benzenoid [SG77a]. benzo [HKP79, KPPH78, LPP78, LP79c]. Benzol [HG75]. Benzol/H [HG75]. benzomorphans [RW79]. benzoquinone [MK72]. benzyl [KKM70, NN74]. benzyne [MPS70]. BeO [SBD^{+}70]. Berechnungen [BKW75]. Bergmann [Cal73b]. BERKELEY [BS78c]. Berlin [Cal73a, Cal79a, Cal77b, ZL78]. Berlin-Heidelberg-New [Cal77b]. Bernard [Shu77a]. Berthier [RCE79]. Beryllium [BS78b, GBV77, TTS71, Tat73]. berylliumlike [JL74]. Bessel [SW77, SW78]. Best [Jak76a, Jak76b, Jak76c, Jak76d]. Between [Gal72b, Sim77a, AMSS79, Bas78, Cas79, Cha71, CL78, Col79, Con70, Csa77b, Dah78, DGC77, DD72, Dea72, ED79, FM72, Fin73, GJP79, Hal71, HPK79, Höl77, Kib75b, KC74, Koc74, Koł74, Koł75, KL73, Kum77, LPP78, LB77, LR77, MM72a, NY76, PLYM78, PČL78, PW79, Rob71, SF76, SRE^{+}79, SDJ79, SE79, Sin73a, SKS77, SS78b, SP71, Suz77, SO77, TTS71, Tic79, WK72, WB76, WV74, WMV77, WP79, ZG79]. beyond [Wil74]. BH [LL72, BS74a, BS76a, EL73, GG79a, MHA70, Mul71b]. Bibliography [RWFS71, Sab75a]. bicyclic [BBG79]. bielectronic [Rue79]. Bifluoride [LS79]. big [KSP78]. bimolecular [Chr77]. binary [DV79]. Binding [LK78, SLJ78, CZNL70, CD74, CT79, CV73, DAC^{+}79, Eis77, GJLL75, Han79, KPD78, Lak75, NKKY79, NTN^{+}79, Pul74, PBG76, PPB78, PD79, RKE79, SS79, ST78, TNB79, UP71, WRF73, ZL78, KD77]. biocatalytic [JM76a]. Biochemistry [Pul78b, Sti79, BP70, Cal73b]. bioenergetics [Dav79]. biogenic [KM72]. biological [Bai74, BP73, BL79, Cho79, Chr79c, GH79, Hos78, MPMK79, Pau78, PBC74b, SJ79, SLOKP79, Tal77, Urr78, WCK^{+}76, WOEG78]. biologically [BBG79, DSMM79, Kla78]. biology [Cal76b, Com74, Hal79a, Mar73b, Pul78b, Rei70, Ros74, Sti79, Sun75a]. biomacromolecules [AMJS79, SAM79]. biomembrane [Urr77]. biomolecules [KPP78a, NS79a, PP74, Pul70, RC78]. Biophysical [Har75]. biopolymer [Kho79]. biopolymers [AMJS79, FB72, Lad76, LSS78, Pul78a, Sun74b, Sun75b, ST78, ST79b, Dun78]. Biorthonormal [Dah78, Lat77]. biphenylene [JPTM78]. bipolar [Sal73, SF75]. birefrigence [Loc72]. biréfringence [Loc72, DQ77]. birthday [Cal75b]. bisglycinato [RRS75]. bisglycinato-Cu [RRS75].

BISON [WBKL70]. Bloch [Kun78a, KOW78, Kun78b, Wan79]. blocking [KVSG79]. BN [SBD^{+}70, ZF77]. body [Bal78, BS74a, BS76a, BWS77, BP78, BFH78, CJCM73, DDD70, Feu77, KBA76, Lev78a, LR77, Lin78, LRJ73, MMM75, Ner75, NY75, Per78, Ros78b, SBM70]. Bogoliubov [Cli73]. Bond [Bur77, Del73, Gal72b, Par72, SS71b, Ada77, Bas78, DBBJ76, DMS^{+}78, GN77, Gin78, Hal75a, HR79, JJ78, KHI79, Kle79, Kle70b, LS72, LS79, MPS70, NG73, OK75, Pet77, PNC^{+}78, RTS74, RS76, Sri77, TRS73, TS77, TS78, TS79, TTN76, Wój76, Woj78, Lar76]. bond-length [TTN76]. bond-order [TTN76]. bonded [AM72b, BS75c, CP70, Cho79, Loc77c, TP77a, Wel77a]. bonding [AMC73, ACG74, Bai74, BL75, CP70, Cur77, DS76a, DMS^{+}78, ES73b, GH72b, Har71a, HR79, Meh79, Nir75, PDE79]. bonds [BVO72, BK75b, Coo79, Gin78, Loc77d, Sab71, SDD^{+}77]. Book [Ahl77a, Ano74c, Ano78c, Cal73a, Cal76a, Cal79b, Cal79a, Cal73b, Cal75c, Cal76b, Cal77b, Com77a, Com77b, Dun78, Gos78b, Gos78a, Lar76, Lar78c, Lin73, Mår78, Sab75a, Shu77b, Shu77a, Sti79]. Books [Ano70d, Ano70e]. borides [Mat77]. Born [CL78, Col79, Ess77, WN71]. boron [LH74, MVN75, PKK75, PK77a]. Bose [MBMM79, MPMK79]. Boston [Mår78]. both [Koe74]. bound [Bur77, ČP77, Dmi75b, GC75, JP77, KSP78, LTR74, MKT77]. boundaries [SP75b]. boundary [WO75]. Bounds [Abd71a, Abd71b, Bur73, LL70, TL79, BR75a, BBK76, BR72, Gra77, JW76, JP73, Lin70, RE70, Vog79, WW71]. Boys [Par73]. Brackets [Wel70, Wel71]. brain [GT75]. Branch [CB72, BF78]. branched [KHI79]. Branching [Ber75, Pau77, BT78b]. Braun [WW71]. breakdown [Ada75b]. Breit [RL77]. Breit-type [RL77]. Brillouin [Ahl70, SC76a, Ahl70, Gra77, LWA74, RCE79]. broadened [BMD70]. Brueckner [LH76]. building [Kau70a]. bulk [SEW73]. Buridan [MKT77]. Burton [Cal75c]. butadiene [TS78]. butatriene [TS79]. Butterworths [Mår78]. bytes [Spa78].

C [Com77a, Com77b, SBD^{+}70, BDT77, CV73, ETF79, ED79, GG79a, GBT79, HLYC77, HT75, Jak76b, KP71b, NN72, Ram78, Ros78a, SMP77, Tru73a]. C-C [SMP77]. Ca [KD79b]. cage [Ave75]. Calais [Shu77b]. calcium [AD77, LB79b, NTN^{+}79]. Calculated [KW76, AB73, BL77c, LVDC74, Sin73a, WM79]. calculating [Gin78, HS77, ML76, Nam75, PLYM78, Spe73]. Calculation [BFH74, BW76b, But76, BW76c, Cas79, Con72, FB76, HH77, HNT78, KPÖ76, Li76, LP74c, LA76, LW76, Loc77a, Loc77b, LL70, Mon77, RR72, SK71, SP75a, SLB77, SH74, Sve79, ANR75, Ahl70, AD77, AFD^{+}78, BNLS72, BN76, BSM72, BM73, BS75c, CZNL70, CJ73a, CSFM72, CW77, Chi77, ČP72, CL78, CWE71, CE73, CAMC71, CH74, CH75, DRCT78, DD72, Del74, Ema74, FRM74, GZ72a, GF76, Goł79, GML77, GBM74, GRM75, Hal73, Har71c, Hos72, ICS74, JR74, Joh75, JPTM78, JS70, JNŽ72, KR71, KTO^{+}78, KBA76, KA76, KT76, LD73, LL75, LGP75, Lin77, Loc76, Loc77d, MCST72, Mes73a, MM72b, MH72, Muf76, MSM70, MBM71, Muk74, MMM75, MS71c, Ner75, NB78, Nir75, NY76, PPK78, Pie71, PK71, PK76a, RBDV77]. calculation [RGM77, RPK78, SK72, Sel75, SLC78, Sin73b, SP75d, SSH78, Tru73b, WC75b, Wei71, WFS77, YL74, BCL79, FG79, MMM71, NS73]. Calculational [CM72a]. Calculations [Ave79b, AB79, Ave79a, BL77b, BB78, Eld73, Gal72b, Har71b, JE79, LCP76, LZ77, Mak71, PNJB78, PK76b, Sas74, Abs72, Alm74, Als73, ABD^{+}79, Ano78c, ACD^{+}74, Ave79c, Bai74, Bar70, BL77a, BBP71, BBL77, BB72c, Bil72, BKW75, BEH77a, Bri73, BB73, BWRD79, CP70, ČZ79, Chr73, Cla70, CM76, Csi76, Cur77, Dan72, DCM74, DMR74, Dau74, DDH78, DEE^{+}75, DK77, ESR74, EH73, ED74, FTN77, FBL79, FB72, FMZ70, FL79, Fro74, Gar78, GG79a, GG79c, GG79b, GV77, GBS70, GPP73, GBV77, GMG75, GH72b, GJ76, GHJ77, Hal75a, HS76, Han79, HPK79, Har70b, HLGC73, HSBD78, HPK72, HU74, HCIE75, HS70a, HS70b, Höj79a, Höj79b, Höl77, HP77b, Ive74, JW76, JŻ79a, Jor73, JL70, Jun78, KL78, KS73a, KT77, KTO^{+}78, KK72]. calculations [KPKC73, Kau73, KK76a, KK77a, KPP77, KSP78, KPP78a, KPPH78, KPP78b, Kau79b, Kel75, KP71b, KM74b, Kle70b, KD77, KD79b, KKK79, KL79, LD74, Lan77, Lar75, LFF79, LAV72, Lev78a, Lev70, LKH77, Loc77c, LH79, LJBN75, LS74b, Lun79, MMB70, MW74a, MEW70, MSKM77, Meh79, MLB76, Mey71, MLB72, MR72, Moh77, MMS70, MDA78, MM70, NN72, NN74, Nee75, NAR72, Nis77, NMS77, NS78, OMC76, Ove73, Pai73, PPC72, PCMS77, PS70, PK75, PK77a, PK77b, PGPP79, PLMB73, PK73, PKK75, PK77c, Pul70, RS76, RTM77, RB77, RU76, RKOU77, Røe75, RL79, RS79, RK77, Ros78b, Rou79, San72, SM73a, SD78, SN79, Sas77, Sha77a, Sha78a, Sim72, SJ79, SDB74, SC78b, Spa78, SF75, TRS73, TS77, TS78, TS79, TP78]. calculations [TL71, Tru73a, Van72, Vla70, WJ79, WMB75, WH75, Wei77, WN78, Wel77a, WJC70, Wit72, Wol71, WMV77, YG77, DPGC79, LH77, May74a, NS72, SDS73, Wes79]. calculus [Pal75]. calibrate [Kau73]. CAMSEQ [WH75]. can [Fri72, Hal79a]. cancer [Abd76, DEC78, LL76, SG76a, SG77b]. candidate [LPP78]. Canonical [Wit72, BW76a, LD71, Lat76, Shu77c]. canonically [Wes73]. carbenes [Har71c]. Carbon [NTN^{+}79, BWS77, DPGC79, PGPP79, SD70, WK72, Wil77b]. Carbon-13 [NTN^{+}79]. carbonic [PD79, SS79, SS74]. carbonyl [KJ75, ST73, VG78]. carboranes [PK77a]. carboxy [RL79]. carboxylic [KPP78b]. carcinogen [HPK79, LPP78]. carcinogenesis [BS79, Dau77, Her74, Kau74a, Kau79b, LSS78, LSB^{+}79, Löw77, Mem75, Mem77, Mem79, Pul79a, Ros75, TPC77]. carcinogenic [HKP79, KPPH78]. Carcinogenicity [SR78, SS78b]. carcinogens [HPK79, KPP78a, NKKY79, Nee75]. carcinoma [WLS77]. career [Ano76n]. Carlo [Cas78b, Col77a, CL78, Col79, RC78]. Cartesian [BM71a, GM73a, KOW77, SC78b]. case [APČ77, BS78b, Dmi75a, Dmi75b, Kib75b, MK75b, NRS79, PAČ77, ST79a, DRCT78]. casino [Fox75]. catalysis [DS76a, Gol70, Sal79]. catalysts [KGB76]. catalytic [JM76a, BCR^{+}78]. catalyzed [LDP79, PDEB78, WEH74]. catastrophic [KKP77]. catechol [SS78a]. catechols [SS78a]. Cation [PBG76, Hin77, LB79b, MEW70, NI70b, PGDM77, SLOKP79, WO77, LB79b]. Cation-ligand [PBG76, LB79b, PGDM77]. Cation-water [LB79b]. cations [KR73b, PPB78, Urr77]. CC [BVO72]. cell [Jeh73, MHH74]. cells [Tal77]. cellular [FRM74, LFF79, Kel75]. Center [Eng72, Thu75, AM74, AT75, BA78, CM76, HR79, HNT78, Jet73, LPT^{+}74, Mag79, MH72, PAM75, SR72b, Sha76, ST70, SBM72, MM74]. centered [MVM75, MLM75, Mil74, SDM74]. centers [MVM75, Fan70]. centre [AM72a, AM77b, AM77a, KZ73]. CEPA [Bur77]. CEPA-PNO [Bur77]. certain [BL79]. CF [Cur77]. CH [Cur77, FJP70, JŻ79a, JR74, Cur77, FMZ70, FJP70, GG79a, GV77, HPK79, HE77, KC72, MHA70, RMB78, WK72]. chain [BBA73, DDH78, KKA78, KG77, LR77, MOV75, TGSL70]. chaine [TGSL70]. chains [KG76, KG77, Ukr72]. chalcogenides [SKS77]. chalcone [BTD79]. champ [Gou73]. change [Amo72, HY78]. changes [BZSD79]. channel [MM72a]. channels [GRB74, SLOKP79, USK^{+}79, VU79]. character [GK78, KE79, NT79]. Characteristics [ÖL78, ÖL79a, Sla79, WCK^{+}76]. characterization [BT78b, Chr73, Har77b, Jan76, USK^{+}79]. characters [TTS^{+}79]. Charge [CM75b, KPP78b, SRE^{+}79, Sin75, ZL78, BT71, BKW75, BW73, Cho79, Csa77a, DC73, DY76, DT75, HRBB76, JMP78, Kar70a, Kar70b, Kar70c, KFY71, Kau73, KKA78, KKKP79, Kib75b, Lad76, LH77, OMC76, OSL77, Pai73, PK75, PS75a, RM72, Sch73a, Sha77a, SLC78, Van72]. charge-conserving [PK75]. charge-dependent [Kar70a, Kar70b, Kar70c, KFY71]. charge-transfer [Van72]. charged [Csa79a, NG73]. Charges [Hal75a, BT71]. Chemical [FL79, Gos78a, GJLL75, SLOKP79, BS79, BCL77, CSR79, Chr79c, ČP71, Dau77, ES73a, Ema74, HG75, INY^{+}71, Jan76, JJ78, KK72, Kau74a, KKK74, KK75, KK76a, Kau79a, Kau79b, LSS78, Lar76, LPP78, LP79a, LJBN75, MPN79, Mos73, NKKY79, Pul79a, Sal79, SDD^{+}77, ST73, SJ79, TP77a, TWC73, TPC77, Wei75]. chemicophysical [Kau76]. chemin [Lei74]. Chemisorption [DS76a, Sch77c]. Chemistry [Cal73b, Lar78c, Ost79, Pul78b, Sti79, And79, BP70, Cal79a, Cas78b, Csi77, DAC^{+}79, GM73a, GM73b, Gos78b, Hal75a, Hal79a, DV79, KG76, KG77, KB77, Mat72a, MF73, MSP73, MCP73, MPW75, MW75a, Mat76a, Mat76b, MWY77, MW77, MK78, Par77, PHL^{+}76, PS75a, Ran79, SP75c, Sin73c, Sla79, Wil77a, Wyb73a, MN79, Cal75c, Lar78c]. chimique [JJ78]. chloramphenicol [BSB77]. chloranil [DT75]. chloranil/para [DT75]. chloranil/para-phenylenediamine [DT75]. chloro [GBS70]. chloro-substituted [GBS70]. chlorofluoromethanes [PK77b]. chlorophyll [Mag79]. chlorophyllide [OMC76]. Chlorophylls [SCYB77]. chlorpromazine [PPK78, PK76a]. choice [SP79]. choosing [HC72]. chromophores [Kli79]. Chrysene [BS79]. CI [Hil79, APČ77, BNLS72, BS78b, KT77, KSP78, KPÖ76, KMLB77, KL79, Lar78a, MW74a, MLB76, ML79, Mey71, NRS79, PAČ77, PCMS77, SH75a, SHR76, Sin73b, WJ79, WMB75]. CIBS [Mcd79]. Circular [Tin79, BCL79, Dub72, MSSF79]. circumvent [LAV72]. cis} [TS78, WMB75]. cis}-butadiene [TS78]. Cl [Cur78, FM78a, MMR74, FM78a, PK76b]. class [KK74a]. Classical [FM78a, Dub72, Sch71c, Sch76b, WO74]. classification [KR76, MCM73, MLM75]. Clebsch [KG76, KG77]. Cliffs [Cal76a]. ClO [FM78a, GH72b]. Closed [Hil79, ČP72, DK77, Gor71, Gor72, JBP76, KC77, Lai73, MMM75, SH73, Sin75, Spe71b, SO77, PAČ77, SS73a]. Closed-shell [Hil79, ČP72, DK77, JBP76, KC77, Lai73, MMM75, Spe71b, SO77, SS73a]. closure [MMD78]. Cluster [Wes73, FCL79, GbK77a, Har77c, Kle77, KGB77, LH76, Lin78, Mat79a, Mon77, Pai73, Rob71, Rob72b, SPČ79, SLC78, Živ77]. cluster-type [SLC78]. clusters [HP77a, Muf76, NY75, Sin73b, Sla79]. CN [GML77, KJ78]. CNDO [CS74, DMS^{+}78, Duk79, Kau70b, KPKC73, SL73, SSH78]. CNDO/2 [DMS^{+}78, KPKC73, SL73]. CNDO/S3 [Duk79]. CNS [BWRD79, KK75]. CNS-active [BWRD79]. CO [LB79a, FM72, Eis77, LK78, PD78a, PD79, RC72, SBD^{+}70, Tru72, VG78, KTO^{+}78]. co-ordination [FM72]. Co-porphine [KTO^{+}78]. Coarse [MF73]. cobalt [KTO^{+}78]. code [Coo78a, FM79]. coefficient [DH73, JW78]. coefficients [Ahl74, Bar79, DH78, KG76, KG77, LV71a, SD79b]. coherence [Kle74b]. Coherent [BK76, Bow72, DMM75, FM78b, Nis77, SK72]. cohérente [TGSL70]. Cohesive [Cal75a, SP75d, ZF77]. coil [Por77, Por78]. collagen [CBV79]. collection [WEH74]. Collective [Wen79, MOV75]. Collinear [vDR79]. Collision [Chr77, Fla79, Kle74b, MFRE76, vDR79]. collisions [Ano79o, Ber75, BK75a, CL76, DLG79, MM72a, Mic74, Nes75, RM74, Tru76]. Colloquium [Sim79]. combination [Com77c]. combinations [Sol71]. Combined [SH75a, SHR76, ZF76, Nes75, RCE79]. Comment [Coo78a, Nov73, Pos70, WRS70, Amo71, HMC71, Sab71]. Comments [Bar77, KSL76, Kle74a, CR72, LB79a]. commonality [Kau74a]. comorosan [Bas75]. compact [BW76b, But76, BW76c, KE79, SH75a]. Company [Ahl77a, Lar78c, Sti79, Ano78c]. Comparative [Lec76a, PLYM78, TGG79, Ada78, KL78, Muk74, PD77, TWC73, ABM78, BB74]. Comparison [KPKC73, Kib75b, Lun79, MCST72, SRP76, SS70, Sla75, Con70, Gin78, Har77b, MK76a, Min79, ML76, PSK77, Tho76, Con72, Fau71, May78, OMC76, RL79, Sch76b, SRP77, SC78b]. Comparisons [Lar78a]. Compatibility [Fan70]. complementarity [RCE75]. complete [Aba72, CLP79a, CLP79b, Sol73]. completely [Sal73]. Completeness [KB77, Dal79, KK76b, LCK76]. Complex [Bar78, CHH^{+}78, Com78, Doo78, Hen76, Jun78, MCW78, Bar70, BFH78, CG78, DT75, FHB77, Hen74, Hen77, Hil79, Joh73, JG72, KTO^{+}78, Kib75a, MHH74, NI70a, NI70b, NB78, Nut78, PLYM78, Rei76, RRS75, Sim78a, SP71, Sun74b, Sun75b]. Complex-coordinate [MCW78]. complexes [AMC73, BR78, BV73, BVS76, CSFM72, ENRM78, HČZ79, INY^{+}71, JW72, LL72, Lew78, LH77, Loc77c, Loc77d, LH79, MCM73, MLM75, Pet78, SC78a, Van72, WO77, Wel77a]. complicated [PP71]. component [LD71]. components [Boy70a, GC77, Har79a, Pul74, SMB^{+}79]. Composition [RK79]. compound [LKH77]. compounds [BS77a, BP70, FM72, Gin78, KKK70, KPP78b, KJ75, Kot76, MYTC77, PN71, PKK75]. Compton [BY76, Dec70, Kai77, MBM73a]. Computation [ADFL71, DD77a, GMG75, KPJM70, KK76c, Spe72b, LF77, ML79, SBB79]. Computational [RW77, ADDLG76, Ano79o, BDT77, Duf71]. Computations [SK73, BP74, BP75a, Ber76, BCR^{+}78, Bur77, GCP79, Pul77, TK78, TG76, YD75]. Computed [Bor74]. computer [BF71]. computers [Ost79]. computing [SB78]. Concentration [Fog71, BF75]. concept [KK77b, SW79]. Concepts [LH74, Atk74, Atk91, Sha77b, Gos78b]. concerned [Sch78]. concerning [Jeh73, Mem77, Sab71]. concerted [TC70]. condensation [MBMM79, MPMK79]. condensed [Chr79c, Sti78]. condensed-phase [Sti78]. condition [CM72b, Kij72]. Conditional [Hun74, Hun75a, Hun75b]. Conditions [Kho79, Erd78a, Kum77, Lai73, RT77, WO75, WZ70, Fan70]. Condon [Sla73b, Lon73, Nis77]. conductance [VU79]. Conduction [Lew78, Lad74b, Pet78, SKS77]. conductivity [BK75b, FB72]. Conference [Cal76b, Ano78d]. Configuration [CS76, Fro70, HPK72, LD74, PW79, Sas74, Spe72b, TS71, TTS^{+}79, WP79, YW73, BS77b, BS78a, BS78c, ČPH74, Cof74b, DD72, FTN77, GHS79, Hac74, HY78, HT75, JG72, JSK75a, JSK75b, KR73a, Kau74b, Kau79b, KKM70, LED73, MM70, NI70a, NI70b, PSK77, Røe75, SD79a, Sas77, SRP76, SRP77, SDJ79, SDB76, SBD^{+}70, Tat73, Tru73b, Wei76, WV74]. Configuration-Interaction [TTS^{+}79, BS78c, SD79a, SDJ79, Tru73b]. Configurations [ENRM78, Cha77, Csa77b, DN75, Dav74, KZ73, MS71a, Muk75, PH77, SM71a, SM71b, SM71c, YW73]. confined [Wei76]. Confirmation [Ram78]. Conformation [AMC73, Kan74, CC79, DS79, FSV79, KD79b, LL72, LCPP79, MGT79, PPC72, PGPP79, PS79b, Pul70, PS75b, PC76, RRNS75, BP73, RRNS75, SMP77]. conformation-energy [KD79b]. Conformational [AJ79, AMJS79, BZSD79, BWRD79, IZZ^{+}79, KK77a, KP79a, KVSG79, ME78, MDA78, RRS75, RU76, RSK^{+}78, Rou73, Rou74, SD77, Sri77, VU79, WTB^{+}76, Wei77, WN78, HPW79, JE79, KM72, PPC72, PBC74b, RKOU77, Sun74a, VGS79a, KL77a]. conformations [DSMM79, Har75, Kar78, Pop79, Pul76, SY77, ULM76, Shu77a]. conformer [Bur77]. congeners [WOEG78]. conjugate [RE70, RPK78]. Conjugated [LL73, BVO72, DM72, ICS74, MCST72, Muk74, NMS77, NS78, SN79, SG77a, TTN76, TSE76]. conjugation [MOV75]. Connection [SF76, LP74a, Tic79]. conservation [BBT71]. conserving [PK75]. considerations [HBB77]. considering [HPW79, ST79a]. consistence [KS75]. consistency [BS77b, BS78a]. Consistent [ÖL78, Sla70a, Abs72, BKS79, BBP71, Bor70, BK75c, CWE71, DS74, Gou73, Hac74, Hal78, JL74, Kar70b, Kar70c, KFY71, KKW73, Kle70b, LG77, Meh78, MM72b, MSM70, MBM71, MMM71, ÖL79a, Sco74, Sel75, SLC78, TGSL70, Vla70, VK74, ZF76]. constant [AB73, BR72, IM73, KL77c, Men71, PGPP79, SEW73]. Constants [TPM70, Csa76, Csa78b, GPP73, JE79, LZ77, Laz79, LG71, Mak71, SP75d, TB74, Tic79]. constituent [KR76, Kla78]. constituents [PGPP79]. constrained [Mor78, MI78, NI74]. constraint [Sch78, CM72b]. constraints [Mor75]. constructed [Ra76b]. constructing [PW74a]. construction [Blo72, Blo73, Jor75, MS75, Pau77, Sab73, MMB70]. contact [HB79, Sim78b]. contacts [Jeh73]. contained [Sur71]. containing [DCS77, Mar79, MM74, SN79, TK78, WRF73]. content [HL78a]. continuation [Jor75]. continuity [MK76b]. Continuous [Eng71, FHB77, WEH74]. Continuum [CG79a, Dmi75b, Nir75, RM79]. contour [HKP79, Kau73, SBB79]. contracted [OK75]. contraction [Dav79]. contribute [Hal79a]. Contribution [Pul77, CG79a, Men71, KKK70, WN71]. Contributions [GBT79, GTBB77, KP76, Par73]. control [BCL77, KKP77]. convenient [Jet73]. Convergence [Ahl77b, BG76, BK71, DSY74, Sal73, AM77b, AM77a, CJS77, Csa77a, Euw71, FP77, JSC78, KS75, LAV72, TRPB77, Ahl70]. convergent [PW74a]. converging [SH73]. conversion [Plo79]. convex [Col77b]. convulsant [AJ79]. CoO [WJ79]. Coordinate [Spe72a, AMC73, Jun78, LV78, MCW78, Nal78, Spe71a, Spe71b, SP79]. coordinated [JMP78]. coordinates [CHH^{+}78, KOW77, Rei76, WRS70]. coordination [RBDV77]. Coplanar [GMZ73]. copolyamino [SWF79]. Core [CE73, SHI73, BFS73, KZ73, San72, TP78, YD75]. core-electron [BFS73]. correct [SK71]. correction [Csa79b, Han77a, RB75b, YD75]. Corrections [BMM75, CFHS77, CT77, DC73, Mal71a, Mal71b, PT78, PPI73, YJ76]. correlated [BS70, GL74, Hen78, KM74a, LG75, Meh73a, Meh73b, Pet77, PNC^{+}78, Rob71, Rob72b, Sim72, WV74]. correlating [SBD^{+}70]. Correlation [BFS73, BS74a, BS76a, BL77d, ČP71, ČPH74, ČPP75, Com77a, GTBB77, Nes71, Nes77, PČH74, SPČ79, Spe72a, Aro78, BB70, BBB73, BP78, BN74, BS76b, CEK72, FU76, Fro75, GL74, Hur76a, Hur77, JMM71, JMP79, JŻ79a, eAJ76, KM74a, Kau70a, KP78, KKK79, Lan73, LG77, LG75, Lud71, MW75a, MVN75, MMM75, Nes70, Ove73, Pal75, PT78, Pet77, PBS76, PKSB78, Res78, Rob71, Rob72b, Rue79, SDD72, Sha77b, Sha78b, SJH79, Spe70, SL70, SMW71, SO77, Tat76, WK72, LB73b, RKOU77]. correlations [BC75, Ran78, Sin73b, SS78b]. corresponding [Bur72, RC78]. Cosmic [Fox75]. Cotton [AB73]. Coulomb [AM77b, BK74, Eng72, KKK79, LH74, Lar78b, LC77, Sil71, Sur71, TL79]. Coulombic [AFD^{+}78, PD78a]. couplage [TGSL70]. Coupled [BSM72, BM73, GBM74, GRM75, Har77c, HDG71, MM72a, RGM77, BP78, BMM75, ČP71, HMV77, HMV78, HBE78, Hun74, JS77, Lin78, Mat79a, Mon77, SPČ79, Vog79, Živ77]. Coupled-channel [MM72a]. Coupled-cluster [Har77c, Lin78, Mat79a, Mon77, SPČ79, Živ77]. coupled-pair [BP78, ČP71]. Coupling [GÖ68, HL79, BOP76, BBA73, GPP73, Hal75b, IMN79, JE79, KK74b, KK77c, Laz79, LRD^{+}77, LRD^{+}78, Mak71, Men71, Shu77c, TGSL70, TB74, GÖJ^{+}73]. courte [JM76b, TBB70]. covalent [Ada77, FCL79]. CR [BFH74]. CrF [KL78]. Crick [Sun77]. Cristal [DQ77]. cristallins [JM76b]. criterion [JTR77, LCK76]. Critical [BG70, Hub71, BBK76, BS78b, ČZ79]. cross [Bar79, BG78, Chi77, Fae79, ML76, SSM75]. cross-sections [ML76]. crossing [NB72, Pop71]. crossings [BR75b, BF78]. Crystal [Ave79b, AB79, Ave79a, Bal71, Loc76, SW71, WN71, AD77, Ave79c, Bow71, Ell73, HP77a, LFF79, MME79, MCM73, RW74, RT77, SLC78, SW70, Sol73, SS77b, Wol71]. crystal-field [RW74]. crystalline [BL70, CKI75, Euw71, JM76b, Kar78, VK74]. Crystals [Jul78, Sla70a, BF71, Cal75a, DS70a, Dav72, DF76, GMG75, HKM71, JMP78, Kho79, Kun78a, KOW78, Kun78b, Lip72, Loc77b, MCM73, Sch71d, Spe73, SBM70, TW77, Wój76, Fan70, Cal79a]. Csizmadia [Ano78c, Lar78c]. Cu [KD77, RRS75]. Cubic [KOW77, Kun78a, KOW78, Kun78b, SS77b]. CuCl [FWJ^{+}79, KWC79]. cumulene [Ukr72]. current [Löw77]. curve [Liu71a, MVN75]. Curves [HL79, Gub79, GBK77b, Jor75, May78, NS73, PK71, Wil77b]. cusp [CM72b]. cut [LC77]. cut-off [LC77]. Cyanide [Mof79]. cyanogen [Mof79]. cyclic [Bur77, Kar78, LV71b, LV72, ME78, Ra76a, Sun74b, Sun75b, ST78, ST79b]. Cyclo [RSK^{+}78]. cyclobutane [KC79]. cyclohexapeptide [RSK^{+}78]. cyclopropane [LG77]. cysteinyl [PPC72]. cyt [Han79]. cytochrome [Har77a, LHK77, PL79, RL79]. cytosine [CLP79b, KD77, PZP79, ZR78].

D [AS78, Cal73a, Cal73b, Sti79, AC70, AJK78, MS71b, Rit76, Ros78a, TP77b, Wei76]. D- [AJK78]. Dancoff [HU74, Lar78a]. dans [DQ77, TGSL70, TBB70]. Darboux [BG72]. data [CFMO73, FKS76, SKW72, WEH74, Ahl77a]. dC [PPP79]. decade [Van71]. decay [BYB79, HLMV77, HF72]. decaying [HMV77, HMV78]. decomposition [Jør77, KM76]. Decoupled [NS72, NS73]. Decoupling [WZ70]. defects [FCL79, Lar75, PBPL78, Rou78, SC79]. defining [Fin73]. definite [KR73a, KR76]. definition [Wil79b]. Definitions [Spe71a]. Deformation [Ave79b, AB79, Ave79a, Ave79c]. degeneracies [LG71]. degeneracy [Eng71]. degenerate [BEF78, BY76, Csa78a, Dmi75a, Dmi75b, IS75]. délocalisation [JM76b]. delocalization [EL73, JM76b, KPP78b]. delocalized [Loc77d]. delta [CB72]. delta-function [CB72]. demonstrating [RU76]. demonstration [RR72]. Densities [BHM75, BW73, CG79b, Hen78, HYL78, Lim75a, NN72, Pai73, Pau78, RR72, Sch73a, Sha77a, Sha78a, Sin75, SBM70]. Density [Ave79b, AB79, Ave79a, Gar78, HS79, Har71b, KL75, LS71, LL70, MW75b, SLM70, SJH79, Spe72a, Spe72b, SS77a, Ada72a, Ada79, AM72b, Ano78d, Ave79c, BG75, Bun77, CM72b, Cli73, CFMO73, Col78, CH77, CESW70, Csa77a, CM75b, DC73, Del74, DD77a, DR75, FM78b, GF76, GH78, Gás71, GY75, GP79, GJLL75, GJ76, Han75, Har70b, Har77b, Har79a, ICS74, Ive74, Kau73, KKA78, KP71b, Kry77, Kry78, KR73b, LS70, Lar73, Lar78a, Lar78b, Lar79, MSL73, Pea74b, PS75a, Ros78b, SP75c, SLC78, Spe71a, Spe71b, SC73, SMW71, TB78, TW77, YL74, ZF76, ZF77, KOW78, Kun78b, SDB76]. density-matrix [Cli73]. dépendance [TBB70]. Dependence [Bow71, DRCT78, FN73, MDC79, AZ72, Csa75, Fog71, HF72, KMLB77, MP71b, NB78, Nor75, RS78a, RS78b, RS78c, Sco71, Tal77, Tan74, TBB70]. Dependent [MK78, Sch71e, Ahl74, AMSS79, Bre73, Cli73, CESW70, CEK72, Gol73, Han70b, Har77a, HU74, JL70, Kar70a, Kar70b, Kar70c, KFY71, Moc73, RGM75, Sch71b, Sch71c, Sch71d, SM73b, VU79, WA71, Jam70]. depolarization [Urr77]. deprotonation [HL78b]. depsipeptides [Pul74]. derivation [Ada75b, Man77]. Derivative [PKSB79, Col79, PW74a]. derivatives [BSZ78, BS79, MPN79, SB75]. derived [RE70]. Desargues [Ran79]. describing [CSDV78, Gos75, PP79]. Description [Koc74, Ada77, Car75, ESG79, GSE78, Kau74b, KN79, LH74, Lar78b, SM77]. descriptions [FM72]. design [Kau79a, ME78, WH75]. detachment [SL79]. detected [MSSF79]. Determination [BS77b, BS78a, BCL79, Boy70a, HE77, JW72, JMP78, KZ73, OJB77, Ben72, LP74b, Pal73, RK77, WB73, BJ73, BJ73]. determining [CD74, CSFM72, MCF^{+}70, MCM73, Sol73]. deux [TBB70]. Development [SC78b]. developments [Mes73b, Mic74, Mic76b, Pai73, PHL^{+}76, Pul79a, Sal79, Tri77]. dG [PPP79]. d’hydrogène [Lei74]. diagonal [Erd78a, ST79a]. Diagram [KE79, Man77, Pau77]. Diagrammatic [SL70, Wil77b, BS75a, WS79]. diagrams [PČH74]. diamagnetic [Ema74, GZ72a, GZ72b, Hal73]. diamagnetism [SM73b]. diamond [SEW73, WESW73, ZF77]. dianion [MW74a]. diatom [KL77b, LKH77]. Diatomic [Jai70, AM77b, AM77a, BR75b, BFR78, Dav72, EB74, FW75, Gia76, Gin78, GJ76, Hal75b, HE77, Hen79, KD79a, Li76, Nal78, NS73, PK75, SM73a, SS71b, SBD^{+}70, WO76, YG77, vDR79]. Diatomics [NS79b, Dun78, Moh77, Pul78a]. Diatomics-in-molecules [NS79b]. diatoms [CB72]. dibenzo [CT79]. dibenzo-p-dioxins [CT79]. dibenzofurans [CT79]. dichroism [BCL79, Dub72, MSSF79, Tin79]. dielectric [Bre73, Bre75, Csa76, Csa78b]. diene [Par72]. Difference [Fro79, FW75, FU76, HPK79, TTS71, WK72]. differences [Hal71, JS70, Sim77a]. different [CL73, HMV77, Kut78, MK76a, May78, SRP76, SRP77, TB74]. Differential [HTG77a, BKW75, HBE78, PK75, SSM75]. Differentiation [AJK78]. Difficulties [ML79, BK71, LAV72]. diffraction [CFMO73, FM78b, MPW75]. diffusion [Tal77, WRF73]. difluoride [RTM77]. dihydrides [LB73a]. dihydrodiol [LPP78]. dihydrouracil [GJP79]. dilated [Bal78]. dilemma [Kat72, KD74]. dimension [Ada72b]. dimensional [CK73, FY79, RW75, Red79]. dimer [ACG74, BS78b, Dau74, MO77, MHW78]. dimerization [EL73]. dimers [Sch71c, WZ70, Cur77, Wel77a]. dimethyl [LCPP79]. dimethylphosphate [PGPP79]. dinitrogen [Hal79b]. diol [LP79a, LP79c, PDE79]. dioxins [CT79]. dioxygen [RBDV77]. dipalmitoyl [VH75]. dipeptide [PS79b]. Dipole [CAMC71, DD71, SBB79, Als73, BSM72, ETF79, GJP79, GBM74, HH77, KMT77, Kob78, LED73, LD73, Lar78a, LCP76, Li76, MCST72, RGM77, SLJ78, Sin73a, TSE76, Wei76]. dipole-dipole [LED73, LD73]. Dipole-optimized [SBB79]. Dirac [ERW75, GY75, Goo76, KL77c, YG77]. Direct [Cal79c, ICS74, KS73a, KBK72, Pal73, Wel75, Wel76, Wel77b, YL74, BJ73, ČP72, Col79, GHS79, PPB78, Jak76c]. directe [BJ73]. directed [BKW75]. direction [LW77]. directional [Kai77]. directions [DM72]. dirheniumoctachloride [MMR74]. Discontinuous [SWP77, HHR70, Lud75]. Discovery [MS71b]. discrete [CS78, Pai71, ZF76]. discretization [RM79]. Discussion [Kau74a, CS73, Feu77, WN71, Pos70]. diseases [ULM76]. dislocations [De 78]. disordered [Fau71, Fau76, Oni79, Sto71]. disorders [Kau76]. Dispersion [Ahl74, BHM75, SE79, Abd71a, Abd71b, BR72, DY73, Koc74, Lan74, Liu71b, MP71b]. Display [NGR^{+}77, WS75]. Dissipative [Pri75, GP74]. dissociation [BS77a, CSR79, PT78]. dissociative [Gub79]. dissolved [AMJS79]. distance [JM76b, TBB70]. distances [CODM77, Mul74]. distinguishable [CSY^{+}74]. distorted [Ada73]. distortion [INI71]. distribution [AM72b, Ben72, BT71, GH72a, Hea71, JMP78, LS72, OMC76, WESW73, ZL78]. distributions [BS70, Ber76, BKW75, DD77a, GN77, Gás71, HRBB76, KR73b, Tru76]. disulfide [Boy74]. dithienyls [AeAs78]. diuretics [SS74]. divalent [AD77]. Divergence [You75]. divers [JM76b]. diverse [Urr78]. diverses [eAJ76]. DMP [LCPP79]. DNA [Abd76, ANR75, CC79, Coo79, ESR74, FSV79, INY^{+}71, KGP79, KKK79, KS79, Lad70, Lad74b, Lad75b, LSOC77, MGT79, PP79, PPP79, Por77, Por78, PZP79, RKE79, Sey76, SR78, Vol79]. DNA-dye [INY^{+}71]. DNA-type [KGP79]. do [PS75a]. DODS [KU70, MK76a, RJ73, WV74]. dominant [GBTB78]. donation [LB79a]. donor [DBBJ76, ZV79]. d’optimisation [BPTLVM73]. Dordrecht [Sti79]. double [BS77b, BS78a, BV73, BVS76, CK73, CC79, GG79a, GG79c, IZZ^{+}79, KP79a, Kle70b, KMLB77, LG77, Lim77a, Ros78a, TP77b]. double-helical [IZZ^{+}79, KP79a]. double-well [CK73]. double-zeta [GG79a, GG79c]. doublet [BMD70, Hal75b]. doublets [Bow71, GHS79]. doubly [AS78]. Dreikörperproblem [KBA76]. driven [FMRG75]. Drug [Ano76e, Dav77b, DSMM79, Kau79a, KVSG79, ME78, Wei75, WO77, WH75, Ano76e]. Drug-Enzyme [Ano76e]. Drug-Receptor [Ano76e]. drugs [BSB77, DH78, Gab76, KK72, KM72, KK74a, KK75, KPP78a, Wei77]. dT [PPP79]. Dual [NI74]. Dual-basis [NI74]. due [HK74, ML79]. d’une [Mar73a]. during [CSR79]. dye [INY^{+}71]. dyes [BTD79, FTN77]. Dynamic [CL77, LP74c, Ahl77b, HS77, KKP77, Moc74, MMS70, MM72b, MSSF79]. Dynamical [Lun75, SIM75, Kun78b, Oda73, PP79, Sin73c]. Dynamics [Eis77, Sim78b, CJCM73, Enz73, FM78a, Hub71, MHRS76, MHRS77, OCRMLR^{+}71, Pri75, RM74, Fan70].

E. [Sla73b]. earth [PBG76, RW74, LB73a]. easier [Col77a]. Easy [Gia73]. Eckart [Jør78]. economic [Bil72]. ecosystem [BCL77]. Ed [Lar78c, Sti79, Cal79b, Dun78]. edifices [RS78c]. Edited [Cal73b, Cal75c, Cal76b, Shu77b, Shu77a]. Editorial [Sab75b]. Edward [Lon73]. Effect [JMM71, KGP79, LRJ73, MK72, NMS77, Nor75, San71, SSM75, SS77a, SH75b, Tat73, Ave75, Bas75, BN74, BMM75, CHH^{+}78, ČPH74, DY76, Ham71, HT75, JVM78, KN79, KBKČD78, LCPP79, MSSF79, MG77, PK72, Rei76, SHO71, VH75, Vol79, YR76, GZ72b, Sch71e]. Effective [BW76a, Bra75, GCJ73, MW75c, ST79a, TP78, TOED79, Wes79, Bra79a, CW77, Jør77, KK74a, Sla79, Wes73, WJC70]. Effectiveness [Sas77, Jak76c]. Effects [ST73, BBB73, BS79, BL77d, ČPH74, ČPP75, CJCM73, DN75, DBJ74, Han79, HPW79, KC73, KC77, Kle74b, LCPP79, Lun77, Mat73a, Mat76c, MVN75, ME78, Nes77, PČH74, PP74, RKE79, RM72, TP77a, TDB77, WC75a, Wen79]. efficient [GJ76, Koe74, PD78a, Rou79]. EHF [May74a, MK75b, MK75a, May78]. eigenfunction [HL78a, KD79a]. eigenfunctions [BB73, GTG76, MS75, RB75a]. Eigenproblem [Kun78a]. Eigenspace [ACH72, Car74]. eigenstates [Bri73, Goł79]. eigenvalue [Kar70b, Kar70c, KFY71, KA76, LA70, LWA74, Nor75, RE70, WAL71, Woo72, KBA76]. Eigenvalues [LL70, Cas78a, Gra77, Koe74, LC77, Ste74, Wel70, Wel71]. eigenvectors [Koe74]. Eigenwertberechnung [KA76, KBA76]. Eikonal [CL76, GLS79]. Einfluß [BKW75]. einsamen [BKW75]. Einstein [MBMM79, Wil79a]. Elaboration [RCE75]. elastic [BG78, SSM75]. elastin [RU76, RSK^{+}78]. Electric [Als73, LGP75, BSM72, CE73, Gra77, HH77, Kob78, Lar78a, LCP76, Loc72, MCST72, PPI73, TSE76]. électrique [Loc72]. electrolyte [Tal77]. Electron [BL77b, BH71, Bra75, Com77b, ES73b, ED79, Gal72a, GN77, GOK79, GLS79, Goo76, GP79, Hur76a, JMP79, KM74a, LS72, LRD^{+}77, LRD^{+}78, Mic76b, Mil71, PKSB78, Spe70, Spe72a, VGS79a, WESW73, Aba72, AC76, Abs78, Ada72a, AB73, AM77b, AM77a, AM72b, AT75, Ano79o, ACK^{+}75, BFS73, BB70, BP78, BTD79, BN74, BL77a, BM74, BBP71, BDP73, BG78, Ben72, BL70, Bil72, BK75a, BK76, BW73, Bra79a, BY76, BB73, BB78, Cal72, Cal73c, CM72a, CHD71, CB72, CSY^{+}74, ČP71, ČP77, CJCM73, CM72b, CFMO73, CH77, CESW70, Coo78a, CS72, Csa77b, Dah78, DL75, DSC^{+}77, DT72, Dec70, Des72, Feu77, FU76, Gal74, Gás71, GBT79, GC77, Gru70, GL74, HR79, HLGC73, Hen79, Hos72, HP77b, Hur76b, HYL78, HR76, Jak76a, Jak76b, Jak76c]. electron [Jak76d, JMM71, JŻ79a, JPTM78, JM73, Jun78, KL75, Kle74b, KP78, Kun78a, KOW78, Kun78b, KS72, KK77d, KKK79, KPÖ76, Lad74a, LH74, LP74a, LG77, Lim76b, LE72, LS74a, LÖ77, LBÅF78, LB73a, LG75, LPT^{+}74, Lun77, LAM71, MMB70, MW74a, MS71a, Man77, Mar79, MM74, Mat76b, Mat79a, May74a, May74b, MK75b, MK75a, May78, MLB72, Moh77, NN74, Nes77, NB78, PČL78, Pau78, Pet77, Pet79, PN76, PBS76, PK77c, PÖ77, RL77, Rat78, RC70, RC71, RC72, Rob71, Rob72b, RJ73, RB75b, RK76, SHI73, San72, SM73a, SM76b, SM76a, SD78, Sar75, SPČ79, SM71a, SM71b, SM71c, SLJ78, SMH77, SG77a, Sha77b, Sha78b, SS73a, SS73b, SIM75, Sil71, ST70, Sin73b, SS71a, Sla70b, SJH79, SR72c, SBM70, SMW71, TS71, TNB79]. electron [TOED79, WMB75, WRS70, WV74, YL74, ZV79, SL79, SDS73, Com77a]. electron-atom [CM72a]. Electron-conformational [VGS79a]. electron-density [CFMO73]. Electron-intramolecular [LRD^{+}77, LRD^{+}78]. electron-lattice [CS72]. electronegativity [Joh77, MR72]. Electronic [ÅFL78, BT78a, CT79, CSR79, Chr79a, Duk79, EP71, Fau71, Fau76, FWJ^{+}79, GPC^{+}71, GML77, GL74, KMO77, KWC79, LSS78, LF77, LGM77, LM72, LG75, LV71c, MW74a, MHWC78, MHW78, MOV75, Nes70, Nes71, Pet78, Pop71, Rit76, RL79, SMB^{+}79, SPČ72, SCYB75, SM79, Sto71, Ukr72, AeAs78, ANR75, Ada75a, AMSS79, BF71, Boy72a, BS78c, Bur77, Car75, CSFM72, CS76, ČPP75, CR72, CL78, CK74, CS78, De 78, DD77a, DM72, Fin73, FBL79, GP79, GK75, Har78, HS70a, Hin77, Hop78, HT75, HR76, JR74, Joh75, JW72, JG72, Kan74, KR71, Kel75, KJ75, KL77a, KC74, KD79a, KU70, KS79, Lad75b, LL76, LVDC74, LP79a, LFF79, LKK^{+}79, LE72, LÖS73, LK78, LS79, Lud71, Lud75, MPN79, MS72a, MS74, MCF^{+}70, MCM73, Mic76b, ME78]. electronic [MR72, MK72, NS75, NT79, Pai71, Pal75, PH77, Pet77, PNC^{+}78, RM79, SHO71, SP74, SP75b, SLB77, SLC78, Spe73, SWP77, SG76a, TB74, Tho76, Tru73b, Tru73a, VG78, WS79, ZF76, LL72, PLSL78]. Electronic-structure [KMO77]. Electronically [FMRG75, BS75c]. électronique [LL72]. Electrons [JBG70, TW77, AL71, BKW75, Col77b, GGJ77, Hil79, KMLB77, Lin71, MK76a, SIM75, SKS77, SM73b, Tat76, TL71, Wan79]. Electrophilic [Mem75, Mem79, HG75]. Electrostatic [CM75a, CFHS77, HKP79, NS79a, SS73c, DRCT78, GPP79, Kib75b, OMC76, PPP79, PT75, PPK78, PD77, PZP79, SBB79, GJP79]. Elektronenpaars [BKW75]. elektrophilen [HG75]. element [Sha78b]. elementary [Bra79b, ČZ79, Lan74, Rei70]. Elements [Gal72b, HL79, Sas74, DD72, Dav74, HBB77, HLGC73, KR73a, LS74a, Nis77, Oda73, PW79, PN76, Pos78, RCE75, Sur71, Tru72, WG74, Wit76, WP79]. Eleventh [Pul78b, Sti79]. elliptical [LGR73]. Elsevier [Ahl77a, Ano78c, Lar78c]. elucidation [Pul77]. embedded [Dmi75b]. embedding [BHLM78]. emergence [Pea74a]. emission [GK75]. emitted [Hea71]. empirical [BBL77, CAMC71, HLYC77, LÖS73, NMS77, San71, San72, SMH77, SG77a, WJC70]. End [RBDV77]. End-on [RBDV77]. endowed [KMT77]. energetic [KC77]. energetics [VU79]. énergie [BJ73, BPTLVM73, eAJ76]. Energies [OJB77, Abd71a, Abd71b, BM74, BJ73, BP75b, Cal72, Cal73c, Cal74, CM72a, Cas79, CJS77, ČP72, CR72, CAMC71, CV73, DGC77, DCM74, DMR74, Dau74, GJP79, GBT79, GLS79, Gin78, Hal75a, Har77c, HE77, Jak76d, JMP79, JSC78, KZ73, Kau70a, KPKC73, Mcd79, Mey71, MDS75, Muk74, MMM75, Nir75, PT78, PNC^{+}78, Rit76, SHO71, SBM72, SP75d, SO77, Tal77, TNB79, Tru73b, Tyu75, Tyu78, WJ79, ZR78, ZBH^{+}78]. Energy [BG75, CD76, DL75, Dav72, DF76, FY79, HF72, Jai70, KKA78, Lad70, Lad76, LL70, LB73b, OK75, Sch71a, SW71, TC70, Wan79, AF73, BFS73, BA78, BS74a, BS76a, BNLS72, BZSD79, Bor70, Bor74, BPTLVM73, BEF78, BFH74, BWRD79, CZNL70, CB72, CJ73a, CJ73b, Cha71, CL76, Cof74a, CESW70, CEK72, CH74, CH75, CZ75, DMS^{+}78, DEC78, EL73, EH73, Eng71, Euw71, FN73, Gar74, GTBB77, Gra77, GBK77b, HS76, Han77a, HS79, HS75, Hen79, Höl77, JTR77, JŻ79a, Jor75, eAJ76, KS73a, KP71a, KP76, KM76, KD79b, KBK72, KP78, KKK79, KMLB77, Lad74b, Lar78a, LV71b, LV72, LW77, Lin77, LG70, Lip72, MP71b, Moh78, MDA78, Muf76, NY76, PT78, PPC72, Pet78, PK71, PD78b, Pop71, RU76, RKOU77, RE70, Rue79]. energy [San71, SP75a, Sco71, SDD72, SW70, SB78, Spe73, Sri77, SP77, SK79, SL73, SO77, TB78, TTS71, TDB77, Vog79, WK72, WS75, Wei77, WN78, Wel76, Wel77b, WS79, WFS77, WM79, ZSR79, CT77]. Energy-band [Lad76, Pet78]. Energy-based [TC70]. energy-dependent [CEK72]. energy-localization [Eng71]. Energy-shifted [CT77]. energy-spectrum [WFS77]. England [Com77a, Com77b]. Englewood [Cal76a]. Enhanced [Ave75]. enhancement [Gol70]. enhancements [BC76]. enol [ZR78, ZBH^{+}78]. ensemble [BHLM78]. enthalpy [BCR^{+}78]. entropy [HČZ79, Pri75, Sch76a]. environment [ST73]. Environmental [TP77a, PP74]. environments [Boy70a]. enzymatic [BCR^{+}78]. enzymatically [LHK77]. enzyme [Gol70, LP79b, VGS79a, WEH74, Ano76e]. enzyme-catalyzed [WEH74]. enzymes [AMSS79, LKK^{+}79]. ephaptic [Wal77]. Epoxide [PDEB78, LPP78, LP79c]. epoxides [LP79a, PDE79]. Equation [Lib76, TN73, Ada73, Amo71, Bie75, Csa79a, ESG79, Fro79, Hen77, Koe74, LA70, LWA74, NS75, Nor75, PW74b, PW74a, RL77, SHI73, Sil71, VB72, WA71, Wel70, Wel71, WZ70, WAL71, Woo72, YL74]. Equations [FG79, HY78, Ada78, BEH77b, Cal79c, Cli73, DSY74, DR75, FW75, HS75, HBE78, JM73, Kar70a, Kar70b, Kar70c, KFY71, Lai73, Lan73, Lim74a, MLB73, May74b, MK75b, Mes79, Ner76, SA77, SIM75, SS71a, SD74, Živ77, Sim77a]. Equations-of-motion [FG79, DR75, SIM75]. equilibrium [DN75, Sim78b, SP75d]. Equivalence [JD74, HT75, SHO71, SDB76]. Equivalence-restricted [JD74]. Equivalent [GÖ68, GÖJ^{+}73, Sch77b, BS74b, DN75, PW74b]. Erich [Ano77l, Par77]. Errata [GZMM73, GÖJ^{+}73]. Erratum [May75, PA77, SRP77]. Error [LTR74, BS77b, BS78a, Col77a, JW76, KA76]. Errors [HS76, SP77]. ESCA [BL77c, Con72, WK73]. ESR [MS71b]. esters [Kla78]. estimate [CS74]. estimated [JŻ79b]. Estimation [BYB79, LSOC77, PT78, DMS^{+}78, LB73b]. état [BJ73]. états [BJ73]. ethers [DBBJ76, GBS70, GBS70]. ethidium [PL77, RKE79]. ethyl [MEW70, OMC76]. ethylating [HPK79]. ethylcarbonium [TWC73]. ethylene [CSR79, FWFM77, FJP70, HMC71, MMD78, PDEB78, Rat71, Rou74, TS78, WC75b, Wes79]. etiology [Kau74a]. Etude [DQ77]. etudié [TGSL70]. europium [Fog71, Liu71b, SKS77]. evaluate [PD78a]. evaluated [SHI73, Sla79]. evaluating [Jet73]. Evaluation [BR78, Che71, DDH78, BOP76, BSG77, Col79, eAJ76, Sha78b, ST70, PAM75, Pos78]. Evaluations [Sha76]. even [Raf75, Sar75]. even-tempered [Raf75]. evidence [Feu77, Oda73, Pet78]. evidenced [Lar75]. Evolution [Sun74a, BF75, DS79]. evolventibus [BP79]. Exact [BS71, CJ73a, HKM71, LC77, Nis77, HL78a]. Examination [Dub72, Jak76c]. examples [BG72]. Exchange [GMZ73, SW71, AM77b, BYB79, CJ73a, CJ73b, CESW70, CEK72, Csa75, Dan72, ES70, Hir70, Ive74, Kle70a, Lin71, LPT^{+}74, Ove73, Pie71, Sco71, Sil71, SW70, Sla71a, SC78a, Suz77, Tat73, SJH79]. exchange-correlation [CEK72]. excimer [AZ72, Fla79, MHWC78]. Excitation [CK74, OJB77, Abd71a, BS77b, BS78a, BK76, BZSD79, Bow79, CW77, CS76, ČP72, DC75, Gos75, Har77c, NM77, Oni79, Tru73b]. excitations [BP78, Bra79b, CKI75, Gre71, Mil71, MOV75, MMS70, Ner76]. Excited [LS79, SCYB77, WMB75, Abd71b, AS78, Bar77, BN76, BJ73, BS75c, Cha77, CD74, ČPH74, CKD73, DS76b, DY73, DC75, GB75, GL74, HNT78, Hea71, HT75, Hos72, KKW73, KSP78, Lin77, LG75, LAM71, MW74a, MCST72, MMS70, MM72b, MG77, MSM70, MBM71, Muk74, Mul71b, PČH74, PLMB73, RB77, Ros78b, SMB^{+}79, SC76a, Tan74, TTS^{+}79, TSE76, You72, YD75, SPČ72]. excited-state [BN76]. excités [BJ73]. exciton [Par72]. Excitons [MBMM79]. Exclusion [Lim76a]. Existence [Živ77, RT77, TK78]. expanded [Ber73]. Expansion [Dal79, Ada72a, Ada79, Csa77a, Ema74, FB76, GPP79, Gra77, HNT78, Lin70, Man77, MHPS70, Mat71a, Mcd79, MW75c, Rob71, Rob72b, RK79, SR72a, Sal73, SR72b, SP79, WS79, You75]. expansions [BG72, CJS77, GbK77a, Gor71, Gor72, JSC78, KL75, Kle77, KGB77, MCST72, Meh73b, NY75, PAM75, SRP76, SRP77, SF75, Wes73]. expectation [BSG77, BL77d, SK71]. experiment [BP72, Ber75, Gin78, IZZ^{+}79]. Experimental [Rob72a, Con70, Coo78a, Nes75, Sin75, Feu77]. Experimentalist [WBKL70]. explicitly [Sim72]. exploration [PBC74b]. explore [SDD^{+}77]. exponent [LS74b]. exponential [CM76, Höj79a, Höj79b, Sch72]. exponential-type [Höj79a, Höj79b]. exponentials [Sol71]. exponents [BPTLVM73, Vla70]. exposants [BPTLVM73]. expression [Gar74, KD79a]. expressions [KP71a, LG70]. Extended [BRS^{+}75, BEF78, EB74, FMZ70, JL74, KT76, AM72b, Als73, BFH78, Bur72, CFHS77, Chi77, FHB77, JS70, JNŽ72, KPKC73, MLB73, May74b, Mes73b, Pel77, Han79]. Extension [KP72, MDS75, TB75]. extensions [Lun75]. external [Sim78b]. Extrapolations [BG70].

F [Cur78, FJP70, GH72b, KJ78, Lar76, Par73, TS77, FJP70, GH72b, LL72, Min79, PK76b, RM74, RW75, ST72, SBD^{+}70, VG78, WK73, ZL78]. F- [FJP70, FJP70]. factor [Gol70, Nam75, SK72]. Factors [Kar78, BS76b, BM73, CT79, GBM74, GRM75, MS71a, MMM71, SM71a, SM71b, SM71c, Sur71, TS71]. family [BSB77]. Faraday [Ave75]. Farnell [Sab75a]. fast [TS71]. FCC [KOW78, Kun78b]. FCO [TB74]. Fe [BMD70, GML77, MSSF79, Har77a]. features [RU76]. Feenberg [GS70b]. Fehlerminimisierung [KA76]. Fermi [Csa79a, GY75, Goo76, YG77, Bie75, Csa77b, Csa79b, DH76, GGJ75, Han70a, Han77a, HB79, KW76, Lun75, PN71, SKW72, SC73, WC75a]. Fermi-like [GGJ75]. Fermion [Col71, Col70, GH78, Per78]. fermions [Bow72]. ferrimagnetic [Fog71]. ferroelectric [CHHS74, HCIE75]. ferromagnetism [MCP73]. ferrous [LK78, RL79]. Feshbach [LKH77]. few [Per78]. few-body [Per78]. Feynman [CD71, NRS79, TG76]. FH [Lei74, ST72]. Fibrous [GK78]. Field [GÖ68, GÖJ^{+}73, Sla70a, AD77, Bal71, BKS79, BM73, Bow71, CE73, DS74, DLG79, Ell73, Gou73, Gra77, GL74, Hac74, Hal78, Har71c, HP77a, HDG71, Hor71, JL74, KKW73, KK74b, KK77c, LGP75, LG77, LG75, Nal78, Nir75, PS75a, RW74, SF76, Sch71a, Sch71b, Sol73, TGSL70, WN71, Wan79, Wol71]. fields [DL75, Gás71, LL73, MGM78, Nal78]. Fifth [Ano78d, Boy70b, BM71a]. Fifth-rank [Boy70b, BM71a]. Filled [Bee76]. filling [Kau79b]. final [Rou73, Tru76]. Fine [BK77]. Finite [HB79, BF70, Blo72, Blo73, ES79, Fae79, FW75, FU76, KE79]. finite-difference [FW75, FU76]. First [CH74, JBP76, Lip72, NRS79, Bun77, Cal74, Coo79, CH75, DGC77, Goł79, GL74, HS76, Lan73, LF77, Laz78, LZ79, MCST72, Moc74, NG73, RB77, ST72, SC73, Van71, Wel77b, Raf75]. first- [Moc74]. First-Order [JBP76, NRS79, Coo79, HS76, Lan73, LF77, Laz78, LZ79, SC73]. first-principle [CH75]. First-principles [CH74]. first-row [Cal74, DGC77, NG73, ST72, Raf75]. five [AMC73]. five- [AMC73]. fixation [NPL^{+}79]. fixed [KM74b, Loc77a]. flexible [DSMM79, LJBN75]. Floating [AM74, HP77b, RMB78, TL71, AF73, AM72a, ADDLG76, RR77]. fluctuation [Bra79b]. fluorescence [AZ72]. fluoride [AD77, Lei74, Sti78, VG78]. fluorides [PK77c]. fluorinated [KR73b]. fluorine [SDD72, WRF73]. fluorine-containing [WRF73]. fluoro [BL77c]. fluoro- [BL77c]. fluoroacetic [BL75]. fluoronium [Lei74]. fluorosulphate [HS70b]. fluorouracil [GJP79]. fluorure [Lei74]. fluxional [TB73]. FNO [HR79, MG77]. Fock [BR78, Bre73, Cli73, MBM73a, SS71a, SDS73, BSM72, BM73, BMM75, BEH77b, BK71, Bur72, Bur73, DSG74, FW75, Fro70, Fro74, GHS79, GMG75, GBM74, GRM75, HKM71, Ive74, Jam70, Jor73, JL70, JL74, Kar70a, KKA78, KM76, KT76, KBK72, KMO77, Lai73, LS70, LS71, LL75, Laz78, MHPS70, MN79, MLB73, May74b, Mcd79, Mes73b, MW74b, Mes78, Mes79, NN72, NN74, NS72, NS73, OKM74, PČL78, PS73, PS70, PKSB79, RS79, RGM77, SA77, SH73, SLB77, SDJ79, SDB74, SDB76, SB78, Spe71b, SP75d, SEW73, TTN76, Tru73a, Van70, Vog79]. fold [Lét74]. folding [PF79]. follow [BCR^{+}78]. fonction [BPTLVM73]. fonctions [eAJ76]. fondamental [BJ73]. forbidden [TH78]. Force [MGM78, CJCM73, KP71a, Nal78]. forces [Abd71a, AM77b, Brä70, Cla71, FY71, GK78, Jeh73, KKW73, Mal71a, Mal71b, Pie71, RM72, RM73, RMP71, WB76]. forcing [JSC78]. form [Duf71, Gor71, Gor72, Hen77, JW78, RSS75]. Formal [Bra75, Bra79a]. formaldehyde [DRCT78, LED73, Tan74]. formaldiminoxyl [MMB70]. formalism [ADDLG76, ACD^{+}74, Car74, CDEO76, Com77c, CE73, Feu77, GP74, GJLL75, GJ76, KP72, Min79, Mor75, Sac72, TC70, ZF76, ZF77]. formalisms [Mic76a]. formally [NG73]. Formamide [PK76b, LBÅF78, OSL77, OSL77]. formation [BB70, HČZ79, Kho79]. forming [Mor75]. forms [GPC^{+}71, Kob78, KL79, LH76]. formula [CA78, JW78, MHRS76, Tyu75, Tyu78, WW71]. formulae [Har71a]. Formulas [BSG77, BSZ78, CM78, Gor71, Gor72, Har77a, Hos72, KE79, Spe71a]. formulation [Cli73, HS71, KPJM70, Lar79, Mat74a, Mat76b, Mat79a, MW74b, Nel79]. formulations [KC74]. formycin [SMP77]. found [ZBH^{+}78]. foundation [CD71]. Four [NY75, HR79, HP77b, LH74, LPT^{+}74, May74a, PK72, SR72b, Sey76, SHR76, SDS73]. four-atomic [PK72]. Four-body [NY75]. four-center [SR72b]. four-electron [HR79, HP77b, LH74, LPT^{+}74, May74a, SDS73]. Fourier [AFD^{+}78, GMZ73, GMG75, MH72, Sha71, Thu75, ZF76]. Fourth [WS79, KP78]. Fourth-Order [WS79, KP78]. fractional [KK77b]. Fraga [Ahl77a]. fragment [GG79c, GG79b]. fragments [CSM71, Chr73, Kau70a, OMC76, PCMS77]. frame [Jør78]. framework [ABM78, LW76, Loc77a, Loc77b, Loc77d, PD78a, WJ79]. Franck [Nis77]. Free [BTD79, Mat76b, MW77, MN79, SD79a, AB73, BL70, DMS^{+}78, GPC^{+}71, KPJM70, KC74, Lim74b, Lim75a, Lim76b, Lim77b, LB73a, MW74a, Mat72a, MF73, MSP73, MCP73, MPW75, MW75a, Mat76a, MWY77, MK78, SS73a, SS73b, SM73b]. Free-electron [BTD79]. French [BJ73, BPTLVM73, Gou73, JM76b, eAJ76, JJ78, LL72, Lei74, Loc72, Mar73a, TGSL70, TBB70]. Frequency [Hea71, RGM75, Ahl74, Bow71, BR72, Bre73]. Frequency-dependent [RGM75, Ahl74, Bre73]. Friedrich [Ano77m, Par77]. frozen [Lar75, Lar78a]. frozen-K-shell [Lar75]. Fues [SS71b]. Fukui [Cal77b]. Full [KKM70, Lar78a]. Function [LV71b, Ada73, BSS78, BKS79, BS75b, BPTLVM73, Bre75, CB72, ČP72, CSH74, CS72, DRCT78, De 76, Duf71, FMZ70, HSBD78, Jol79, KM74a, Kut78, LD71, LV72, LA76, Meh73b, Meh78, NS75, NRS79, PBPL78, Pau78, PGPP79, Rat71, Sch72, SB75, SD70, SDB76, SL70, WFS77, Sim77a]. Function-Level [Sim77a]. functional [BG75, Fin73, Gia73, GJLL75, GJ76, HS79, KS73a, PS75a, RSS75, SP75c, SJH79, SMW71, TB78]. functionals [HS79]. functioning [AMSS79, VGS79a]. Functions [Gal72b, Jai70, Aba72, Ada73, ANR75, Ada75a, AB73, Als73, ADFL71, Ban74, BA78, BN76, BDP73, BS71, Ben72, BSG77, BC79, BBA73, BG72, Bre73, Bur77, Cas78a, Cha77, CL78, CII79, DS74, DMM75, Eng72, Gal74, GN77, GM73a, GM73b, Gre71, GH72a, Gru70, GB72, GCJ73, HO73, HHR70, HKP79, Höj79a, Höj79b, Hur77, HR76, Jak76a, Jak76b, Jak76c, Jak76d, JTR77, JW78, Jor73, eAJ76, KC72, KC77, KK76c, LF77, Lau79, Lec76a, LTR74, Li76, Lim76a, Loc77a, Lud75, Lud77, Meh73a, Meh73b, MDC79, Moc74, MK76b, MM72b, Mos73, MSM70, MBM71, Nal78, Ner75, Pau77, Pol79, Pop71, RWFS71, Rob71, Rob72b, RK76, Sab75a, SDD^{+}77, SH73, Sch78, Sch72, SC76a, Sco74, Sha76, SS73a, SS73b, Sim72, SH75a, Sla70b, SC78b, SW77]. functions [SW78, SWP77, SSH78, Tan74, Tru72, Tru73a, Urr78, VB72, VF76, Wel76, Wel77b, Wit76, WV74, Mos73, Pos77]. fundamental [HPK79, Kau79b, Pul74]. fundus [JG74]. Future [Sla73a, Ost79].

G [Ano78c, Cal76a, Cal79b, Lar78c, Mår78, Sab75a, IMN79, Nam75]. Galerkin [Jan76]. gallium [Fog71]. gap [Bor74, Cal79c]. garnet [Fog71, Liu71b]. gas [BM74, BY76, ČZ79, Chr77, Csa77b, FBL79, Fla79, GOK79, LW77, LRJ73, LPT^{+}74, MHW78, MS76b, NY75, RJ73, Sti78, TW77]. gas- [Sti78]. gas-phase [ČZ79]. gas-solid [FBL79, LW77]. Gauge [Hal73, Kob78]. Gaussian [AF73, ADDLG76, BBP71, BP74, BSG77, DPGC79, Euw71, GM73a, GM73b, Hal73, HS70a, HP77b, JP73, KC72, KC77, KSP78, Lec76b, Lo75, LA76, MEW70, Mat71a, Mat73a, Mat76c, MYTC77, NY76, OK75, PAM75, Pos77, Raf75, RR77, RMB78, RM72, RS78a, RS78b, RS78c, SBB79, SBM72, SC78b, TB74, TL71, Van74, Van70]. Gaussian-Lobe [RS78a, RS78b, RS78c]. Gaussian-type [Lec76b, Mat73a, Mat76c, NY76]. Gaussians [FHP70]. Geminal [SPČ72]. geminals [Lya78, LB73b]. genealogical [GTG76]. General [ADDLG76, KSP78, Gar74, Hac74, Jak76a, JG72, JNŽ72, KR73a, Kau74b, Kib75b, PW74a, Rob72b, SP75a, ACD^{+}74, BSZ78, BS78c, Car74, Kle77, Lar79, NI70a, Rob72b]. Generalization [GHS79, KS73b, SL79, DSG74, Ste77]. Generalized [CGRC74, Car72, Hal78, Hal79b, Joh70, LG70, RT77, RK76, Sch76a, Suz77, Ahl77b, CW76, FM79, LA70, LWA74, Lev78b, MW74b, Pet77, PNC^{+}78, RC70, RC71, RC72, SB75, SL70, Tyu75, Tyu78, Van74, WAL71, Woo72, SC76a]. Generalized-molecular-orbital [Hal78]. generated [BW76a, HKP79]. Generating [Wit76]. Generation [GTG76, BL77a]. generator [LV78]. generators [Lar75, PW79, WP79]. genetic [Coo78a, Coo78b, FM79]. Geometric [BB71, HTG77b, Laz79, Har77b, HTG77a]. géométrie [TBB70]. geometries [ZR78, ZBH^{+}78]. Geometry [Har79a, JŻ79b, AF73, AMSS79, BSZ78, CD71, CLP79a, CLP79b, FWFM77, JR74, LK78, TBB70]. gerichteter [BKW75]. German [Ahl70, BKW75, FJP70, HG75, KBA76, KA76, PTUK69, Sch73b]. giant [Cal79a, Jul78]. Given [GÖ68, GÖJ^{+}73, KMLB77, May78]. GIy [VU79]. Glauber [GLS79]. globular [MME79]. Gly [BWRD79, CBV79, RSK^{+}78]. Gly-Pro-Pro [CBV79]. glycine [DEC78, RRNS75, RRS75, RC78]. Glycosyl [SY77]. glyoxal [DEC78, DAP79, ED79, LDP79, OSL77]. glyoxalase [SWF79]. glyoxals [Abd76]. glyoxylase [LP79b]. Goldstone [LH76]. Gordan [KG76, KG77]. Goscinski [Shu77b]. Gradient [Gar74, CE73, LGP75]. gradients [RPK78]. Graph [Ran78, RW79, Sha77b, Ran79]. Graph-theoretical [Ran78]. graphical [APČ77, LH74, Lar78b, PAČ77, PW79]. graphs [Ran79]. green [CSH74, BSS78, BKS79, BBA73, ČP72, HR76, LD71, LF77, Lau79, MHRS76, MHRS77, Ner75, PBPL78, Rat71, Sco74, VB72, WFS77]. Ground [Bra79b, ÖL78, SDS73, Spe72a, ADFL71, BWS77, BJ73, BS70, BP75b, Bun77, Cal72, Cal74, DRCT78, Fro75, GL74, Ive74, Koł74, KÖ71, Kut78, LH74, Lar78a, Lec76a, Lec76b, Lin77, LG75, Man77, MCST72, Mul71a, ÖL79a, PS70, RB77, RS79, Ros78b, SK71, SK72, SMB^{+}79, SH75a, Sin73b, SD70, SDB74, Tan74, Tat76, WK73, WRS70, Wil77b, WB73, Hal79b]. ground-state [Ive74, Sin73b, WB73]. Group [BDT77, CSFM72, Hur77, KR76, MCF^{+}70, MCM73, TB75, BTD79, BF70, Boy72b, Boy72c, Cor72, ES73a, FMZ70, JS70, Kle70b, LP74a, Mat74a, Mat76b, Mat79a, Meh73b, Meh78, MM77b, Nel79, Nov73, Oda73, Pal75, PW79, PTUK69, PS79b, Rei72b, SDD^{+}77, Sha77b, Sha78b, ST73, TP77b, WP79]. group-function [Meh73b]. Group-theoretical [CSFM72, MCF^{+}70]. Groups [Gal72a, Blo72, Blo73, BW76b, But76, BW76c, Flu72, Gal74, GP72, KG76, KG77, KE79, KK77b, LP79c, Lét74, Lim77b, Ros78a, SAM79, Sin73c, SS73c, SLOKP79]. growing [Lak76]. growth [BC75, MDA78]. gruppen [Wyb73b]. Gruppentheorie [Cal73a, PTUK69]. guanine [HPK79, KD77, LPP78, PLYM78, PZP79, ZBH^{+}78]. GVB [PNC^{+}78]. Györgyi [Lak76]. Gyrotropique [DQ77].

H [Bur77, DY73, GG79a, HG75, HLYC77, JŻ79a, KP71a, Koc74, Lar76, LB79a, RR77, RW75, RKOU77, Sab75a, Sla71b, TG76, SY77, ÅSW77, Abd71a, Ada78, Ada73, Ada77, AM72a, AM74, AM77b, And79, BB74, BC79, BDT77, CJ73a, CJ73b, Doo78, EH73, FMZ70, GGJ77, HNT78, HG75, HP77b, JTR77, JE79, JP77, KR71, Kat72, KSP78, Koc74, Kut78, Lec76a, Lec76b, LTR74, LG77, Liu71a, LGR73, MS71b, MM72a, MG77, OSL77, RW75, RS79, ST72, SHR76, Spe72a, SW77, TS77, TBB70, TG76, Tru72, Tru73b, Tru73a, TS72, WB73, You75, TBB70]. H-like [AM72a, AM74]. haemoglobin [Har71a]. Hafner [Lar76]. Half [SDS73, SDB74, SB78, SDB76]. Half-Projected [SDS73, SDB74, SB78, SDB76]. halide [Fla79, Moh79]. halides [BB72c, CKD73, JM76b, LVDC74, LRJ73, LPT^{+}74]. Halkjaer [ÖL79b]. Hall [Cal76a]. halogenated [DBBJ76]. halogens [Hal75b]. Halogénures [JM76b]. HAM [ÅFL78, LBÅF78]. Hameka [Lar76]. Hamilton [TL72]. Hamiltonian [AC76, Abs78, CK74, DD72, Feu77, HL78a, Jør77, KS72, KK77d, MLM75, MMM75, WJC70]. Hamiltonians [BW76a, Bra75, Bra79a, CS78, Han70b, Rei72a, Sac72, Wes73, Wes79]. Hammerstein [PW74b, PW74a]. Handbook [FKS76, Atk74, Atk91, Gos78b, Ahl77a]. handed [SY77]. handedness [DS79, SY77]. hard [GH72a, Wei76]. hard-sphere [GH72a, Wei76]. hardcover [Gos78b]. harmonic [Bas78, DR77, Hun74, RB75a, WO75, ST79a]. harmonics [KOW77, Sch77b, SR72c, SF75]. Hartree [BR78, Har78, Jam70, Jor73, KKA78, KMO77, PS70, SDS73, BBK76, BSM72, BM73, BMM75, BEH77b, BK71, Bre73, Bur72, Bur73, Cli73, DSG74, FW75, Fro70, Fro74, GHS79, GMG75, GBM74, GRM75, HKM71, Ive74, JL70, JL74, Kar70a, KM76, KT76, KBK72, Lai73, LS70, LS71, LL75, Laz78, Lim74b, Lim75a, Lim75b, Lim76b, Lim77b, MHPS70, MN79, MLB73, May74b, Mcd79, MBM73a, Mes73b, MW74b, Mes78, Mes79, NN72, NN74, NS72, NS73, OKM74, PČL78, PS73, PKSB79, RS79, RGM77, SA77, SH73, SLB77, SDJ79, SS71a, SDB74, SDB76, SB78, Spe71b, SP75d, SEW73, TTN76, Tru73a, Van70, Vog79]. Hartree-like [Har78]. having [Lét74, SP75b]. HBeCH [RMB78]. HBF [TB74]. HCl [PK76b]. HCO [TB74]. He* [RM79]. He-like [BS71]. He-sequence [MMS70]. heats [GTBB77, GBT79]. heavier [HLGC73]. heavy [AZ72, LZ77, MFRE76, PK77c]. HeH [AM77a, BB70, Hop78, JP77, Koł76, SW78]. HeHe [DMR74]. Heidelberg [Cal77b]. Heisenberg [GbK77a, SHI73, SK73, SL70]. Heitler [Aba72, Ada77, Ano77-28, LP74a, LP74b, Wil77a]. Heitler-London [Ada77, LP74a, LP74b, Wil77a]. held [BP70, BP71, BP72, BP73, Pul78b, Sti79]. helical [IZZ^{+}79, KP79a, SY77, Urr78, VU79]. helices [VBC78]. Helium [BG70, LP74c, AC70, BK74, CL78, DD71, HG78, LA70, MCW78, SK71, SK72, SE79, Spe70, Tat76, WAL71, YD75, Pel77]. heliumlike [Lin70]. helix [CC79, DS79, Meh79, PPP79, Por77, Por78, PZP79, SY77]. helix-coil [Por78]. Hellmann [CD71, NRS79, PF74, TG76]. Hellmann-type [PF74]. Helmholtz [SF75]. Heme [DAC^{+}79, Eis76, Eis77, LK78]. hemoglobin [Sal72]. Hemoproteins [MSSF79, KN79]. HeNe [SBD^{+}70]. Herman [HC72]. Hermitian [Gou73, Han75, Hen77]. hermitien [Gou73]. heteroatoms [SN79, SG77a]. heterocycles [KRKŠ78, Rou73, Rou74]. heterocyclic [BP70]. heterogeneous [Sal79]. Heteronuclear [AM77a, Tyu78]. heteroorganic [Hil79]. Heuristic [HSBD78]. HF [BS74a, BS76a, HR79, LZ79, LS74b, Meh78, PK76b, TS77]. HFS [HC72]. HgI [HLGC73]. High [Mal71a, Mal71b, KTO^{+}78, Lib76, Sch76b]. high-spin [KTO^{+}78]. higher [RB75b, SB75, GbK77a]. highest [FG79]. Highly [Hen78, KHI79, SC76a]. Hilbert [Lan77]. Hill [Lin73]. hindered [DMS^{+}78]. Hirschfelder [CJS77]. histamine [MHH74]. HKS [Har79b]. HL [CSFM72, FM72]. HNO [HR79]. Hoffmann [Mat76a]. hole [BFS73, BK74, Csa75, LD71, LH74, Lar75, Lar78b, MWY77, Rat71, Wen79]. holes [Csa78a, Res78]. Holland [Sti79]. homogène [Mar73a]. homogeneous [Mar73a]. homonuclear [vDR79, Tyu75]. homopolymers [FB72]. homopolynucleotides [SL73]. honour [Cal75b]. hopping [Fro79]. hormone [MDA78]. Hovi [Cal75b]. HP [LL72]. Hubbard [FN73, Mat76a]. Hückel [Mat76a, AM72b, Als73, Ano77l, BRS^{+}75, CFHS77, EB74, Han79, KPKC73, Par77, SDM74, SS74, Tri77]. Hulthén [Cha77, GB72]. human [Har75, WLS77]. Humanities [Cal73b]. Hund [Ano77m, Par77]. Hurley [Com77a, Com77b]. HY [SH75a, SHR76]. Hybrid [Lec76b, Loh76, DF76, Eng72, MI78, Pol70, Sil71]. Hybridization [Ran74, AAC76]. hydration [LCPP79, LDP79, PP74, PA74]. hydrazines [GPC^{+}71]. hydrazyl [GPC^{+}71]. hydride [Bur72, DCS77, MVN75, SDB74]. hydrides [DGC77, DSC^{+}77, KC77, NG73, PK77a, PK77c, Swi71]. hydrocarbon [GBTB78, Her74, KT76, NN72, GG79c]. hydrocarbons [BOP76, BK71, ČPH74, DBBJ76, DEE^{+}75, GBT79, LP79a, Mak71, Mem77, Mem79, Muk74, NMS77, NS78, PČH74, PD77, Pul79a, SG77a, SR78, TP78, TTN76, TSE76, UP71]. hydrodynamic [SE79]. Hydrogen [ACG74, BL75, BK75b, PPI73, YR76, AM74, AM72b, Bai74, BRS^{+}75, Bas78, BS75c, BFH74, Cho79, Coo79, CSH74, Cur77, DBBJ76, DMS^{+}78, Dea72, GZ72a, GZ72b, GGCK79, HKM71, Koł74, Koł75, KMO77, KGB76, Lei74, Lib76, LS72, Loc77c, LS79, Mal71a, Meh79, MO77, MS76b, Nir75, OKM74, OK75, Pie71, PDE79, PGPP79, Rei76, Rit76, Sab71, Sch77c, SLC78, Sti78, SH74, TP77a, Wel77a, Wil77a, Wój76, Woj78, Woo72, Zer74, CP70, GGCK79]. hydrogen-bond [DMS^{+}78]. hydrogen-bonded [BS75c, Cho79]. Hydrogen-like [PPI73]. hydrogen-molecule [BRS^{+}75, MS76b]. hydrogenic [JP73, Lau79, WG74]. hydronium [MS71b]. hydroquinone [SP71]. hydroxyapatite [WRF73]. hydroxyindole [WO77]. hydroxyl [MK72, WRF73]. hydroxylation [PL79]. hydroxymethylcarbonium [TWC73]. Hylleraas [BS71, CL78, Jol79]. Hylleraas-type [BS71, CL78]. Hyper [LS71, BSM72, LS70]. Hyper-Hartree [LS71, LS70]. hyper-polarizabilities [BSM72]. hypercycle [Eig78]. hyperfine [BNLS72, IMN79, LL75, Lun79, Men71, NN72, NN74, PPI73, TB74]. hypersurface [HG75]. hypervalent [ST72]. hypervirial [Cas78a, Gop76, ST79a]. hypervirial-scaling [Cas78a]. hypochromic [Vol79]. hypothesis [HS71]. hypothetical [PS70].

I. [KA76]. Icosahedral [BO73, Boy72a, BS74b]. ideal [SKL79]. ideas [Tho75a]. idempotent [Hen78]. idempotents [Sch71a]. Identification [DSMM79]. identity [Lat77]. IETS [Feu77, Feu77]. II [Har77a, KD77, MSSF79, RRS75, Abs78, APČ77, AM77a, AB79, Bar79, Blo73, But76, CW78, CD78, CSFM72, CSR79, Coo78b, CH75, DSC^{+}77, DMR74, Dmi75b, DDD70, Fau76, GZ72b, GL74, HH77, HS79, HR79, Höj79b, Jak76b, KC73, KG77, KBA76, KOW78, LZ79, LP74b, MMB70, Mal71b, MM74, May74b, MK75b, MHRS76, MVN75, NI70b, NS73, PČH74, PW79, PHL^{+}76, Pet77, Pos79, PS75a, Ran79, RC71, Rob72b, Røe73, RS78b, Sch71c, SG77a, SPČ72, SDM74, SS73b, SH75a, SDB74, SCYB77, Spe71b, SC73, ST78, Tyu78, Wel77b, WAL71]. III [AM72b, Ave79c, Ave79a, BB72c, BW76c, ČP71, Dau74, DH78, FB72, Gás71, Har70a, Har79a, Jak76c, Kun78b, LBÅF78, LG75, MW74a, MK75a, MHRS77, Mcd79, MCM73, MLB72, MDS75, PP74, RC72, RS78c, SP75c, Sch71d, Spe72a, Spe73, ST79b]. imaginary [BR72]. imidazole [KRKŠ78]. imidazolium [WO77]. imino [ZR78, ZBH^{+}78]. imino-enol [ZBH^{+}78]. immunity [Ram78]. Impact [Wil77a, BK76, PS79b]. impacting [HG78]. Implications [KKK70, Kob78, Cho79, Van75]. Importance [EH73, BP78, CW77, JM76a, RR72]. improve [Kau74b]. Improved [Jol79, KS75, Moh79, Vog79, PF74, RDH79]. improvements [ESR74, Sla74]. improves [Ram78]. impurities [JVM78]. Impurity [Fan70, Ada72a, AD77, BHLM78, CD78, Csa76, Csa77a, Csa78b, DT72, KGP79, MCM73, SLC78]. inadequacy [Cla71, Feu77]. incluant [JM76b]. including [FGV77, JŻ79a, KK74a, SBM70, SSH78]. incoherent [BK76]. incomplete [Mcd79]. Incompletely [Kun78a, KOW78, Kun78b]. incorporating [PBS76]. Increased [Har71a]. Increased-valence [Har71a]. increases [MK76a]. independence [KB77]. Independent [Meh73a, Coo74, GGJ75, GK76, Lud75, Man77, Sol73, TGG79]. independent-particle [GK76, TGG79]. indicate [Kau79b]. indices [SS78b]. indirect [Cal79c]. INDO [BL77c, FGV77, Kau70b, KP71a, KP72, KPKC73]. indoleacetic [BC75]. Indolealkylamines [WCK^{+}76]. induced [BC76, INY^{+}71, Lar73]. inelastic [Feu77, MM72a]. Inequalities [GH78, Gol73]. infection [Ram78]. infinite [DDH78, MCP73, PD78a, Res78, RT77]. infinitesimal [Han70b]. Influence [FSV79, eAJ76, MGT79, NT79, BKW75, Sal72, SP77, eAJ76]. influences [Kau79a, NG73]. Information [Ber76, HL78a, Lar73]. Information-theoretic [Ber76]. infrared [Wój76, Woj78]. Inhibition [SWF79]. inhibitor [MDA78]. inhibitors [SS74]. inhibitory [Gab76]. inhomogeneous [GOK79]. initial [GPC^{+}71]. initio [Alm74, Bai74, BP74, BP75a, BCR^{+}78, BVR78, ČZ79, Chr73, Cla70, CDEO76, Cur77, DCM74, DMR74, Dau74, DEE^{+}75, ED74, FBL79, GG79a, GG79c, GG79b, GV77, GCP79, HO73, Hac74, Hal73, HKP79, HPK79, HSBD78, HS70a, HS70b, HL78b, JŻ79a, JR74, Joh75, KR71, KTO^{+}78, KPKC73, Kau73, Kau74b, KPP77, KPP78a, KPPH78, KPP78b, Kau79b, Koc74, KPD78, KP79b, LSOC77, LS74b, Meh79, MDC79, OMC76, OSL77, PCMS77, PT75, PPK78, PN76, PK71, PK75, PK76a, PK77b, PP74, PGPP79, PLMB73, PK73, Pul74, PA74, PBG76, PB77, PPB78, RTS74, RWFS71, RBDV77, RL79, Rou74, Sab75a, Sel75, SDJ79, SLC78, Spa78, TRS73, TS77, TS78, TS79, TB74, Tru73a, WJ79, ZSR79, BS75c, LD73, MBM73b, Tru73b]. injection [VH75]. Inner [GB71, Löw71a, Ahl77b, BÖ78, GBT79, GBL^{+}73, Löw70]. inorganic [MS76a]. insertion [JŻ79a]. instabilities [FMRG75, TTN76]. Instability [Mes79, Mes78, NS78, SN79]. Instrument [WBKL70]. insulating [Lip72]. insulator [DD77b]. insulator-to-metal [DD77b]. insulators [Bra76]. Integral [Ban74, BL77b, Kas76, TG76, BS75b, GCJ73, JW76, KT77, Lan74, PGQ77, PK75, PW74b, PW74a, RSS75, Rou79, San72, Tat73]. Integrals [Ave79b, AB79, Ave79a, BH71, GMZ73, Mos73, RB75a, Thu75, TL79, Abs72, AT75, Ave79c, BL77a, BBP71, Bil72, CHS77, Che71, Cof74a, Eng72, Gru70, HLGC73, Jet73, Kar70a, KZ73, KM74b, Lo75, LA76, MM74, Mat71a, Mat73a, Mat76c, PAM75, PPI73, RW74, SHI73, SR72b, Sha76, Sil71, SWN78, ST70, SR72c, SP77, Van74, GM73b]. integrated [Bow71]. integration [BEH77a, DD72, Hur77]. intense [Sch71b]. intensities [BL77c, MS71c, Sch76b, Sve79, WJ79]. intensity [BS71]. Inter [JS70].

Inter-level [JS70]. interacting [CS72, DT72, DDD70, GHS79, Mal71a, Meh78]. Interaction [BB74, Bas78, CC79, DEC78, Höl77, RM74, Sas74, SO77, TTS^{+}79, TL79, Abd77, AMJS79, ACK^{+}75, BS77b, BS78a, BP74, BK75c, BR72, BS78c, CJS77, Cha71, ČPH74, Cof74b, Dah78, DD72, Dav77b, DMS^{+}78, Dea72, DY73, DAP79, ESR74, ED79, FY79, Fro70, GHS79, HB79, HSBD78, HPK72, JMM71, JSC78, JSK75b, Kap79, Kau74b, Kau79b, KL77b, Kob78, Koł74, Koł75, Koł76, KD77, KKM70, KKK79, LED73, LD74, LL75, LB79b, LSB^{+}79, MCP73, MCF^{+}70, MSSF79, Moh79, NY76, OK75, PW79, PSK77, RSR72, Røe75, SD79a, Sas77, SRP76, SRP77, SDJ79, SE79, SBM72, SR78, SDB76, SS77a, TS71, TH78, Tru73b, Wei76, WV74, WP79, YW73, You75, Mal71b]. Interactions [BKW75, KL73, SKS77, TBB70, Abd76, AFD^{+}78, BRS^{+}75, BB72a, BB72b, BP75a, CW78, CW77, DY73, DY76, Dun78, FL79, GJP79, GCP79, Höl77, KM76, Kle70a, KMT77, KR78, KPP78c, LB79b, Lim74a, LW76, Loc77a, Loc77d, LRJ73, MME79, MGT79, MHRS76, MHRS77, MP75, MS76b, OCRMLR^{+}71, PLYM78, PD78a, PDEB78, Pol79, PBG76, PGDM77, Pul78a, RKE79, RM73, SRE^{+}79, SC78a, SBM70, Sun74b, Sun75b, ST78, ST79b, Suz77, VGS79a, VGS79b, Wei75, WO77, Wes73, WMV77, Ano76e]. interative [HR76]. Interatomic [RMP71, DY76, FY71, GK78, Kle70a]. intercalation [PL77]. interest [KG76, KG77, Mos73, Ran79]. interface [FBL79]. interfaces [Joh77, PBPL78]. interference [HLMV77]. Interfragment [VGS79b]. interhalogen [SM73a]. Intermediate [Rei72a, Ste74, BL79, BKW75, CL76, GLS79, Koł74, Koł75, Koł76, KK74b, Sac72]. intermediate-energy [CL76]. intermediate-field [KK74b]. intermediate-range [Koł74, Koł75, Koł76]. intermediates [Tho76, TPC77]. intermetallic [Gin78, PN71]. Intermolecular [Aro78, BB72a, Cho79, GCP79, MME79, MS76b, Pul78a, AM77b, BB72b, Brä70, Cla71, DCM74, DMR74, Dau74, FL79, HSBD78, Jeh73, Kap79, Kle70a, KR78, LW76, Loc77a, Loc77d, MGT79, Pol79, PGDM77, RKE79, RM72, RM73, SC78a, WB76, Dun78]. Internal [Bor71, GP72, MEW70, TBB70, WC75b]. International [Cal76b, BP70, BP71, BP72, BP73, Pul79b, Ede70]. interne [TBB70]. internuclear [CODM77, GZ72b, Mul74, Zer74]. Interpolation [Thu75, Sha71]. Interpretation [LBÅF78, DBJ74, Feu77, INY^{+}71, JJ78, Kot76, Woj78, JJ78, BEF78]. Interpretative [SLM70]. Interrelations [AMSS79, HYL78]. Interscience [Cal75c, Gos78a]. intramolecular [Bri73, LRD^{+}77, LRD^{+}78, Pol79, Rat78]. intrinsically [BMD70]. Introduction [Ano78e, Chr79b, Hur76b, Löw71b, Löw78, Sla71b, Sla73b, Bar78, JW76, Com77b]. Introductory [Nau77, Nau78, Wij78]. Invariance [SP79, FGV77, Kob78, NN77]. Invariant [CA72, Hal73]. invariants [BM71a]. inverse [Ste77]. inverses [Ra76a]. inversion [SKW72, SMMA73]. Investigation [Csa77a, ÅSW77, Coo78a, Ema74, KJ75, LÖS73, LAM71, MP71a, MPN79, MM72a, SLOKP79, UP71, VBC78]. investigations [TPC77]. involvement [WLS77]. involving [Ald75, DY73, RM73, Wel77a]. Ion [GRB74, CJ73a, CSH74, GZ72a, GZ72b, GGCK79, Goo76, HS70a, INI71, JS71, KKK70, KP71a, KSP78, Koł76, KD77, KD79b, LH74, Lar75, Lar78a, Lar78b, Lei74, LS79, MFRE76, MBM73b, PK71, RR77, Sun74b, Sun75b, Tal77, TWC73, Urr75, Woo72, Zer74, CJ73b, DT75, ST78]. ion-molecule [PK71]. ionic [Cal75a, GPC^{+}71, Höl77, RKE79, SM76b, SM76a]. Ionization [Hun75b, Mey71, ÅSW77, CSDV78, Con72, Gos75, Hal79b, HLGC73, HYL78, MVN75, MK72, Ner75, PT78, PK77c, PÖ77, RM79, SMH77, SG77a, TS79, DH78, Jak76d, MDS75, TS78]. ionizations [Ner76]. ionized [KTO^{+}78, PLMB73]. ions [AMJS79, BFS73, BT78a, Bar79, BS71, BM73, Cal73c, Cal74, Col79, CS74, Csa75, Csa76, Csa77a, Csa77b, Csa78b, Csa79a, Fro79, GGJ75, Joh75, KR71, Lin70, MHW78, MMS70, MM72b, PA74, PBG76, RW74, RGM77, Sim77b, SS71a, Tal77]. IRDO [BKW75]. iron [Fog71, Han79, JS70, Liu71b]. Irradiation [BC76]. irrational [Sur71]. Irreducible [KK74b, KK77c, BF70, Blo72, Blo73, BO73, BS74b, Sol73]. irregular [SR72c]. isethionic [CS74]. isocyanate [Mof79]. Isocyanide [Mof79]. isocytosine [CLP79b]. Isoelectronic [BG70, AC70, AS78, BK74, BG72, Mar73a, PD78b, SK71, SK72, SA77, TTS71, Tat73, Jak76b]. isoélectronique [Mar73a]. isolated [Dmi75a, NJF^{+}77]. isomer [GML77, Kot76]. isomeric [JE79]. isomerism [Sla79]. isomerization [BBG79, GP78, Mat78, Mof79]. isomers [Mof79, PLMB73]. isopotential [NS79a]. isotopic [Hal71]. isotropic [BM71a, Mal71b, TB74]. isovalent [AAC76]. Israel [Pul78b, Sti79]. iteration [BEH77b, Gia73]. iterations [KÖ71]. Iterative [Cas78a, JW76, AM72b, Als73, BEH77a, Goł79, TRPB77]. IUCHF [Vog79]. IV [BL77c, CP70, CT77, Hir70, Hos72, Jak76d, LCPP79, May78, PK77a, RTM77, Sch71e, SHR76]. IX [KPPH78, TS78].

J [Cal73a, Cal79b, Cal75c, Gos78a, RKOU77, Shu77b, Sla71b, Mak71]. J.-L [Shu77b]. Jacobi [TL72]. Jahn [MSSF79, Ham71, INI71]. Jansen [Suz77]. January [Sla71b, Sla73b]. Jeremy [Ano76n, Cus75]. Jersey [Cal76a]. Jerusalem [BP70, BP71, BP72, BP73, Cal73b, Pul78b, Sti79]. John [Cal79b, Cal75c, Dun78, Lin73, Shu77a]. Johnson [LB79a]. Julg [Cal79a]. June [Cal76b, Mar73b]. justification [Nov73].

K-shell [CV73, Lar78a]. K. [LB79a]. Karger [Cal76b]. Karwowski [Ahl77a]. Kashiwagi [Ste77]. Kenichi [Cal77b]. ket [Mor75]. kets [Kle74a]. keV [BW73]. kind [Lau79]. kinetic [Han77a, Nal78, TB78]. Kinked [VBC78]. KNiF [WN71]. Knight [DV79]. Kondo [JVM78]. Konvergenz [Ahl70]. Koopmans [Chi77, CT77, MDS75]. Kr [ES70]. Kramers [BMD70, Bow71]. KrF [SD78]. Kronecker [Lét74]. Kronig [RT77, WO76]. krypton [Bor74].

L [Cal75c, Sab75a, Shu77b, RRS75, AJK78, BP75a, Fae79, Mem77]. l-alanine [RRS75]. L-region [Mem77]. L-shell [Fae79]. L-Thyroxine [AJK78]. LaB [ACK^{+}75]. Lactam [Boy77]. ladder [HBB77]. Ladungsverteilungen [BKW75]. Laguerre [Mos73]. Landau [Csa78a]. lanthanide [NTN^{+}79]. lanthanide-substituted [NTN^{+}79]. Laplace [MKF77, SF75]. Laplace-type [SF75]. Large [ACK^{+}75, CSM71, Chr73, DM72, DEE^{+}75, Duk79, Hat75, KHI79, KPP77, NS79a, OMC76, PLYM78, PCMS77, PK75, Ra76a, RSR72, Rob72a, Sas77, Tin79, WCK^{+}76]. laser [DLG79, WS75, WLS77]. laser-Raman [WS75, WLS77]. laserlike [ESG79]. lasers [Fla79, MHWC78]. Latin [BP79]. Lattice [Enz73, MHRS76, MHRS77, OCRMLR^{+}71, CJCM73, CS72, KOW78, Kun78b, SP75c, Sch71e, SP75d, SEW73, TTN76]. LCAO [Abs72, Kau70b, LS74b, ADFL71, BR78, BDP73, BBL72, BBL77, Bil72, Del73, Del74, DCS77, FTN77, KS73a, Kar70a, KT77, KTO^{+}78, LSOC77, LAV72, Min79, NI70a, NI70b, PPK78, PD78a, PK76a, PK77b, SLC78, Spe72a, Spe73, SW78, Vla70]. LCAO-calculations [FTN77]. LCAO-CO [PD78a]. LCAO-MO [SLC78, SW78]. LCAO-MO-SCF [PK77b]. LCAO-X [DCS77]. LDH [BC76]. LDH-pyruvate [BC76]. lead [YR76]. learn [Fri72]. least [Ada73, BT71, HS75]. least-squares [HS75]. Lebenslauf [Ano77m]. Lecture [Cal79a]. left [SY77]. left-and [SY77]. length [Kob78, Sin73a, TTN76]. lente [BP79]. Lenz [Loh76]. Leu [BWRD79]. leur [Mar73a, TBB70]. Level [BF78, HU74, SH73, JS70, Sim77a]. Level-dependent [HU74]. levels [CH74, CH75, Csa78a, Dmi75a, HP77a, San71, TDB77, Wen79, WM79]. Levi [Ran79]. Levy [RCE79]. Lewis [LL72]. Li [Bar79, BS70, BS76b, Fro75, GS70a, Jak76b, KP71a, LL75, LE72, MM72b, MDS75, RM74, RS79, SP75d, SBD^{+}70, ZL78]. Li-like [Bar79]. liaison [JJ78]. Liapunov [SB75]. LiCH [RMB78]. Lie [Wyb73a]. LiF [Pai71, SBD^{+}70, ZF77]. life [Con73, SG76b]. lifetimes [BNLS72, BOP76, HE77, Naz72]. ligand [Bal71, Han79, LB79b, PBG76, PGDM77, SF76, Sch71a, Van72, WN71, LK78, PD79]. ligands [AMSS79]. Light [Sch76b, Sch71c, TS71]. LiH [BS74a, BS76a, BS76b, HH77, LG77, MM70]. like [AM72a, AM74, Bar79, BS71, GGJ75, Har78, PPI73, YD75]. limit [Kat72, OKM74, Ra76a]. Limitations [CAA^{+}75]. limited [APČ77, ČPH74, KMLB77, NRS79, PAČ77, WJ79]. limits [Wei71]. Linda [ÖL79b]. Linderberg [Shu77b, KSS73]. line [BMD70, Fog71, HMV77, HMV78, Kot76, WEH74]. linéaire [TGSL70]. Linear [Han77a, BBA73, BBT71, Com77c, DN75, GbK77a, Hur77, KB77, KL73, LR77, Liu71a, Nor75, Sol71, TK78, TGSL70, KA76]. Lineare [KA76]. linearly [Sol73]. linkage [Boy74, Har77a]. linked [LH76, Man77, SMP77]. linked-cluster [LH76]. linked-diagram [Man77]. LiNO [HR79]. Linus [Ano78f, Ano78u, Eyr78]. LiOH [NS79b]. LiON [HR79]. lipophilicity [Tic79]. lipoprotein [Har75]. liquid [Col79, HK74, DV79, Kho79, Nir75]. liquid- [DV79]. lithium [AS78, Bur72, CS73, RB77, Rei72a, SA77, SC73, Spe73, SDB74]. little [Gia73]. living [MBMM79, SGM78]. Lobe [Pos77, RS78c, Rou78, SC78b, RS78a, RS78b, RS78c]. Local [CKI75, Ive74, SMW71, ZF77, EH73, GSE78, HS75, HK74, JTR77, MR73, Sco71, TW77, ZF76, LB73b, PS75a]. local-mode [GSE78]. Localization [SPČ72, BO79, Eng71, LB77, Lud71, RC72, Wen79]. Localized [De 78, GH72b, SS77b, Car75, Cof74b, DGC77, DSC^{+}77, DD72, DS70b, Del73, DC75, GbK77a, GC77, Kau74b, Kle77, KGB77, LAV72, Lin71, Lud77, MMB70, MI78, Pol70, RPK78, SS77a, SK79]. localized-site [GbK77a, Kle77, KGB77]. logarithmic [PW74a]. loge [ACD^{+}74, Lud71, Lud77, MDC79, SDD^{+}77]. London [Com77a, Com77b, Lar76, Lin73, Mår78, Sab75a, Shu77a, Aba72, Ada77, LP74a, LP74b, Mal71a, Mal71b, Moh78, Wil77a]. lone [BKW75, DSC^{+}77, Kau74b]. Long [AFD^{+}78, DY73, Kle79, Koł74, Koł75, Koł76, SBM72, Abd71a, Cha71, Dea72, FY71, GPP73, KP71a, KL73, MOV75, PD78a, SZ76, Ukr72, WMV77]. Long- [Koł74, Koł75, Koł76]. Long-range [AFD^{+}78, DY73, Kle79, Cha71, Dea72, FY71, GPP73, PD78a, SZ76]. longitudinal [HDG71]. look [Tho75a]. Low [Hen79, AMJS79, ACK^{+}75, Boy70a, ČPH74, DM72, GLS79, Han75, KTO^{+}78, KGP79, MSSF79, PČH74, PCMS77, Tan74]. Low-energy-electron [Hen79]. low-lying [PČH74, PCMS77, Tan74]. low-molecular [KGP79]. low-order [Han75]. low-spin [KTO^{+}78]. low-temperature [ACK^{+}75, MSSF79]. Löwdin [Sch77a, CGLÖ76, Cro77, De 76, Duf71, Gra77, Jan77, JW78, KS73b, LCK76, Sco73, Ste77]. Lower [JP77, Lin70, LL70, BBK76, JP73, LTR74, SMB^{+}79, TTS^{+}79, Wei71, Gra77]. lower-bound [LTR74]. lowest [FG79, JŻ79b, Koł76, Mul71b, Pie71, SHR76, TTS^{+}79]. LSD [JG74, WCK^{+}76, WOEG78]. LSD/serotonin [WCK^{+}76]. luminescence [BYB79]. lying [ČPH74, PČH74, PCMS77, Tan74].

M [Ahl77a, Cal73a, Cal79b, Cal75c, Cal76b, SS78a]. m- [SS78a]. macromolecules [BVO72, BL79, KGP79, NGR^{+}77]. macroscopic [Fog71]. Magee [Cal75c]. magic [CA78]. Magnesium [LB79b]. magnetic [Ada75b, BMD70, Bow71, DL75, Fog71, HDG71, Hor71, JVM78, LL73, LZ77, Loc72, MF73, MSSF79, NTN^{+}79, PGPP79, Pul78b, PPI73, RKOU77, SKS77, Sti79, Wan79]. magnétique [Loc72]. Magnetism [Mat72a, Mat79b]. magnetization [Hub71]. Magneto [DR77, IM73, SH74, SH75b]. Magneto-optical [DR77, IM73, SH75b]. Magnon [Liu71b]. main [MOV75]. Mainly [BB73]. mammary [WLS77]. manganese [Han79]. manifolds [Dal79]. manipulation [ACH72, Car74, NGR^{+}77]. Many [BP78, CJCM73, DDD70, Feu77, Gás71, Lev78a, Lim76b, MVM75, AC76, Abs78, BS74a, BS76a, BWS77, BFH78, BB73, CHD71, ČP71, Dav72, DF76, FP77, Gru70, JMM71, JBP76, KL75, Lim74a, LR77, LÖ77, Lin78, LRJ73, Lun77, MR73, Man77, Mat76b, Mat79a, MK75b, MMM75, Ner75, NB78, RK76, Sha77b, SS73a, SS73b, Sla70b, WV74]. Many-body [BP78, DDD70, Feu77, Lev78a, BS74a, BS76a, BWS77, BFH78, CJCM73, LR77, Lin78, LRJ73, MMM75, Ner75]. Many-centered [MVM75]. Many-electron [CJCM73, Gás71, Lim76b, AC76, Abs78, BP78, BB73, CHD71, ČP71, JMM71, KL75, LÖ77, Lun77, Man77, Mat76b, Mat79a, MK75b, NB78, RK76, Sha77b, SS73a, SS73b, Sla70b, WV74]. many-neighbor [Dav72, DF76]. many-particle [Lim74a]. many-points [MR73]. many-shell [FP77]. map [Sri77]. mapping [TC70]. maps [HKP79, Kau73, NS79a, PPC72, SBB79]. March [BP70, BP71]. Markoffian [SM77]. Marois [Cal76b]. mast [MHH74]. master [ESG79]. Masthead [Ano70l, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano71i, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano72h, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano73j, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano77n, Ano77o, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano77u, Ano77v, Ano77w, Ano77x, Ano77y, Ano77z, Ano77-27, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano78s, Ano78t, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m]. Masthead [Ano79n]. Mathematical [CA78, SW78, KN79]. Matrices [Gal72a, LL70, SLM70, Spe72a, Ano78d, DR75, GH78, GP79, Han70a, Han75, Har77b, Har79a, ICS74, KP71b, Kry77, Lar73, Lar78a, Lar78b, Ra76a, Spe71a, Spe71b, Wel70]. Matrix [Dav74, Gal72b, HBB77, HL79, Har71b, KR73a, LS71, LS74a, MW74b, PN76, Sas74, Sha78b, Spe72b, Ada72a, Ada79, BG75, Bun77, CM72b, Cli73, CFMO73, DD72, Del74, Fox75, FM78b, GF76, Gar78, Goł79, Har70b, JW78, KL75, Kij72, Kry78, LS70, MS71b, MHRS77, MW75b, MSL73, Nis77, PW79, Pea74b, Pos78, Ros78b, Sco73, SDB76, SC73, Sur71, Tru72, WG74, Wit76, WP79]. matter [Lay77, PS75a]. Mavroyannis [Cha71]. maxima [TK78]. Maximum [MI78, DL75, Mak71, Ran74, Sch78]. Mayer [Cal79b]. MBPT [BS78b]. MC [KS72, KS75, KK77d, Lev70]. MCD [JPTM78, MSSF79]. McGraw [Lin73]. McGraw-Hill [Lin73]. MCHF [Fro75]. McIver [TB75]. MCSCF [BS76b, Gar78, RDH79, RCE79]. mean [TL72]. means [ETF79, Fin73, LG71, MI78, Nis77, Nor75, Pol79]. Measurement [BW73, Com74]. measurements [KKK70]. mechanical [CBV79, Che71, Coo74, DAP79, FJP70, GJ76, HS79, HSBD78, Hos78, KKK79, MP71a, MKT77, PP74, PGPP79, PS75b, PC76, Pul77, RW75, TPC77, Wal77, Wei71, WO74]. mechanics [Col77a, Con73, GP74, GT75, Hun75a, MM77a, NB72, PTUK69, PP71, Pul76, SE79, Shu77a, TL72, Van71, Wil79b, Wil79a, Cal79b]. Mechanism [PF79, BYB79, Dav79, HG75, KKK74, KK75, LP79b, PL79, Pul79a, Sal79, SM79, SLOKP79, TPC77, UP71]. Mechanisms [KKKP79, Pet79, BDT77, LSS78, TB75, Urr75, WO77, WOEG78]. Mechanismus [HG75]. media [BBT71, Chr79c]. medium [HLMV77, KPP77, Sim78b]. mediums [HMV77]. Meeting [Pul79b]. melting [KGP79]. membrane [Sun74b, Sun75b, ST78, ST79b, Urr77, Van75]. membrane-active [ST79b]. membranes [BK75b, Fro79, GH79, Hos78, Mar79, Urr75]. memory [Spa78]. meperidine [LJBN75]. MERGE [HKP79, HPK79, KPPH78]. Metabolic [NKKY79]. metabolism [ME78]. metabolites [HKP79, KPPH78, NKKY79]. Metal [Ber73, Swi71, BR78, BFS73, CS73, CHHS74, CS74, DD77b, DCS77, FN73, JW72, Joh77, KD77, KGB76, LVDC74, MS72a, MS74, Moh77, MM77b, SC73, Spe73, Sto71, Van72]. metal-ferroelectric [CHHS74]. Metal-to-nonmetal [Ber73]. metallic [KMO77, OKM74]. metalloporphyrins [DAC^{+}79]. metals [CJCM73, CM75b, De 78, FMRG75, KW76, LB73a, MGM78, PK77c, Sch77c, SKW72]. metastable [DD71]. metastables [KSP78]. Methadone [LJBN75]. Methane [Dan72, DPGC79, HSBD78]. methionyl [PPC72]. Method [Hen74, LS71, Rei76, Aba72, AFD^{+}78, BR78, BB75, BBL72, Bil72, BK75c, BFH74, Bur72, CS73, CK75, Cas78a, Cas78b, CD74, CODM77, CSY^{+}74, Cla71, CSH74, DC73, DRCT78, Eps77, FFCB79, FG79, FRM74, FHB77, GHS79, GB75, GML77, GL74, GJ76, GHJ77, Hal78, HS77, Har77c, HC72, Hen76, Hen77, HC78, Jan76, Jet73, JW72, Joh73, JM76a, Joh77, JM73, JS70, JNŽ72, KKW73, Kau73, KSP78, KA76, KR78, KU70, KK77d, LS70, LVDC74, LV71a, LV78, LP74a, LP74b, LFF79, LG77, Lim74b, Lim75a, Lim75b, Lim76b, Lim77b, LG75, Loc76, MKF77, Mak71, Mat71a, May74a, MK75b, MK75a, May78, MCF^{+}70, MS76a, Mes73b, Mes78, MW75c, Mon77, Mor78, MMR74, MS75, Nam75, NI74, Pai71, PD78a, Ran74, Rob71, Rob72b, RDH79, RK77]. method [Rue79, RPK78, Sac72, Sam75, SF76, SM76a, SH73, SC76a, SIM75, SKL79, Shu77c, SBM72, Sla73a, Sla74, SC76b, SC73, Spe73, Sri77, TGSL70, Tru73b, Van70, VK74, WEH74, WK73, WA71, Wei74, Wei76, YJ76, ZR78, ZBH^{+}78, ZF76, DCM74, KS72, LB73b, Wel76]. méthode [TGSL70]. Methodology [BW76b]. Methods [Chi77, CFMO73, GMZ73, Shu77b, SBD^{+}70, Tur74, APČ77, Ano79o, BG76, Bar70, BNLS72, BBL77, BS78b, BL77c, BS78c, CGLÖ76, CSFM72, ČP71, Cof74a, DS76b, DCS77, Ell73, Ema74, Fau71, Gar78, GV77, Har77b, Har78, Kar70b, Kar70c, KFY71, Kau70b, KK74b, KK77c, KVSG79, Lar78a, Lev78a, Lun79, MK76a, May78, MCM73, MW74b, Min79, ML76, NS72, NS73, PLYM78, PAČ77, PČL78, PBPL78, PSK77, Pos79, RW77, Sin73b, SS70, Sla75, Vog79, WFS77, Cal76a]. methyl [BS79, Kla78, MLB72, MK72, RRNS75, Sha77a]. methyl-substituted [Sha77a]. methylacetamide [DEC78]. methylamine [Abd77]. methylating [HPK79]. methylation [SS78a]. Methylbenz [SS78b]. methylene [LED73, LD73, TRS73]. methylglyoxal [Abd77, ED79, LDP79, Lew78, Pet78]. methylglyoxal-protein [Lew78]. methylimidazoles [KRKŠ78]. methyltransferase [SS78a]. Metric [Mor78]. Metric-constrained [Mor78]. Mg [PPB78]. MgH [BL70]. MgO [KGB76]. microelectronics [Ost79]. microscopic [Coo78a]. microwave [Web74, WLS77]. migration [FY79]. Milne [YL74]. MINDO [Boy77, SL73, ZR78, ZBH^{+}78]. MINDO/2 [SL73, ZBH^{+}78]. MINDO/3 [Boy77]. minicomputer [Spa78]. minimal [Cur78, KPÖ76, Mal71a, RPK78]. minimisation [BJ73, Muk75]. minimization [BJ73, HS75, Jol79, KS73a, KA76, KBK72, Wel76, Wel77b]. Minimum [GY75, HHR70, JŻ79a, MI78, You72]. mixed [ACG74, Euw71]. mixing [LV71a]. mixtures [GH72a]. MO [Ano78c, BBL72, BBL77, Csi77, KT77, Kau70b, KL79, Lar78c, LAV72, MW74a, NN77, PCMS77, PPK78, PK76a, PK77b, Vla70, WMB75, MM77b, Abs72, Bil72, CA78, CSFM72, Cof74a, Csi76, Del73, FTN77, FM72, Hil79, KHI79, KS73a, Kar70a, Kar70b, Kar70c, KFY71, KTO^{+}78, LS74b, MSKM77, MR72, Moh77, NI70a, NI70b, Sam75, SD78, SMH77, SG77a, SS74, SLC78, SW78, WRF73, ZR78, ZBH^{+}78, BFH74]. mode [DBBJ76, GSE78]. Model [KL77b, LPP78, LP79a, PD79, SPČ72, AF73, ADDLG76, ABD^{+}79, BS77b, BS78a, BL70, BFR78, BK71, Bra79b, CB72, CD78, Cho79, ČPP75, CE73, Coo74, Coo78a, Coo78b, Coo79, CZ75, Dav77b, Dec70, Doo78, Duk79, ERW75, ED79, FCL79, FM78b, GbK77a, GOK79, Goo76, GGJ75, Hal75b, Han77a, HB79, HC78, ICS74, JS71, JVM78, JL70, JM76b, Kib75b, Kli79, KBK72, Lar75, Lar78a, LP79c, LP79b, LW77, LB73a, LK78, LPT^{+}74, Lud75, Mal71a, MCP73, May74a, Mic76b, PČL78, Par72, Per78, Pet77, PNC^{+}78, PP74, Por77, PGPP79, RSS75, RR77, RMB78, RC70, RC71, RC72, RL79, SM76b, SPČ79, Sch71a, SP75a, Sha77a, Sha78a, SB78, SL70, Sti78, Tal77, TP77a, WN78, WO76, Wei76, YJ76, Sch71d, SPČ72, YD75]. modèle [JM76b]. Modeling [SKL79]. modelling [FHP70]. Models [LHK77, Pau78, WO77, AM77a, ČPH74, GK76, Höl77, JPTM78, Lad74a, LS74a, Löw77, Mag79, PL79, PČH74, PBS76, TW77, Urr77, WOEG78, SK73]. Modern [Kap79]. modes [BDT77, Hor71, SLOKP79]. Modification [Bee79]. Modified [GM73a, GM73b, KK77d, KL79, Rue79, BSG77, Csa77b, JW78, PS73]. MODPOT [HKP79, HPK79, KPP78a, KPPH78, KPP78b, PK76b, PK77a, PK77b]. MODPOT/VRDDO [HPK79, KPP78a, KPP78b, PK77b]. MODPOT/VRDDO/MERGE [HKP79]. MODPOT/VRDO [PK76b]. modulus [SEW73]. moiety [RKOU77]. MOLE [PHL^{+}76]. moléculaire [LL72]. moleculaires [TBB70, BJ73, Mar73a]. Molecular [AeAs78, Ald75, BB70, BL77b, BH71, CD78, Cur78, DH73, DLG79, DK77, Eis76, Eld73, FWFM77, FB72, FHP70, GPP73, GMZ73, GHJ77, JG74, KC73, Kau76, KPPH78, KPP78b, KL77a, Kot76, KD77, LV78, Lev70, LGR73, MS72a, MS72b, Mat73a, Mat76c, MHA70, Mul74, MHH74, NPL^{+}79, PPC72, PPP79, PK76b, PK77a, PK77b, Pul70, PZP79, SMP77, SS78a, SJ79, SR72c, TPM70, Urr75, Van72, Van74, AF73, AM74, Ano79o, BTD79, BNLS72, BL77a, BJ73, BBP71, BV73, BVS76, BT71, BT78b, BVR78, Boy70b, BS78c, BSB77, Cal76a, CM75a, CW76, Cas79, CSDV78, CJS77, CSM71, Chr73, ČP71, CR72, CDEO76, CM76, Com77c, Con72, CAMC71, CSH74, CD71, CZ75, CA72, DGC77, DSC^{+}77, DRCT78, DT75, DEE^{+}75, Ell73, Eng71, FTN77, FCL79]. molecular [Fla79, Flu72, GGCK79, GG79a, GG79c, GG79b, Gia73, GC77, GPP79, HO73, Hac74, HH77, HHR70, Hal78, Hal79b, HLMV77, HKP79, HLGC73, Hen74, Hen76, Hen77, HLYC77, HL78b, Hos72, HTG77a, HTG77b, IS75, JSC78, Joh70, JS71, JP77, KR76, KTO^{+}78, Kau70a, Kel75, KM76, KK77b, KGP79, KPD78, LL72, Lan77, LGAM79, LP74a, LP74b, LFF79, Lev78a, Lib76, Lo75, Loc77b, Loc72, MMB70, MEW70, MS74, Mat71a, MLB76, MDC79, Mic74, Mic76a, MHWC78, Mil74, Nal78, NB72, Nee75, Ner75, OMC76, PH77, PCMS77, PT75, PPK78, PK75, PGPP79, PS75a, PBC74b, Pul76, Pul78b, Pul79b, PÖ77, RR72, RWFS71, RK77, Rou78, Sab75a, SM73a, SD79a, SPČ79, SP75a, Sch72, SLB77, SLJ78, Sel75, SMMA73, SDM74, SBB79, Shu77a, Sil71, Sim72]. molecular [Sla71a, SC76b, SC78a, SC78b, SF75, SP77, SS73c, Sti79, SWP77, TBB70, TNB79, TP77b, TH78, Tur74, Tyu75, Tyu78, ULM76, VBC78, WO77, WOEG78, WJC70, Wil79b, Wit72, Wój76, Woo72, Fan70, KD79b, Mar73a]. Molecular-orbital [Van72, BTD79, CSDV78, Joh70, JS71, KPD78, KD79b]. Molecule [OJB77, Ada78, AF73, BRS^{+}75, BWS77, BG78, DN75, ES73a, FSV79, GZ72a, GZ72b, GGCK79, GLS79, GS70a, HH77, Hen79, HK74, IM73, JŻ79b, Jør78, KP71a, KSP78, KD79a, KL73, LD74, Laz79, LE72, MO77, MS76b, Min79, Nes75, Pie71, PK71, Rit76, SH74, TDB77, TOED79, WO76, Wil77b, Wil77a, Zer74, vDR79]. Molecules [Jai70, KM74b, Nes71, AM72a, AM74, AM77b, AM77a, Als73, AMJS79, Ave72, Bar77, BB74, BJ73, BBP71, BP74, BP75a, BSZ78, BP73, BR75b, BFR78, BS76b, BBG79, BW73, Boy74, Boy72a, CK75, Cha71, CSFM72, CW77, CSM71, Chr73, CR72, Csi76, DN75, DSMM79, DM72, Duk79, DCS77, ERW75, EB74, FW75, Fox75, GP72, GP78, GB75, GJLL75, GJ76, Hal71, Hal73, HE77, HF72, HRBB76, Hil79, Hur76a, Hur76b, Hur77, IM73, IMN79, ICS74, JBP76, Joh73, Jul78, Kla78, Kle70b, Koc74, KBK72, KL73, LL72, LL73, LCP76, LZ77, Li76, LR77, LM72, Lin73, LGR73, MMB70, Mal71a, Mal71b, MCST72, Meh78, MCF^{+}70, MLM75, MLB72, MOV75, Moh79, Mor78, MI78, Muf76, NI70a, NI70b, NS73, NS79b, Nes70, OMC76, PLYM78, Pau78]. molecules [PK72, PCMS77, PS79a, PS70, Plo79, PF74, PK75, Pos70, RR72, RSR72, RTS74, RM73, Rob72a, Rob71, Rob72b, Sam75, SM73a, SM76b, SM76a, SD79a, Sas77, SP74, SP75b, Sch73b, SG77a, Sim79, Sim78b, Sla75, SZ76, SBD^{+}70, Suz77, SH75b, TGSL70, Tin79, TB73, Tyu75, Tyu78, VGS79a, VGS79b, WCK^{+}76, Wen79, WS79, WMV77, BJ73, LL72, Loc72, Kle70b, Com77a, Com77b, Ano78c, Cal79a]. Molecules-in-molecules [KM74b]. Moleküle [Sch73b]. Møller [BP75b, PKSB79]. Moment [Ell77, AAC76, Als73, CA72, DSC^{+}77, Fae79, KP71a, MMD78, Tru72]. Moment-polarized [Ell77]. Moments [OJB77, CAMC71, DH76, ETF79, HH77, Lar78a, Lau79, Li76, RR72, SMW71]. momentum [BS70, BBT71, BW73, But76, Cro77, CM75b, GN77, WESW73].

Monday [Sla71b]. mono [Har77a, INI71]. mono-negative [INI71]. Monocyclic [Boy77]. monoelectric [BSG77]. Monographs [Gos78a]. monolayers [VH75]. monomer [AZ72]. monomolecular [Chr77]. monopole [KP71a]. monopole-transition [KP71a]. monosubstituierten [FJP70]. monosubstituted [FJP70]. Monosubstitutions [MR72]. monoxide [BWS77, SD70, Wil77b, WFS77]. Monte [Cas78b, Col77a, CL78, Col79, RC78]. more-electron [Sin73b]. Mössbauer [GML77, Kot76]. motion [DR75, FG79, GP79, LG71, SHI73, SIM75, SD74, SLOKP79, SH75b, Web74, Sim77a]. motions [Bri73]. Mott [Bra76]. Mouton [AB73]. moving [DR77]. MS [GML77, BFH74, Con72, CG79b, KC79, Kau73, PKK75, PK77c, WC75b]. MSW [SP75a]. MSX [RK79]. MT [CD78, KC79, RK77]. muffin [DC73, FFCB79, FRM74, GHJ77, Pai73, Sla74, Wil74]. muffin-tin [FFCB79, GHJ77, Sla74, Wil74]. Mulliken [CM75a, CA78, Kau70b]. Multi [MM70, Hac74, NI70a, NI70b]. Multi-configuration [MM70, Hac74, NI70a, NI70b]. multibody [NY75]. Multicenter [GMZ73]. Multichannel [RM79]. Multiconfiguration [Har71b, SC76a, BG76, BSS78, BKS79, Fro74, GB75, Hal78, Har70b, JS77, KKW73, PNC^{+}78]. Multiconfigurational [GBV77, Gre71, KK76c]. multidimensional [BF78, LG71]. Multimolecular [Sla79]. Multiple [AD77, Dav77a, Wil74, BL77c, CW76, CW78, Dan72, DC73, FFCB79, GGCK79, Gin78, KL78, Kel75, LVDC74, Li76, MMR74, RK77, WK73]. multiple-scattering [Dan72, FFCB79, Li76]. multiple-scattering-X [DC73]. multiplet [BFS73]. multiplets [LP74b]. Multipole [CJ73b, DH76, GPP79, MM72b, SP79]. muons [Fri72]. muscarinic [Kan74]. muscle [Dav79, NTN^{+}79]. Musher [Ano76n, Cus75]. mutation [GRE^{+}77]. mutations [Coo74, Coo78b, RCE75]. Mutually [Abs72]. my [SG76b]. myosin [BZSD79].

N [PK71, RKOU77, SY77, AAC76, Bur77, DEC78, Har70a, HLYC77, KPÖ76, MLB76, PL79, PK71, RRNS75, RC71, SBD^{+}70, Tru72, WK73]. N-acetyl [RRNS75]. n-h [SY77]. N-hydroxylation [PL79]. N-methyl [RRNS75]. N-methylacetamide [DEC78]. N3 [ME78]. Nakajima [HLMV77]. nalorphine [KK77a]. Nanosecond [GC75]. naphtalene [BOP76]. naphthalene [KT77]. narcotic [KK72, KKK74, KK75]. narcotics [KM72, KKK74, KK75, LJBN75]. native [FSV79]. Natural [BS70, BC79, KÖ71, SD74, Eig78, RCE79, SDB76, TRPB77, ICS74]. Nature [KBKČD78, Sim78c, BVO72, BM71b, MPS70]. Naturel [DQ77]. Nb [MM77b]. near [De 78]. nearest [SS77a, SS77b]. nearly [BL70, BEF78, LB73a, SM73b]. necessary [Kum77]. necessity [Kau79b]. need [Del74]. negative [BT78a, Csa77b, Csa79a, INI71, KSP78, Sim77b]. neighbor [Dav72, DF76, MHRS76, MHRS77, OCRMLR^{+}71, SS77b, WJC70]. neighborhood [KL73]. neighbour [SS77a]. Neon [Fae79, NY76]. nerve [BK75b, NM77]. nested [MHPS70]. networks [KHI79]. neuroleptics [KK74a]. neurotransmitter [Kau76, KKP77]. neurotransmitters [GRB74]. neutron [MPW75]. next [SS77a, SS77b]. next-nearest [SS77a, SS77b]. NH [Bai74, BWRD79, CE73, FMZ70, HNT78, KC72, LV71c, RC70, RC71, Bai74]. Ni [JR74, LB79a]. nickel [CZNL70, DCS77, SDJ79, SLC78]. NiH [GBK77b]. nitric [KÖ71]. nitrite [HCIE75]. Nitrobenzene [PK76b]. nitrogen [BKW75, HK74, LD74, Mul74, NPL^{+}79]. n’l [JSK75b, JSK75b]. NMR [JE79, KVSG79, Pul77]. NO [PK71, Cal73a, Cal76a, Cal77b, Lin73, Sab75a, Cal79a, KKK70, LK78, MSSF79, Cal75b]. noble [BM74, Csa77b, MHW78, NY75, Sto71]. noble-gas [MHW78]. Non [DC73, Gal72b, HL79, JBG70, KP71a, Sun77, BBT71, BMD70, Bow71, Boy70a, DT72, FRM74, GP72, Gou73, HLYC77, JG72, Kun78a, KOW78, Kun78b, NB72, Pai73, Sch72, Sin75, TK78]. Non-Adiabatic [HL79]. Non-applicability [KP71a]. non-Bloch [Kun78a, KOW78, Kun78b]. non-closed-shell [Sin75]. non-crossing [NB72]. non-empirical [HLYC77]. non-existence [TK78]. non-exponential [Sch72]. non-Hermitian [Gou73]. non-interacting [DT72]. non-Kramers [BMD70, Bow71]. non-linear [BBT71, TK78]. Non-muffin-tin [DC73, FRM74, Pai73]. Non-Orthogonal [Gal72b, JBG70, JG72]. non-rigid [GP72]. Non-Watson [Sun77]. non-zero [Boy70a]. Nonadditive [Wei76]. Nonadiabatic [YD75, BC79]. nonalternant [Lai73]. nonbenzenoid [SG77a]. Nonbonded [Sun74b, Sun75b, ST78, ST79b]. noncanonical [GP74]. noncharacteristic [MFRE76]. Nonclassical [Urr78]. nonclosed [SS73b]. noncompact [Sin73c]. Nondegenerate [Hor71]. Nondiagonal [RM73, Oni79]. Nonempirical [HPK79, ABD^{+}79, KPP78a, KPPH78, KPP78b]. nongeometrically [DN75]. nonlinear [CK75, Wel75]. Nonlocal [Ove73]. nonmetal [Ber73]. nonmetallic [JMP78]. Nonorthogonal [JSK75a, JSK75b, Lya78, Meh78, Wel76, Wel77b]. nonorthogonalized [KR73a]. nonplanarity [RKOU77]. nonpolar [SAM79]. nonradiative [HE77, HF72]. nonreactive [Ber76, Tru76]. Nonrigid [DN75, TDB77, ES73a, GP78, PS79a]. nontensor [Bri73]. Nonzero [PBC74a]. Normal [KKP77]. normalized [Nal78]. notation [KL77c]. Note [Csa79b, Sla70b, Bee72, CW77, FHB77, FY71, HK74, Lad75a, LL76, Sab75b, Sco73, Sim77a]. Notes [FM78b, Cal79a]. nouvelle [JJ78]. novel [FBL79]. Nuclear [RKOU77, TB73, Boy72b, Bri73, GP79, Laz79, RMP71, Pul78b, Sti79]. nuclei [LZ77, PGPP79]. nucleic [AMJS79, DS79, FB72, IZZ^{+}79, KD77, KL79, LP79c, LSB^{+}79, NKKY79, PL77, Pol79, PGPP79, SD77, SMB^{+}79, SRE^{+}79, Sun74a, UP71, VBC78]. nucleic-acid [FB72]. nucleophile [PDEB78]. nucleophilic [PLSL78]. nucleoside [KP79b, PB77, SMP77]. nucleosides [GPP73, JE79]. nucleotide [Lak75, ME78, SW79]. nucleotide-binding [Lak75]. nucleus [HB79]. number [KL77c, KMLB77, MW77, MK76a, Res78]. numbers [GCJ73, Har79b, ICS74]. Numerical [GMZ73, KP71b, Thu75, BEH77a, CHS77, DD72, GJ76, RS79, Sur71, WA71, BBL77, DD72, ES79, TW77]. Nutrition [Web74].

O [BMD70, FM78a, Har70a, Har77a, LGP75, LGM77, Shu77b, ÅSW77, BZSD79, Bur77, CV73, DMS^{+}78, FM78a, FMZ70, Gub79, HL79, Han79, HNT78, HPK72, LG77, LE72, LK78, MS71b, MG77, PK71, Ros78a, ST72, SS78a, TTS^{+}79, Tru73b, WK73]. O-methyltransferase [SS78a]. O. [Jan77]. Obituary [Löw73]. observables [Kun78b, ML79]. observations [GK76]. obtained [BB75, CESW70]. obtaining [Cas78a, CFMO73, CH74]. occasion [Cal75b]. occupation [Har79b, ICS74]. occurring [DDH78]. octahedral [LVDC74]. octupolar [MMS70]. Odd [May74b]. Odd-electron [May74b]. odorant [LM72]. off [LC77, Mil71, ST79a]. off-diagonal [ST79a]. OH [Cur77, LV71c]. Öhrn [Shu77b]. old [Tho75a]. Ole [ÖL79b]. Olov [CGLÖ76]. on-line [WEH74]. One [BSS78, BDP73, BH71, Eng72, SBM70, TNB79, Ada72b, AM72a, AM74, AM77b, AM77a, AT75, CB72, CD78, ČP77, CM76, Dah78, DSG74, Dec70, FY79, Fri72, GC77, HNT78, Hen78, Hos72, HR76, Jak76a, Jak76b, Jak76c, Jak76d, JM73, KZ73, KS72, KK77d, Lét74, May78, MVM75, MLM75, SS73b, ST70, SBM72, SR72c, MM74]. One- [BH71]. one-band [CD78]. One-body [SBM70]. One-Center [Eng72, AM74, AT75, CM76, HNT78, MM74]. one-centered [MVM75, MLM75]. one-centre [AM77a, KZ73]. one-densities [Hen78]. one-dimensional [FY79]. One-electron [BDP73, TNB79, AM77a, CB72, ČP77, Dah78, Dec70, GC77, Hos72, HR76, Jak76a, Jak76b, Jak76c, Jak76d, JM73, KS72, KK77d, May78, ST70, SR72c]. One-particle [BSS78, DSG74]. only [Lét74, MPW75]. onto [MHPS70]. Open [CGRC74, APČ77, BS78c, GHS79, GBM74, GRM75, HP77b, JD74, KBK72, Lin78, MMB70, MW75b, MLB72, MM72b, MDS75, MMM71, Muk75, Nel79, Nes77, PT78, RGM75, RGM77, SPČ79, SM77, CDEO76, Hen76]. open-shell [BS78c, GHS79, HP77b, JD74, Lin78, MMB70, MLB72, MDS75, Muk75, Nel79, Nes77, PT78, RGM75, RGM77, SPČ79, CDEO76, Hen76]. Opening [Ede70, Lon74, Lon75, Lon73]. Operator [BL72, Mic76a, Mor75, BBK76, CE73, Dal79, DC75, HBB77, KK74b, KK77c, Ner76, PS75a, Sur71, WG74, Sim77a]. Operator-Level [Sim77a]. Operatoren [KA76]. operators [Bal78, BSG77, BFH78, Col78, CEK72, Cro77, Gos75, KA76, Lar75, Lar79, Lim75b, MK78, MW75c, SL79, Sch71a, Sol73]. opiate [LJBN75]. Oppenheimer [CL78, Col79, Ess77]. Optical [PS79a, BBT71, DR77, GML77, INY^{+}71, IM73, Lip72, Mes73a, Sch71e, SH74, SH75b]. optically [Cha71]. Optimal [Raf75, Ra76b, Abs78]. optimisation [Muk75]. Optimization [BT71, Gia73, ABM78, BSZ78, Bur77, CLP79a, CLP79b, Gar78, Jak76a, Jak76c, LS74b, MK75b, RCE79, Sun75a, Wel75, BPTLVM73]. Optimized [Sol71, Jak76a, Jak76b, Jak76c, Jak76d, JP73, LTR74, SBB79, SBD^{+}70]. optimizing [KK76c]. Optimum [Cas78b, Pol70, ZR78, ZBH^{+}78, CHD71].

OPW [VK74]. Orbit [BOP76, Lo75, MSSF79, BP74, BP75a, BK75c, Hal75b, JSK75b, LA76, MM74, MCF^{+}70, SM71a, SM71b]. Orbit-orbit [Lo75]. Orbital [ABM78, Ave79b, AB79, Ave79a, BL72, GMZ73, HRBB76, MVN75, Muk75, Abs78, AeAs78, AF73, Ald75, ADDLG76, Ave79c, BTD79, BBP71, BT71, BC79, BVR78, BSB77, Cal76a, Cas79, CSDV78, CR72, Cof74b, CDEO76, CM76, Com77c, CAMC71, Cur78, DH73, DEE^{+}75, FB72, Gar78, GPP73, GC77, GH72b, GHJ77, HO73, Hal78, Hal79b, HLGC73, Hen76, HLYC77, HL78b, HP77b, Joh70, JS71, JG74, Joh77, KC73, KTO^{+}78, KPKC73, Kau79b, KSL76, KL77a, Kot76, KD77, KPD78, KD79b, KÖ71, LV71a, LB77, LGR73, LS74b, Mar73a, MEW70, MHA70, MH72, MG77, MHH74, NPL^{+}79, Nee75, NG73, PPC72, PK75, Pop71, Pul70, PBC74b, Raf75, RR77, RMB78, SM73a, SMP77, Sch73b, SRP76, SRP77, SDM74, SS78a, SM73b, TRPB77, Tur74, Van72, VBC78, Vla70]. orbital [WJC70]. orbitales [BJ73]. Orbitals [Bee79, Gal72b, GMZ73, Kas76, Thu75, AC76, Abs78, BS74a, BS76a, BJ73, BP74, BS70, BSZ78, Bur72, CW77, Cof74b, CM76, CA72, CA73, DGC77, DSC^{+}77, DD72, Des72, DT75, Eng71, FGV77, Gia73, GB72, Hal73, HLGC73, Hen74, Hen77, Hos72, HTG77a, ICS74, JP73, JP77, JG72, JSK75a, JSK75b, KR73a, KT77, KP72, Kau74b, Lec76b, LAV72, LGR73, Lo75, Loh76, MMB70, Mar73a, MS72b, Mat73a, Mat76c, MK75b, Mcd79, MI78, Mul74, NN72, NN74, NS72, Pal73, PS73, Pol70, PAM75, Ra76b, Rob72a, Ros78a, RS78a, RS78b, RS78c, Rou78, RCE79, RPK78, SR72b, SD79a, SBB79, ST70, SC76b, SDB76, SD74, SR72c, SW77, SW78, Tat73, TP77b, Tyu75, Tyu78, Van72, Wyb73a, JBG70]. Orbitalstruktur [Sch73b]. orbiting [SSM75]. Order [JBP76, OJB77, SLM70, WS79, Abd71b, BBB73, BWS77, Bun77, Coo79, HS76, Han75, HBE78, KD74, Kle79, KP78, Lan73, LF77, Laz78, LZ79, Moc74, Mos73, NRS79, PLMB73, RM73, SK71, SKL79, SC73, TTN76, Vog79, You75]. ordered [PN71]. ordering [AS78]. orders [KHI79, Mal71a, Mal71b]. ordination [FM72]. organic [Ano78c, Csi76, Csi77, DT75, Lar78c, RD79]. Organization [GH79, Eig78, Pea74a]. organo [MYTC77]. organo-silicon [MYTC77]. organometallic [RMB78]. orientatiion [TBB70]. Orientation [Fuk70, Jør78, Cal77b, Eno76, Fuk75, TBB70]. Origin [Sur71, GRE^{+}77, CA78]. origins [PL77]. Orthogonal [CH77, Gal72b, JBG70, JG72, MS75]. orthogonality [RS76]. Orthogonalization [KS73b, Ra76b, LB77, Sco73, Ste77]. orthogonalized [CWE71, ML76, NN72, NN74, NS72]. orthonormalisation [Lat76]. orthonormality [MI78, NI74]. orthonormality-constrained [MI78, NI74]. Orthonormalization [Jør77, Jør78, Sch77a]. Oscillator [WG74, AC70, BHLM78, CM78, GBV77, Lar78a, RB75a, ST79a, Sin73a, WO75, Wit76, WM79]. oscillators [Bas78, BBA73, BHLM78, HMV77, HMV78, Hun74]. other [Boy74, Har77b, Jan77, KKK74, KK75, LDP79, RU76, SM71a, SM71b]. outer [Des72, GBT79]. Overlap [Thu75, BB72a, BKW75, DL75, DY76, GPP79, KP76, Kib75b, LV71b, LV72, Mak71, MI78, PNC^{+}78, PK75, Ra76a, RR72, Ran74, RM72, Sch71a, Van70, WW71]. overlapping [CG79b, Ra76b, SC76b, YJ76]. overlapping-sphere [SC76b, YJ76]. overlaps [Nis77]. overview [Sim78a]. Oxford [Ahl77a, Ano78c, Gos78b, Sab75a]. oxide [KÖ71, PDEB78]. oxocomplexes [BL77c]. oxopyridine [KPP78c]. oxopyridine-water [KPP78c]. oxopyridines [CLP79a]. oxopyrimidines [CLP79a]. oxy [RL79]. oxycumulenes [LÖS73]. oxygen [ÅSW77, DAC^{+}79, Har71a, HK74, IM73, KC77, LGAM79, Sal72, TH78]. oxygenases [Har77a]. oxyhemoglobin [Har77a]. ozone [Bur77]. ozonides [Rou73, Rou74].

P [Gos78b, LL72, Sab75a, Sim78c, AMSS79, CT79, LHK77, MK72, NT79, SS78a, AC70, GL74, Har77a, LL75, LG75, MSM70, MBM71, SHR76, TTS71, Tat73]. p-450 [LHK77]. P-450-dependent [Har77a]. p-benzoquinone [MK72]. p-methylation [SS78a]. p-quinones [NT79]. P. [Jan77]. P.-O. [Jan77]. P450 [Han79, PL79, RL79]. Padé [GBL^{+}73, ČPP75, CM78, GB71, PW74a, Røe73, Røe75]. Padé-type [PW74a]. Pages [Sab75a, Cal73a, Cal76a, Cal77b, Lin73, Shu77b, Shu77a]. Pair [GH72a, Nes71, Ada78, BP78, BKW75, ČP71, ENRM78, JMP79, KP76, Lan73, LE72, Mag79, Meh73a, MSL73, MMM75, NS75, Nes70, PNC^{+}78, PZP79, RC70, RC71, RC72, SE79, Sin73b, Sri77, WV74, SPČ72]. pair-correlated [WV74]. pair-correlation [Lan73, MMM75]. pair-potential [Meh73a]. pairing [JPTM78, Kle79, MWY77, Mat79b]. pairs [BM74, DSC^{+}77, GRE^{+}77, Jan76, Kau74b, KKK79, Sun77, ZV79]. pairwise [Pet77, PNC^{+}78]. pairwise-correlated [PNC^{+}78]. panel [WN71]. papain [BCR^{+}78]. paper [Wil77a]. paperback [Gos78b]. papers [LB79a]. para-phenylenediamine [DT75]. parallel [BCL79]. paramagnetism [BVO72]. parameter [Amo72, BS71, Höj79a, Höj79b, Jol79, MSKM77, RK79, SD79a]. parameters [AD77, BYB79, EB74, FN73, Har71c, KP72, LED73, LD73, NMS77, San71, Sel75, Sol73, TK78, Wel75, Wel76, Wel77b, Wol71]. parametric [Kar70b, Kar70c, KFY71]. Parametrization [LV71a, Kib75a]. parametrizations [Gia73]. parent [KKK70]. parentage [KK77b]. Paris [Cal76b]. Pariser [JL70]. Parr [JL70]. Parseval [Lat77]. part [GBS70, Sla70b, Sur71, Wel70, Wel71]. partenaires [TBB70]. Partial [Bar79, AM77b, AM77a, RK79, Wel71]. partially [Sim72]. Particle [Lar75, WA71, Ada79, BSS78, Col78, DR77, DSG74, GGJ75, GK76, Hen78, LD71, Lim74a, Lud75, Mat74a, Rat71, TGG79, MWY77]. Particle-hole [Lar75, LD71, Rat71, MWY77]. Particles [GÖ68, GÖJ^{+}73, KL77c]. particularly [Abd71b]. partition [DH78]. Partitioning [Mil74, MSL73, SDM74, BB71, BF78, GGCK79, GS70b, LG71, MMM75, Ner76, Røe73]. partners [TBB70]. parts [KR76, Kla78]. parvalbumins [NTN^{+}79]. path [JŻ79a, Lei74, PGQ77]. path-integral [PGQ77]. pathways [TH78]. pattern [KP76, Pal75]. Pauli [Kij72, RL77]. Pauling [Ano78f, Ano78u, Eyr78, Ram78]. PCILO [Kau74b, KK77a, KVSG79, Loc76, LW76, Loc77a, LH77, Loc77c, Loc77d, LH79, Wel77a]. PCILOCC [Loc77b]. PCILON [DD72]. PCN [Tho76]. PDC [DT75]. peak [SKL79]. peculiar [Mes79]. peculiarities [LKK^{+}79]. Peierls [Cal77a]. penetration [YD75]. Penney [RT77, WO76]. Penning [RM79]. pentapeptide [RU76]. pentofuranosyl [JE79]. peptide [BWRD79, RKOU77, Sri77, VGS79b]. peptides [NJF^{+}77, PP74, RRNS75, RKOU77]. Per-Olov [CGLÖ76]. perfect [Kun78a, KOW78, Kun78b]. periodic [Lad74a, Oda73]. permanganate [MBM73b]. permeation [Urr75, Urr77]. permitted [CSFM72, MCF^{+}70, MCM73]. permutation [LP74a, BDT77]. Permutational [Kry77, Rou79]. permutations [PW79, WP79]. peroxotitaniumporphyrin [RBDV77]. perspective [Bas75, SW79]. pertaining [Sal79]. Perturbation [ANR75, Ada72a, Ada79, CW78, CT77, DY76, Eps77, IS75, PGQ77, SA77, SS71a, Abd71b, Ahl70, AC70, AM72a, AM74, Amo72, Bar70, BS74a, BS75a, BS76a, BWS77, BP78, BB72b, BB74, BV73, BVS76, BK75c, Brä70, BO79, Bur73, Car72, CJS77, Cla71, Coo79, CSH74, CZ75, Dmi75a, Dmi75b, ES79, ETF79, GZ72a, GS70b, Gra77, Hir70, Hos72, JS77, JBP76, JSC78, JM73, Kle70a, KP78, Kut78, LWA74, Lin78, LPT^{+}74, LG71, MCST72, MMM71, PSK77, PLMB73, SS73a, SS73b, SS70, SC78a, Suz77, TS72, Vla70, Vog79, WRS70, WB76, Wil77b, WS79, Dmi75a, Dmi75b, MLB72]. perturbation-variation [TS72]. Perturbational [Pie71]. perturbations [Bow71, HS76, Mil74, SDM74]. Perturbative [DD72, DCM74, DMR74, Dau74, Kau74b, Koc74]. perturbed [LA70, Laz78, LZ79, WAL71, Woo72]. pest [Wyb73b]. Petraschen [Cal73a]. Petrov [Jan76]. PGF [KD79b]. PH [BT78a, SR77]. pharmacological [ZG79]. pharmacology [Kau77, PT75, Ric77, Van75, Mår78]. pharmacon [Höl77]. Phase [LP74c, ÖL78, SKW72, ČZ79, DBJ74, Enz73, Lar75, Nel79, NMS77, ÖL79a, Sab73, Sti78, SO77, Tru73b]. phenanthrene [Chr73]. phenethylamines [PBC74b]. phenolic [KPP78b]. phenols [DMS^{+}78]. phenomena [BEF78, BO79, CK73, LH75, Loc72]. phénomènes [Loc72]. phenoxyacetic [BC75]. phenylene [INI71]. phenylenediamine [DT75]. PHF [Lun79]. Phonon [Men71, ACK^{+}75, HDG71, LW77]. phonon-plasmon [HDG71]. phonons [MPMK79, TW77]. phosphate [LCPP79, PPB78, SLOKP79]. phosphatidylethanol [VH75]. phospholipid [USK^{+}79]. phospholipids [FL79, GH79, PS75b]. phosphorus [Cha77]. photoabsorption [Fae79]. photobiological [CS76, SCYB77]. photochemical [Kli79]. photochemistry [KBKČD78]. photodetachment [ML76]. photodimerization [SM79]. photoelectric [Ave75]. Photoelectron [DH78, Gos78a, KRKŠ78, GML77, Hal75b, KKK70, Kla78, Lun77, PLSL78, Rab77, WJ79]. photoemission [Dav77a, Mes73a]. Photoinduced [Mar79]. Photoionization [LGAM79, Bar79, Chi77, Lan77]. photon [MKF77]. photons [Ave72, Lin73]. photoreactivity [SCYB75]. photoreceptor [Eno76]. photosynthesis [Chr79a, CD76, KKKP79]. phthalocyanine [MS72b]. phthalocyanines [MS72a, MS74]. Physical [GB71, GBL^{+}73, HL78a, ML79, Pea74a, Pet77, Coo78b]. Physicochemical [KK75, KKK74, Kau79a]. Physics [Gos78a, Cal76b, Ess77, Fri72, DV79, Mar73b, Sin73c]. phytochrome [SCYB77]. pi [NN72]. picoline [BKW75]. Picolins [BKW75]. picture [CSDV78]. pigment [Kli79]. pigments [MP71a]. pilocarpine [Kan74]. Pioneers [Par77]. planar [Tan74, CSR79]. plane [CWE71, Euw71, KMO77, ML76, OKM74, WFS77]. plane-wave [CWE71, KMO77]. plane-wave-Gaussian [Euw71]. plant [BC75]. plasma [AJK78, Har75, Hor71]. plasmon [HDG71]. Plenum [Shu77b]. Plesset [BP75b, PKSB79]. plethysm [MCM73]. plethysms [MVM75, MLM75]. plots [SY77]. PNC [Tho76]. PNO [Bur77, Mey71]. PNO-CI [Mey71]. Point [BG70, FCL79, MM77b, SC79, ACD^{+}74, CB72, ČPP75, Flu72, Kib75b, KK77b, LP74a, Lét74, PBPL78, RCE75]. point-charge [Kib75b]. Point-group [MM77b]. points [BF78, MR73]. Pointwise [Lev78b]. Poisson [Sil71]. polar [Jør77]. polarisability [Sri77]. polarity [PS79b]. polarizabiities [HTG77b]. polarizabilities [ANR75, Ahl74, Ahl77b, Bar77, BSM72, BM73, CG79a, ETF79, GBM74, HH77, HS77, HTG77a, KKW73, LCP76, LG70, Mal71b, MM72b, MMM71, RB77, RB75b, SSH78, TSE76]. Polarizability [LP74c, BR72, MCST72, NAR72, RGM75, Zer74]. Polarization [BKS79, DM72, Hal71, HCIE75, OJB77, Sti78, Boy70b, CJ73a, CJ73b, KC77, Kle74b, KBKČD78, LW77, MLB72, SSH78, YD75]. polarizations [RGM77]. polarized [Bor70, Ell77, Fro70]. pole [CG78]. Poly [PPP79, VU79]. poly- [VU79]. polyamino [NJF^{+}77]. polyatomic [Als73, BBP71, BP74, BP75a, CSFM72, HF72, Jør78, Kle70b, LCP76, LZ77, MMB70, MCF^{+}70, MLB72, Muf76, Plo79, RR72]. polycyclic [Mem77, Mem79, PD77, SR78, UP71]. polycytosine [LSOC77]. polyelectronic [KPJM70]. Polyene [AL71]. polyenes [KHI79, LV71b, LV72]. polyglycine [KL77a]. polyhedral [PKK75]. Polyhedric [RS78c]. polyions [RKE79]. polymals [Pos78, Pos79]. polymer [DDH78]. polymers [ADFL71, ADDLG76, AFD^{+}78, ABD^{+}79, BP73, Del74, DD77a, Par72, PD78a, SKL79]. polymethines [FTN77]. polynomials [Ra76b, Ra76b]. polynuclear [DEE^{+}75, NT79]. polynucleotide [SY77]. polynucleotides [KP79a]. Polypeptide [Urr77, ULM76]. polypeptides [Kar78, KL77a, PPC72, Urr78]. polypolar [SF75]. Pople [JL70]. population [Kau70b, KPKC73]. populations [KP76]. porphin [Alm74]. porphine [KTO^{+}78]. porphins [MW74a]. porphyrins [Han79, MW74a]. portée [TBB70]. Position [Csa75]. positive [BT78a, Csa77b, Sco73]. possibilities [IZZ^{+}79]. Possible [Tic79, WLS77, Coo78a, Gol70, HPK79, HK74, Kau74a, LSS78, SY77, Wei74]. potassium [CD74, Sun75b, Sun74b]. Potential [BR75b, Cof74a, Gub79, GBK77b, HL79, HG75, Jai70, Liu71a, ÅSW77, ABD^{+}79, BA78, BNLS72, CL77, CL78, Col79, CG79b, CH74, Csa76, Csa77b, Csa78b, DR77, DRCT78, DSG74, DEC78, Gás71, HS79, HKP79, HSBD78, HC78, Jor75, KP71a, KD79a, LC77, May78, Meh73a, MVN75, MO77, MS76b, MK72, Moh78, Moh79, NS73, Nee75, OMC76, PPP79, PT75, PPK78, PK71, PD78b, Pol79, PKSB78, PZP79, Rue79, SP75a, SBB79, SB78, SS71b, TN73, TP78, TOED79, WJC70, WM79, MVN75, Wil77b]. potential-energy [BA78, Moh78, PD78b, SB78]. Potentialhyperfläche [HG75]. potentials [Ban74, CM75a, CK73, Con72, DMS^{+}78, DH78, Ell77, FRM74, GPP79, Hal79b, HLGC73, Hun75b, HYL78, KL77b, Nal78, Ner75, Nis77, Ove73, Pai73, PT78, PD77, PK77c, PÖ77, RM74, Sab73, SMH77, SG77a, SJH79, SBM70, TGG79, TS78, TS79, Wei74, WO75]. pour [Mar73a]. Pouvoir [DQ77]. Power [Lau79, BG72]. powers [Boy72c, RB75a]. pp [Ahl77a, Com77a, Com77b, Sti79, Cal79a, Dun78, Lar78c, Mår78]. PPP [BK71, ČPH74, ČPP75, FTN77, KBK72, Mic76b, PČH74]. PPP-type [FTN77]. practical [SMMA73]. practice [Ano78c, Csi76]. prealbumin [AJK78]. precision [Del74]. Predicting [MP75]. prediction [BDT77, IMN79, KK74a]. predictions [Tru76]. predictors [DH78]. predissociating [HE77]. predissociation [BFR78, BEF78]. Preface [CG77, Löw79]. Preliminary [BCR^{+}78, BS75c, BSB77, LDP79, BCL77, Meh79]. Prentice [Cal76a]. Prentice-Hall [Cal76a]. presence [Abd71a]. Present [DS70a, Ost79]. Press [Com77a, Com77b, Lar76, Sab75a, Shu77b]. pressure [Bor71, Lib76]. Price [Ahl77a, Cal76a, Cal79b, Cal76b, Cal77b, Com77a, Com77b, Dun78, Gos78a, Lar78c, Lin73, Sab75a, Shu77b, Shu77a, Sti79, Ano78c, Gos78b, Mår78]. primary [MP75]. Primitive [Hir70, Kut78]. principal [GCJ73]. principle [CH75, Csa76, Csa78a, DR75, Eig78, HHR70, Lim76a, Lip72, Moc73, PBC74a, RCE75, Sre75, You72]. Principles [Rab77, CH74, GY75, Lud75, Sun74a, Gos78a]. prior [Cof74b]. Pro [CBV79, RSK^{+}78, BWRD79]. Pro-Leu-Gly-NH [BWRD79]. probabilities [BR75a]. probability [Ada75b, AC70, Ave75, Hun74, Hun75a]. Probing [SY77]. Problem [Gal72a, LL70, Aba72, AC76, Abs78, Ada72a, AFD^{+}78, AMJS79, BP78, Col78, Com74, DMR74, ES73a, FU76, Gop76, KPJM70, KBA76, KK76b, LR77, Lud71, Mal71a, MHPS70, Mat74a, Mat76b, Mcd79, MVM75, MLB76, MFRE76, Nor75, Pal75, Per78, RS76, RE70, Sha77b, Sha78b, SMMA73, Shu77c, Sla70b, Ste77, Wes79, ZF76]. Problems [Loc77d, AM74, ČP71, Com78, DL75, DSY74, FBL79, Gás71, Kau79b, SPČ79, Ste74]. procedure [ADDLG76, BB71, BK71, But76, CM72a, Fro75, GCP79, HS70b, Jak76a, Kau74b, Koc74, Lan77, PD78a, PK75]. Procedures [PK76b, BB74, KPPH78, KPP78b, KS75, LB77, MR73, PK77a, PK77b, TRPB77, Wei75]. Proceedings [Cal76b, Sti79, BP70, BP71, BP72, BP73, Csi77, Mar73b, Pul78b]. processes [CSDV78, Chr79c, Fla79, GP78, Kle74b, Plo79, Rat78, Rei70]. processing [Bil72]. prodines [LJBN75]. produce [KKK70]. Product [Meh73b, Ada73, Bri73, Lya78]. Production [BS77a]. products [Lét74, MH72]. Professor [Sla71b]. profile [KK77a]. profiles [Kai77, MBM73a]. program [BF71, CAMC71, HC72, Jeh73, KSP78, Lin77, SP75c]. Progress [Lar78c, Dau77]. project [Nes75]. Projected [May74b, SDS73, Hur77, LL75, MLB73, May74a, MK75b, MK75a, May78, PČL78, SLB77, SDB74, SDB76, SB78, Van70, SDS73]. projected-unrestricted [SLB77]. Projection [BL72, Fin73, CHD71, Cro77, LKH77, CE73, KGB77]. Projections [GB71, Löw71a, Ahl77b, BÖ78, GBL^{+}73, Löw70, Sar75]. promazine [PPK78, PK76a]. proof [Rei72b, You72]. propagation [LW77]. Propagator [OJB77, KPÖ76, PÖ77]. propagators [LÖ77]. Properties [BH71, GB71, HO73, Löw71a, Ahl77b, AMJS79, BBB73, BN74, BN76, BDP73, BL79, Boy74, Cal75a, CJS77, Chr79a, CS78, DR77, DBBJ76, DMS^{+}78, Del73, DDD70, DD71, GZ72a, GZ72b, GC77, GBL^{+}73, Hin77, JSC78, KMT77, KWC79, Lad74b, Lip72, Löw70, LG71, Mat71a, ME78, Moc74, Mon77, NRS79, Ove73, Pau77, Pet78, PBC74b, PBG76, PD79, Ra76a, Ran79, RL79, SD77, Sch76a, SP79, SZ76, Wei71, WV74, WB73, ZF77]. property [Boy70a, De 76, Ran78]. Proposed [Kau73]. propranolol [KVSG79]. prospects [Ost79, Sla73a]. prostaglandins [KD79b]. Protein [SG77b, AJK78, BCL79, ED79, Lad74a, Lak75, Lew78, Mar79, Meh79, Pet78, PF79]. protein-containing [Mar79]. protein-methylglyoxal [Pet78]. proteins [Abd76, AMJS79, BCL79, Boy74, DAC^{+}79, Eis76, Eis77, GC75, LK78, MME79, PPC72, Pet79, Pop79, PP74]. protobiological [Fox75]. Proton [Coo79, Coo78a, DBBJ76, DMS^{+}78, GPP73, Gol70, TP77a, NN72, NN74]. proton-acceptor [DMS^{+}78]. proton-proton [GPP73]. protonated [KJ75]. protonation [FJP70, Kap73, Lei74]. Protonisierung [FJP70]. protons [Gia76, HG78, Sal72, TL71]. prototype [FTN77]. proximate [LPP78]. pseudo [BP71, CH74, Kar70b, Kar70c, KFY71, Mil74, SDM74]. pseudo-aromaticity [BP71]. pseudo-eigenvalue [Kar70b, Kar70c, KFY71]. pseudonatural [Pal73]. Pseudopotential [RTM77, Sch73a, CH75, LB79b, PF74]. pseudopotentials [GG79a, GG79c, GG79b, PGDM77]. pseudorotation [BDT77]. pseudospectra [Fae79]. psoralens [SCYB75]. psychoactivities [DH78]. psychotropic [DH78, KK74a, KK75, Pur75]. PtCl [BBL77, Mes73a]. pteridines [Kap73]. Publications [Ano78u]. Published [Cal73a, Cal76a, Cal73b, Cal75c, Cal76b, Cal77b, Lar76, Lin73, Sab75a]. Publishing [Ahl77a, Lar78c, Sti79, Ano78c]. puff [AM74]. Pullman [Cal73b, Dun78, Shu77a, Sti79, Mem75, Mem79]. pure [Lim74a, WRF73]. Purine [GRE^{+}77, GJP79, GPP73, KP79b, ME78, SMB^{+}79]. Purine-purine [GRE^{+}77]. Purines [BP72]. pyrazine [HO73, Hin77, PCMS77]. pyrazole [KPKC73, PK73]. pyrene [AZ72, HKP79, KPPH78, LPP78, LP79c]. pyrene-7 [LPP78]. Pyridine [PK76b, AM72b, HO73, JL70]. pyridoxal [AMSS79, WTB^{+}76]. pyridoxal-p-dependent [AMSS79]. pyrimidine [CLP79b, GJP79, GPP73, SMB^{+}79, SM79]. pyrimidinic [KL79]. Pyrrole [PK76b, AM72b, KPKC73, PK73]. pyruvate [BC76].

quadruple [BP78]. quadrupolar [MMS70, SR72a]. Quadrupole [ETF79, BM73, Boy72b, Kot76, LK78, RR72, Tru72]. Qualitative [ČPH74]. quality [DPGC79, GG79a, GG79c, MYTC77, MDC79]. Quanta [Atk74, Atk91, Gos78b]. quantal [Ber76]. Quantenchemische [HG75]. Quantenmechanik [PTUK69, Cal73a]. quantique [Loc72, DQ77]. Quantitative [Pop79, Pur75, BBL77, DBJ74]. quantities [BW73, GMG75]. quantization [KL77c, SL79]. quantizing [HDG71, Hor71]. Quantum [And79, Bul71, CBV79, CGLÖ76, Cal73b, DAP79, FL79, GP74, GT75, Her74, Hos78, KK72, KKK74, KK76a, Kau79a, LH75, LW77, LJBN75, MPN79, PP74, PS75b, Pul76, PC76, Pul78b, Ric77, Sti79, SLOKP79, TPC77, Wal77, WMV77, Ave72, Cas78b, Che71, ČP71, Col77a, Con73, Fox75, Gol73, GM73a, GM73b, GCJ73, GJ76, Hal75a, Hal79a, HS79, HSBD78, Hat75, DV79, INY^{+}71, Jan76, KK75, KG76, KG77, KB77, KKK79, LPP78, LP79a, Lin73, MP71a, Mat72a, MF73, MSP73, MCP73, MPW75, MW75a, Mat76a, Mat76b, MWY77, MW77, MK78, MKT77, Mos73, NB72, Par77, PHL^{+}76, PGQ77, PTUK69, PT75, PGPP79, PS75a, PP71, Pul77, Pul79b, RW75, Red79, Rei70, Ros74, SP75c, Sch71b, Sch71c, Sch71d, Sch71e, Sch76b, SE79, SJ79, Sla79, Sre75]. quantum [TP77a, TL72, TWC73, Tho75a, Van71, Wei71, Wei75, WO74, WO76, Wil79b, Wil79a, Wyb73a, BP70, HG75, Loc72, MN79, Sch71c, Lar76, Mår78, Shu77b, Shu77a]. Quantum-Chemical [FL79, SLOKP79, MPN79, LPP78, SJ79]. Quantum-mechanical [DAP79, Hos78, PP74, PS75b, PC76, TPC77, HSBD78, KKK79, MKT77, PGPP79, Pul77, Wei71]. quartets [GHS79]. Quasi [BBL72, BBL77, MMB70]. quasi-localized [MMB70]. Quasi-relativistic [BBL72, BBL77]. Quasiclassical [Tru76, Ber76]. quasidegeneracy [Wes79]. quasidegenerate [Dmi75a]. quasimolecular [MFRE76]. quenching [AZ72]. quinone [SP71]. quinones [NT79].

R [Sab75a, MM74, WRS70, FWFM77, You75]. Rabalais [Gos78a]. Radial [AAC76, Gru70, JBG70, BS70, Ben72, CA72, JG72, JSK75a, JSK75b, RW74]. radiation [Cal75c, Hea71, Kob78, Sch71b]. radiationless [Plo79]. radiative [Bar79, BOP76, LH75, SZ76]. radical [Cla70, DT75, HS70b, KR73b, KKM70, Sha77a, Sha78a, TS78]. radicals [CDEO76, GPC^{+}71, KT76, KU70, MMB70, MS71b, MLB72, NN72, NN74, SG77b]. radii [Ra76b]. Raman [DH73, MS71c, Sve79, WS75, WLS77]. Random [LP74c, ÖL78, And79, Lar75, Nel79, NMS77, ÖL79a, SO77, Tru73b]. random-phase [Nel79, NMS77]. randomness [NJF^{+}77]. range [Abd71a, AFD^{+}78, Cha71, Dea72, DY73, FY71, GPP73, JBP76, KP71a, Kle79, Koł74, Koł75, Koł76, MCP73, PD78a, RMP71, SKL79, SBM72, SZ76, WMV77]. rank [Boy70b, BM71a]. Rapid [Euw71, SBB79]. rapidly [PW74a]. rare [Fla79, LRJ73, LPT^{+}74, MS76b, RW74, TW77]. rare-earth [RW74]. rare-gas [Fla79, LRJ73, MS76b]. rare-gas-halide [Fla79]. rat [JG74]. rate [Bar79, BC76, ČZ79]. rates [Mat78]. rational [Gab76]. ratios [Ber75]. ray [BS71, CFMO73, FM78b, GK75, GML77, HYL78, Mes73a, TS71, PNJB78]. Rayleigh [Amo72, Car72, CJS77, Cla71, Sre75, WB76]. rays [Dec70, MFRE76]. Re [MMR74]. reaction [BC76, BCR^{+}78, BDT77, GTBB77, GBT79, JŻ79a, LPP78, Mag79, Nal78, PK71, PDEB78, PKSB78, Sal79, TB75, TH78]. reactions [ČZ79, CSR79, Chr77, Chr79c, FM78a, GBTB78, RW75, TC70, WEH74]. reactive [Ber75, Ber76, Hil79, Mic74, Mic76a, Nes75, RW77, Red79, Tru76]. Reactivity [WCK^{+}76, CW78, ES73a, Joh77, LP79a, LP79c, Sal72, SS78b]. real [Bre75, Fau76, Goł79, Lad75b, Wel77b]. realistic [FRM74]. reality [Živ77]. realization [SP75c]. Rearrangement [RB75b, BMM75]. rearrangements [BVR78, BDT77, CSR79, TG76]. Rebane [WW71]. received [Ano70d, Ano70e]. Receptor [Ano76e, CT79, Dav77b, Höl77, JG74, Wei75, WCK^{+}76, WO77, WOEG78]. receptors [CS76, SCYB77]. recherche [Mar73a]. recombination [Bar79, Gub79]. reconstructed [DS70a]. recurrence [JW78]. Rederivation [ČP71]. redox [LKK^{+}79, DAC^{+}79]. Reduced [AC76, Abs78, Bun77, Col78, JTR77, Kry78, LL70, SLM70, Spe72a, SW78, Ano78d, BB71, KP71b, Kry77, Lar79, MW74a, NS75, Spe71a, Spe71b, SW77]. Reduced-density-matrix [Kry78]. reduction [Per78]. reference [BSS78, BKS79, Lar73, Lar78b]. Refined [RD79]. Reflections [Ano76n]. region [BB72a, Mem77, PZP79]. regions [ZL78]. Regularities [Plo79]. Regularity [CS78]. regulation [BC75]. regulations [SG77b]. related [BW73, NN74, OMC76, Wol71]. relation [Abd76, CBV79, Cas79, DEC78, KSS73, KC74, LD71, Liu71b, ST79a, WV74, Rob71]. Relations [BHM75, JW78, Kau79a, Lan74, ST79a]. Relationship [LB77, ZG79, BCL77, DGC77, Kum77, PČL78, PW79, WB76, WP79]. relationships [PD78b, Pur75, SJ79]. Relative [AS78, BKW75, CS74, JM76b, WJ79, ZR78, ZBH^{+}78]. Relativistic [CW76, DS70b, Des72, ES70, MBM73a, RW74, Ros78a, TP77b, WC75a, BBL72, BBL77, BK75c, Cha71, CWE71, Ell77, Koe74, Mat73a, Mat76c, PPI73]. relaxation [MG77, SM77, Sch71d, Sch71e, Wen79]. relay [TP77a]. relevance [Urr77]. relevant [TPC77]. Remark [INY^{+}71, Bie75, Brä70, BF75]. Remarks [Cus75, Rei70, Sin73c, Sis74, SC76b, Wei71, AT75, BVS76, Bry74, Han77b, KSS73, Kau74a, Mar76, Nau77, Nau78, San72, Sch78, SG77c, Exe74, Wij78]. Reminiscences [Con73, Pul79c, Tho75b, Van71, SG76b]. Removal [CHS77, HT75, SHO71]. Renner [PK72]. repeat [RU76, RKOU77]. repetitive [CBV79]. replacement [Mil74, SDM74]. replication [Coo74]. report [BS75c, BSB77]. Representability [Col71, Erd78b, LL70, Col70, Kum77, LD71]. Representable [Lar73, Lar78b]. Representation [Spe72a, BS76b, BT71, Bow72, LH74, NM77, Nis77, PGQ77, PS75a, SP75c, Spe71a, Spe71b, SL70, Sti78]. representations [BF70, Blo72, Blo73, Boy72c, Cor72, HS79, Lét74, Rei72b, RM72, SS73c]. Representing [Gal72a, Lar79]. Reproduction [GCP79]. repulsion [AT75, HLGC73, JBP76, MM74, Sil71, WN71]. repulsive [Csa77b, GBTB78]. repulsive-dominant [GBTB78]. requirements [NN77]. research [Nes75]. residues [PPC72]. Resolvent [Røe75, BÖ78]. resonance [BMD70, Bow71, Boy72b, DL75, EL73, FY79, FB76, GBV77, HE77, KM74b, NTN^{+}79, Pul78b, RSS75, RKOU77, San72, SSM75, Sti79]. Resonances [Sim78a, BFH78, Doo78, FHB77, Jun78, KSP78, LKH77, MCW78, NB78]. resonant [CG78]. respect [KMLB77, SP79]. response [Han77a]. responsible [BL79]. Restricted [Mes78, CDEO76, JD74, Spe71b]. restriction [HT75, Kij72]. restrictions [Kau79b, SHO71]. Results [Col71, Fau76, SK72, BCL77, CS73, Col70, Doo78, GCP79, HKM71, JL74, Min79, PT75, RTM77, Red79, Sch76b, SBD^{+}70, Wei76]. retarded [MP71b]. retention [BKW75, PK75]. retinal [WMB75]. retinol [GC75]. reunification [Bul71]. reversal [VF76]. Review [Ahl77a, Ano78c, Cal73a, Cal76a, Cal79b, Cal79a, Cal73b, Cal75c, Cal76b, Cal77b, Com77a, Com77b, Dun78, Gos78b, Gos78a, Lar76, Lar78c, Lin73, Mår78, Sab75a, Shu77b, Shu77a, Sti79, DS70b, CS73]. Reviews [Ano74c]. revisited [Kle77]. RHF [SP77]. rhodium [AMC73]. ribonucleic [Sun77]. Riccati [PW74b]. Richards [Mår78, Sab75a]. Riedel [Sti79]. right [SY77]. right-handed [SY77]. rigid [GP72, Gra77, Meh78, MO77, SW79]. rigid-rotor [MO77]. rigorous [FRM74, KPP77, Sim78a, KMO77, SP75c]. Ring [Spe72a, NT79]. RNA [MP75]. Roby [CAA^{+}75]. Role [HČZ79, KD79b, KKK79, GK78, Har79b, KBKČD78, Lar79, Lim77a, Mem77, Per78, PDE79, Ros74, LW77]. roles [ULM76]. root [Sco73]. rotamers [BL75]. rotating [DLG79]. rotating-wave [DLG79]. rotation [CK75, Cor72, IM73, KD79a, MEW70, WC75b]. rotation-vibration [KD79a]. Rotational [FGV77, Ber76, HP77a, SH75b, Web74]. rotationally [HE77, MM72a]. rotations [Gia73]. Rotatoire [DQ77]. rotator [Gra77]. rotor [MO77]. row [Cal74, CP70, DGC77, KP72, NG73, Raf75, ST72, Tho76, Wel77a, Wel77b]. Rowe [SD74]. RPA [Bre75, Jor73, Lar78a, NS78, SN79]. rule [NB72]. rules [ČPH74, CD71, HBB77, Hur77, Mat76a, Res78, TB75]. Runge [Loh76]. Rydberg [FWFM77, Hos72, Mil71, Rob72a, TTS^{+}79].

S [Ahl77a, AS78, Cal76b, OSL77, RS79, BS70, BM73, BK74, DD71, LG75, LAM71, SHR76, Spe70]. S-state [BM73]. S. [Par73]. S3 [Duk79]. Sagamore [Smi79]. salts [AMJS79, DT75, Sal72]. same [HMV78]. sampling [EH73]. Sanibel [Sla71b, Sla73b]. Sasaki [Ste77]. Satellite [BL77c]. satisfying [Kij72, Sch78]. saturated [Rou73, Rou74]. scaling [Bar78, BFH78, CG78, CG79a, Cas78a, Cas78b, Com78, Doo78, Gás71, Gop76, HS71, HTG77a, NB78, NG73, Nut78, Sim78a]. scattered [BB75, Joh70, JS71, JW72, Joh73, JM76a, Wei74, YJ76]. scattered-wave [Joh70, JS71, JW72, Joh73, JM76a, Wei74, YJ76]. Scattering [Ada72b, Bal78, Nut78, AD77, Ano79o, BG78, Ber76, BW73, BL77c, BY76, CM72a, CW76, Com78, DH73, Dan72, DC73, Dav77a, Dec70, FFCB79, GGCK79, Gia76, GLS79, Hen79, HYL78, KL78, Kel75, KL77b, LVDC74, Lev78a, Li76, LKH77, Mic76a, Mil71, MMR74, Nes77, PBPL78, RW77, Red79, RK77, SK72, SMMA73, TS71, TOED79, WK73, Wil74]. scattering-theoretic [PBPL78]. SCF [KT77, KTO^{+}78, KL79, LS74b, ZBH^{+}78, Abs72, ACH72, BB72b, BDP73, BBL72, BBL77, BPTLVM73, CGRC74, CP70, Car74, Cas79, DS76b, DSY74, DK77, Eld73, FP77, GML77, Hil79, HS70b, JD74, JW76, JS77, JS71, JW72, Joh73, JM76a, Kau70b, KPKC73, Kau73, KSP78, KS72, KS75, KK77d, LSOC77, LAV72, Lev70, Lim74a, Lud77, MW74a, MS76a, MR72, Moh77, MG77, MM70, NN77, PCMS77, PPK78, PS70, PK76a, PK77b, PK73, PBG76, PPB78, Sam75, Vla70, WRF73, WMB75, YJ76, ZR78]. SCF-MO [MR72]. SCF-MO-CI [PCMS77, WMB75]. SCF-MO-LCAO [BBL72]. SCF/CI [KSP78]. scheme [Aro78, BEH77a, Goł79, HR76, JG72, KT77, KM76, KK76c, MG77, Nov73, RK79]. schemes [CSFM72, Kau74b, Kut78, MSL73, RC72]. Schiff [WTB^{+}76]. Schrödinger [Ada73, Amo72, Bal78, BFH78, Car72, CJS77, Cla71, KA76, VB72, WA71, Wel70, Wel71, WB76]. Schrödinger-Operatoren [KA76]. science [CGLÖ76, Pul79b, Shu77b]. Sciences [Cal73b, Bul71]. Scientific [Ahl77a, Lar78c, Ano76n, Ano78c, Jan77, Par73]. scientist [SG76b]. SCOPW [Sco71]. Scott [Sab75a]. screened [AM74, TN73, WM79]. Screening [TPM70, Csa77a, Nee75, Sel75, SIM75, SLC78, YW73]. screening-charge [Csa77a]. SCRPA [Lar78a]. Se [Sch73a]. search [Mar73a]. Second [KL77c, OJB77, SLM70, Abd71b, BBB73, CP70, HS76, HBE78, KP72, Lau79, MHRS76, MHRS77, Moc74, OCRMLR^{+}71, PLMB73, RM73, SK71, SL79, Vog79, Wel77a, You75, KD74]. second-neighbor [MHRS77]. Second-Order [SLM70, Abd71b, BBB73, HS76, HBE78, Moc74, PLMB73, Vog79, You75, KD74]. second-row [KP72]. secondary [BCL79, MP75, Meh79, RU76]. section [BG78]. sections [Bar79, Chi77, Fae79, ML76, SSM75]. secular [Amo71, Hen77, Koe74, Nor75]. Selected [Hat75, Moh77, SM73a, SD78, DGC77, SM76b, SM76a]. selection [DLG79, HBB77]. selective [Tal77]. selenide [YR76]. selenium [Ald75, VK74]. Self [BBP71, Bor70, CWE71, Kar70b, Kar70c, KFY71, Kle70b, Meh78, MM72b, MSM70, MBM71, Sla70a, BKS79, BK75c, DS74, Eig78, Gou73, Hac74, Hal78, JL74, KKW73, KS75, LG77, MMM71, Sco73, Sco74, Sel75, SLC78, Tal77, TGSL70, Vla70, VK74, ZF76, GL74, LG75]. self-adjoint [Sco73]. self-consistence [KS75]. Self-Consistent [Sla70a, Bor70, CWE71, Kar70b, Kar70c, KFY71, Kle70b, Meh78, MSM70, MBM71, BK75c, DS74, Hac74, JL74, KKW73, MMM71, Sco74, Sel75, SLC78, TGSL70, Vla70, VK74, ZF76]. self-consistent-field [BKS79, Hal78, LG77]. Self-correlated [GL74, LG75]. self-diffusion [Tal77]. self-organization [Eig78]. Semi [BBL77, WJC70, CAMC71, Dub72, LÖS73, NMS77, PK75, RT77, San71, San72, Sch76b, SMH77, SG77a]. semi-ab [PK75]. semi-classical [Dub72, Sch76b]. Semi-empirical [WJC70, BBL77, CAMC71, LÖS73, NMS77, San71, San72, SMH77, SG77a]. semi-infinite [RT77]. Semi-quantitative [BBL77]. Semiclassical [Ada75b, FL79]. semiconductors [CD78, CJCM73, Csa76, Csa77a, Csa78b, DT75, FCL79, RD79]. Semiempirical [HLGC73, JPTM78, KHI79, PL79, PT78, ABM78, AT75, BOP76, Cof74a, NS78, PLYM78, RL79, SN79]. Semiorthogonalized [Kas76, KT77]. semiquantitative [Hal75b]. seniority [MLM75]. sensitizing [SCYB75]. separability [Ada75a]. separable [Ban74]. separated [LE72, RC70, RC71, RC72, Sal73, SPČ72]. separated-electron [LE72]. Separated-Pair [SPČ72]. Separation [Lan73, GZ72b, GP79, KKKP79, Ra76a, Zer74]. separations [RMP71]. Sequence [BG70, Ell73, AC70, AS78, BSM72, BK74, MP75, MMS70, MM72b, PL77, SK71, SK72, SA77, TTS71, Tat73, Van72]. Sequence-adapted [Ell73]. sequences [BG72, RSK^{+}78]. sequential [Muk75, Urr78]. Serber [Pau77]. Serber-type [Pau77]. série [Mar73a]. series [BG72, ČZ79, DDH78, Gos78b, Jak76b, Mar73a, PD78b, ZG79, Cal75b]. Serotonin [WO77, WOEG78, WCK^{+}76]. session [KKK70, Pos70]. set [ÅSW77, Aba72, Bur77, CHS77, CG79b, Cur78, DRCT78, DLG79, Euw71, FGV77, HS70a, Mal71a, Mcd79, RPK78, Sol73, SSH78, RS78b]. sets [Abs78, Dah78, DPGC79, ES79, Fin73, KB77, LCK76, MYTC77, Pel77, SWN78, SBM72, SC78b, TB74, Van74]. Several [Bar70, Kau79b, MCST72]. Sexena [Ahl77a]. SH [KPD78]. shape [BMD70, CA73, HMV77, HMV78, RR77, SC73]. shaped [WO75]. shapes [Fox75, ST72]. Sharma [JW78]. Shell [CGRC74, APČ77, BW76a, BS78c, ČP72, CDEO76, CV73, DK77, Fae79, FP77, GBT79, GHS79, GBM74, GRM75, Hen76, Hil79, HP77b, JD74, JBP76, KC77, KBK72, Lai73, Lar75, Lar78a, Lin78, MMB70, MW75b, MLB72, MM72b, MDS75, MMM71, Muk75, MMM75, NN74, Nel79, Nes77, PAČ77, PT78, PAM75, RGM75, RGM77, SM73a, SM76b, SM76a, SD78, SH73, SPČ79, SS73a, SS73b, Sin75, Spe71b, SO77]. Shells [Bee76, JBG70]. shielding [BM73, GBM74, GRM75, LZ77, MMM71, PGPP79]. shift [DV79, MP71a, SKW72]. shifted [CT77]. Shifting [SH73]. shifts [BK77, Ema74, GML77, Kot76, Sab73]. short [HK74, JBP76, JM76b, LS72, SKL79, TBB70]. Short-distance [TBB70]. short-range [JBP76, SKL79]. showdomycin [SMP77]. SHS [Sab71]. side [RBDV77]. side-on [RBDV77]. sigma [CP70, Koł76]. sigma-bonded [CP70]. sigmatropic [BVR78]. SiH [BT78a]. silane [DPGC79]. Silbey [CJS77]. silicon [DPGC79, MYTC77]. similarities [KM72]. Similarity [KP76, RW79, Sch78]. Simple [PS73, Pel77, PP71, Spe72a, TS72, APČ77, ČP77, FTN77, GLS79, GMG75, Han77a, KM74a, Lad74a, LDP79, NMS77, PČL78, PGQ77, Pet77, PKSB78, RTS74, SPČ79, Sch72, Sol71, SS77b, TW77, Hal78]. simplest [HPK79, VB72]. Simplification [Lét74]. Simplifications [BL77a, BL77b]. Simplified [CP70, GV77, BFR78, Rei72b]. simulate [CEK72]. Simulated [DEE^{+}75, CDEO76]. Simulation [Bai74, Pol79, NGR^{+}77, RC78]. simultaneous [Goł79, Vla70, Wel76, Wel77b]. sine [WO75]. sine-shaped [WO75]. Singer [Pos78, Pos79]. Single [Rat71, Ada72a, Ada79, BS77b, BS78a, BM74, CC79, CM76, Gre71, Mil74, Muk75, PAM75, SD79a, SDM74, SLC78, SC79]. single-center [PAM75]. single-centered [Mil74, SDM74]. single-electron [Ada72a, BM74]. single-parameter [SD79a]. single-particle [Ada79]. Singlet [BB75, TSE76, BP75a, MW74a, MCST72, NT79, PCMS77, SMB^{+}79, SDB74, TH78, SDS73]. singlets [GHS79]. singly [Csa79a, LAM71]. Singularities [BG70, CHS77, CB72]. SiO [BEF78]. Site [Flu72, AMSS79, CW78, GbK77a, Höl77, Kap73, Kle77, KGB77, PD79, SS79]. sites [LK78]. six [AMC73, Jol79]. six-coordinate [AMC73]. six-parameter [Jol79]. sixtieth [Cal75b]. size [BS77b, BS78a]. size-consistency [BS77b, BS78a]. sizes [DGC77, Kau73, MDC79]. Skillman [HC72]. skin [SCYB75]. skin-sensitizing [SCYB75]. Slater [BR78, BS74a, BS76a, BSZ78, BPTLVM73, Bre73, Cha77, ERW75, GMZ73, LS70, LS71, Lec76b, MH72, SR72b, Sam75, ST70, SS71a, SR72c]. Slater-type [BS74a, BS76a, SR72b, ST70, SR72c]. slow [Gia76]. Small [Com77a, Com77b, Vla70, BB72a, CODM77, GJLL75, HS70a, Hur76a, Hur76b, KPP77, LGR73, Mul74, Pos70, Sim79]. Smooth [Thu75, Sha71]. SN [KKA78]. SOCL [GH72b]. sodium [Ber73, Bor70, BK75b, CD74, Col79, HCIE75]. software [WH75]. solid [FBL79, DV79, LW77, Muf76, Nam75, Sla71a, SR72c, SP75d]. solid-state [DV79, Sla71a]. solides [KMO77]. solids [ADFL71, Bor71, Bre75, Car75, Dec70, Fau76, Joh73, JMP78, PBPL78, Sco74, Sla75, Wen79, ZF77]. Solitons [Dav79]. Solution [MHPS70, Ada73, BEH77b, Csa79a, DSY74, FW75, HS75, JM73, Koe74, KK76b, Kun78a, Ner76, Nor75, Pul77, RRNS75, SD77, WA71, ZF76]. solutions [Cal79c, ENRM78, FG79, Fro79, KKA78, Kho79, Lai73, Mes79, Tal77, TTN76, Živ77]. Solvation [SAM79, WN78]. Solvent [AZ72, FSV79, HPW79, LCPP79, RC78]. Somatostatin [MDA78]. Some [AT75, BS75a, BVS76, ESR74, GSE78, GK76, HC78, Jai70, KSS73, LKK^{+}79, Löw70, Löw71a, Löw77, Mes73b, Nal78, Oda73, PD78b, San72, Sch78, Spe72a, SG76b, Vol79, Wei75, AMSS79, BG76, BYB79, BM74, BS75c, CK75, Cha77, CL73, DDH78, Doo78, DT75, JL74, KMT77, KR73b, Kum77, Lai73, LM72, MS76a, MR72, Mor75, Muk74, NN72, Nis77, PGQ77, PS70, PD77, PDE79, PK77c, RMB78, Ros78b, RD79, Sha77a, Sha78a, SKW72, Sim77b, Sto71, SL73, SBD^{+}70, Tho75a, TP78, WN78, DEC78, ED79, RTM77, SW78]. Sons [Cal79b, Cal75c, Dun78, Shu77a]. Space [GGCK79, Spe72b, Bee72, BS70, BS76b, Cli73, GHS79, HY78, KC74, Lan77, NM77, SP75c, Sla70b]. space-dependent [Cli73]. space-spin [KC74]. spanning [Bee72]. spatial [AMSS79, Gal74, MPN79, Rou79, ZG79]. spatially [Csa76, Csa78b]. special [Mag79, Pau77, WV74]. special-pair [Mag79]. Specific [Gol70, AMSS79, BL79, Jeh73, KMT77, Lar79]. specificity [PL77]. spectra [AeAs78, BCL79, BL77c, BEF78, CW77, DH78, GGJ77, GPC^{+}71, GK75, Hal75b, Kib75a, KJ75, KS79, LV78, LGM77, LBÅF78, Lun77, MS71b, MR72, MK72, NT79, PNJB78, SLB77, Sch71e, SM79, SK79, Van72, WK73, Wój76, Woj78, LBÅF78]. Spectral [BHM75, Boy74, BF75, CS72, Ra76b, BFH78]. Spectroscopic [Kli79, TPM70, KKK70, WB73]. spectroscopy [Bas78, Feu77, GC75, KKK70, Kla78, KRKŠ78, MKF77, Mic76b, Pul78b, Rab77, Sti79, WS75, Gos78a]. spectrum [BHLM78, CS78, FHB77, GML77, HG78, HDG71, Oni79, VG78, WJ79, Wan79, WFS77]. Speed [DMS^{+}78]. sphere [GH72a, Kau73, SC76b, Wei76, YJ76]. spherical [AF73, ADDLG76, FHP70, HP77b, Kat72, RR77, RMB78, RK76, Sch77b, SR72c, TL71]. spherically [Csa77b]. spheroidal [SP74]. sphingomyelins [PC76]. Spin [BP74, BOP76, BK75a, Joh77, Kau79b, Kle74b, KT76, KR73b, Lim75a, MS71a, MF73, MSP73, MCP73, MPW75, MW75a, Mat76a, Mat76b, MWY77, MW77, MN79, MLB73, May74a, May74b, MK75b, MK75a, May78, MK78, Sar75, SD79a, SM71a, SM71b, SM71c, SK73, SS73a, SS73b, TH78, APČ77, BP75a, Bor70, BK75c, CL73, CDEO76, DL75, Dav74, Fro70, GTG76, GJLL75, GJ76, Hal75b, JMM71, JSK75b, KR73a, KTO^{+}78, KPJM70, KC74, KGB77, Kle79, LED73, LD73, Laz79, Lim74a, Lim74b, Lim76b, Lim77b, MM74, Mat72a, MCF^{+}70, MLM75, Mes73b, MLB72, MS75, NN72, PAČ77, Pau77, Sha77a, Sha78a, SL70, TB73, VF76, Har79a, MSP73, MPW75, MW75a]. spin-adapted [Dav74, VF76]. spin-density-functional [GJLL75, GJ76]. Spin-Dependent [MK78]. Spin-Extended [KT76, Mes73b]. Spin-forbidden [TH78]. Spin-Free [Mat76b, MW77, MN79, SD79a, MF73, MSP73, MCP73, MPW75, MW75a, Mat76a, MWY77, SS73a, SS73b, KPJM70, KC74, Lim74b, Lim75a, Lim76b, Lim77b, Mat72a, MK78]. Spin-only [MPW75]. Spin-Orbit [BOP76, BP74, BP75a, BK75c, Hal75b, JSK75b, MM74, MCF^{+}70, SM71a]. Spin-orbital [Joh77]. Spin-other-orbit [SM71a, SM71b]. spin-polarized [Bor70, Fro70]. Spin-Projected [May74b, MK75b, MK75a, May78]. Spin-spin [MS71a, SM71c, MM74]. spin-wave [CL73]. spinfree [Lim75b]. spinless [DR77]. spinors [CW76]. spins [HLMV77]. split [SM73a, SM76b, SM76a, SD78]. split-shell [SM73a, SM76b, SM76a, SD78]. splitting [Har71c, HP77a, Kot76, MS71a, SM71a, SM71b, SM71c, WN71, BP75a]. splittings [BFS73, BB75, LK78, NN72, NN74]. Springer [Cal79a, Cal77b]. Springer-Verlag [Cal79a, Cal77b]. Square [RSS75, And79, Sco73]. squares [BT71, HS75]. stabilité [JM76b]. Stability [Lai73, MK76b, Bur77, Coo78b, JM76b, LÖS73, LRJ73, LPT^{+}74, MHA70, Lai73]. stabilization [HK74, KSP78]. stabilized [MS71b]. stable [Hat75, RS79]. staggered [Hub71]. standard [RS78a]. Stanton [TB75]. Stark [BK77, CHH^{+}78, Rei76]. starting [ACD^{+}74]. State [LÖ77, NM77, ÖL78, Abd77, Bar77, BWS77, BN76, BJ73, BS70, BM73, BP75b, BCL77, Bow72, BK74, Bra79b, Bun77, Cal72, Cal74, CR72, CKD73, Dav77b, DD71, FWFM77, Fro75, GBV77, GL74, Hal79b, HE77, DV79, HMC71, Ive74, JŻ79b, KR73a, Kap79, Kar78, Koł74, Koł75, KÖ71, Kut78, LH74, Lar75, Lar78a, Lar78b, Lec76a, Lec76b, Lib76, LKH77, Lin77, LG75, LHK77, MME79, Man77, Mat79b, MM72b, MSM70, MBM71, Muk74, Mul71a, Nis77, ÖL79a, RS79, Ros78b, SK71, SK72, SA77, SH75a, Sin73b, SD70, Sla71a, SGM78, Tan74, Tat76, TSE76, Tru73a, Tru76, WRS70, Wil77b, WB73, PČH74]. statement [AFD^{+}78, Lon75]. States [Spe72a, Abd71b, AM77a, AS78, ADFL71, BSS78, BKS79, BP75a, Ber73, BL79, BF71, BS75c, Boy72a, BS78c, CSFM72, ČPH74, ČP77, CESW70, DRCT78, DS76b, DS70a, DS70b, DT72, De 78, DD77a, DC75, Dmi75b, DMM75, ES73b, Fau71, Fau76, GV77, GB75, GL74, Hal75b, HNT78, Hat75, HS70b, HT75, Hos72, IS75, KR71, KR76, KKW73, KTO^{+}78, KSP78, Koł76, Kun78a, KOW78, Kun78b, LD71, Lim74a, LV71c, LS79, LAM71, MW74a, MCST72, MCF^{+}70, MCM73, MLM75, Mil71, MMS70, MG77, Muf76, Mul71b, PČH74, PCMS77, PS70, Pie71, Pop71, PLMB73, RB77, Rit76, Ros78b, RT77, SMB^{+}79, SP75a, SC76a, Sim78c, SH75a, SHR76, Sin75, SDB74, SCYB75, SCYB77, Spe70, SS77a, SS77b, SK79, Sto71, Tan74, TTS71, Tat73, TTS^{+}79, TNB79, TS72, Urr78, WS75, WMB75]. states [WK73, You72, YD75, SPČ72, SDS73]. Static [INI71, Moc74, GBM74, GRM75, HS71, MCST72, RGM77, WO76]. Stationarity [CG78]. stationary [Com78]. Statistical [MM77a, SW70, SW71, Gol73, Han75, KR78, Lin71, Cal79b]. statistics [TB73]. status [DS70a]. steady [BCL77, Dav77b]. steady-state [Dav77b]. steering [KSL76]. stentorin [SCYB77]. Stephen [Cha71]. stereoregular [DD77a]. stereoselection [Fuk70, Fuk75, Cal77b]. steric [Tic79]. Stickstoff [BKW75]. Stieltjes [Lan74, Lan77]. Stieltjes-integral [Lan74]. STO [CHS77, JW78, Mar73a, Pel77, ZSR79]. STO-3G [ZSR79]. stochastic [ESG79]. storage [Bil72, MHH74]. Störungsrechnung [Ahl70]. STOS [Wel77b]. STP [Eng72]. strain [BMD70]. strength [AC70]. strengths [CM78, GBV77, Lar78a, Sin73a, WM79]. stretch [SS71b]. strip [JG74]. Strong [Lun77, KK77c, LL73, Loc77d, SRE^{+}79, Shu77c, Pet78]. strong-field [KK77c]. strongly [Kle74a]. Structural [Boy74, Chr79a, Enz73, Mof79, RW79, RU76]. Structure [BC75, LL72, Lei74, ÅFL78, AF73, AMSS79, AL71, BT78a, BK77, BCL77, Bur77, CBV79, CGLÖ76, Car75, CJ73b, CR72, Col77b, Duk79, EP71, Erd78a, Fin73, FBL79, FWJ^{+}79, FSV79, FRM74, Fro74, GSE78, GK75, GML77, Har78, HS70a, Hin77, HK74, Hop78, HT75, JR74, Joh75, eAJ76, KK74a, Kau79a, Kel75, KKA78, KJ75, KL77a, KC74, KU70, KVSG79, KMO77, Lad74b, Lad75a, Lad75b, Lad76, LSOC77, LSS78, LVDC74, LP79a, Lay77, LFF79, Lev78a, LE72, LM72, LÖS73, LK78, Lud75, Lun79, MW74a, MPN79, MS72a, MS74, MF73, Meh79, MHWC78, MHW78, MR72, MOV75, Ove73, Pai71, Pai73, PN71, Pea74a, Pet77, PNC^{+}78, Pul77, Pur75, Ran78, RL79, SMB^{+}79, SP74, SP75b, Sch73b, Sim79, SJ79, SJH79, SR77, SCYB75, Spe71b, SM79, SK79]. structure [Swi71, TB74, Tho76, Ukr72, Van75, Wil79b, Wit72, ZG79, ZF76, LL72, Lei74, Shu77b]. structure-activity [Ran78]. structure-property [Ran78]. Structures [Boy77, ABD^{+}79, BNLS72, BCL79, Con70, FY79, GK78, Har70a, JW72, JNŽ72, Kan74, Kry78, Lad74a, NG73, Pri75, SL73, USK^{+}79]. studied [ÅFL78, BF78, HLMV77, KVSG79, LBÅF78, MSSF79, TGSL70]. Studies [Cal75b, Hos72, LCPP79, LG71, TB78, Ald75, BK77, BCL77, BVR78, Bur77, Cur78, GZ72a, GZ72b, GL74, HR79, Har75, HG75, JE79, JS71, JG74, KKK74, KK75, KN79, Kli79, KMO77, LPP78, LP79a, LG75, LGR73, Meh79, MHA70, MS76b, MCW78, MHH74, PL79, PLSL78, PKSB79, PP74, Pul74, PS75b, PBG76, PC76, PD79, RW75, RRNS75, RSK^{+}78, Rou74, SMP77, SS79, SS74, SS78a, Sim77b, SH75a, SHR76, Sti78, TGG79, TW77, PCMS77]. Study [CR72, DPGC79, FFCB79, Jai70, KU70, Res78, AeAs78, ABM78, BB74, BM74, BBA73, BSB77, CBV79, CS73, CEK72, CLP79a, CLP79b, DGC77, DAP79, ES70, Fro75, GC77, HLYC77, HL78b, KM74a, KL78, KC79, KHI79, KD74, KL77a, KPD78, KS79, KJ78, KP79b, LED73, LDP79, LP79c, LP79b, Lec76a, LCK76, LE72, MMS70, MBM73b, NPL^{+}79, NI70a, NI70b, NTN^{+}79, NS79b, Nor75, OSL77, PLYM78, PKSB78, PGPP79, PA74, PB77, PGDM77, PPB78, RW79, RMB78, Rob72a, SD79a, SD70, SR77, SC78a, SC73, Spe73, SP71, TS71, TP77a, TTS^{+}79, TWC73, TB74, TRPB77, VG78, WRF73, Web74, WLS77, WK73, Boy77, GML77, LCPP79]. Sub [Sim78c]. Subduction [SD79b]. submolecular [Sun75a]. subspace [Sre75]. subspaces [Jan76]. Substituent [BS79, KC73, Mil74, SDM74, Tic79]. substituents [MK72]. Substituted [Boy77, DH78, Fog71, GBS70, Kap73, MR72, NTN^{+}79, PLSL78, Rou73, Sha77a, SS78a]. Substitution [HG75]. substitutions [KMLB77]. substrate [SS79]. Subsymposium [Ano79o]. subsystem [Mil74, SDM74]. subsystems [FTN77]. subunits [Fin73, Pol79]. Successive [GS70b, MK75b]. Sudden [Bow79, KBKČD78]. Suggested [Pea74b, Kau74b]. suggestion [Mem77]. suitable [BB73, Gia73]. sulfide [ES73b]. sulfides [Mat76d]. sulfur [Ald75, BP75a, KC77, TNB79]. sulphate [HS70a, JS71]. sulphur [Cha77]. Sum [RB77, Res78]. Sum-over-states [RB77]. summary [KM72]. sumrule [BB71]. super [Amo71, Nor75]. super-secular [Nor75]. super-secular-equation [Amo71]. superconductant [PP79]. superconducting [Mat76d]. Superconductivity [Mat71b, Mat72b, Mat73b, Mat74b, Mat75, Mat77, Mat79b, ACK^{+}75, Cli73, MM77b]. superdelocalizability [Mem75, Mem79]. supermolecule [GCP79, OSL77]. superoperator [BÖ78]. Supplement [RWFS71, Sab75a]. support [BDT77]. supported [KGB76]. Surface [SK79, BR75b, DS70a, DT72, DEC78, Kai77, KP71a, KW76, MO77, RT77, SKW72, SC73]. surfaces [BNLS72, BFH74, KKW73, Lin77, Moh78, PBPL78, PN71, PKSB78, Sim79, SB78]. survey [Sim77b]. susceptibilities [BBT71, Ema74]. susceptibility [Hal73, PLSL78]. suspected [KPP78a]. SW [MS76a]. symbols [BW76b, But76, BW76c]. Symmetric [Gal72a, Csa77b, DMM75, Goł79, Jør78, Kun78a, KOW78, Kun78b, Lim77b, Liu71a, Sch77a, Sch73b, Sla70b]. symmetrical [Moh78]. Symmetries [Gal74, Mat76d, KGB77, MWY77, Oda73]. Symmetrischer [Sch73b]. symmetrized [Boy72c, VK74, Wyb73a]. Symmetry [Bee76, BBT71, BH71, BB73, De 76, Eld73, GP78, GÖ68, JSC78, Kle70a, KC74, LP74a, LP74b, Ran79, Amo71, BF71, Boy70a, DM72, DK77, Flu72, GSE78, GP72, Han75, JPTM78, Kat72, KD74, Kau79b, Kle74a, Kry77, Kut78, Lar79, Laz78, LZ79, Lim76a, LG71, MK76b, MM77b, Ros78a, Rou78, Rou79, GÖJ^{+}73]. symmetry-adapted [Kut78]. symmetry-lobe [Rou78]. Symposia [Cal73b]. Symposium [Pul78b, Sla71b, Sla73b, Sti79, BP70, BP71, BP72, BP73, Lon73, Ede70]. synaptic [Wal77]. synthesis [Gab76]. System [Spe72a, AM72b, BZSD79, BS75c, Bra79b, BS78c, CS72, HL79, HBE78, HG75, HP77b, JBP76, KR76, KP71a, Lad75a, Man77, May74a, PČL78, PHL^{+}76, Res78, Sre75, SP79, TBB70, WH75]. systematibus [BP79]. système [TBB70]. systems [ACG74, AM72b, Bai74, BA78, Bar70, BB72c, BEH77a, BEH77b, BFH78, BL77d, CP70, Cho79, ČP71, ČP72, ČPP75, ČP77, CK74, CKI75, CII79, ESG79, GOK79, GP74, GBS70, GL74, GBM74, GRM75, HHR70, HS79, Hat75, Hea71, HG75, HR76, IS75, JMM71, JM76a, JPTM78, JS70, JL74, KL75, KPP77, KKP77, KMT77, KL77c, Lai73, Lar79, Lat77, Lin78, LRD^{+}77, LRD^{+}78, LG75, May74b, MK75a, MS76a, Mic76b, Mil74, MBMM79, MPMK79, MMM71, MMM75, NS75, NB78, Oni79, OSL77, PT78, Pea74a, Per78, PGQ77, RMB78, RGM75, RB75b, SN79, Sar75, SPČ79, SM77, SDM74, SS73a, ST72, Sol71, SO77, TS71, TP77a, TL71, VB72, WOEG78, YG77, SS73b, SDS73]. Szent [Lak76]. Szent-Györgyi [Lak76].

T [Sab75a]. table [Oda73]. Tableau [Gal72b]. taking [MCF^{+}70]. Tamm [DS70a, HU74, Lar78a]. targets [Gia76, GLS79]. taurine [CS74]. tautomer [ZR78, ZBH^{+}78]. Tautomerism [CLP79a, CLP79b, ZSR79]. tautomers [HLYC77]. Tc [MM77b]. Tchebycheff [Lan77]. TCNQ [DT75, ÅFL78, CHHS74, HRBB76, Joh75, Lad75a, LRD^{+}77, LRD^{+}78]. TCNQ-TTF [Lad75a]. Te [Sch73a]. Technique [BL72, BPTLVM73, AFD^{+}78, ABD^{+}79, BF78, CHD71, ČP72, CDEO76, Dan72, GS70b, HLMV77, HPK79, KPPH78, LS74b, MSKM77, Mil74, Muk75, Ner76, PW74a, Røe73, Røe75, SDM74, Sun75a]. techniques [ABM78, Che71, DD72, FB76, GGCK79, JS77, PBC74b, Pul77]. Teller [Ham71, INI71, MSSF79]. Temperature [Tal77, ACK^{+}75, MSSF79, PBC74a]. temperatures [AMJS79]. tempered [Raf75]. Tempering [Pos79]. template [Gab76]. ten [DSC^{+}77]. ten-electron [DSC^{+}77]. tensor [KK74b, KK77c]. tensors [Boy70a, Boy70b, BM71a, BM71b, BO73, BS74b, Ell73, IMN79]. teratogens [KPP78a]. term [Cha71, Koc74, WJC70]. terms [Ada73, AF73, CSFM72, CD71, FTN77, GC77, JPTM78, MCF^{+}70, MCM73, NY75, PS75a, WS79, ZL78]. ternary [Mat77]. Test [KT77, Kij72, PK75, BS78b, ČZ79, NJF^{+}77, PS73, GJ76]. tests [MMD78, Sin75]. tetracyanoquinodimethane [DT75]. Tetrahedral [Pos77, SC78b]. tetrahedrally [JMP78]. TFD [Sla75]. their [Ano79o, BT78a, BO73, DBBJ76, DEC78, GK78, GM73a, GM73b, GPC^{+}71, HBB77, HS79, KL75, Kau70a, KKK74, Kau76, Kla78, Lar78a, Mar73a, Pau77, PKSB78, PK77c, SMB^{+}79, Sin75, SR78, SH75b, TBB70, Tho76, ULM76, Urr77, USK^{+}79, WK73, WCK^{+}76, GK78, Sla79]. theorem [BEF78, BFH78, Chi77, CT77, CD71, FHB77, Gop76, Lan73, LH76, Mil74, MDS75, Nal78, NRS79, Rei72b, RCE79, WO74, WO75, WO76, CAA^{+}75, SC76a]. theorems [Dal79, Lev78b]. theoretic [Ber76, BDT77, KKP77, PBPL78]. Theoretical [BB72c, GZ72a, GZ72b, GBS70, HPW79, IMN79, Kau77, KN79, KJ75, KS79, KJ78, Lar78c, LP79c, PBS76, Rat78, RRNS75, SS79, SP71, VG78, WOEG78, Woj78, WB73, BB75, Boy72b, Cal76b, CSFM72, Con70, CD71, GS70a, JVM78, KR76, KKK74, KK74a, KK75, Kau76, LDP79, LP79b, LAM71, Mar73b, MCF^{+}70, MCM73, Nes75, Nov73, Ran78, RW79, Sha77b, Sim77b, SS78b, SR77, UP71, WMV77, CLP79a, CLP79b]. theoretically [JŻ79b]. Theorie [Mar73a, Loc72]. theories [ANR75, BB73, CA78, Feu77, FGV77, PKSB78, PKSB79, PS75a, WB76]. Theory [BN76, BMD70, BFH78, Bra75, Bra76, CGRC74, Chr79c, Cla71, Com77b, Coo74, Coo78b, Csi76, DSC^{+}77, Fuk70, Fuk75, JBG70, Oni79, RKE79, Sch77c, Wój76, ZV79, AC70, Amo72, AFD^{+}78, ACH72, ACD^{+}74, Ave72, BFS73, Bal71, Bal78, Bar70, BRS^{+}75, BS74a, BS75a, BS76a, BWS77, BP78, BB72b, BV73, BP72, BVS76, BMM75, BP75b, Blo72, Blo73, BDT77, Bow79, Brä70, BO79, Bra79a, Bur73, Car72, CFHS77, Car74, ČZ79, CW76, CT77, CL73, Chr77, ČP71, Csa79b, Csi77, Dav77a, DT72, DH76, DY76, Dmi75a, Dmi75b, Dub72, ES73a, Ell73, Eps77, ES79, ETF79, Fae79, FP77, FY71, Goo76, GS70b, Hal78, Hal79b, Han70a, Han77a, HS79, Har78, Har79b, HBE78, Hen76, Her74, Hir70, Hos78, Hur76b, IS75, JD74, Jam70, JS77, Joh70]. theory [JNŽ72, KL75, Kap79, Kib75a, Kle70a, Kle77, Kle79, KR78, KP78, Kry78, KS72, KS75, KK77d, LL76, LH75, Lar73, Lar76, Lar78c, LWA74, Laz78, LZ79, Lev78a, Lim76b, LÖ77, Lin78, Lin73, Loc72, LPT^{+}74, LG71, Lun75, MKF77, Mar73a, MF73, Mat76a, Mat78, MN79, Mat79a, MCST72, Mem75, Mem79, MW75b, Mic74, MMM71, NI70a, Naz72, PP79, PGQ77, PTUK69, PN76, PSK77, RR72, RSR72, Ram78, Red79, RCE75, Rob71, Rob72b, Ros74, RK76, Sam75, SF76, SLB77, Sch71b, Sch71c, Sch71d, Sch71e, Sch76b, Sel75, SMH77, SG77a, SDM74, SS73a, SS73b, Sin75, Sla71a, SJH79, SMW71, Suz77, SG76a, Tho75a, Tri77, TB75, TOED79, Tur74, Vla70, Vog79, Vol79, WN71, Wil74, Wil77b, Wol71, Ber75, IZZ^{+}79, JMP78, NN77, Cal76a]. Theory [Ano78c, Cal77b]. there [Sey76]. thermal [Mat78, NJF^{+}77, SWF79]. Thermodynamic [CII79]. Thermodynamics [Por78, Van75]. Theta [LV71b, LV72]. Theta-Function [LV71b, LV72]. thioridazine [KK76a]. thiouracils [GJP79]. Third [BWS77, Cal76b].

Third-order [BWS77]. Thomas [Bie75, Csa77b, Csa79a, Csa79b, DH76, GY75, Goo76, GGJ75, Han70a, Han77a, Lun75, WC75a, YG77]. Three [RW75, Bal78, CK73, HR79, Koł76, LG75, LPT^{+}74, PČL78, RL77, Red79, RC72, ST72, ST70, SS71a]. three-body [Bal78]. three-center [HR79, LPT^{+}74, ST70]. Three-dimensional [RW75, CK73, Red79]. three-electron [LG75, PČL78, RL77, SS71a]. threshold [Mes79, Sim78c]. through-water [PPB78]. thulium [AD77]. thymine [GJP79, ZR78]. Thyroxine [AJK78]. Ti [BL77c]. TiC [GK75]. TiCl [EP71]. TiCo [PN71]. TiFe [PN71]. Tight [CZNL70, CD74]. Time [Gol73, JL70, Moc73, NB78, Sch71b, Sch71c, Sch71d, Sch71e, VF76, BF75, Han70b, Man77, SB75, WA71, Jam70]. Time-Dependent [Sch71e, Gol73, JL70, Moc73, Sch71b, Sch71c, Sch71d, Han70b, WA71, Jam70]. time-independent [Man77]. tin [DC73, FFCB79, FRM74, GHJ77, Kot76, Pai73, Sla74, Wil74, GK75]. TiNi [PN71]. TiO [GK75]. tissue [Jeh73]. tissue-specific [Jeh73]. titanium [ES73b]. Tm [DS74]. today [Wyb73b]. Toivo [Cal75b]. tomorrow [Wyb73b]. top [DMM75]. Topological [KM72, BT78b]. Topologie [Sch73b]. topology [Sch73b]. Torsional [BBG79, SY77]. Total [BG78, SW71, BM74, Cas79, KP76, SW70, Spe73]. traditional [ČP71]. trajectories [CG78, CH77]. trans} [TS78, WMB75]. trans}- [TS78]. transfer [BZSD79, Cho79, CD76, DT75, Feu77, Lad76, LKK^{+}79, LW77, LH77, OSL77, Pet79, Rat78, SRE^{+}79, Van72, ZV79]. Transferability [CSM71, SDD^{+}77, SWN78]. Transferable [Ave79c, Ave79b, AB79, Ave79a]. Transform [GMZ73, Ban74, BS75b, GMG75, GB72, GCJ73, MKF77, MH72, ZF76]. Transformation [Thu75, WV74, AFD^{+}78, BL77a, Cof74b, FHB77, Jør77, Sch78, Sha71, Shu77c]. transformations [BW76a, Com78, Han70b, SRP76, SRP77, Wit72]. transformed [Wes73]. Transition [AC70, MMD78, OJB77, BR78, BFS73, BR75a, Cal77a, CHHS74, De 78, DR75, DCS77, FN73, GBV77, Gos75, Joh77, KP71a, KGB76, LH75, Lar78a, Lar79, LVDC74, MM77b, Por77, Por78, San71, Sch77c, SKW72, Sin75, SD74, SR77, Sto71, Van72]. transition-metal [BR78, BFS73, Joh77, KGB76]. transitions [BP75a, Ber73, Coo79, DM72, DD77b, Enz73, FWFM77, GML77, KD74, Lar78a, Mes73a, NT79, SIM75]. translation [BBA73, FM72]. Transmembrane [USK^{+}79, VU79]. Transmission [DS79, Wal77]. transport [Fro79, Hos78, KWC79, Mar79]. transversion [GRE^{+}77]. transversion-type [GRE^{+}77]. trapped [BBA73]. treated [BK75c]. treating [FHB77, Pai73]. Treatment [Sla71a, AM72a, BBB73, BTD79, CW78, Chr77, Cla71, CSH74, Csa78b, DS70b, Des72, FJP70, GS70a, Hil79, JBP76, Kau76, KMO77, MM74, MLB72, RTS74, RM79, SA77, SDJ79, Shu77c, SS71a, SW78, WRS70, LG71, SIM75]. treatments [AM74, AM77b, AM77a, CHD71, TS72]. tri [INI71]. tri-phenylene [INI71]. triatomic [KKW73]. tribute [CGLÖ76, Cus75]. Tridiagonal [Wel70]. tridiagonalization [Wel71]. trifluoromethyluracil [PLSL78]. Trifonow [Cal73a]. trigonal [Sch73a]. trigonometric [DDH78]. trimers [Moh78]. trimolecular [CW78]. triple [Kle70b, MPS70, Wei76]. triple-dipole [Wei76]. Triplet [SN79, Abd77, BB75, BP75a, IMN79, JŻ79b, NS78, PCMS77, PS70, San71, SMB^{+}79, SCYB75, TH78, TSE76]. tRNA [Pul77]. tropoelastin [RKOU77]. true [LH79]. Truncated [YJ76, ML79]. Truncated-atomic-sphere [YJ76]. Tryptamine [WOEG78]. tryptamines [JG74]. TTF [CHHS74, HRBB76, Lad75a, LRD^{+}77, LRD^{+}78]. TTF-TCNQ [LRD^{+}78]. Tunneling [CK73, Ada75b, CL77, Eis76, Feu77, Gol70, Wal77]. Turkuensis [Cal75b]. turn [RU76]. Turner [Cal76a]. Twenty [Lak76]. Two [BL77b, BH71, DT72, Gal72b, Höj79a, Höj79b, MSKM77, Thu75, Aba72, BA78, BB74, Bas78, BL77a, BBP71, Bil72, Cal72, Cal73c, ČPP75, CL78, Col79, Dea72, Eng72, GL74, HMV77, HLMV77, HMV78, Har75, Jet73, KM74a, Kij72, KBA76, Koc74, Koł74, LAM71, MKF77, Mal71b, MM74, Meh78, MH72, OK75, Pie71, Rat71, Ros78b, RB75b, Sha76, SDJ79, SS70, SJH79, SL70, TBB70, TTS^{+}79, WRS70, JBG70, Mal71a]. two-body [KBA76, Ros78b]. Two-Center [Thu75, BA78, Eng72, MH72, Sha76]. Two-Electron [BL77b, BH71, DT72, Aba72, BL77a, BBP71, Bil72, Cal72, Cal73c, GL74, KM74a, LAM71, MM74, RB75b, SJH79, WRS70]. two-matrix [Kij72]. Two-parameter [Höj79a, Höj79b, MSKM77]. two-particle [Rat71]. two-photon [MKF77]. two-point [ČPP75]. two-spin [SL70]. Type [Spe72a, BS74a, BS76a, BS71, Cha77, CL78, FTN77, GRE^{+}77, Höj79a, Höj79b, KGP79, LS72, LM72, LGR73, Mat73a, Mat76c, NY76, OK75, Pau77, PF74, PAM75, PW74a, RL77, Røe73, Sab71, SR72b, ST70, SLC78, SR72c, SF75, Lec76b].

U [Lon73, Sla73b]. U. [LB79a]. UHF [Cla70, Lun79, Sha77a, Sha78a]. ultimate [HPK79]. ultrasonic [PP71]. Ultraviolet [Gos78a, PLSL78, Ros75, Rab77]. unbounded [BFH78]. unconstrained [Muk75]. unconventional [Lec76a]. uncoupled [Vog79]. understand [KKK74]. Understanding [HP77a, Kau76]. uniaxial [Hub71]. Unified [MM74, CL73, LÖ77]. Uniform [MYTC77, BM73]. unit [CBV79]. Unitary [PH77, Mat74a, Mat76b, Mat79a, Nel79, Pal75, PW79, Sha77b, Sha78b, WP79]. unitary-group [Mat79a, Nel79]. united [CSH74]. units [HR79, Sri77]. Universal [SWN78]. universe [Lay77]. Universitatis [Cal75b]. University [Sab75a]. unperturbed [AM74]. Unrestricted [NN72, NN74, PČL78, SLB77, Jor73, KKA78, PS70]. unsaturated [NI70a, NI70b]. unstable [Tho76, TPC77]. Untersuchungen [HG75]. upon [DRCT78, LCPP79, Sco71]. Upper [BBK76, LTR74, Wei71]. upper- [LTR74]. uracil [CLP79b, GJP79, KD77, ZR78, ZSR79]. uracils [PLSL78]. urea [ED74]. Uridine [PB77]. Use [Eld73, GPP79, Jor73, LAV72, BW73, Car75, CM76, Ema74, Euw71, GM73a, GM73b, HHR70, Kau73, Laz78, LZ79, MCM73, Nee75, Ner76, PT75, PAM75, PGDM77, RK77, Rou79, RCE79, Sch72, Sim72, Sol73, Sti78, Wei74, Wol71]. used [BC75, KB77, San72]. useful [LH74]. Using [JBG70, SDD^{+}77, AM77a, BS74a, BS76a, Bri73, Bur72, Chi77, Chr73, ČP71, ČP72, CL78, CK74, CSH74, Cur78, Dan72, DD72, Goo76, GHJ77, HS70a, HR76, JG72, JSK75b, KPPH78, KP71b, KMO77, LF77, LGR73, Loc76, Loc77a, OMC76, OK75, PCMS77, SD79a, Sha76, SBM72, Sri77, TB74, Vla70, WN78, WRS70, WFS77]. usual [Cla71]. utility [PS73]. utilization [ST79a].

V [FTN77, HMC71, IM73, Kla78, Kle70b, KMO77, LGP75, LGM77, PPP79, PNC^{+}78, PK77b]. vacancy [SC79]. Väinö [Cal75b]. Valence [CESW70, Gal72b, Gin78, Goo76, Har70a, Har77a, HR79, WJ79, AM72b, BW76a, Bra79b, GN77, GV77, Har71a, HLYC77, Kle79, Lad74a, LS74a, MW74a, Mil71, Moh77, NN74, NG73, OK75, Pet77, PNC^{+}78, RTS74, RS76, SM73a, SM76b, SM76a, SD78, SBD^{+}70, TRS73, TS77, TS78, TS79, WMB75, TTS^{+}79]. Valence-Bond [Gal72b, HR79, GN77, OK75]. Valence-Electron [Goo76, SD78, WMB75]. valence-fluctuation [Bra79b]. valence-shell [BW76a, NN74]. valency [CR72]. Validity [LPT^{+}74, CD78, TL72]. valinomycin [Sun75b, ST78, ST79b, Sun74b]. valinomycin-potassium [Sun75b]. Values [Spe72b, BBK76, BS71, BSG77, BL77d, RK77, SK71, SG77a, SP77]. Vanadium [WFS77, PNJB78]. Variable [ES79, MR72, Sel75, Amo72, Csa76, Csa78b, PK75]. variance [HS75]. Variation [BL72, BEH77b, Zer74, Eps77, GZ72a, Mor78, MI78, MMM71, NI74, PGPP79, SY77, TS72, Wel76, Wel77b]. variation-perturbation [GZ72a, MMM71]. Variational [BR75a, BR72, Bri73, Csa78a, GbK77a, Kle77, KGB77, LFF79, Lud75, PSK77, ACD^{+}74, CK75, CSH74, Csa76, Csa78b, Csa79a, CZ75, DR75, GF76, HS77, Moc73, Moc74, MK76b, NS75, Pai71, PBC74a, TK78, WW71, Wei76, ZF76]. variationally [Csa77b, LTR74]. various [BB74, Gal74, JM76b, eAJ76]. varying [GZ72b]. vascular [ULM76]. vector [Bre73, Loh76]. vectors [LÖ77, Mor75]. velocity [Kob78, Sin73a]. verdazyl [MPN79]. Verdet [IM73]. Verfahren [KA76]. Verlag [Cal73a, Cal79a, Cal77b]. Vers [JJ78]. Versailles [Cal76b, Mar73b]. versions [BBL77]. versus [AM77b, Ber75, ES79, HPK79, Ive74, Kob78, PPB78, RBDV77]. Vertebrate [Eno76]. vertical [VG78, SG77a, TS79]. VI [CDEO76, Kle77, Smi79, Tho76]. Via [GB71, Als73, FB76, GBL^{+}73, SS77b, YL74]. vibrantibus [BP79]. Vibration [CK75, BBA73, KD79a, LRD^{+}77, LRD^{+}78]. vibration-translation [BBA73]. Vibrational [Tru72, BS77a, Bow79, GSE78, Sch71d, SR77, Ste77, Tru76]. vibrationally [MM72a]. vibronic [CII79, WZ70, TGSL70]. vibronique [TGSL70]. vibrotors [SS71b]. VII [GbK77a, Lai73, SPČ79, TS77]. VIII [Com77a, FWFM77, KPP78b, KGB77]. vinyl [Cla70, MMB70, MLB72]. viral [WLS77]. virial [BEF78, BFH78, FHB77, Lev78b, Nal78, WO74, WO75, WO76]. Virtual [Bee79, Cof74b, Mcd79]. visual [Kli79, MP71a]. vitamin [ED79, Ram78]. vivo [Dav77b, Web74, WS75]. Vleck [Sla71b, Gor71, Gor72]. Vol [Cal75c, Lar78c]. volatile [DBBJ76]. voltage [VU79]. voltage-dependent [VU79]. VRDDO [HKP79, HPK79, KPP78a, KPPH78, KPP78b, PK75, PK77a, PK77b]. VRDO [PK76b].

W [Gos78b, Gos78a, Mår78, Sab75a]. Waals [Mal71a, Mal71b, HČZ79, KKW73, LH79, WMV77]. Wahlgren [LB79a]. walk [And79]. Walker [Sab75a]. wall [ULM76]. Waller [Lim74b, Lim75a, Lim75b, Lim76b, Lim77b]. Walsh [CD71]. Walter [Ano77-28]. water [AF73, ACG74, AMJS79, CC79, Dau74, Hal79b, HSBD78, HK74, KPP78c, Laz79, LB79b, Mey71, Nir75, PDEB78, PPB78, RC78, Tal77]. Watson [Sun77]. Wave [Bre73, Ada73, ANR75, Ada75a, AB73, AM77b, AM77a, Als73, ADFL71, BB75, BN76, BDP73, BS71, BC79, Cas78a, CL73, CL78, CWE71, DS74, DRCT78, DLG79, DMM75, Euw71, Gal74, GN77, Gre71, HO73, HHR70, HKP79, Hun75a, Jak76a, Jak76b, Jak76c, Jak76d, JTR77, Joh70, JS71, JW72, Joh73, JM76a, Jol79, Jor73, KM74a, KC72, KKA78, KMO77, KK76c, LF77, LTR74, Li76, Lim76a, Loc77a, Lud75, Lud77, Meh73a, Meh73b, Moc74, ML76, MK76b, MM72b, MSM70, MBM71, NS75, NRS79, OKM74, Pau78, Pop71, RWFS71, Rob71, Rob72b, RK79, Sab75a, SH73, Sch72, SC76a, SS73a, SS73b, Sim72, SH75a, SD70, Sla70b, SWP77, Tan74, Tru73a, VF76, Wei74, WFS77, WV74, YJ76, Coo74, FJP70]. wave-functions [Jak76b, Jak76c, Jak76d, JTR77, KK76c, Lim76a, Loc77a, MK76b, SWP77, VF76]. Wave-mechanical [FJP70]. Wave-vector- [Bre73]. wavefunctions [OK75]. Wavenumber [SM73b]. waves [Ada72b, MSP73]. weak [Lya78]. weakly [MKT77]. Wechselwirkungen [BKW75]. weighted [DL75]. welcome [Sis74]. Welcoming [Bry74, Han77b, Mar76, SG77c, Exe74]. well [CK73, RSS75]. Wellenmechanische [FJP70]. Weyl [HBE78]. which [CEK72, KKK70]. wide [RMP71]. width [Fog71]. widths [FB76]. Wigner [Ahl70, Ahl70, LWA74]. Wiley [Cal79b, Cal75c, Dun78, Gos78a, Shu77a]. Wiley-Interscience [Cal75c, Gos78a]. windows [BHLM78]. Within [Spe72a, BMD70, Eps77, KM76, LB79b, LW76, Loc77a, Loc77b, Loc77d, PD78a, SDM74, WJ79, ZF76]. Woodward [Mat76a]. work [PS75a].

X [BFH74, Com77b, Con72, GML77, JW72, JM76a, KC79, Kau73, MS76a, PKK75, PK77c, RK77, SS71a, WC75b, YJ76, FM78b, HYL78, AD77, BB75, BS71, CFMO73, Dan72, DC73, Dec70, DCS77, GK75, GML77, HPK79, Joh73, Joh77, Li76, Mes73a, Min79, MFRE76, MMR74, PNJB78, PPC72, RK77, Sam75, SF76, Sla73a, Sla75, TS71, TS79, WK73]. x-ray [FM78b, HYL78, PNJB78, BS71, CFMO73, GK75, GML77, Mes73a, TS71]. X-rays [Dec70, MFRE76]. Xe [ES70]. XeF [SD78, TDB77]. XII [AF73, MF73]. xiii [Cal76a, MSP73]. XIV [LG71, MCP73]. XIX [MWY77]. XV [CR72, MPW75]. XVI [MW75a]. XVII [Mat76a]. XVIII [Mat76b]. XX [MW77]. XXI [MN79]. years [Lak76]. yesterday [Wyb73b]. Yield [HG78, GGJ77]. York [Ahl77a, Ano78c, Cal79b, Cal75c, Cal77b, Dun78, Gos78a, Lar76, Lar78c, Shu77b]. Young [Lim75b].

Zero [Col77a, Boy70a, Har71c, Nal78]. zeta [GG79a, GG79c]. zinc [MS72b, MHH74, RTM77]. zinc-histamine [MHH74]. ZN [JMP79, KPD78]. ZnSe [CESW70, CWE71]. zum [BKW75, HG75]. zur [KA76, HG75]. Zwanzig [HLMV77]. zwitterion [RC78].

References

Aldrich:1976:RMA

[AAC76] Haven S. Aldrich, John Aiken, and L. Chopin Cusachs. Radial moment analysis of isovalent hybridization in N_{2}. International Journal of Quantum Chemistry, 10(6):955–960, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adnan:1973:CMC

[AB73] Syed Sajjad Zaheer Adnan and Sadhan Basu. Cotton–Mouton constant of benzene calculated by free electron wave functions. International Journal of Quantum Chemistry, 7(2):405–406, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1979:TIDb

[AB79] John Avery and Erik Berg. Transferable integrals in a deformation density approach to crystal orbital calculations. II. International Journal of Quantum Chemistry, 16(6):1279–1299, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abarenkov:1972:CSF

[Aba72] I. V. Abarenkov. On the complete set of functions in the Heitler–London method for two-electron problem. International Journal of Quantum Chemistry, 6(1):93–100, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abdulnur:1971:BEE

[Abd71a] Suheil F. Abdulnur. Bounds to excitation energies of H in the presence of long range dispersion forces. International Journal of Quantum Chemistry, 5(S5):169–174, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Abdulnur:1971:BSO

[Abd71b] Suheil F. Abdulnur. Bounds to second-order perturbation energies (particularly dispersion) for excited states. International Journal of Quantum Chemistry, 5(5):525–537, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abdulnur:1976:IGP

[Abd76] S. F. Abdulnur. The interactions of glyoxals with proteins and DNA in relation to cancer. International Journal of Quantum Chemistry, 10(S3):59–64, January 12–14, 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Abdulnur:1977:IMM

[Abd77] Suheil F. Abdulnur. The interaction of methylglyoxal with methylamine and the triplet state. International Journal of Quantum Chemistry, 12(S4):217–223, January 9–12, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Andre:1979:NMP

[ABD^{+}79] J. M. André, L. A. Burke, J. Delhalle, G. Nicolas, and P. H. Durand. A nonempirical model potential technique for calculations of band structures of polymers. International Journal of Quantum Chemistry, 16(S13):283–291, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Adnan:1978:OOT

[ABM78] S. S. Z. Adnan, S. Bhattacharya, and D. Mukherjee. Orbital optimization techniques: Comparative study in a semiempirical framework. International Journal of Quantum Chemistry, 14 (3):289–297, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Absar:1972:MCI

[Abs72] Ilyas Absar. Mutually consistent integrals for approximate SCF–LCAO–MO calculations. International Journal of Quantum Chemistry, 6(S6):445–450, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Absar:1978:RHO

[Abs78] Ilyas Absar. Reduced Hamiltonian orbitals. II. Optimal orbital basis sets for the many-electron problem. International Journal of Quantum Chemistry, 13(6):777–790, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ali:1970:TPO

[AC70] M. Asgar Ali and Richard J. S. Crossley. Transition probability and oscillator strength by perturbation theory: 1s3p ^{1,3}P – 1s3d ^{1,3}D transition in helium isoelectronic sequence. International Journal of Quantum Chemistry, 4(5):541–543, September 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Absar:1976:RHO

[AC76] Ilyas Absar and A. J. Coleman. Reduced Hamiltonian orbitals. I. A new approach to the many-electron problem. International Journal of Quantum Chemistry, 10(S10):319–330, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Aslangul:1974:LTS

[ACD^{+}74] C. Aslangul, R. Constanciel, R. Daudel, L. Esnault, and E. V. Ludeña. The loge theory as a starting point for variational calculations. I. General formalism. International Journal of Quantum Chemistry, 8(4):499–522, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Aldrich:1974:HBM

[ACG74] Haven S. Aldrich, L. Chopin Cusachs, and L. P. Gary. Hydrogen bonding in mixed dimer systems with water. International Journal of Quantum Chemistry, 8(S1):55–64, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Arnau:1972:EMS

[ACH72] Catalina Arnau, Ramón Carbó, and Sigeru Huzinaga. Eigenspace manipulation in SCF theory. International Journal of Quantum Chemistry, 6(5):843–847, September 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Arko:1975:LEP

[ACK^{+}75] A. J. Arko, G. Crabtree, J. B. Ketterson, F. M. Mueller, P. F. Walch, L. R. Windmiller, Z. Fisk, R. F. Hoyt, A. C. Mota, R. Viswanathan, D. E. Ellis, A. J. Freeman, and J. Rath. Large electron–phonon interaction but low-temperature superconductivity in LaB_{6}. International Journal of Quantum Chemistry, 9(S9):569–578, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Alves:1977:MSX

[AD77] J. L. Alves and M. L. De Siqueira. Multiple scattering Xα calculation of crystal field parameters for a divalent thulium impurity in calcium fluoride. International Journal of Quantum Chemistry, 12(S11):75–79, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Adawi:1972:PES

[Ada72a] I. Adawi. Perturbation expansion of the single-electron density matrix for the impurity problem. International Journal of Quantum Chemistry, 6(S6):443–444, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Adawi:1972:SWO

[Ada72b] I. Adawi. Scattering of waves in one dimension. International Journal of Quantum Chemistry, 6(S6):453–454, January 17– 22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Adams:1973:SSE

[Ada73] William H. Adams. On the solution of the Schrödinger equation for H_{2} in terms of a wave function least distorted from a product of atomic wave functions. International Journal of Quantum Chemistry, 7(S7):127–133, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Adams:1975:SEW

[Ada75a] William H. Adams. On the separability of electronic wave functions. International Journal of Quantum Chemistry, 9(S9):367– 373, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Adawi:1975:SDT

[Ada75b] I. Adawi. Semiclassical derivation of the tunneling probability in magnetic breakdown. International Journal of Quantum Chemistry, 9(S9):477–480, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Adams:1977:HLD

[Ada77] William H. Adams. The Heitler-London description of the covalent bond in H_{2}. International Journal of Quantum Chemistry, 12(S11):193–199, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement:

Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Adamowicz:1978:CAP

[Ada78] Ludwik Adamowicz. A comparative analysis of pair equations for the H_{2} molecule. International Journal of Quantum Chemistry, 13(3):265–270, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adawi:1979:PES

[Ada79] I. Adawi. Perturbation expansion of the single-particle density matrix. International Journal of Quantum Chemistry, 16(S13): 81–92, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Andre:1976:FSG

[ADDLG76] J. M. André, J. Delhalle, C. Demanet, and M. E. Lambert-Gerard. A floating spherical Gaussian orbital model for polymers: I. general formalism and computational procedure. International Journal of Quantum Chemistry, 10(S10):99–105, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Andre:1971:CLW

[ADFL71] J. M. Andre, J. Delhalle, J. G. Fripiat, and G. Leroy. Computation of LCAO wave functions for ground states of polymers and solids. International Journal of Quantum Chemistry, 5(1):67– 84, January 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abu-eittah:1978:MOS

[AeAs78] R. Abu-eittah and F. Al-sageir. Molecular orbital study of the electronic absorption spectra of dithienyls. International Journal of Quantum Chemistry, 13(5):565–577, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Afzal:1973:FSG

[AF73] M. Afzal and A. A. Frost. A floating spherical Gaussian orbital model of molecular structure. XII. Analysis of energy terms affecting the geometry of the water molecule. International Journal of Quantum Chemistry, 7(1):51–62, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Andre:1978:LRC

[AFD^{+}78] J. M. André, J. G. Fripiat, C. H. Demanet, J. L. Brédas, and J. Delhalle. Long-range Coulombic interactions in the theory of polymers: A statement of the problem and a method for calculation by the Fourier transformation technique. International Journal of Quantum Chemistry, 14(S12):233–247, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

AAsbrink:1978:EST

[ÅFL78] L. Åsbrink, C. Fridh, and E. Lindholm. Electronic structure of TCNQ, studied with HAM /3. International Journal of Quantum Chemistry, 13(3):331–347, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ahlrichs:1970:KBW

[Ahl70] Reinhart Ahlrichs. Die Konvergenz der Brillouin–Wigner Störungsrechnung. (German) [Convergence of the Brillouin– Wigner perturbation calculation]. International Journal of Quantum Chemistry, 4(2):149–172, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ahlberg:1974:DCF

[Ahl74] R. Ahlberg. Dispersion coefficients and frequency-dependent polarizabilities. International Journal of Quantum Chemistry, 8(S8):363–377, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Ahlberg:1977:BRB

[Ahl77a] Ragnar Ahlberg. Book review: Handbook of atomic data. S. Fraga, I. Karwowski, and K. M. S. Sexena. Elsevier Scientific Publishing Company, Amsterdam, Oxford, New York, 1976. 551 pp. Price: $74. International Journal of Quantum Chemistry, 12(1):231, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [FKS76].

Ahlberg:1977:CPG

[Ahl77b] Ragnar Ahlberg. Convergence properties of generalized inner projections to dynamic polarizabilities. International Journal of Quantum Chemistry, 12(3):443–447, September 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Andrews:1979:CAC

[AJ79] P. R. Andrews and G. P. Jones. Conformational analysis of convulsant and anticonvulsant barbiturates. International Journal of Quantum Chemistry, 16(S6):439–452, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Andrea:1978:DDT

[AJK78] Tariq A. Andrea, E. C. Jorgensen, and Peter A. Kollman. Differentiation of D- and L-Thyroxine by the plasma protein prealbumin. International Journal of Quantum Chemistry, 14(S5): 191–200, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Andre:1971:AEB

[AL71] Jean-Marie André and Georges Leroy. All-electrons band structure of Polyene. International Journal of Quantum Chemistry, 5(5):557–563, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Aldrich:1975:MOS

[Ald75] Haven Scott Aldrich. Molecular orbital studies involving sulfur and selenium. International Journal of Quantum Chemistry, 9 (S2):271–278, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Almlof:1974:ICP

[Alm74] Jan Almlöf. Ab initio calculations on porphin. International Journal of Quantum Chemistry, 8(6):915–924, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Alster:1973:EDM

[Als73] J. Alster. Electric dipole moment calculations of polyatomic molecules via iterative extended Hückel wave functions. International Journal of Quantum Chemistry, 7(S7):311–318, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Ali:1972:FOC

[AM72a] M. K. Ali and W. J. Meath. A floating one centre perturbation treatment for H-like molecules. International Journal of Quantum Chemistry, 6(5):949–966, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Almlof:1972:VED

[AM72b] Jan Almlöf and Olle Mårtensson. The valence electron density distribution of hydrogen bonded systems in the iterative extended Hückel approach. III. The pyrrole–pyridine system. International Journal of Quantum Chemistry, 6(3):491–500, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ali:1974:FOC

[AM74] M. K. Ali and William J. Meath. Floating one-center perturbation treatments for H-like molecules based on screened hydrogen atom or molecular puff unperturbed problems. International Journal of Quantum Chemistry, 8(1):119–136, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ali:1977:COCb

[AM77a] M. K. Ali and W. J. Meath. On the convergence of onecentre partial wave treatments for diatomic molecules. II. Heteronuclear one-electron diatomic molecules using the 1sσ, 2sσ, 3sσ, 2pσ, 3pσ, and 3pσ states of HeH^{++} as models. International Journal of Quantum Chemistry, 12(5):897–914, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ali:1977:COCa

[AM77b] M. K. Ali and William J. Meath. On the convergence of one centre partial wave treatments for diatomic molecules. H (1sσ_{g}) and Coulomb versus electron exchange intermolecular forces. International Journal of Quantum Chemistry, 12(1): 35–59, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Aldrich:1973:CBF

[AMC73] H. S. Aldrich, J. T. Mague, and L. C. Cusachs. Conformation and bonding of five- and six-coordinate rhodium complexes from ARCANA. International Journal of Quantum Chemistry, 7(S7):239–247, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Andronikashvili:1979:CPB

[AMJS79] E. L. Andronikashvili, G. M. Mrevlishvili, G. Sh. Japaridze, and V. M. Sokhadze. Conformational properties of biomacromolecules (proteins and nucleic acids) at low temperatures and the problem of interaction of biopolymers with water molecules and ions of dissolved salts. International Journal of Quantum Chemistry, 16(2):367–377, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Amos:1971:CSS

[Amo71] A. T. Amos. A comment on the super-secular-equation in a symmetry adapted basis. International Journal of Quantum Chemistry, 5(6):699–702, November 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Amos:1972:CVP

[Amo72] A. T. Amos. A change of variable for the perturbation parameter in Rayleigh–Schrödinger perturbation theory. International Journal of Quantum Chemistry, 6(1):125–130, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Almazov:1979:IBE

[AMSS79] V. P. Almazov, Yu. V. Morozov, F. A. Savin, and B. S. Sukhareva. Interrelations between electronic structure and spatial geometry of specific ligands in the functioning active site of some pyridoxal-p-dependent enzymes. International Journal of Quantum Chemistry, 16(4):769–775, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anderson:1979:QCR

[And79] James B. Anderson. Quantum chemistry by random walk: H_{4} square. International Journal of Quantum Chemistry, 15(1): 109–120, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:Aa

[Ano70a] Anonymous. Announcements. International Journal of Quantum Chemistry, 4(1):121, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:Ab

[Ano70b] Anonymous. Announcements. International Journal of Quantum Chemistry, 4(4):451–452, July 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:Ac

[Ano70c] Anonymous. Announcements. International Journal of Quantum Chemistry, 4(6):631, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:BRa

[Ano70d] Anonymous. Books received. International Journal of Quantum Chemistry, 4(1):123–124, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:BRb

[Ano70e] Anonymous. Books received. International Journal of Quantum Chemistry, 4(2):219–221, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:Ma

[Ano70f] Anonymous. Masthead. International Journal of Quantum Chemistry, 4(1):fmi–fmii, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:Mb

[Ano70g] Anonymous. Masthead. International Journal of Quantum Chemistry, 4(2):fmi–fmii, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:Mc

[Ano70h] Anonymous. Masthead. International Journal of Quantum Chemistry, 4(3):fmi–fmii, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:Md

[Ano70i] Anonymous. Masthead. International Journal of Quantum Chemistry, 4(4):fmi–fmii, July 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:Me

[Ano70j] Anonymous. Masthead. International Journal of Quantum Chemistry, 4(5):fmi–fmii, September 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:Mf

[Ano70k] Anonymous. Masthead. International Journal of Quantum Chemistry, 4(6):fmi–fmii, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1970:M

[Ano70l] Anonymous. Masthead. International Journal of Quantum Chemistry, 5(S4):i, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Anonymous:1971:Aa

[Ano71a] Anonymous. Announcement. International Journal of Quantum Chemistry, 5(1):115, January 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1971:Ab

[Ano71b] Anonymous. Announcements. International Journal of Quantum Chemistry, 5(4):475, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1971:Ma

[Ano71c] Anonymous. Masthead. International Journal of Quantum Chemistry, 5(1):fmi–fmii, January 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1971:Mb

[Ano71d] Anonymous. Masthead. International Journal of Quantum Chemistry, 5(2):fmi–fmii, March 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1971:Mc

[Ano71e] Anonymous. Masthead. International Journal of Quantum Chemistry, 5(3):fmi–fmii, May 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1971:Md

[Ano71f] Anonymous. Masthead. International Journal of Quantum Chemistry, 5(4):fmi–fmii, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1971:Me

[Ano71g] Anonymous. Masthead. International Journal of Quantum Chemistry, 5(5):fmi–fmii, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1971:Mf

[Ano71h] Anonymous. Masthead. International Journal of Quantum Chemistry, 5(6):fmi–fmii, November 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1971:M

[Ano71i] Anonymous. Masthead. International Journal of Quantum Chemistry, 5(S5):fmi, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Anonymous:1972:Aa

[Ano72a] Anonymous. Announcements. International Journal of Quantum Chemistry, 6(1):197–198, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1972:Ab

[Ano72b] Anonymous. Announcements. International Journal of Quantum Chemistry, 6(4):799, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1972:Ma

[Ano72c] Anonymous. Masthead. International Journal of Quantum Chemistry, 6(1):fmi–fmii, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1972:Mb

[Ano72d] Anonymous. Masthead. International Journal of Quantum Chemistry, 6(2):fmi–fmii, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1972:Mc

[Ano72e] Anonymous. Masthead. International Journal of Quantum Chemistry, 6(3):fmi–fmii, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1972:Md

[Ano72f] Anonymous. Masthead. International Journal of Quantum Chemistry, 6(4):fmi–fmii, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1972:Me

[Ano72g] Anonymous. Masthead. International Journal of Quantum Chemistry, 6(5):fmi–fmii, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1972:M

[Ano72h] Anonymous. Masthead. International Journal of Quantum Chemistry, 6(S6):fmi, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Anonymous:1973:Aa

[Ano73a] Anonymous. Announcement. International Journal of Quantum Chemistry, 7(1):135, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1973:Ab

[Ano73b] Anonymous. Announcements. International Journal of Quantum Chemistry, 7(3):637, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1973:Ac

[Ano73c] Anonymous. Announcements. International Journal of Quantum Chemistry, 7(6):1227–1228, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1973:Ma

[Ano73d] Anonymous. Masthead. International Journal of Quantum Chemistry, 7(1):fmi–fmii, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1973:Mb

[Ano73e] Anonymous. Masthead. International Journal of Quantum Chemistry, 7(2):fmi–fmii, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1973:Mc

[Ano73f] Anonymous. Masthead. International Journal of Quantum Chemistry, 7(3):fmi–fmii, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1973:Md

[Ano73g] Anonymous. Masthead. International Journal of Quantum Chemistry, 7(4):fmi–fmii, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1973:Me

[Ano73h] Anonymous. Masthead. International Journal of Quantum Chemistry, 7(5):fmi–fmii, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1973:Mf

[Ano73i] Anonymous. Masthead. International Journal of Quantum Chemistry, 7(6):fmi–fmii, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1973:M

[Ano73j] Anonymous. Masthead. International Journal of Quantum Chemistry, 7(S7):i–ii, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Anonymous:1974:Aa

[Ano74a] Anonymous. Announcement. International Journal of Quantum Chemistry, 8(2):315–316, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1974:Ab

[Ano74b] Anonymous. Announcement. International Journal of Quantum Chemistry, 8(4):633, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1974:BR

[Ano74c] Anonymous. Book Reviews. International Journal of Quantum Chemistry, 8(6):995, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1974:Ma

[Ano74d] Anonymous. Masthead. International Journal of Quantum Chemistry, 8(1):fmi–fmii, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1974:Mb

[Ano74e] Anonymous. Masthead. International Journal of Quantum Chemistry, 8(2):fmi–fmii, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1974:Mc

[Ano74f] Anonymous. Masthead. International Journal of Quantum Chemistry, 8(3):fmi–fmii, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1974:Md

[Ano74g] Anonymous. Masthead. International Journal of Quantum Chemistry, 8(4):fmi–fmii, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1974:Me

[Ano74h] Anonymous. Masthead. International Journal of Quantum Chemistry, 8(5):fmi–fmii, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1974:Mf

[Ano74i] Anonymous. Masthead. International Journal of Quantum Chemistry, 8(6):fmi–fmii, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1974:Mg

[Ano74j] Anonymous. Masthead. International Journal of Quantum Chemistry, 8(S1):fmi, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1974:Mh

[Ano74k] Anonymous. Masthead. International Journal of Quantum Chemistry, 8(S8):fmi, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Anonymous:1975:A

[Ano75a] Anonymous. Announcements. International Journal of Quantum Chemistry, 9(3):557, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1975:Ma

[Ano75b] Anonymous. Masthead. International Journal of Quantum Chemistry, 9(1):fmi–fmii, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1975:Mb

[Ano75c] Anonymous. Masthead. International Journal of Quantum Chemistry, 9(2):fmi–fmii, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1975:Mc

[Ano75d] Anonymous. Masthead. International Journal of Quantum Chemistry, 9(3):fmi–fmii, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1975:Md

[Ano75e] Anonymous. Masthead. International Journal of Quantum Chemistry, 9(4):fmi–fmii, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1975:Me

[Ano75f] Anonymous. Masthead. International Journal of Quantum Chemistry, 9(5):fmi–fmii, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1975:Mf

[Ano75g] Anonymous. Masthead. International Journal of Quantum Chemistry, 9(6):fmi–fmii, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1975:Mg

[Ano75h] Anonymous. Masthead. International Journal of Quantum Chemistry, 9(S2):fmi, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1975:Mh

[Ano75i] Anonymous. Masthead. International Journal of Quantum Chemistry, 9(S9):i, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Anonymous:1976:Aa

[Ano76a] Anonymous. Announcements. International Journal of Quantum Chemistry, 10(1):187–189, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Ab

[Ano76b] Anonymous. Announcements. International Journal of Quantum Chemistry, 10(2):381, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Ac

[Ano76c] Anonymous. Announcements. International Journal of Quantum Chemistry, 10(4):717, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Ad

[Ano76d] Anonymous. Announcements. International Journal of Quantum Chemistry, 10(6):1089–1090, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:DRD

[Ano76e] Anonymous. Drug-receptor and Drug-Enzyme Interactions: A Basic Approach. International Journal of Quantum Chemistry, 10(3):555–557, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Ma

[Ano76f] Anonymous. Masthead. International Journal of Quantum Chemistry, 10(1):fmi–fmii, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Mb

[Ano76g] Anonymous. Masthead. International Journal of Quantum Chemistry, 10(2):fmi–fmii, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Mc

[Ano76h] Anonymous. Masthead. International Journal of Quantum Chemistry, 10(3):fmi–fmii, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Md

[Ano76i] Anonymous. Masthead. International Journal of Quantum Chemistry, 10(4):fmi–fmii, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Me

[Ano76j] Anonymous. Masthead. International Journal of Quantum Chemistry, 10(5):fmi–fmii, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Mf

[Ano76k] Anonymous. Masthead. International Journal of Quantum Chemistry, 10(6):fmi–fmii, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1976:Mg

[Ano76l] Anonymous. Masthead. International Journal of Quantum Chemistry, 10(S3):fmi, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1976:Mh

[Ano76m] Anonymous. Masthead. International Journal of Quantum Chemistry, 10(S10):iv, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1976:RSC

[Ano76n] Anonymous. Reflections on the scientific career of Jeremy Musher. International Journal of Quantum Chemistry, 10(1): 1–3, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Ab

[Ano77a] Anonymous. Announcement. International Journal of Quantum Chemistry, 11(2):380, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Af

[Ano77b] Anonymous. Announcement. International Journal of Quantum Chemistry, 12(2):431, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Ag

[Ano77c] Anonymous. Announcement. International Journal of Quantum Chemistry, 12(2):432, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Ak

[Ano77d] Anonymous. Announcement. International Journal of Quantum Chemistry, 12(6):1049, December 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Aa

[Ano77e] Anonymous. Announcements. International Journal of Quantum Chemistry, 11(1):187, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Ac

[Ano77f] Anonymous. Announcements. International Journal of Quantum Chemistry, 11(3):524, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Ad

[Ano77g] Anonymous. Announcements. International Journal of Quantum Chemistry, 11(4):703, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Ae

[Ano77h] Anonymous. Announcements. International Journal of Quantum Chemistry, 11(4):704, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Ah

[Ano77i] Anonymous. Announcements. International Journal of Quantum Chemistry, 12(3):591, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Ai

[Ano77j] Anonymous. Announcements. International Journal of Quantum Chemistry, 12(4):784, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Aj

[Ano77k] Anonymous. Announcements. International Journal of Quantum Chemistry, 12(5):969, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:EH

[Ano77l] Anonymous. Erich Hückel. International Journal of Quantum Chemistry, 12(S11):3, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Anonymous:1977:LFH

[Ano77m] Anonymous. Lebenslauf von Friedrich Hund. International Journal of Quantum Chemistry, 12(S11):5–10, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Anonymous:1977:Ma

[Ano77n] Anonymous. Masthead. International Journal of Quantum Chemistry, 11(1):fmi–fmii, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Mb

[Ano77o] Anonymous. Masthead. International Journal of Quantum Chemistry, 11(2):fmi–fmii, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Mc

[Ano77p] Anonymous. Masthead. International Journal of Quantum Chemistry, 11(3):fmi–fmii, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Md

[Ano77q] Anonymous. Masthead. International Journal of Quantum Chemistry, 11(4):fmi–fmii, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Me

[Ano77r] Anonymous. Masthead. International Journal of Quantum Chemistry, 11(5):fmi–fmii, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Mf

[Ano77s] Anonymous. Masthead. International Journal of Quantum Chemistry, 11(6):fmi–fmii, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Mg

[Ano77t] Anonymous. Masthead. International Journal of Quantum Chemistry, 12(1):fmi–fmii, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Mh

[Ano77u] Anonymous. Masthead. International Journal of Quantum Chemistry, 12(2):fmi–fmii, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Mi

[Ano77v] Anonymous. Masthead. International Journal of Quantum Chemistry, 12(3):fmi–fmii, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Mj

[Ano77w] Anonymous. Masthead. International Journal of Quantum Chemistry, 12(4):fmi–fmii, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Mk

[Ano77x] Anonymous. Masthead. International Journal of Quantum Chemistry, 12(5):fmi–fmii, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Ml

[Ano77y] Anonymous. Masthead. International Journal of Quantum Chemistry, 12(6):fmi–fmii, December 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1977:Mm

[Ano77z] Anonymous. Masthead. International Journal of Quantum Chemistry, 12(S4):fmi, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1977:Mn

[Ano77-27] Anonymous. Masthead. International Journal of Quantum Chemistry, 12(S11):fmi, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Anonymous:1977:WH

[Ano77-28] Anonymous. Walter Heitler. International Journal of Quantum Chemistry, 12(S11):1–2, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Anonymous:1978:Aa

[Ano78a] Anonymous. Announcement. International Journal of Quantum Chemistry, 13(1):160, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Ab

[Ano78b] Anonymous. Announcement. International Journal of Quantum Chemistry, 13(6):811, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:BRB

[Ano78c] Anonymous. Book review: Theory and practice of MO calculations on organic molecules. I. G. Csizmadia. Elsevier scientific publishing company, Amsterdam, Oxford, New York, 1976. price: $38.50. International Journal of Quantum Chemistry, 13(1):159, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Csi76].

Anonymous:1978:FCR

[Ano78d] Anonymous. Fifth conference on reduced density matrices. International Journal of Quantum Chemistry, 13(6):693–695, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:I

[Ano78e] Anonymous. Introduction. International Journal of Quantum Chemistry, 14(5):iii, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:LP

[Ano78f] Anonymous. Linus Pauling. International Journal of Quantum Chemistry, 14(S5):3–4, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1978:Ma

[Ano78g] Anonymous. Masthead. International Journal of Quantum Chemistry, 13(1):fmi–fmii, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Mb

[Ano78h] Anonymous. Masthead. International Journal of Quantum Chemistry, 13(2):fmi–fmii, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Mc

[Ano78i] Anonymous. Masthead. International Journal of Quantum Chemistry, 13(3):fmi–fmii, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Md

[Ano78j] Anonymous. Masthead. International Journal of Quantum Chemistry, 13(4):fmi–fmii, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Me

[Ano78k] Anonymous. Masthead. International Journal of Quantum Chemistry, 13(5):fmi–fmii, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Mf

[Ano78l] Anonymous. Masthead. International Journal of Quantum Chemistry, 13(6):fmi–fmii, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Mg

[Ano78m] Anonymous. Masthead. International Journal of Quantum Chemistry, 14(1):fmi–fmii, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Mh

[Ano78n] Anonymous. Masthead. International Journal of Quantum Chemistry, 14(2):fmi–fmii, August 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Mi

[Ano78o] Anonymous. Masthead. International Journal of Quantum Chemistry, 14(3):fmi–fmii, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Mj

[Ano78p] Anonymous. Masthead. International Journal of Quantum Chemistry, 14(4):fmi–fmii, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Mk

[Ano78q] Anonymous. Masthead. International Journal of Quantum Chemistry, 14(5):fmi–fmii, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Ml

[Ano78r] Anonymous. Masthead. International Journal of Quantum Chemistry, 14(6):fmi–fmii, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1978:Mm

[Ano78s] Anonymous. Masthead. International Journal of Quantum Chemistry, 14(S5):fmi, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1978:Mn

[Ano78t] Anonymous. Masthead. International Journal of Quantum Chemistry, 14(S12):fmi, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Anonymous:1978:PLP

[Ano78u] Anonymous. Publications of Linus Pauling. International Journal of Quantum Chemistry, 14(S5):5–8, March 3–11, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1979:Ma

[Ano79a] Anonymous. Masthead. International Journal of Quantum Chemistry, 15(1):fmi–fmii, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Mb

[Ano79b] Anonymous. Masthead. International Journal of Quantum Chemistry, 15(2):fmi–fmii, February 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Mc

[Ano79c] Anonymous. Masthead. International Journal of Quantum Chemistry, 15(3):fmi–fmii, March 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Md

[Ano79d] Anonymous. Masthead. International Journal of Quantum Chemistry, 15(4):fmi–fmii, April 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Me

[Ano79e] Anonymous. Masthead. International Journal of Quantum Chemistry, 15(5):fmi–fmii, May 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Mf

[Ano79f] Anonymous. Masthead. International Journal of Quantum Chemistry, 15(6):fmi–fmii, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Mg

[Ano79g] Anonymous. Masthead. International Journal of Quantum Chemistry, 16(1):fmi–fmii, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Mh

[Ano79h] Anonymous. Masthead. International Journal of Quantum Chemistry, 16(2):fmi–fmii, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Mi

[Ano79i] Anonymous. Masthead. International Journal of Quantum Chemistry, 16(3):fmi–fmii, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Mj

[Ano79j] Anonymous. Masthead. International Journal of Quantum Chemistry, 16(4):fmi–fmii, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Mk

[Ano79k] Anonymous. Masthead. International Journal of Quantum Chemistry, 16(5):fmi–fmii, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Ml

[Ano79l] Anonymous. Masthead. International Journal of Quantum Chemistry, 16(6):fmi–fmii, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1979:Mm

[Ano79m] Anonymous. Masthead. International Journal of Quantum Chemistry, 16(S6):33–45, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1979:Mn

[Ano79n] Anonymous. Masthead. International Journal of Quantum Chemistry, 16(S13):fmi, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Anonymous:1979:SES

[Ano79o] Anonymous. Subsymposium on electron scattering, molecular collisions, and their computational methods. International Journal of Quantum Chemistry, 16(S13):473, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Adams:1975:PTW

[ANR75] Steven Adams, Shlomo Nir, and Robert Rein. Perturbation theories and wave functions for calculation of electronic polarizabilities application to DNA bases. International Journal of Quantum Chemistry, 9(4):701–710, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adams:1977:AGM

[APČ77] B. G. Adams, J. Paldus, and J. Čížek. Application of graphical methods of spin algebras to limited CI approaches. II. A simple open shell case. International Journal of Quantum Chemistry, 11(5):849–867, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Aronowitz:1978:ICN

[Aro78] S. Aronowitz. Intermolecular correlation in a new approximation scheme. International Journal of Quantum Chemistry, 14 (3):253–269, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ali:1978:ROD

[AS78] M. A. Ali and S. R. Samanta. Relative ordering of ^{4}S ^{e}, ^{4}D e, and ^{4}D^{o} doubly excited states of lithium isoelectronic sequence. International Journal of Quantum Chemistry, 14(S12): 523–527, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

AAgren:1977:BSI

[ÅSW77] H. Ågren, S. Svensson, and U. I. Wahlgren. A basis set investigation for the oxygen 1s ionization potential in H_{2}O. International Journal of Quantum Chemistry, 11(2):317–324, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anno:1975:SRS

[AT75] Tosinobu Anno and Hirohide Teruya. Some remarks on the semiempirical one-center electron repulsion integrals. International Journal of Quantum Chemistry, 9(1):51–65, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Atkins:1974:QHC

[Atk74] P. W. Atkins. Quanta: a handbook of concepts, volume 21 of Oxford chemistry series. Clarendon Press, New York, NY, USA, 1974. ISBN 0-19-855493-1, 0-19-855494-x. 309 pp. LCCN 2/28.

Atkins:1991:QHC

[Atk91] P. W. Atkins. Quanta: a handbook of concepts. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 1991. ISBN 0-19-855572-5, 0-19-855573-3. 434 pp. LCCN ????

Avery:1972:QTA

[Ave72] John Avery. The quantum theory of atoms, molecules and photons. European chemistry series. McGraw-Hill, New York, NY, USA, 1972. ISBN 0-07-094178-5, 0-07-084013-X (paperback). xv + 378 pp. LCCN QD462 .A9.

Avery:1975:EPP

[Ave75] John S. Avery. Enhanced probability for the photoelectric effect in a Faraday cage. International Journal of Quantum Chemistry, 9(3):445–457, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1979:TIDc

[Ave79a] J. Avery. Transferable integrals in a deformation density approach to crystal orbital calculations. III. International Journal of Quantum Chemistry, 16(S13):403–412, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Avery:1979:TIDa

[Ave79b] John Avery. Transferable integrals in a deformation density approach to crystal orbital calculations. I. International Journal of Quantum Chemistry, 16(6):1265–1277, December 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1979:TID

[Ave79c] John Avery. Transferable integrals in a deformation density approach to crystal orbital calculations. III. International Journal of Quantum Chemistry, 16(S13):403–412, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Abu-Zeid:1972:SDP

[AZ72] Mohyi-Eldin M. Abu-Zeid. Solvent dependence of pyrene monomer and excimer fluorescence quenching by heavy atoms. International Journal of Quantum Chemistry, 6(S6):279–293, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Barnsley:1978:APE

[BA78] Michael F. Barnsley and Jacques G. Aguilar. On the approximation of potential-energy functions for two-center systems. International Journal of Quantum Chemistry, 13(5):641–677, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Baird:1974:SHB

[Bai74] N. Colin Baird. Simulation of hydrogen bonding in biological systems: Ab initio calculations for NH_{3} − NH_{3} and NH_{3} − NH_{4}^{+}. International Journal of Quantum Chemistry, 8(S1):49–54, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ballhausen:1971:CLF

[Bal71] C. J. Ballhausen. Crystal and ligand field theory. International Journal of Quantum Chemistry, 5(S5):373–377, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Balslev:1978:STD

[Bal78] Erik Balslev. Scattering theory of dilated three-body Schrödinger operators. International Journal of Quantum Chemistry, 14(4):361–370, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bandrauk:1974:ITF

[Ban74] A. D. Bandrauk. Integral transform functions and separable potentials. International Journal of Quantum Chemistry, 8 (2):209–221, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barr:1970:SPT

[Bar70] Tery L. Barr. Several perturbation theory methods for accurate calculations on complex systems. International Journal of Quantum Chemistry, 5(S4):239–250, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Barnett:1977:CES

[Bar77] Gene Barnett. Comments on excited state polarizabilities of molecules. International Journal of Quantum Chemistry, 12 (2):427–429, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bardsley:1978:CSI

[Bar78] J. N. Bardsley. Complex scaling: An introduction. International Journal of Quantum Chemistry, 14(4):343–352, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barfield:1979:PPC

[Bar79] W. D. Barfield. Partial photoionization cross sections and radiative recombination rate coefficients for Li-like ions. II. International Journal of Quantum Chemistry, 16(S13):683–693, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Bass:1975:CET

[Bas75] George E. Bass. The comorosan effect: Toward a perspective. International Journal of Quantum Chemistry, 9(S2):321– 324, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Basu:1978:IBT

[Bas78] Sadhan Basu. Interaction between two harmonic oscillators and spectroscopy of hydrogen bond. International Journal of Quantum Chemistry, 13(2):261–262, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Banyard:1970:MFE

[BB70] K. E. Banyard and C. C. Baker. Molecular formation and electron correlation in HeH^{+}. International Journal of Quantum Chemistry, 4(4):431–450, July 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bartlett:1971:GSR

[BB71] Rodney J. Bartlett and Erkki J. Brändas. Geometric sumrule and the reduced partitioning procedure. International Journal of Quantum Chemistry, 5(S5):151–159, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Basilevsky:1972:IIR

[BB72a] M. V. Basilevsky and M. M. Berenfeld. Intermolecular interactions in the region of small overlap. International Journal of Quantum Chemistry, 6(1):23–45, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Basilevsky:1972:SPT

[BB72b] M. V. Basilevsky and M. M. Berenfeld. SCF perturbation theory and intermolecular interactions. International Journal of Quantum Chemistry, 6(3):555–574, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bhattacharyya:1972:TCS

[BB72c] Subirnath Bhattacharyya and Amiyansu Bhaumik. Theoretical calculations on σ-systems. III. Alkyl halides and alkanes. International Journal of Quantum Chemistry, 6(5):935–948, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bunge:1973:SES

[BB73] Carlos F. Bunge and Annik Bunge. Symmetry eigenfunctions suitable for many-electron theories and calculations. I. Mainly atoms. International Journal of Quantum Chemistry, 7(5): 927–944, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Basilevsky:1974:ITH

[BB74] M. V. Basilevsky and M. M. Berenfeld. Interaction of two H_{2} molecules: Comparative study of various perturbation procedures. International Journal of Quantum Chemistry, 8(3):467– 489, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bagus:1975:STS

[BB75] P. S. Bagus and B. I. Bennett. Singlet–triplet splittings as obtained from the X α-scattered wave method: A theoretical analysis. International Journal of Quantum Chemistry, 9(1): 143–148, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bunge:1978:CAE

[BB78] Carlos F. Bunge and Annik Vivier Bunge. Calculations of atomic electron affinities. International Journal of Quantum Chemistry, 14(S12):345–355, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Blaisten-Barojas:1973:AGF

[BBA73] E. Blaisten-Barojas and M. Allavena. An application of Green’s functions to the study of the vibration-translation coupling of trapped oscillators in a linear chain. International Journal of Quantum Chemistry, 7(2):195–213, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bartlett:1973:TCE

[BBB73] Rodney J. Bartlett, John C. Bellum, and Erkki J. Brändas. The treatment of correlation effects in second-order properties. International Journal of Quantum Chemistry, 7(S7):449– 462, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Bolotin:1979:TIB

[BBG79] Adolph B. Bolotin, Victor A. Bolotin, and Victoria L. Gineityte. Torsional isomerization of biologically active bicyclic molecules. International Journal of Quantum Chemistry, 16 (4):839–845, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Behling:1976:ULB

[BBK76] R. Behling, A. Bongers, and T. Küpper. Upper and lower bounds to critical values of the Hartree operator. International Journal of Quantum Chemistry, 10(6):985–992, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bersuker:1972:QRA

[BBL72] I. B. Bersuker, S. S. Budnikov, and B. A. Leizerov. Quasirelativistic approximation in the SCF-MO-LCAO method. International Journal of Quantum Chemistry, 6(5):849–858, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bersuker:1977:SQS

[BBL77] I. B. Bersuker, S. S. Budnikov, and B. A. Leizerov. Semi-quantitative and semi-empirical versions in the quasirelativistic SCF–MO–LCAO methods: Numerical calculations for (PtCl_{6})^{2−}. International Journal of Quantum Chemistry, 11(4):543–559, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bendazzoli:1971:SCM

[BBP71] G. L. Bendazzoli, F. Bernardi, and P. Palmieri. Self consistent molecular orbital calculations on polyatomic molecules: Gaussian approximation for two-electron integrals. International Journal of Quantum Chemistry, 5(5):497–503, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bloembergen:1971:SNL

[BBT71] N. Bloembergen, W. K. Burns, and C. L. Tang. Symmetry of non-linear optical susceptibilities in absorbing media and conservation of angular momentum. International Journal of Quantum Chemistry, 5(S5):555–562, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Block:1975:SAC

[BC75] Arthur McB. Block and Richard G. Clements. Structure– activity correlations for phenoxyacetic acids and indoleacetic acids used for plant growth regulation. International Journal of Quantum Chemistry, 9(S2):197–202, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bass:1976:IIR

[BC76] George E. Bass and James E. Chenevey. Irradiation induced rate enhancements for the LDH-pyruvate reaction. International Journal of Quantum Chemistry, 10(S3):247–250, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bishop:1979:NOA

[BC79] David M. Bishop and Lap M. Cheung. Natural orbital analysis of nonadiabatic H_{2}^{+} wave functions. International Journal of Quantum Chemistry, 15(5):517–532, May 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Block:1977:ASA

[BCL77] Arthur McB. Block, Elvira Cuevas, and Ram S. Lamba. Auxin structure–activity relationship. Preliminary results of studies of chemical control of an ecosystem in its steady state. International Journal of Quantum Chemistry, 12(S4):127–133, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bolotina:1979:DSS

[BCL79] I. A. Bolotina, V. O. Chekhov, and V. Yu. Lugauskas. Determination of the secondary structures of proteins by circular dichroism spectra. Calculation of the protein basic circular dichroism spectra for antiparallel and parallel β-structures and β-bends. International Journal of Quantum Chemistry, 16(4): 819–824, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bolis:1978:PAF

[BCR^{+}78] G. Bolis, E. Clementi, M. Ragazzi, D. Salvaderi, and D. R. Ferro. Preliminary attempt to follow the enthalpy of an enzymatic reaction by ab initio computations: Catalytic action of papain. International Journal of Quantum Chemistry, 14(6): 815–838, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bendazzoli:1973:OEP

[BDP73] G. L. Bendazzoli, M. Dixon, and P. Palmieri. One-electron properties from approximate LCAO–SCF wave functions. International Journal of Quantum Chemistry, 7(2):223–232, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bouman:1977:GTR

[BDT77] Thomas D. Bouman, Charles D. Duncan, and Carl Trindle. Group theory and reaction mechanisms: Permutation theoretic prediction and computational support for pseudorotation modes in C_{2}H and C_{5}H rearrangements. International Journal of Quantum Chemistry, 11(3):399–413, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beebe:1972:NSS

[Bee72] Nelson H. F. Beebe. A note on space spanning. International Journal of Quantum Chemistry, 6(S6):439–441, January 17– 22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Beebe:1976:SFS

[Bee76] Nelson H. F. Beebe. On the symmetry of filled shells. International Journal of Quantum Chemistry, 10(6):1007–1010, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beebe:1979:MVO

[Bee79] Nelson H. F. Beebe. Modification of virtual orbitals. International Journal of Quantum Chemistry, 15(6):589–600, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandas:1978:EVT

[BEF78] Erkki Brändas, Nils Elander, and Piotr Froelich. Extended virial theorem applied to predissociation phenomena. Interpretation of the nearly degenerate G^{1}Π ∼ I^{1}Π energy spectra in SiO. International Journal of Quantum Chemistry, 14(4):443– 456, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Blakemore:1977:NIS

[BEH77a] M. Blakemore, G. A. Evans, and J. Hyslop. A numerical integration scheme for iterative calculations on atomic systems. International Journal of Quantum Chemistry, 11(2):325–339, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Blakemore:1977:VIS

[BEH77b] M. Blakemore, G. A. Evans, and J. Hyslop. Variation–iteration solution of the Hartree–Fock equations for atomic systems. International Journal of Quantum Chemistry, 12(2):305–315, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Benesch:1972:DRE

[Ben72] Robert Benesch. On the determination of radial electron– electron distribution functions. International Journal of Quantum Chemistry, 6(1):181–192, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berggren:1973:MNT

[Ber73] K. F. Berggren. Metal-to-nonmetal transitions in expanded states of sodium. International Journal of Quantum Chemistry, 7(S7):597–603, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Bernstein:1975:BRR

[Ber75] R. B. Bernstein. Branching ratios in reactive collisions: Theory versus experiment. International Journal of Quantum Chemistry, 9(S9):385–395, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Bernstein:1976:ITA

[Ber76] R. B. Bernstein. Information-theoretic analysis of rotational distributions from quantal and quasiclassical computations of reactive and nonreactive scattering. International Journal of Quantum Chemistry, 10(S10):267–276, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Blokker:1970:AFG

[BF70] Esko Blokker and Stig Flodmark. The arbitrary finite group and its irreducible representations. International Journal of Quantum Chemistry, 5(S4):463–472, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Blokker:1971:CPS

[BF71] Esko Blokker and Stig Flodmark. A computer program for symmetry adaptation of electronic states in crystals. International Journal of Quantum Chemistry, 5(S5):569–578, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Brandas:1975:RTE

[BF75] Erkki Brändas and Piotr Froelich. A remark on time evolution and spectral concentration. International Journal of Quantum Chemistry, 9(S9):457–459, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Brandas:1978:LCB

[BF78] Erkki Brändas and Piotr Froelich. Level crossings and branch points studied by the multidimensional partitioning technique. International Journal of Quantum Chemistry, 13(5):619–626, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brunn:1974:CAE

[BFH74] Wolfgang Brunn, Lothar Fritsche, and Klaus Hermann. Calculation of the adsorption energy of hydrogen on CR and Mo (001) surfaces by the MS X_{α} method. International Journal of Quantum Chemistry, 8(S8):483–490, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Brandas:1978:TRM

[BFH78] Erkki Brändas, Piotr Froelich, and Michael Hehenberger. Theory of resonances in many-body systems: Spectral theory of unbounded Schrödinger operators, complex scaling, and extended virial theorem. International Journal of Quantum Chemistry, 14(4):419–441, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berrondo:1978:SMP

[BFR78] M. Berrondo and A. Flores-Riveros. A simplified model for the predissociation of diatomic molecules. International Journal of Quantum Chemistry, 13(3):391–410, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bagus:1973:CMH

[BFS73] P. S. Bagus, A. J. Freeman, and F. Sasaki. Correlation and multiplet hole theory of core-electron energy splittings in transition-metal ions. International Journal of Quantum Chemistry, 7(S7):83–92, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Brandas:1970:CPS

[BG70] E. Brändas and O. Goscinski. Critical point, singularities, and extrapolations in the helium isoelectronic sequence. International Journal of Quantum Chemistry, 4(6):571–578, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandas:1972:DFP

[BG72] E. Brändas and O. Goscinski. Darboux functions and power series expansions with examples from isoelectronic sequences. International Journal of Quantum Chemistry, 6(S6):59–72, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Berrondo:1975:EFD

[BG75] M. Berrondo and O. Goscinski. Energy as a functional of the density matrix. International Journal of Quantum Chemistry, 9(S9):67–73, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Banerjee:1976:CBS

[BG76] A. Banerjee and F. Grein. Convergence behavior of some multiconfiguration methods. International Journal of Quantum Chemistry, 10(1):123–134, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Benedict:1978:TCS

[BG78] M. G. Benedict and I. Gyemant. Total cross section of the SF_{6} molecule for elastic electron scattering. International Journal of Quantum Chemistry, 13(5):597–603, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brailsford:1971:SPO

[BH71] D. F. Brailsford and G. G. Hall. Symmetry properties of oneand two-electron molecular integrals. International Journal of Quantum Chemistry, 5(6):657–668, November 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Breton:1978:WAS

[BHLM78] J. Breton, A. Hardisson, R. Lefebvre, and F. Mauricio. On the windows in the absorption spectrum of an ensemble of oscillators embedding an impurity oscillator. International Journal of Quantum Chemistry, 13(2):255–260, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandas:1975:DRS

[BHM75] Erkki Brändas, Michael Hehenberger, and Harold V. McIntosh. Dispersion relations and spectral densities. International Journal of Quantum Chemistry, 9(1):103–117, January 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Biedenharn:1975:RTF

[Bie75] L. C. Biedenharn. A remark on the Thomas–Fermi equation. International Journal of Quantum Chemistry, 9(S9):31– 33, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Billingsley:1972:ESP

[Bil72] Frank P. Billingsley II. An economic storage and processing method for two-electron integrals in LCAO–MO calculations. International Journal of Quantum Chemistry, 6(4):617–623, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Benard:1973:DOM

[BJ73] Marc Benard and André Julg. Détermination des orbitales moléculaires par minimisation directe de l’énergie de l’état fondamental et des états excités des molécules. (French) [Determination of molecular orbitals by direct minimization of the ground state and excited state energies of molecules]. International Journal of Quantum Chemistry, 7(5):945–950, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonacic:1971:CDH

[BK71] Vlasta Bonačić and Jaroslav Koutecký. Convergence difficulties in the Hartree–Fock procedure for the PPP model of alternant hydrocarbons. International Journal of Quantum Chemistry, 5(S5):137–142, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Boyd:1974:CHS

[BK74] R. J. Boyd and J. Katriel. The Coulomb hole in the 2^{3} S state of the helium isoelectronic sequence. International Journal of Quantum Chemistry, 8(2):255–261, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Blum:1975:SAE

[BK75a] K. Blum and H. Kleinpoppen. Spin analysis of electron atom collisions. International Journal of Quantum Chemistry, 9(S9): 415–424, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Born:1975:HBS

[BK75b] Sidney Born and Kazuo Kitahara. Hydrogen bonds in the sodium conductivity of nerve membranes. International Journal of Quantum Chemistry, 9(S2):237–257, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Boyer:1975:SCR

[BK75c] L. L. Boyer and B. M. Klein. A self-consistent relativistic APW method with the spin-orbit interaction treated as a perturbation. International Journal of Quantum Chemistry, 9(S9):511– 518, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Blum:1976:CIE

[BK76] K. Blum and H. Kleinpoppen. Coherent and incoherent excitation of atoms by electron impact. International Journal of Quantum Chemistry, 10(S10):231–238, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Beyer:1977:FSS

[BK77] H-J. Beyer and H. Kleinpoppen. Fine structure and Stark shifts from anticrossing studies. International Journal of Quantum Chemistry, 12(S11):271–287, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Banerjee:1979:PGF

[BKS79] Ajit Banerjee, John W. Kenney III, and Jack Simons. Polarization Green’s function with multiconfiguration self-consistent-field reference states. International Journal of Quantum Chemistry, 16(6):1209–1237, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Birner:1975:WGL

[BKW75] P. Birner, H.-J. Köhler, and C. Weiss. Wechselwirkungen gerichteter ladungsverteilungen. I. IRDO (intermediate retention of differential overlap)-berechnungen zum einfluß des einsamen elektronenpaars am stickstoff auf die relative acidität des β-picolins. (German) [interactions of directed charge distributions. I. IRDO (intermediate retention of differential overlap) calculations on the influence of the lone pair of electrons on nitrogen on the relative acidity of the β picoline]. International Journal of Quantum Chemistry, 9(5):917–922, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berggren:1970:NFE

[BL70] K.-F. Berggren and P. Lindner. A ‘nearly free electron’ model for crystalline MgH_{2}. International Journal of Quantum Chemistry, 5(S4):431–436, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Beebe:1972:POT

[BL72] Nelson H. F. Beebe and Sten Lunell. A projection operator technique for orbital basis variation. International Journal of Quantum Chemistry, 6:1149–1155, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bak:1975:HBF

[BL75] B. Bak and P. A. Lund. Hydrogen bonding in 1-fluoroacetic acid rotamers. International Journal of Quantum Chemistry, 9(S9):285–287, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Beebe:1977:SGT

[BL77a] Nelson H. F. Beebe and Jan Linderberg. Simplifications in the generation and transformation of two-electron integrals in molecular calculations. International Journal of Quantum Chemistry, 12(4):683–705, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beebe:1977:STE

[BL77b] Nelson H. F. Beebe and Jan Linderberg. Simplifications in the two-electron integral array in molecular calculations. International Journal of Quantum Chemistry, 12(4):683–705, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Braga:1977:SIE

[BL77c] Manuel Braga and Sven Larsson. Satellite intensities in ESCA spectra of Ti(IV) fluoro- and oxocomplexes calculated by the multiple scattering and INDO methods. International Journal of Quantum Chemistry, 12(S11):61–69, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Brown:1977:CEE

[BL77d] Richard E. Brown and S. Larsson. Correlation effects to the expectation values of atomic systems. International Journal of Quantum Chemistry, 12(1):83–92, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Biczo:1979:ISR

[BL79] G. Biczó and I. Lukovits. Are intermediate states responsible for certain specific properties of biological macromolecules? International Journal of Quantum Chemistry, 16(1):31–42, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Blokker:1972:TCI

[Blo72] Esko Blokker. A theory for the construction of the irreducible representations of finite groups. International Journal of Quantum Chemistry, 6(5):925–933, September 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Blokker:1973:TCI

[Blo73] Esko Blokker. A theory for the construction of the irreducible representations of finite groups. II. International Journal of Quantum Chemistry, 7(6):1091–1097, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boyle:1971:IIF

[BM71a] L. L. Boyle and P. S. C. Matthews. The isotropic invariants of fifth-rank Cartesian tensors. International Journal of Quantum Chemistry, 5(4):381–386, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boyle:1971:NAT

[BM71b] L. L. Boyle and P. S. C. Matthews. On the nature of axial tensors. International Journal of Quantum Chemistry, 5(S5):579– 590, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Bhattacharya:1973:CHF

[BM73] A. K. Bhattacharya and P. K. Mukherjee. Coupled Hartree– Fock calculation of the uniform field quadrupole polarizabilities and shielding factors of S-state ions. International Journal of Quantum Chemistry, 7(3):491–499, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bellum:1974:SSE

[BM74] John C. Bellum and David A. Micha. A study of single-electron and total energies for some pairs of noble gas atoms. International Journal of Quantum Chemistry, 8(S8):229–240, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Bowden:1970:TIS

[BMD70] C. M. Bowden, H. C. Meyer, and P. L. Donoho. Theory of the intrinsically strain broadened line shape for magnetic resonance absorption within the non-Kramers doublet of Al_{2}O_{3}:Fe^{2+}.

International Journal of Quantum Chemistry, 5(S4):395–406, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Bhattacharya:1975:CCH

[BMM75] A. K. Bhattacharya, R. K. Moitra, and A. Mukherji. Corrections to coupled Hartree–Fock theory: the rearrangement effect. International Journal of Quantum Chemistry, 9(3):385– 392, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beck:1974:EEC

[BN74] Donald R. Beck and Cleanthis A. Nicolaides. The effect of electron correlation on atomic properties. International Journal of Quantum Chemistry, 8(S8):17–28, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Beck:1976:TCE

[BN76] Donald R. Beck and Cleanthes A. Nicolaides. Theory and calculation of excited-state wave functions and properties. International Journal of Quantum Chemistry, 10(S10):119–134, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Beck:1972:CMC

[BNLS72] D. Beck, C. Nicolaides, W. Luken, and O. Sinanoğlu. CI methods for the calculation of hyperfine structures, lifetimes, and molecular potential energy surfaces. International Journal of Quantum Chemistry, 6(S6):1–7, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Boyle:1973:IIT

[BO73] L. L. Boyle and Z. Ożgo. Icosahedral irreducible tensors and their applications. International Journal of Quantum Chemistry, 7(2):383–404, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Born:1978:AIP

[BÖ78] Gregory Born and Yngve Öhrn. Alternative inner projections of the superoperator resolvent. International Journal of Quantum Chemistry, 14(S12):143–151, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Brandas:1979:APT

[BO79] E. J. Brändas and C. H. Obcemea. Analytic perturbation theory and localization phenomena. International Journal of Quantum Chemistry, 16(5):993–1007, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bendazzoli:1976:SOC

[BOP76] G. L. Bendazzoli, G. Orlandi, and P. Palmieri. Spin-orbit coupling in aromatic hydrocarbons: A semiempirical evaluation of the radiative lifetimes of naphtalene and anthracene. International Journal of Quantum Chemistry, 10(4):659–668, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boring:1970:SCS

[Bor70] A. Michael Boring. Self-consistent spin-polarized energy bands of sodium. International Journal of Quantum Chemistry, 4(3): 315–332, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Borelius:1971:IPS

[Bor71] Gudmund Borelius. Internal pressure in solids. International Journal of Quantum Chemistry, 5(S5):591, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Boring:1974:CEG

[Bor74] Michael Boring. Computed energy gap in krypton. International Journal of Quantum Chemistry, 8(S8):451–455, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Bowden:1971:DIM

[Bow71] Charles M. Bowden. Dependence of the integrated magnetic resonance absorption on resonance frequency and crystal field perturbations for non-Kramers doublets. International Journal of Quantum Chemistry, 5(S5):673–683, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Bowden:1972:CSR

[Bow72] Charles M. Bowden. Coherent state representation for fermions. International Journal of Quantum Chemistry, 6(S6): 133–143, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Bowman:1979:SAT

[Bow79] Joel M. Bowman. Sudden approximation theory of vibrational excitation. International Journal of Quantum Chemistry, 16 (S13):487–500, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Boyle:1970:DNZ

[Boy70a] L. L. Boyle. Determination of the non-zero components of property tensors in environments of low symmetry. International Journal of Quantum Chemistry, 4(4):389–395, July 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boyle:1970:FRM

[Boy70b] L. L. Boyle. Fifth-rank molecular polarization tensors. International Journal of Quantum Chemistry, 4(4):413–425, July 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boyle:1972:ESI

[Boy72a] L. L. Boyle. The electronic states of icosahedral molecules. International Journal of Quantum Chemistry, 6(5):919–924, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boyle:1972:GTA

[Boy72b] L. L. Boyle. The group theoretical approach to nuclear quadrupole resonance. International Journal of Quantum Chemistry, 6(2):313–325, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boyle:1972:SPG

[Boy72c] L. L. Boyle. The symmetrized powers of group representations. International Journal of Quantum Chemistry, 6(4):725– 746, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boyd:1974:SSP

[Boy74] Donald B. Boyd. Structural and spectral properties of the disulfide linkage in proteins and other molecules. International Journal of Quantum Chemistry, 8(S1):13–19, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Boyd:1977:MSS

[Boy77] Donald B. Boyd. MINDO/3 Study of N-substituted monocyclic β-lactam structures. International Journal of Quantum Chemistry, 12(S4):161–167, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bergmann:1970:QAH

[BP70] Ernst David Bergmann and Bernard Pullman, editors. Quantum aspects of heterocyclic compounds in chemistry and biochemistry: proceedings of an international symposium held in Jerusalem, 31 March–4 April 1969, volume 2 of The Jerusalem symposia on quantum chemistry and biochemistry. Israel Academy of Sciences and Humanities; [distributors: Academic Press, New York], Jerusalem, Israel, 1970. ISBN 0-12-567050-8. LCCN QD462 .Q34.

Bergmann:1971:APA

[BP71] Ernst David Bergmann and Bernard Pullman, editors. “Aromaticity, pseudo-aromaticity, anti-aromaticity”: proceedings of an international symposium held in Jerusalem, 31 March–3 April, 1970, volume 3 of The Jerusalem symposia on quantum chemistry and biochemistry. Israel Academy of Sciences and Humanities; [distributors: Academic Press, New York], Jerusalem, 1971. ISBN 0-12-091040-3. LCCN QD331 .A75.

Bergmann:1972:PTE

[BP72] Ernst David Bergmann and Bernard Pullman, editors. The Purines: theory and experiment; proceedings of an international symposium held in Jerusalem, 4–8 April 1971, volume 4 of The Jerusalem symposia on quantum chemistry and biochemistry. Israel Academy of Sciences and Humanities; [distributed by Academic Press, New York], Jerusalem, 1972. LCCN QD401 .P979.

Bergmann:1973:CBM

[BP73] Ernst David Bergmann and Bernard Pullman, editors. Conformation of biological molecules and polymers; proceedings of an international symposium held in Jerusalem, 3–9 April 1972, volume 5 of The Jerusalem symposia on quantum chemistry and biochemistry. Israel Academy of Sciences and Humanities; [distributed by Academic Press, New York], Jerusalem, 1973. ISBN 0-12-091065-9. LCCN QP501 .C66.

Bendazzoli:1974:SOI

[BP74] G. L. Bendazzoli and P. Palmieri. Spin-orbit interaction in polyatomic molecules: Ab initio computations with Gaussian orbitals. International Journal of Quantum Chemistry, 8(6): 941–950, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bendazzoli:1975:ICS

[BP75a] G. L. Bendazzoli and P. Palmieri. Ab initio computations of spin-orbit interactions in polyatomic molecules. Splitting of sulfur L_{II}, III states and singlet–triplet transitions in SO_{2}. International Journal of Quantum Chemistry, 9(3):537–544, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Binkley:1975:MPT

[BP75b] J. S. Binkley and J. A. Pople. Møller–Plesset theory for atomic ground state energies. International Journal of Quantum Chemistry, 9(2):229–236, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bartlett:1978:MBP

[BP78] Rodney J. Bartlett and George D. Purvis. Many-body perturbation theory, coupled-pair many-electron theory, and the importance of quadruple excitations for the correlation problem. International Journal of Quantum Chemistry, 14(5):561–581, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Buchler:1979:VSL

[BP79] J. R. Buchler and J. Perdang. De vibrantibus systematibus lente autonomiterque evolventibus. (Latin) []. International Journal of Quantum Chemistry, 16(S13):183–187, March 11– 17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Bouscasse:1973:TOE

[BPTLVM73] L. Bouscasse, R. Phan-Tan-Luu, E. J. Vincent, and Et. J. Metzger. Technique d’optimisation de l’énergie SCF en fonction des exposants de Slater. (French) [Optimization technique for the SCF energy as a function of Slater exponents]. International Journal of Quantum Chemistry, 7(2):357–364, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Braun:1972:VBI

[BR72] P. A. Braun and T. K. Rebane. Variational bounds for imaginary frequency polarizability and dispersion interaction constant. International Journal of Quantum Chemistry, 6(4):639– 646, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barnsley:1975:VBT

[BR75a] Michael F. Barnsley and Peter D. Robinson. Variational bounds on transition probabilities. International Journal of Quantum Chemistry, 9(3):479–487, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berrondo:1975:PSC

[BR75b] M. Berrondo and D. H. Rojas. Potential surface crossings for diatomic molecules. International Journal of Quantum Chemistry, 9(1):119–131, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Baerends:1978:ELH

[BR78] E. J. Baerends and P. Ros. Evaluation of the LCAO Hartree– Fock–Slater method: Applications to transition-metal complexes. International Journal of Quantum Chemistry, 14(S12): 169–190, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Brandas:1970:RPT

[Brä70] Erkki Brändas. A remark on perturbation theory for intermolecular forces. International Journal of Quantum Chemistry, 5(S4):285–287, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Brandow:1975:FTE

[Bra75] B. H. Brandow. Formal theory of effective π-electron Hamiltonians. International Journal of Quantum Chemistry, ??(??):??, ?? 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandow:1976:TMI

[Bra76] B. H. Brandow. Theory of Mott insulators. International Journal of Quantum Chemistry, 10(S10):417–434, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Brandow:1979:FTE

[Bra79a] B. H. Brandow. Formal theory of effective π-electron Hamiltonians. International Journal of Quantum Chemistry, 15(2):

207–242, February 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandow:1979:GSE

[Bra79b] B. H. Brandow. Ground state and elementary excitations of a model valence-fluctuation system. International Journal of Quantum Chemistry, 16(S13):423–444, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Brener:1973:WVF

[Bre73] N. E. Brener. Wave-vector- and frequency-dependent dielectric functions in the Hartree–Fock–Slater approximation. International Journal of Quantum Chemistry, 7(S7):559–567, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Brener:1975:RDF

[Bre75] N. E. Brener. RPA dielectric function for real solids. International Journal of Quantum Chemistry, 9(S9):555–561, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Brickmann:1973:VCE

[Bri73] J. Brickmann. Variational calculations of eigenstates for intramolecular nuclear motions using nontensor product bases. International Journal of Quantum Chemistry, 7(S7):427–436, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Barriel:1975:EHT

[BRS^{+}75] J. M. Barriel, J. M. Riera, F. Sanz, R. Caballol, M. Vidal, and R. Carbó. Extended Hückel theory of hydrogen-molecule interactions. International Journal of Quantum Chemistry, 9 (6):1021–1031, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bryan:1974:WR

[Bry74] Robert A. Bryan. Welcoming remarks. International Journal of Quantum Chemistry, 8(S8):3, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Benesch:1970:NOM

[BS70] Robert Benesch and Vedene H. Smith, Jr. Natural orbitals in momentum space and correlated radial momentum distributions: I. The ^{1}S ground state of Li^{+}. International Journal of Quantum Chemistry, 5(S4):131–138, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Benesch:1971:EAX

[BS71] Robert Benesch and Vedene H. Smith, Jr. Exact and asymptotic X-ray intensity values from 20-parameter Hylleraas-type wave functions. He and He-like ions. International Journal of Quantum Chemistry, 5(S5):35–45, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Bartlett:1974:CEL

[BS74a] Rodney J. Bartlett and David M. Silver. Correlation energy in LiH, BH, and HF with many-body perturbation theory using Slater-type atomic orbitals. International Journal of Quantum Chemistry, 8(S8):271–276, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Boyle:1974:AEI

[BS74b] L. L. Boyle and C. E. Schäuffer. Alternative equivalent icosahedral irreducible tensors. International Journal of Quantum Chemistry, 8(2):153–161, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bartlett:1975:SAD

[BS75a] Rodney J. Bartlett and David M. Silver. Some aspects of diagrammatic perturbation theory. International Journal of Quantum Chemistry, 9(S9):183–198, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Bishop:1975:NIT

[BS75b] David M. Bishop and Barbara E. Schneider. A new integral transform basis function. International Journal of Quantum Chemistry, 9(1):67–74, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Born:1975:ICS

[BS75c] Gregory Born and John R. Sabin. Ab Initio calculation of some electronically excited states of a hydrogen-bonded system: A preliminary report. International Journal of Quantum Chemistry, 9(S2):259–262, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bartlett:1976:CEL

[BS76a] Rodney J. Bartlett and David M. Silver. Correlation energy in LiH, BH, and HF with many-body perturbation theory using Slater-type atomic orbitals. International Journal of Quantum Chemistry, 10(1):185–186, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Besson:1976:ASC

[BS76b] M. A. Besson and M. Suard. Analysis of the space correlation factors in a MCSCF representation of LiH and Li_{2} molecules. International Journal of Quantum Chemistry, 10(1):151–162, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bak:1977:PNC

[BS77a] Børge Bak and Henrik Svanholt. Production of new compounds by vibrational dissociation. International Journal of Quantum Chemistry, 12(S11):259–262, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Bartlett:1977:DSC

[BS77b] Rodney J. Bartlett and Isaiah Shavitt. Determination of the size-consistency error in the single and double excitation configuration interaction model. International Journal of Quantum Chemistry, 12(S11):165–173, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Bartlett:1978:DSC

[BS78a] Rodney J. Bartlett and Isaiah Shavitt. Determination of the size-consistency error in the single and double excitation configuration interaction model. International Journal of Quantum Chemistry, 14(S12):543–544, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Blomberg:1978:BDC

[BS78b] Margareta R. A. Blomberg and Per E. M. Siegbahn. Beryllium dimer, a critical test case of MBPT and CI methods. International Journal of Quantum Chemistry, 14(5):583–592, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brooks:1978:BSI

[BS78c] Bernard R. Brooks and Henry F. Schaefer III. The BERKE-LEY system. III. General configuration-interaction methods for open-shell molecular electronic states. International Journal of Quantum Chemistry, 14(5):603–612, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berger:1979:SEC

[BS79] Gregory D. Berger and Paul G. Seybold. Substituent effects in chemical carcinogenesis: Chrysene and its methyl derivatives. International Journal of Quantum Chemistry, 16(S6):305–311, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Brown:1977:MOS

[BSB77] Richard E. Brown, Alfredo M. Simas, and Roy E. Bruns. A molecular orbital study of the chloramphenicol family of drugs: A preliminary report. International Journal of Quantum Chemistry, 12(S4):357–362, January 9–12, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Biondi:1977:FEE

[BSG77] Franco Biondi, Oriano Salvetti, and A. Gołebiewski. Formulas for evaluation of expectation values of monoelectric operators with modified Gaussian functions. International Journal of Quantum Chemistry, 12(1):29–34, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bhattacharya:1972:CHF

[BSM72] A. K. Bhattacharya, S. Sengupta, and A. Mukherji. Coupled Hartree–Fock calculation of the electric dipole hyperpolarizabilities: He sequence. International Journal of Quantum Chemistry, 6(2):337–345, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Banerjee:1978:OPG

[BSS78] Ajit Banerjee, Ron Shepard, and Jack Simons. One-particle Green’s function with multiconfiguration reference states. International Journal of Quantum Chemistry, 14(S12):389–404, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Beran:1978:AGO

[BSZ78] S. Beran, Z. Slanina, and D. C. Zidarov. Automatic geometry optimization for molecules with d orbitals. I. General analytic formulas for derivatives of Slater orbitals. International Journal of Quantum Chemistry, 13(2):227–233, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Billingsley:1971:OAO

[BT71] Frank P. Billingsley II and Carl Trindle. Optimization of atomic orbital charges by least squares representation of the molecular charge distribution. International Journal of Quantum Chemistry, 5(S5):317–327, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Ball:1978:ESS

[BT78a] John R. Ball and Colin Thomson. Electronic structure of SiH_{2}, PH_{2}, and their positive and negative ions. International Journal of Quantum Chemistry, 14(1):39–53, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonchev:1978:TCM

[BT78b] D. Bonchev and N. Trinajstić. On topological characterization of molecular branching. International Journal of Quantum Chemistry, 14(S12):293–303, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Basu:1979:FEM

[BTD79] Biswajit Basu, R. D. Tiwari, and Arun K. Dey. Free-electron molecular-orbital treatment of chalcone group of dyes. International Journal of Quantum Chemistry, 16(6):1337–1346, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bulloff:1971:QRS

[Bul71] Jack J. Bulloff. Quantum reunification of the sciences. International Journal of Quantum Chemistry, 5(S5):777–786, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Bunge:1977:RFO

[Bun77] Carlos F. Bunge. Reduced first order density matrix for the Be ground state. International Journal of Quantum Chemistry, 12 (2):343–353, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Burden:1972:EHF

[Bur72] F. R. Burden. An extended Hartree–Fock method using corresponding orbitals: Application to lithium hydride. International Journal of Quantum Chemistry, 6(4):647–650, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Burrows:1973:BHF

[Bur73] B. L. Burrows. Bounds for Hartree–Fock perturbation theory. International Journal of Quantum Chemistry, 7(2):345– 355, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Burton:1977:BFC

[Bur77] Peter G. Burton. Bond functions in CEPA-PNO computations of electronic structure. Basis set optimization studies in N_{2} H_{2}. Application to ozone and the stability of the bound cyclic conformer of O_{3}. International Journal of Quantum Chemistry, 12 (S11):207–213, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Butler:1976:CSAb

[But76] P. H. Butler. Calculation of j and jm symbols for arbitrary compact groups. II. An alternate procedure for angular momentum. International Journal of Quantum Chemistry, 10(4): 599–613, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beretta:1973:DPT

[BV73] E. Beretta and F. Vetrano. A double perturbation theory for molecular complexes. International Journal of Quantum Chemistry, 7(2):333–344, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berlin:1972:NPM

[BVO72] A. A. Berlin, G. A. Vinogradov, and A. A. Ovchinnikov. On the nature of paramagnetism in macromolecules with conjugated C − C bonds. International Journal of Quantum Chemistry, 6(2):263–269, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bouma:1978:IMO

[BVR78] Willem J. Bouma, Mark A. Vincent, and Leo Radom. Ab initio molecular orbital studies of sigmatropic rearrangements. International Journal of Quantum Chemistry, 14(6):767–777, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bertetta:1976:SRD

[BVS76] E. Bertetta, F. Vetrano, and F. Solimano. Some remarks on a double perturbation theory for molecular complexes. International Journal of Quantum Chemistry, 10(1):175–179, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonham:1973:MQR

[BW73] R. A. Bonham and H. F. Wellenstein. Measurement of quantities related to the charge and momentum densities of atoms and molecules by use of keV electron scattering. International Journal of Quantum Chemistry, 7(S7):377–394, January 21– 27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Bradford:1976:EVS

[BW76a] E. G. Bradford and Paul Westhaus. Effective valence-shell Hamiltonians generated by canonical transformations. International Journal of Quantum Chemistry, 10(S3):231–236, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Butler:1976:CSAa

[BW76b] P. H. Butler and B. G. Wybourne. Calculation of j and jm symbols for arbitrary compact groups. I. Methodology. International Journal of Quantum Chemistry, 10(4):581–598, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Butler:1976:CSAc

[BW76c] P. H. Butler and B. G. Wybourne. Calculation of j and jm symbols for arbitrary compact groups. III. Application to SO_{3} ⊃ T ⊃ C_{3} ⊃ C_{1}. International Journal of Quantum Chemistry, 10(4):615–628, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Butt:1979:CEC

[BWRD79] Glenn Butt, Roderich Walter, V. Renugopalakrishnan, and M. E. Druyan. Conformational energy calculations on the CNS-active peptide Pro-Leu-Gly-NH_{2}. International Journal of Quantum Chemistry, 16(S6):453–458, March 4–10, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bartlett:1977:TOM

[BWS77] Rodney J. Bartlett, Stephen Wilson, and David M. Silver. Third-order many-body perturbation theory for the ground state of the carbon monoxide molecule. International Journal of Quantum Chemistry, 12(4):737–757, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Buchler:1976:CSD

[BY76] Jean-Robert Buchler and William R. Yueh. Compton scattering in a degenerate electron gas. International Journal of Quantum Chemistry, 10(S10):277–282, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bazhin:1979:ESP

[BYB79] N. M. Bazhin, B. I. Yakobson, and A. I. Burshtein. Estimation of some parameters of the exchange mechanism of luminescence decay. International Journal of Quantum Chemistry, 16(4): 931–935, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bogach:1979:CCE

[BZSD79] P. G. Bogach, V. L. Zyma, A. M. Shakhovsky, and V. M. Danilova. Conformational changes and excitation energy transfer in myosin–auramine O system. International Journal of Quantum Chemistry, 16(1):43–50, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cusachs:1972:IAO

[CA72] Louis Chopin Cusachs and Haven Scott Aldrich. Invariant atomic orbitals by radial moment analysis of accurate molecular orbitals. International Journal of Quantum Chemistry, 6 (S6):221–229, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Cusachs:1973:SAO

[CA73] Louis Chopin Cusachs and Haven Scott Aldrich. The shape of atomic orbitals. International Journal of Quantum Chemistry, 7(S7):231–237, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Carbo:1978:MBA

[CA78] Ramon Carbo and Caterina Arnau. Mathematical basis of approximate MO theories: Origin of Mulliken’s magic formula. International Journal of Quantum Chemistry, 14(2):209–212, August 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cusachs:1975:LRT

[CAA^{+}75] L. C. Cusachs, H. S. Aldrich, J. G. Aiken, H. B. Jonassen, and K. S. Wheelock. Limitations of Roby’s Theorem. International Journal of Quantum Chemistry, 9(S9):265–267, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Callan:1972:AGS

[Cal72] Edwin J. Callan. Analytic ground state energies of two-electron atoms. International Journal of Quantum Chemistry, 6(S6): 431–434, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Calais:1973:BRB

[Cal73a] Jean-Louis Calais. Book review: Anwendung der Gruppentheorie in der Quantenmechanik. Authors: M. J. Petraschen and E. D. Trifonow. Published by: Akademie-Verlag, Berlin, 1969. No. of pages: 240. International Journal of Quantum Chemistry, 7 (6):1221, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [PTUK69].

Caldwell:1973:BRB

[Cal73b] Dennis Caldwell. Book review: The Jerusalem Symposia on Quantum Chemistry and Biochemistry. Edited by E. D. Bergmann and B. Pullman. Published by: The Jerusalem Academy of Sciences and Humanities. Jerusalem 1970, 1971, and 1972. International Journal of Quantum Chemistry, 7(6): 1221–1225, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [BP70, BP71, BP72, BP73].

Callan:1973:ETE

[Cal73c] Edwin J. Callan. On the energies of two-electron atoms and ions. International Journal of Quantum Chemistry, 7(S7):111– 115, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Callan:1974:GSE

[Cal74] Edwin J. Callan. On the ground state energies of first-row atoms and ions. International Journal of Quantum Chemistry, 8(S8):29–36, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Calais:1975:CPI

[Cal75a] Jean-Louis Calais. Cohesive properties of ionic crystals. International Journal of Quantum Chemistry, 9(S9):497–509, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Calais:1975:SHV

[Cal75b] Jean-Louis Calais. Studies in honour of Väinö Toivo Hovi on the occasion of his sixtieth birthday. Annales Universitatis Turkuensis, Series A No. 162, 1973. International Journal of Quantum Chemistry, 9(5):941, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Caldwell:1975:BRB

[Cal75c] Dennis Caldwell. Book review: Advances in radiation chemistry. Vol. 4. Edited by: M. Burton and J. L. Magee. Published by: Wiley-Interscience, John Wiley & Sons, Inc. New York, 1974. International Journal of Quantum Chemistry, 9(4):755– 757, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Calais:1976:BRB

[Cal76a] Jean-Louis Calais. Book review: Methods in molecular orbital theory. Author: Almon G. Turner. Published by: Prentice-Hall, Englewood Cliffs, New Jersey, 1974. Price: £9.45. No. of pages: xiii + 225. International Journal of Quantum Chemistry, 10 (3):552–553, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Tur74].

Caldwell:1976:BRB

[Cal76b] Dennis Caldwell. Book review: From theoretical physics to biology. Proceedings of the Third International Conference ‘From Theoretical Physics to Biology’, Versailles, June 1971. Edited by: M. Marois, Paris. Published by: S. Karger AG, Basel, 1973. Price: $43.00. International Journal of Quantum Chemistry, 10(3):551–552, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Mar73b].

Calais:1977:PTT

[Cal77a] Jean-Louis Calais. Is the Peierls transition a transition? International Journal of Quantum Chemistry, 12(S11):547–553, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Caldwell:1977:BRB

[Cal77b] Dennis Caldwell. Book review: Theory of orientation and stereoselection. Author: Kenichi Fukui. Published by: Springer-Verlag, Berlin-Heidelberg-New York, 1975. Price: $18.10. No. of pages: 134. International Journal of Quantum Chemistry, 11(1):185, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Fuk70, Fuk75].

Calais:1979:BRBb

[Cal79a] Jean-Louis Calais. Book review: Crystals as giant molecules. Lecture notes in chemistry, no. 9. by A. Julg. Springer-Verlag, Berlin, 1978. 142 pp. International Journal of Quantum Chemistry, 16(5):1157, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Jul78].

Calais:1979:BRBa

[Cal79b] Jean-Louis Calais. Book review: Statistical mechanics. J. E. Mayer and M. G. Mayer. 2nd ed. John Wiley & Sons, Inc., New York, 1977. Price: £17.50/$29.00. International Journal of Quantum Chemistry, 15(2):259, February 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Calais:1979:DIS

[Cal79c] Jean-Louis Calais. Direct and indirect solutions of gap equations. International Journal of Quantum Chemistry, 16(S13): 387–393, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Corrington:1971:DMO

[CAMC71] Joyce H. Corrington, H. S. Aldrich, C. W. Mccurdy, and L. C. Cusachs. Dipole moments and orbital energies from ARCANA:

A semi-empirical molecular orbital calculation program. International Journal of Quantum Chemistry, 5(S5):307–316, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Carbo:1972:GRS

[Car72] Ramón Carbó. Generalized Rayleigh–Schrödinger perturbation theory. International Journal of Quantum Chemistry, 6(4):609– 611, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Carboa:1974:EMS

[Car74] Ramóan Carbóa. Eigenspace manipulation in SCF theory: General formalism. International Journal of Quantum Chemistry, 8(3):423–426, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cartling:1975:ULD

[Car75] Bo G. Cartling. On the use of a localized description of the electronic structure of solids. International Journal of Quantum Chemistry, 9(S9):117–121, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Castro:1978:IHS

[Cas78a] Eduardo A. Castro. Iterative hypervirial-scaling method for obtaining wave functions and eigenvalues. International Journal of Quantum Chemistry, 14(3):231–237, September 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Castro:1978:OSM

[Cas78b] Eduardo A. Castro. Optimum scaling in a Monte Carlo method for quantum chemistry. International Journal of Quantum Chemistry, 13(3):455–456, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Castro:1979:CTM

[Cas79] Eduardo A. Castro. Calculation of total molecular energies from an approximate relation between SCF orbital energies and total SCF energies. International Journal of Quantum Chemistry, 15(3):355–358, March 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Certain:1972:BPS

[CB72] P. R. Certain and W. Byers Brown. Branch point singularities in the energy of the delta-function model of one-electron diatoms. International Journal of Quantum Chemistry, 6(1): 131–142, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cabrol:1979:QMS

[CBV79] D. Cabrol, H. Broch, and D. Vasilescu. Quantum mechanical study of (Gly-Pro-Pro) repetitive unit in relation to collagen structure. International Journal of Quantum Chemistry, 16(S6):365–375, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Clementi:1979:IWD

[CC79] E. Clementi and G. Corongiu. Interaction of water with DNA single and double helix in the B conformation. International Journal of Quantum Chemistry, 16(4):897–915, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coulson:1971:TFW

[CD71] C. A. Coulson and B. M. Deb. On the theoretical foundation of Walsh’s rules of molecular geometry in terms of the Hellmann–Feynman theorem. International Journal of Quantum Chemistry, 5(4):411–434, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chaney:1974:AMT

[CD74] Roy C. Chaney and Franklin Dorman. Application of the method of tight binding for determining the excited bands of sodium and potassium. International Journal of Quantum Chemistry, 8(S8):465–472, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Colbow:1976:ETP

[CD76] Konrad Colbow and R. P. Dunyluk. Energy transfer in photosynthesis. International Journal of Quantum Chemistry, 10 (S3):151–159, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Chao:1978:MMI

[CD78] K. A. Chao and A. Ferreira Da Silva. Molecular model of impurity bands in semiconductors. II. the validity of the one-band AMO–MT model. International Journal of Quantum Chemistry, 14(S12):461–468, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Collins:1976:SIM

[CDEO76] Michael P. S. Collins, Brian J. Duke, James E. Eilers, and Brian O’Leary. The simulated ab initio molecular orbital technique. VI. Open-shell radicals in the spin restricted formalism. International Journal of Quantum Chemistry, 10(4):629–642, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cook:1973:CMC

[CE73] Earl C. Cook, Jr. and Darrell Ebbing. Core model calculation of the electric field gradient: Projection operator formalism with application to NH_{3}. International Journal of Quantum Chemistry, 7(4):707–715, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Collins:1972:SED

[CEK72] T. C. Collins, R. N. Euwema, and A. B. Kunz. A study of energy-dependent operators which simulate exchangecorrelation operators. International Journal of Quantum Chemistry, 6(S6):459–461, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Collins:1970:VED

[CESW70] T. C. Collins, R. N. Euwema, D. J. Stukel, and G. G. Wepfer. Valence electron density of states of ZnSe obtained from an energy dependent exchange approximation. International Journal of Quantum Chemistry, 5(S4):77–85, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Carbo:1977:ECE

[CFHS77] R. Carbó, J. M. Fornos, J. A. Hernández, and F. Sanz. Electrostatic corrections to extended Hückel theory. International Journal of Quantum Chemistry, 11(2):271–276, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Clinton:1973:MOE

[CFMO73] W. L. Clinton, C. A. Frishberg, L. J. Massa, and P. A. Oldfield. Methods for obtaining an electron-density matrix from X-ray diffraction data. International Journal of Quantum Chemistry, 7(S7):505–514, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Calais:1977:P

[CG77] Jean-Louis Calais and Osvaldo Goscinski. Preface. International Journal of Quantum Chemistry, 11(6):895, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Canuto:1978:SRP

[CG78] Sylvio Canuto and Osvaldo Goscinski. Stationarity of resonant pole trajectories in complex scaling. International Journal of Quantum Chemistry, 14(4):383–391, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Canuto:1979:CCP

[CG79a] Sylvio Canuto and Osvaldo Goscinski. Continuum contribution to polarizabilities and scaling. International Journal of Quantum Chemistry, 16(5):985–991, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Costas:1979:MPS

[CG79b] M. Costas and A. Garritz. The MS potential from a set of overlapping densities. International Journal of Quantum Chemistry, 16(S13):141–162, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Calais:1976:QSM

[CGLÖ76] Jean-Louis Calais, Osvaldo Goscinski, Jan Linderberg, and Yngve Öhrn, editors. Quantum science: methods and structure: a tribute to Per-Olov Löwdin. Plenum Press, New York, NY, USA; London, UK, 1976. ISBN 0-306-30968-8. ix + 595 pp. LCCN QC174.125.

Caballol:1974:GOS

[CGRC74] Rosa Caballol, Ramóan Gallifa, Josep M. Riera, and Ramóan Carbóa. Generalized open shell SCF theory. International Journal of Quantum Chemistry, 8(3):373–394, May 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1974:FPP

[CH74] P. Csavinszky and R. Hucek. First-principles pseudo potential calculation for obtaining the (ns) energy levels of the K atom. International Journal of Quantum Chemistry, 8(S8):37– 47, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Csavinszky:1975:FPP

[CH75] P. Csavinszky and R. Hucek. A first-principle pseudopotential calculation of the (ns) energy levels of the K atom. II. International Journal of Quantum Chemistry, 9(2):215–227, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Collard:1977:OTE

[CH77] Kathleen Collard and George G. Hall. Orthogonal trajectories of the electron density. International Journal of Quantum Chemistry, 12(4):623–637, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chang:1971:MSR

[Cha71] Tai Yup Chang. On the Mavroyannis–Stephen relativistic long-range interaction energy term between optically active molecules. International Journal of Quantum Chemistry, 5(4): 469–473, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chandler:1977:HST

[Cha77] G. S. Chandler. Hulthén and Slater type 2d functions in some excited configurations of sulphur and phosphorus. International Journal of Quantum Chemistry, 11(4):685–694, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cantu:1971:OPT

[CHD71] A. A. Cantu and A. Hart-Davis. The optimum projection technique in many-electron treatments. International Journal of Quantum Chemistry, 5(3):297–309, May 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chen:1971:ETQ

[Che71] Tien Chi Chen. Evaluation techniques for quantum mechanical integrals. International Journal of Quantum Chemistry, 5(S5): 775, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Cerjan:1978:CCS

[CHH^{+}78] C. Cerjan, R. Hedges, C. Holt, W. P. Reinhardt, K. Scheibner, and J. J. Wendoloski. Complex coordinates and the Stark effect. International Journal of Quantum Chemistry, 14(4):393– 418, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cohen:1974:MFT

[CHHS74] Morrel H. Cohen, John A. Hertz, Paul M. Horn, and Vinod K. S. Shante. The metal-ferroelectric transition in TTF– TCNQ. International Journal of Quantum Chemistry, 8(S8): 491–498, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Chipman:1977:MCP

[Chi77] Daniel M. Chipman. Methods for the calculation of photoionization cross sections using the extended Koopmans’ theorem. International Journal of Quantum Chemistry, 12(S11): 365–377, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Chojnacki:1979:ICT

[Cho79] Henryk Chojnacki. Intermolecular charge transfer in model hydrogen-bonded systems — biological implications. International Journal of Quantum Chemistry, 16(2):299–309, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Christoffersen:1973:ICL

[Chr73] Ralph E. Christoffersen. Ab initio calculations on large molecules using molecular fragments characterization of anthracene and phenanthrene. International Journal of Quantum Chemistry, 7(S7):169–179, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Christov:1977:CTT

[Chr77] S. G. Christov. Collision theory treatment of monomolecular and bimolecular gas reactions. International Journal of Quantum Chemistry, 12(3):495–503, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Christoffersen:1979:ESP

[Chr79a] Ralph E. Christoffersen. Electronic structural properties and photosynthesis. International Journal of Quantum Chemistry, 16(3):573–604, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Christoffersen:1979:I

[Chr79b] Ralph E. Christoffersen. Introduction. International Journal of Quantum Chemistry, 16(1):3–4, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Christov:1979:TCR

[Chr79c] S. G. Christov. Theory of chemical reactions in condensed media and its applications to biological processes. International Journal of Quantum Chemistry, 16(2):353–366, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Careless:1977:RSN

[CHS77] P. N. Careless, D. Hyatt, and L. Stanton. Removal of singularities in the numerical integrals for an STO basis set. International Journal of Quantum Chemistry, 12(3):569–578, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Crotov:1979:TFV

[CII79] S. S. Crotov, A. A. Ischenko, and L. S. Ivashkevich. Thermodynamic functions for vibronic systems. International Journal of Quantum Chemistry, 16(5):973–983, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chalasinski:1973:ECE

[CJ73a] G. Chałasiński and B. Jeziorski. Exact calculation of exchange polarization energy for H ion. International Journal of Quantum Chemistry, 7(1):63–73, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chalasinski:1973:MSE

[CJ73b] G. Chałasiński and B. Jeziorski. Multipole structure of exchange polarization energy for H_{2}^{+} Ion. International Journal of Quantum Chemistry, 7(4):745–757, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Claesson:1973:MEM

[CJCM73] A. Claesson, W. Jones, G. G. Chell, and N. H. March. Manyelectron and many-body force effects in the lattice dynamics of metals and semiconductors. International Journal of Quantum Chemistry, 7(S7):629–652, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Chalasinski:1977:CPR

[CJS77] Grzegorz Chałasiński, Bogumił Jeziorski, and Krzysztof Szalewicz. On the convergence properties of the Rayleigh– Schrödinger and the Hirschfelder–Silbey perturbation expansions for molecular interaction energies. International Journal of Quantum Chemistry, 11(2):247–257, February 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Carbonell:1973:TPT

[CK73] R. G. Carbonell and M. D. Kostin. Tunneling phenomena in three-dimensional double-well potentials. International Journal of Quantum Chemistry, 7(2):319–332, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Collins:1974:EHE

[CK74] T. C. Collins and A. B. Kunz. Excitation Hamiltonian of electronic systems using ÔÂÔ. International Journal of Quantum Chemistry, 8(S8):437–445, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Carney:1975:VRA

[CK75] G. D. Carney and C. W. Kern. Vibration–rotation analysis of some nonlinear molecules by a variational method. International Journal of Quantum Chemistry, 9(S9):317–323, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Collins:1973:ESA

[CKD73] T. C. Collins, A. B. Kunz, and J. T. Devreese. An excited state of alkali halides. International Journal of Quantum Chemistry, 7(S7):551–557, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Collins:1975:LEC

[CKI75] T. C. Collins, A. B. Kunz, and J. L. Ivey. Local excitations in crystalline systems. International Journal of Quantum Chemistry, 9(S9):519–527, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Chow:1973:UAS

[CL73] Yutze Chow and C. J. Liu. A unified analysis of some different approaches to spin-wave theory. International Journal of Quantum Chemistry, 7(S7):521–525, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Chen:1976:EAI

[CL76] Joseph C. Y. Chen and Yim Tin Lee. Eikonal approximation for intermediate-energy atomic collisions. International Journal of Quantum Chemistry, 10(S10):251–258, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Chojnacki:1977:DPB

[CL77] Henryk Chojnacki and Zbigniew Laskowski. Dynamic potential barrier and tunneling. International Journal of Quantum Chemistry, 11(6):1017–1020, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coldwell:1978:MCC

[CL78] R. L. Coldwell and R. E. Lowther. Monte Carlo calculation of the Born–Oppenheimer potential between two helium atoms using Hylleraas-type electronic wave functions. International Journal of Quantum Chemistry, 14(S12):329–341, March 12– 18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Claxton:1970:IUC

[Cla70] T. A. Claxton. Ab initio UHF calculations on the vinyl radical. International Journal of Quantum Chemistry, 4(3):337– 339, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Claverie:1971:TIF

[Cla71] Pierre Claverie. Theory of intermolecular forces. I. On the inadequacy of the usual Rayleigh–Schrödinger perturbation method for the treatment of intermolecular forces. International Journal of Quantum Chemistry, 5(3):273–296, May 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Clinton:1973:DMF

[Cli73] William L. Clinton. A density-matrix formulation of the Hartree–Fock–Bogoliubov equations for space-dependent superconductivity. International Journal of Quantum Chemistry, 7(S7):479–489, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Czerminski:1979:TOO

[CLP79a] R. Czermiński, B. Lesyng, and A. Pohorille. Tautomerism of oxopyridines and oxopyrimidines: Theoretical study with complete optimization of geometry. International Journal of Quantum Chemistry, 16(5):1141–1148, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Czerminski:1979:TPB

[CLP79b] R. Czermiński, B. Lesyng, and A. Pohorille. Tautomerism of pyrimidine bases — uracil, cytosine, isocytosine: Theoretical study with complete optimization of geometry. International Journal of Quantum Chemistry, 16(3):605–613, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Callaway:1972:CPE

[CM72a] Joseph Callaway and John J. Matese. Calculational procedure for electron-atom scattering at arbitrary energies. International Journal of Quantum Chemistry, 6(S6):79–85, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Clinton:1972:CCC

[CM72b] W. L. Clinton and L. J. Massa. The cusp condition: Constraint on the electron density matrix. International Journal of Quantum Chemistry, 6(3):519–523, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Carbo:1975:EMP

[CM75a] Ramón Carbó and Miguel Martin. Electrostatic molecular potentials: Mulliken approximation. International Journal of Quantum Chemistry, 9(2):193–214, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cusack:1975:CMD

[CM75b] S. Cusack and N. H. March. Charge and momentum density in metals and alloys. International Journal of Quantum Chemistry, 9(S9):481–496, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Combs:1976:USE

[CM76] L. L. Combs and Carl P. Miller. The use of single exponential orbitals in one-center molecular orbital calculations. International Journal of Quantum Chemistry, 10(3):455–460, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cohen:1978:PAF

[CM78] M. Cohen and R. P. Mceachran. Padé approximant formulas for atomic oscillator strengths. International Journal of Quantum Chemistry, 14(S12):59–66, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Chao:1977:AMS

[CODM77] K. A. Chao, F. A. Oliveira, R. O. De Cerqueira, and Norberto Majlis. The AMO method at small internuclear distances. International Journal of Quantum Chemistry, 12(1):11–17, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coffey:1974:PEI

[Cof74a] Patrick Coffey. Potential energy integrals in semiempirical MO methods. International Journal of Quantum Chemistry, 8(2): 263–266, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coffey:1974:VOT

[Cof74b] Patrick Coffey. Virtual orbital transformation prior to configuration interaction with localized orbitals. International Journal of Quantum Chemistry, 8(5):777–782, September 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coleman:1970:RRF

[Col70] A. J. Coleman. Recent results on fermion N-representability. International Journal of Quantum Chemistry, 5(S4):355–358, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Coleman:1971:RRF

[Col71] A. J. Coleman. Recent results on fermion N-representability. International Journal of Quantum Chemistry, 4(??):355–??, ?? 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coldwell:1977:ZMC

[Col77a] Robert Lynn Coldwell. Zero Monte Carlo error or quantum mechanics is easier. International Journal of Quantum Chemistry, 12(S11):215–222, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement:

Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Coleman:1977:CSE

[Col77b] A. J. Coleman. The convex structure of electrons. International Journal of Quantum Chemistry, 11(6):907–916, June 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coleman:1978:RDO

[Col78] A. J. Coleman. Reduced density operators and the N-particle problem. International Journal of Quantum Chemistry, 13 (1):67–82, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coldwell:1979:DMC

[Col79] R. L. Coldwell. A direct Monte Carlo evaluation of the derivative of the Born–Oppenheimer potential between two ions in liquid sodium. International Journal of Quantum Chemistry, 16(S13):705–717, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Comorosan:1974:MPB

[Com74] Sorin Comorosan. The measurement problem in biology. International Journal of Quantum Chemistry, 8(S1):221–228, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Combs:1977:BRBa

[Com77a] Leon L. Combs. Book review: Electron Correlation in Small Molecules. A. C. Hurley. Academic Press, London, England, 1976. VIII + 276 pp. Price: £10.50. International Journal of Quantum Chemistry, 12(5):967, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Hur76a].

Combs:1977:BRBb

[Com77b] Leon L. Combs. Book review: Introduction to the Electron Theory of Small Molecules. A. C. Hurley. Academic Press, London, England, 1976. X + 329 pp. Price: £12.00. International Journal of Quantum Chemistry, 12(5):967–968, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Hur76b].

Combs:1977:LCM

[Com77c] Leon L. Combs. A linear combination of molecular orbital formalism. International Journal of Quantum Chemistry, 11 (6):1001–1004, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Combes:1978:CST

[Com78] J. M. Combes. Complex scaling transformations in stationary and scattering problems. International Journal of Quantum Chemistry, 14(4):353–359, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Connolly:1970:CBT

[Con70] John W. D. Connolly. The comparison between theoretical and “experimental” band structures. International Journal of Quantum Chemistry, 5(S4):419–424, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Connolly:1972:CMI

[Con72] John W. D. Connolly. Calculation of molecular ionization potentials: Comparison of MS Xα and ESCA. International Journal of Quantum Chemistry, 6(S6):201–208, January 17– 22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Condon:1973:RLQ

[Con73] E. U. Condon. Reminiscences of a life in and out of quantum mechanics. International Journal of Quantum Chemistry, 7 (S7):7–22, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Cooper:1974:TMW

[Coo74] W. Grant Cooper. Theory of mutations. I. Wave mechanical model for replication independent mutations. International Journal of Quantum Chemistry, 8(S1):135–148, January 17– 19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cooper:1978:CPE

[Coo78a] W. Grant Cooper. Comment on possible experimental investigation of the microscopic proton–electron model of the genetic code. International Journal of Quantum Chemistry, 14(S5): 463–467, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cooper:1978:TMI

[Coo78b] W. Grant Cooper. Theory of mutations. II. Physical model of genetic stability. International Journal of Quantum Chemistry, 14(1):71–89, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cooper:1979:PTH

[Coo79] W. Grant Cooper. Proton transitions in hydrogen bonds of DNA: A first-order perturbation model. International Journal of Quantum Chemistry, 16(S6):171–188, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Corio:1972:RRG

[Cor72] P. L. Corio. On the representations of the rotation group. International Journal of Quantum Chemistry, 6(2):289–296, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cantril:1970:SSC

[CP70] Jerry M. Cantril and Herbert A. Pohl. Simplified SCF calculations for sigma-bonded systems IV: Hydrogen bonding among second row atoms. International Journal of Quantum Chemistry, 5(S4):165–178, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Cizek:1971:CPA

[ČP71] J. Čižek and J. Paldus. Correlation problems in atomic and molecular systems III. Rederivation of the coupled-pair manyelectron theory using the traditional quantum chemical methods. International Journal of Quantum Chemistry, 5(4):359– 379, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cizek:1972:DCE

[ČP72] J. Čižek and J. Paldus. A direct calculation of the excitation energies of closed-shell systems using the Green function technique. International Journal of Quantum Chemistry, 6(S6): 435–438, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Cizek:1977:AAB

[ČP77] J. Čížek and J. Paldus. An algebraic approach to bound states of simple one-electron systems. International Journal of Quantum Chemistry, 12(5):875–896, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cizek:1974:CEL

[ČPH74] J. Čížek, J. Paldus, and I. Hubač. Correlation effects in the low–lying excited states of the PPP models of alternant hydrocarbons. I. Qualitative rules for the effect of limited configuration interaction. International Journal of Quantum Chemistry, 8(6):951–970, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cizek:1975:CEP

[ČPP75] J. Čížek, A. Pellégatti, and J. Paldus. Correlation effects in the PPP model of alternant π-electronic systems: two-point Padé approximant approach. International Journal of Quantum Chemistry, 9(6):987–1007, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Clementi:1972:SES

[CR72] Enrico Clementi and Anne Routh. Study of the electronic structure of molecules. XV. Comments on the molecular orbital valency state and on the molecular orbital energies. International Journal of Quantum Chemistry, 6(3):525–539, May 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Crossley:1977:LPO

[Cro77] Richard J. S. Crossley. On Löwdin’s projection operators for angular momentum. I. International Journal of Quantum Chemistry, 11(6):917–929, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coulter:1972:SFI

[CS72] C. Alton Coulter and Romas A. Shatas. A spectral function for an interacting electron-lattice system. International Journal of Quantum Chemistry, 6(S6):419–427, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Calais:1973:SAM

[CS73] Jean-Louis Calais and Gunnar Sperber. A study of the AMO method as applied to the lithium metal. I. Review, results, and discussion. International Journal of Quantum Chemistry, 7(3): 501–520, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cornelius:1974:CER

[CS74] John Cornelius and John R. Sabin. A CNDO estimate of the relative affinities of taurine and isethionic acid for alkali metal ions. International Journal of Quantum Chemistry, 8(S1):43– 48, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Chin:1976:CAE

[CS76] Chen-An Chin and Pill-Soon Song. Configuration analysis of the electronic excitation in photobiological receptors. International Journal of Quantum Chemistry, 10(S3):89–98, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Combes:1978:RAP

[CS78] J. M. Combes and R. Seiler. Regularity and asymptotic properties of the discrete spectrum of electronic Hamiltonians. International Journal of Quantum Chemistry, 14(2):213–229, August 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1975:PDE

[Csa75] P. Csavinszky. Position dependence of the exchange hole in atoms and ions. International Journal of Quantum Chemistry, 9(S9):123–128, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Csavinszky:1976:VPP

[Csa76] P. Csavinszky. A variational principle for the potential of impurity ions in semiconductors with spatially variable dielectric constants. International Journal of Quantum Chemistry, 10 (S10):305–311, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Csavinszky:1977:ICE

[Csa77a] P. Csavinszky. Investigation of the convergence of an expansion of the screening-charge density for impurity ions in semiconductors. International Journal of Quantum Chemistry, 12 (2):335–342, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1977:VMT

[Csa77b] P. Csavinszky. A variationally modified Thomas–Fermi approach to the repulsive potential between noble gas atoms and positive and negative ions of spherically symmetric electron configurations. International Journal of Quantum Chemistry, 12(S11):105–117, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Csavinszky:1978:VPL

[Csa78a] P. Csavinszky. Variational principle for the Landau levels of holes in degenerate bands. International Journal of Quantum Chemistry, 14(S12):441, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Csavinszky:1978:VTP

[Csa78b] P. Csavinszky. A variational treatment of the potential of impurity ions in semiconductors with spatially variable dielectric constants. International Journal of Quantum Chemistry, 13(2): 221–226, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1979:AVS

[Csa79a] P. Csavinszky. Approximate variational solution of the Thomas–Fermi–Amaldi equation for singly charged negative ions. International Journal of Quantum Chemistry, 16(S13): 115–120, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Csavinszky:1979:NFA

[Csa79b] P. Csavinszky. Note on the Fermi–Amaldi correction for the Thomas–Fermi theory of atoms. International Journal of Quantum Chemistry, 15(5):511–515, May 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cederbaum:1978:AMO

[CSDV78] L. S. Cederbaum, J. Schirmer, W. Domcke, and W. Von Niessen. On the adequacy of the molecular-orbital picture for describing ionization processes. International Journal of Quantum Chemistry, 14(5):593–601, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cherepanov:1972:GTM

[CSFM72] V. I. Cherepanov, A. A. Shchetkov, D. S. Farberov, and A. N. Men. Group-theoretical methods for determining permitted terms of the electronic states of polyatomic molecules and complexes in MO and HL schemes of calculation. II. International Journal of Quantum Chemistry, 6(1):83–91, January 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Corsaro:1974:VPT

[CSH74] Felicia A. Corsaro, Françoise R. Solenberger, and Hendrik F. Hameka. A variational perturbation treatment of the hydrogen molecular ion using the green function method on the united atom. International Journal of Quantum Chemistry, 8(3):319– 333, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csizmadia:1976:TPM

[Csi76] I. G. Csizmadia. Theory and practice of MO calculations on organic molecules, volume 1 of Progress in theoretical organic chemistry. Elsevier, Amsterdam, The Netherlands, 1976. ISBN 0-444-41468-1 (American Elsevier). 378 pp. LCCN QD476 .C78. fl 100.00.

Csizmadia:1977:AMT

[Csi77] I. G. Csizmadia, editor. Applications of MO theory in organic chemistry: [proceedings], volume 2 of Progress in theoretical organic chemistry. Elsevier, Amsterdam, The Netherlands, 1977. ISBN 0-444-41565-3. LCCN QD461 .T43 1976.

Christoffersen:1971:TMF

[CSM71] Ralph E. Christoffersen, Lester L. Shipman, and Gerald M. Macgiora. Transferability of molecular fragments to large molecules. International Journal of Quantum Chemistry, 5 (S5):143–149, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Cheung:1979:ERD

[CSR79] L. M. Cheung, K. R. Sundberg, and K. Ruedenberg. Electronic rearrangements during chemical reactions. II. Planar dissociation of ethylene. International Journal of Quantum Chemistry, 16(5):1103–1139, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chong:1974:ADE

[CSY^{+}74] Delano P. Chong, William R. Scott, Chee P. Yue, Pearl S. C. Wang, Margaret L. Benston, and William E. Palke. Application of the distinguishable electron method. International Journal of Quantum Chemistry, 8(1):137–150, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chong:1977:PCK

[CT77] Delano P. Chong and Yuuzi Takahata. Perturbation corrections to Koopmans’ theorem. IV. Energy-shifted perturbation theory. International Journal of Quantum Chemistry, 12(3):549–559, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cheney:1979:EFA

[CT79] B. Vernon Cheney and Timothy Tolly. Electronic factors affecting receptor binding of dibenzo-p-dioxins and dibenzofurans. International Journal of Quantum Chemistry, 16(1):87– 110, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Curtiss:1977:ICH

[Cur77] L. A. Curtiss. Ab initio calculations on hydrogen bonding in alcohols: Dimers of CH_{3}OH, CH_{3}CH_{2}OH, and CF_{3}CH_{2}OH. International Journal of Quantum Chemistry, 12(S11):459–467, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Curtiss:1978:MOS

[Cur78] L. A. Curtiss. Molecular orbital studies of Al_{2}F_{6} and Al_{2}Cl_{6} using a minimal basis set. International Journal of Quantum Chemistry, 14(6):709–715, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cusachs:1975:RTJ

[Cus75] L. C. Cusachs. Remarks in tribute to Jeremy Musher (1935-1974). International Journal of Quantum Chemistry, 9(S9):75– 81, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Corvilain:1973:KSB

[CV73] S. Corvilain and G. Verhaegen. K-shell binding energies in C and O. International Journal of Quantum Chemistry, 7(S7):69– 81, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Cartling:1976:RMS

[CW76] Bo G. Cartling and Douglas M. Whitmore. Relativistic molecular spinors by generalized multiple scattering theory. International Journal of Quantum Chemistry, 10(3):393–412, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cherng:1977:NIO

[CW77] Sharon Cherng and Paul Westhaus. A note on the importance of d orbitals in the calculation of effective interactions and the excitation spectra of atoms and molecules. International Journal of Quantum Chemistry, 12(1):193–205, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chang:1978:PTM

[CW78] Shih Y. Chang and Harel Weinstein. Perturbation treatment of multiple site reactivity. II. Additivity in trimolecular interactions. International Journal of Quantum Chemistry, 14(6): 801–813, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Collins:1971:SCR

[CWE71] T. C. Collins, G. G. Wepfer, and R. N. Euwema. Selfconsistent relativistic orthogonalized plane-wave calculation of ZnSe. International Journal of Quantum Chemistry, 5(S5):451– 458, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Cullen:1975:AVP

[CZ75] John Cullen and Michael C. Zerner. An approximate variational perturbation model for molecular energy. International Journal of Quantum Chemistry, 9(S9):343–350, January 19– 25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Carsky:1979:ICA

[ČZ79] P. Čársky and R. Zahradník. Ab initio calculations and absolute rate theory. Critical test for a series of elementary gasphase reactions. International Journal of Quantum Chemistry, 16(2):243–256, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Callaway:1970:TBC

[CZNL70] J. Callaway, H. M. Zhang, T. E. Norwood, and J. Langlinais. Tight binding calculation of energy bands in nickel. International Journal of Quantum Chemistry, 5(S4):425–429, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Dolphin:1979:HPM

[DAC^{+}79] David Dolphin, Anthony W. Addison, Max Cairns, Robert K. Dinello, Nicholas P. Farrell, Brian R. James, Donald R. Paulson, and Curtis Welborn. Heme proteins and metalloporphyrins: Redox chemistry and oxygen binding. International Journal of Quantum Chemistry, 16(2):311–329, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dahl:1978:BBS

[Dah78] Jens Peder Dahl. Biorthonormal basis sets and the interaction between one-electron atoms. International Journal of Quantum Chemistry, 14(2):191–207, August 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dalgaard:1979:ECT

[Dal79] Esper Dalgaard. Expansion and completeness theorems for operator manifolds. International Journal of Quantum Chemistry, 15(2):169–180, February 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Danese:1972:MCU

[Dan72] J. Bryan Danese. Methane calculations using the multiplescattering technique with Xα exchange. International Journal of Quantum Chemistry, 6(S6):209–215, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Demoulin:1979:QMS

[DAP79] D. Demoulin, A-M. Armbruster, and B. Pullman. Quantummechanical study of the interaction of glyoxal with arginine. International Journal of Quantum Chemistry, 16(3):631–639, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Daudey:1974:PICc

[Dau74] J. P. Daudey. Perturbative ab initio calculations of intermolecular energies. III. The water dimer. International Journal of Quantum Chemistry, 8(1):29–43, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Daudel:1977:RPC

[Dau77] R. Daudel. Recent progress in chemical carcinogenesis. International Journal of Quantum Chemistry, 12(S4):169–177, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Davison:1972:EBD

[Dav72] Sydney G. Davison. Energy bands of diatomic crystals in the many-neighbor approximation. International Journal of Quantum Chemistry, 6(S6):387–393, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Davidson:1974:MES

[Dav74] Ernest R. Davidson. Matrix elements for spin-adapted configurations. International Journal of Quantum Chemistry, 8 (1):83–89, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davenport:1977:MST

[Dav77a] James W. Davenport. Multiple scattering theory of photoemission. International Journal of Quantum Chemistry, 12(S11): 89–96, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Davies:1977:SSM

[Dav77b] R. H. Davies. A steady-state model of drug–receptor interaction in vivo applied to antagonists of the β-adrenergic receptor. International Journal of Quantum Chemistry, 12(S4):413–442, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Davydov:1979:SBM

[Dav79] A. S. Davydov. Solitons, bioenergetics, and the mechanism of muscle contraction. International Journal of Quantum Chemistry, 16(1):5–17, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davies:1976:HBP

[DBBJ76] R. H. Davies, R. D. Bagnall, W. Bell, and W. G. M. Jones. The hydrogen bond proton donor properties of volatile halogenated hydrocarbons and ethers and their mode of action in anaesthesia. International Journal of Quantum Chemistry, 10 (S3):171–185, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Davies:1974:QIP

[DBJ74] R. H. Davies, R. D. Bagnall, and W. G. M. Jones. A quantitative interpretation of phase effects in anaesthesia. International Journal of Quantum Chemistry, 8(S1):201–212, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Danese:1973:NMT

[DC73] J. B. Danese and J. W. D. Connolly. Non-muffin-tin charge density corrections to multiple-scattering-Xα-method. International Journal of Quantum Chemistry, 7(S7):279–287, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Deutsch:1975:EOA

[DC75] P. W. Deutsch and T. C. Collins. An excitation operator approach to localized excited states. International Journal of Quantum Chemistry, 9(S9):213–220, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Daudey:1974:PICa

[DCM74] J. P. Daudey, P. Claverie, and J. P. Malrieu. Perturbative ab initio calculations of intermolecular energies. I. Method. International Journal of Quantum Chemistry, 8(1):1–15, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dunlap:1977:ALX

[DCS77] Brett I. Dunlap, John W. D. Connolly, and John R. Sabin. On the applicability of LCAO-X α methods to molecules containing transition metal atoms: The nickel atom and nickel hydride. International Journal of Quantum Chemistry, 12(S11): 81–87, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Dutta:1971:DPM

[DD71] C. M. Dutta and N. C. Dutta. Dipole properties of the metastable 2^{3}S state of helium. International Journal of Quantum Chemistry, 5(S5):59–66, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Daudey:1972:PPC

[DD72] J. P. Daudey and S. Diner. PCILON. perturbative configuration interaction using localized orbitals and numerical integration. I. Numerical integration techniques for the calculation of Hamiltonian matrix elements between localized orbitals. International Journal of Quantum Chemistry, 6(3):575–591, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Delhalle:1977:CED

[DD77a] J. Delhalle and S. Delhalle. Computation of the electronic density of states distributions of stereoregular polymers. International Journal of Quantum Chemistry, 11(2):349–358, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dugay:1977:AIM

[DD77b] M. Dugay and G. Debarge. AMOS and insulator-to-metal transitions. International Journal of Quantum Chemistry, 11(6): 1021–1033, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dutta:1970:MBA

[DDD70] N. C. Dutta, C. M. Dutta, and T. P. Das. Many-body approach to the properties of interacting atoms: II. International Journal of Quantum Chemistry, 5(S4):299–305, January 19–24, 1970.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Delhalle:1978:EST

[DDH78] J. Delhalle, S. Delhalle, and F. E. Harris. Evaluation of some trigonometric series occurring in infinite chain polymer calculations. International Journal of Quantum Chemistry, 13(2): 239–254, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DeAViccaro:1976:SPL

[De 76] M. H. De A. Viccaro. Symmetry property of Löwdin’s “alpha” function. International Journal of Quantum Chemistry, 10(6): 1081–1082, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DeHosson:1978:LES

[De 78] J. Th. M. De Hosson. Localized electronic states near dislocations in transition metals. International Journal of Quantum Chemistry, 14(S12):469–482, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Deal:1972:LRI

[Dea72] Walter J. Deal. The long-range interaction between two hydrogen atoms. International Journal of Quantum Chemistry, 6(3): 593–596, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Decicco:1970:CSX

[Dec70] P. D. Decicco. Compton scattering of X-rays from solids in the one-electron model. International Journal of Quantum Chemistry, 5(S4):437–444, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

DelConde:1978:IGG

[DEC78] Guillermo Del Conde, Mirna Estrada, and Alberto Cárdenas. Interaction of glyoxal with glycine and N-methylacetamide: Some aspects of their potential energy surface and its relation with cancer. International Journal of Quantum Chemistry, 14(S5):393–401, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Duke:1975:SIM

[DEE^{+}75] Brian J. Duke, Deidre R. Eilers, James E. Eilers, Sungzong Kang, A. Liberles, and Brian O’Leary. Simulated ab initio molecular orbital calculations of large polynuclear aromatic hydrocarbons. International Journal of Quantum Chemistry, 9 (S2):155–170, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

DelRe:1973:BPL

[Del73] Giuseppe Del Re. Bond properties from a localized MO–LCAO approach. International Journal of Quantum Chemistry, 7(S7): 193–206, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Delhalle:1974:NPC

[Del74] J. Delhalle. On the need of precision in the calculation of the LCAO density matrix of polymers. International Journal of Quantum Chemistry, 8(2):201–208, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Desclaux:1972:RTO

[Des72] Jean-Paul Desclaux. Relativistic treatment of outer electron orbitals. International Journal of Quantum Chemistry, 6(S6): 25–41, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Davison:1976:EBH

[DF76] Syndey G. Davison and E-Ni Foo. Energy bands of sphybrid crystals in the many-neighbor approximation. International Journal of Quantum Chemistry, 10(5):867–872, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Daudel:1977:RBS

[DGC77] R. Daudel, J. D. Goddard, and I. G. Csizmadia. A relationship between the sizes and energies of localized molecular orbitals. I. A study of selected first-row hydrides. International Journal of Quantum Chemistry, 11(1):137–147, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DaRAndrade:1973:MOA

[DH73] P. Da R. Andrade and A. Holz. Molecular orbital admixture coefficient by Raman scattering. International Journal of Quantum Chemistry, 7(2):407–409, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Day:1976:MMT

[DH76] Orville W. Day and George S. Handler. Multipole moments in Thomas–Fermi theory. International Journal of Quantum Chemistry, 10(S10):369–373, January 19–24, 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Domelsmith:1978:PSP

[DH78] L. N. Domelsmith and K. N. Houk. Photoelectron spectra of psychotropic drugs. III. Ionization potentials and partition coefficients as predictors of substituted amphetamine psychoactivities. International Journal of Quantum Chemistry, 14(S5): 257–268, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Dupuis:1977:MSC

[DK77] Michel Dupuis and Harry F. King. Molecular symmetry and closed-shell SCF calculations. I. International Journal of Quantum Chemistry, 11(4):613–625, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dalgaard:1975:EWM

[DL75] Esper Dalgård and Jan Linderberg. Energy weighted maximum overlap in magnetic fields applications to electron spin resonance problems. International Journal of Quantum Chemistry, 9(S9):269–277, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Devries:1979:MCL

[DLG79] Paul L. Devries, Kai-Shue Lam, and Thomas F. George. Molecular collisions in a laser field: Basis set selection and the rotating-wave approximation. International Journal of Quantum Chemistry, 16(S13):541–551, March 11–17, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Downing:1972:PDE

[DM72] J. Downing and J. Michl. Polarization directions of electronic transitions in large conjugated molecules of low symmetry. International Journal of Quantum Chemistry, 6(S6):311– 317, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Dmitriev:1975:APTa

[Dmi75a] Yuri Dmitriev. Adiabatic perturbation theory for the degenerate case. I. Perturbation of isolated degenerate or quasidegenerate levels. International Journal of Quantum Chemistry, 9(6): 1033–1045, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dmitriev:1975:APTb

[Dmi75b] Yuri Dmitriev. Adiabatic perturbation theory for the degenerate case. II. Perturbation of degenerate bound states embedded in the continuum. International Journal of Quantum Chemistry, 9(6):1047–1053, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Doktorov:1975:CSA

[DMM75] E. V. Doktorov, I. A. Malkin, and V. I. Man’ko. Coherent states and asymptotic behavior of the symmetric top wave functions. International Journal of Quantum Chemistry, 9(6): 951–968, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Daudey:1974:PICb

[DMR74] J. P. Daudey, J. P. Malrieu, and Olivia Rojas. Perturbative ab initio calculations of intermolecular energies. II. The He⋯He problem. International Journal of Quantum Chemistry, 8(1):17–28, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davies:1978:SAA

[DMS^{+}78] R. H. Davies, R. C. Mason, D. A. Smith, D. J. McNeillie, and R. James. Speed of action of anesthetics and hydrogen-bond proton-acceptor properties. Estimation of the free energy of interaction of O·H − O bonding in hindered phenols based on CNDO/2 potentials. International Journal of Quantum Chemistry, 14(S5):221–243, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Dalton:1975:NME

[DN75] Bryan J. Dalton and Philip D. Nicholson. Nonrigid molecule effects in molecules with linear or nongeometrically equivalent equilibrium configurations. International Journal of Quantum Chemistry, 9(2):325–377, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Doolen:1978:CSA

[Doo78] Gary D. Doolen. Complex scaling: An analytic model and some new results for e^{+} H resonances. International Journal of Quantum Chemistry, 14(4):523–528, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Daudel:1979:SQG

[DPGC79] R. Daudel, R. A. Poirier, J. D. Goddard, and I. G. Csizmadia. Study of the quality of Gaussian basis sets for carbon and silicon: Calculations on methane and silane. International Journal of Quantum Chemistry, 15(3):261–270, March 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duprez-Quesnoit:1977:EQD

[DQ77] C. Duprez-Quesnoit. Etude Quantique du Pouvoir Rotatoire Naturel et de la Biréfringence Gyrotropique dans un Cristal. International Journal of Quantum Chemistry, 11(3):381–391, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dmitriev:1975:VPT

[DR75] Yuri Dmitriev and Björn Roos. A variational principle for transition and density matrices and approximations to the equations-of-motion. International Journal of Quantum Chemistry, 9(5):825–834, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Datta:1977:MOP

[DR77] Sambhu Datta and F. S. Richardson. Magneto-optical properties of a spinless particle moving in a harmonic potential. International Journal of Quantum Chemistry, 11(4):525–542, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Daudel:1978:DEM

[DRCT78] R. Daudel, H. Le Rouzo, R. Cimiraglia, and J. Tomasi. Dependence of the electrostatic molecular potential upon the basis set and the method of calculation of the wave function. Case of the ground ^{3}A_{1} (π → π*) and ^{1}A_{1} (π → π*) states of formaldehyde. International Journal of Quantum Chemistry, 13(4):537–552, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davison:1970:PST

[DS70a] S. G. Davison and M. Steślicka. Present status of Tamm states in surface reconstructed crystals. International Journal of Quantum Chemistry, 5(S4):455–462, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Davison:1970:RTL

[DS70b] S. G. Davison and M. Steślicka. Relativistic treatment of localized states. A review. International Journal of Quantum Chemistry, 5(S4):445–453, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Damommio:1974:AAS

[DS74] A. Damommio and M. Synek. Accurate analytical selfconsistent field wave functions for Tm^{2+} and Tm^{3+}. International Journal of Quantum Chemistry, 8(1):73–78, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Danese:1976:CBC

[DS76a] J. B. Danese and J. R. Schrieffer. Chemisorption bonding and catalysis. International Journal of Quantum Chemistry, 10 (S10):289–295, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Davidson:1976:SME

[DS76b] Ernest R. Davidson and Larilyn Zeller Stenkamp. SCF methods for excited states. International Journal of Quantum Chemistry, 10(S10):21–31, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Dhingra:1979:TEH

[DS79] M. M. Dhingra and Ramaswamy H. Sarma. Transmission, evolution, and helix handedness of nucleic acid conformation. International Journal of Quantum Chemistry, 16(S6):131–152, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Daudel:1977:TLP

[DSC^{+}77] R. Daudel, M. E. Stephens, I. G. Csizmadia, C. Kozmutza, E. Kapuy, and J. D. Goddard. Theory of lone pairs. II. A moment analysis of localized molecular orbitals in ten-electron hydrides. International Journal of Quantum Chemistry, 11 (4):665–683, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Day:1974:GHF

[DSG74] Orville W. Day, Darwin W. Smith, and Claude Garrod. A generalization of the Hartree–Fock one-particle potential. International Journal of Quantum Chemistry, 8(S8):501–509, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Davies:1979:IBA

[DSMM79] R. H. Davies, D. A. Smith, D. J. McNeillie, and T. R. Morris. Identification of biologically active conformations in flexible drug molecules. International Journal of Quantum Chemistry, 16(S6):203–221, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Dementev:1974:CPS

[DSY74] A. I. Dement’ev, N. F. Stepanov, and S. S. Yarovoi. Convergence problems in the solution of SCF equations. International Journal of Quantum Chemistry, 8(1):107–117, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davison:1972:TET

[DT72] S. G. Davison and W. K. Tan. Two-electron theory of noninteracting surface and impurity states. International Journal of Quantum Chemistry, 6(3):549–554, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dugay:1975:AMO

[DT75] M. Dugay and J. Thomas. Alternant molecular orbitals (AMO’s) for some organic semiconductors: Ion radical salts from 7, 7, 8, 8-tetracyanoquinodimethane (TCNQ) and chloranil/para-phenylenediamine (PDC) charge transfer complex. International Journal of Quantum Chemistry, 9(6):1087–1095, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duben:1972:ESC

[Dub72] Anthony J. Duben. Examination of the semi-classical theory of circular dichroism. International Journal of Quantum Chemistry, 6(4):787–791, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duff:1971:CFL

[Duf71] K. J. Duff. A computational form for Löwdin’s alpha function. International Journal of Quantum Chemistry, 5(1):111– 113, January 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duke:1979:ESL

[Duk79] C. B. Duke. Electronic structure of large molecules: CNDO/S3 model. International Journal of Quantum Chemistry, 16(S13): 267–281, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Dunne:1978:BRB

[Dun78] Lawrence J. Dunne. Book review: Intermolecular interactions: From diatomics to biopolymers. B. Pullman, ed. John Wiley & Sons Inc., New York, 1978. Price: $48.50, 447 pp. International Journal of Quantum Chemistry, 14(3):341, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Pul78a].

Hosson:1979:KSL

[DV79] J. T. H. M. De Hosson and W. Van Der Lugt. The Knight shift in liquid binary alloys: An application of quantum chemistry in liquid- and solid-state physics. International Journal of Quantum Chemistry, 16(S13):367–375, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Deal:1973:LRD

[DY73] Walter J. Deal and Ralph H. Young. Long-range dispersion interactions involving excited atoms: The H(1s)–H(2s) interaction. International Journal of Quantum Chemistry, 7(5): 877–892, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Deal:1976:PTA

[DY76] Walter J. Deal and Ralph H. Young. Perturbation theory with an approximate perturbation: The effect of charge overlap on interatomic interactions. International Journal of Quantum Chemistry, 10(3):419–428, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Julg:1976:ISD

[eAJ76] Alain Pellegatti et André Julg. Influence de la structure de diverses fonctions sur l’evaluation de l’énergie de correlation. (French) [Influence of the structure of various functions on the evaluation of correlation energy]. International Journal of Quantum Chemistry, 10(1):113–121, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Engelke:1974:EHP

[EB74] Sandra Z. Engelke and Charles L. Beckel. Extended Hückel β parameters applied to diatomic molecules. International Journal of Quantum Chemistry, 8(S8):209–217, January 20– 26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Elbert:1974:ICU

[ED74] Stephen T. Elbert and Ernest R. Davidson. Ab initio calculations on urea. International Journal of Quantum Chemistry, 8(6):857–892, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Estrada:1979:EAG

[ED79] Mirna Estrada and Guillermo Del Conde. Electron affinities of glyoxal, methylglyoxal, and vitamin C: Some aspects of the interaction between vitamin C and a protein model. International Journal of Quantum Chemistry, 16(S6):89–92, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Edenman:1970:OAI

[Ede70] Ragnar H. Edenman. Opening address to the International Symposium. International Journal of Quantum Chemistry, 5 (S4):1–2, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Ehrenson:1973:ISL

[EH73] S. Ehrenson and G. D. Harp. Importance sampling in local energy calculations on H_{2}. International Journal of Quantum Chemistry, 7(6):1099–1116, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Eigen:1978:AHP

[Eig78] Manfred Eigen. Abstract: The hypercycle: A principle of natural self-organization. International Journal of Quantum Chemistry, 14(S5):219, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Eisenstein:1976:MTH

[Eis76] Laura Eisenstein. Molecular tunneling in heme proteins. International Journal of Quantum Chemistry, 10(S3):21–27, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Eisenstein:1977:DCB

[Eis77] Laura Eisenstein. Dynamics of CO binding to heme proteins. International Journal of Quantum Chemistry, 12(S4):363–374, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Edmiston:1973:DBA

[EL73] Clyde Edmiston and Peter Lindner. On the dimerization of BH_{3} and the associated delocalization (resonance) energy. International Journal of Quantum Chemistry, 7(2):309–318, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Elder:1973:UMS

[Eld73] M. Elder. Use of molecular symmetry in SCF calculations. International Journal of Quantum Chemistry, 7(1):75–85, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ellzey:1973:SAM

[Ell73] M. L. Ellzey, Jr. Sequence-adapted molecular tensors: Algebraic methods and application to crystal field theory. International Journal of Quantum Chemistry, 7(2):253–284, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ellis:1977:MPR

[Ell77] D. E. Ellis. Moment-polarized relativistic potentials. International Journal of Quantum Chemistry, 12(S11):201–205, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Emanuel:1974:IUE

[Ema74] R. V. Emanuel. An investigation into the use of expansion methods in the calculation of chemical shifts and diamagnetic susceptibilities. International Journal of Quantum Chemistry, 8(5):745–754, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

England:1971:CDE

[Eng71] Walter England. Continuous degeneracy and energy-localization of molecular orbitals. International Journal of Quantum Chemistry, 5(6):683–697, November 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

England:1972:OCC

[Eng72] Walter England. One-center Coulomb, two-center hybrid, and two-center Coulomb integrals over STP functions. International Journal of Quantum Chemistry, 6(3):509–518, May 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Enoch:1976:VPO

[Eno76] Jay M. Enoch. Vertebrate photoreceptor orientation. International Journal of Quantum Chemistry, 10(S3):65–88, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Egan:1978:CBP

[ENRM78] John T. Egan, Shlomo Nir, Robert Rein, and Robert MacElroy. Configurations of base-pair complexes in solutions. International Journal of Quantum Chemistry, 14(S5):433–440, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Enz:1973:LDA

[Enz73] Charles P. Enz. Lattice dynamics above structural phase transitions. International Journal of Quantum Chemistry, 7(S7): 653–657, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Ellis:1971:EST

[EP71] D. E. Ellis and T. Parameswaran. Electronic structure of TiCl_{4}. International Journal of Quantum Chemistry, 5(S5): 443–449, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Epstein:1977:PTW

[Eps77] Saul T. Epstein. Perturbation theory within the variation method. International Journal of Quantum Chemistry, 11(5):

881–884, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Erdahl:1978:SDC

[Erd78a] R. M. Erdahl. On the structure of the diagonal conditions. International Journal of Quantum Chemistry, 13(6):731–736, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Erdahl:1978:R

[Erd78b] R. M. Erdahl. Representability. International Journal of Quantum Chemistry, 13(6):697–718, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ellis:1975:ADS

[ERW75] D. E. Ellis, A. Rosén, and P. F. Walch. Applications of the Dirac–Slater model to molecules. International Journal of Quantum Chemistry, 9(S9):351–358, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Euwema:1970:REA

[ES70] R. N. Euwema and D. J. Stukel. Relativistic exchange approximation study for He, Ne, Ar, Kr and Xe. International Journal of Quantum Chemistry, 5(S4):55–75, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Ellinger:1973:ANM

[ES73a] Y. Ellinger and J. Serre. An application of the nonrigid molecule group theory to a problem of chemical reactivity. International Journal of Quantum Chemistry, 7(S7):217–221, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Ellis:1973:ESB

[ES73b] D. E. Ellis and A. Seth. Electron states and bonding in titanium sulfide. International Journal of Quantum Chemistry, 7 (S7):223–229, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Epstein:1979:VBS

[ES79] Saul T. Epstein and Andrzej J. Sadlej. Variable basis sets in perturbation theory: Numerical finite perturbation versus analytic approach. International Journal of Quantum Chemistry, 15(2):147–167, February 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Englund:1979:SME

[ESG79] John C. Englund, William C. Schieve, and Robert F. Gragg. A stochastic master equation description of laserlike systems. International Journal of Quantum Chemistry, 16(S13):695–704, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Egan:1974:SID

[ESR74] John T. Egan, Thomas J. Swissler, and Robert Rein. Some improvements in DNA interaction calculations. International Journal of Quantum Chemistry, 8(S1):71–79, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Essen:1977:PBO

[Ess77] H. Essén. The physics of the Born–Oppenheimer approximation. International Journal of Quantum Chemistry, 12(4):721– 735, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Espinoza:1979:QMD

[ETF79] Ll. Espinoza, A. Toro, and P. Fuentealba. Quadrupole moments, dipole quadrupole A and quadrupole C polarizabilities by means of perturbation theory. International Journal of Quantum Chemistry, 16(5):939–954, November 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Euwema:1971:RCC

[Euw71] R. N. Euwema. Rapid convergence of crystalline energy bands by use of a plane-wave-Gaussian mixed basis set. International Journal of Quantum Chemistry, 5(S5):471–487, January 18– 23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Wijkman:1974:WR

[Exe74] Executive Vice President Gunnar Wijkman. Welcoming remarks. International Journal of Quantum Chemistry, 8(S1): 1, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Eyring:1978:ALP

[Eyr78] Henry Eyring. About Linus Pauling. International Journal of Quantum Chemistry, 14(S5):11–13, March 3–11, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Faegri:1979:NSP

[Fae79] K. Faegri, Jr. Neon L-shell photoabsorption cross sections from moment theory and finite basis pseudospectra. International Journal of Quantum Chemistry, 15(4):411–421, April 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fankhauser:1970:DCM

[Fan70] Hans R. Fankhauser. On the Dynamics of Crystals with Molecular Impurity Centers. I. Compatibility Conditions. International Journal of Quantum Chemistry, 4(3):342–345, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Faulkner:1971:ESD

[Fau71] J. S. Faulkner. Electronic states in disordered alloys: Comparison of methods. International Journal of Quantum Chemistry, 5(S5):543–553, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Faulkner:1976:ESD

[Fau76] J. S. Faulkner. Electronic states in disordered alloys. II. results for real solids. International Journal of Quantum Chemistry, 10 (S10):405–415, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Flurry:1972:MOC

[FB72] R. L. Flurry, Jr. and D. L. Breen. Molecular orbital calculations on biopolymers. III. Conductivity in nucleic-acid base homopolymers. International Journal of Quantum Chemistry, 6(S6):305–310, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Froelich:1976:CRW

[FB76] Piotr Froelich and Erkki Brändas. Calculation of resonance widths via L^{2} expansion techniques. International Journal of Quantum Chemistry, 10(S10):353–357, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fink:1979:NIA

[FBL79] William H. Fink, Aldona M. Butkus, and Jesus P. Lopez. A novel ab initio approach to problems of electronic structure calculations of the gas-solid interface. International Journal of Quantum Chemistry, 16(S13):331–347, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Fazzio:1979:PDC

[FCL79] A. Fazzio, M. J. Caldas, and J. R. Leite. Point defects in covalent semiconductors: A molecular cluster model. International Journal of Quantum Chemistry, 16(S13):349–361, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Feuchtwang:1977:MBT

[Feu77] T. E. Feuchtwang. Many-body theories of inelastic electron tunneling spectroscopy (IETS): Experimental evidence of the inadequacy of the transfer Hamiltonian formalism of IETS, interpretation and discussion of alternative theories. International Journal of Quantum Chemistry, 12(S11):607–624, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Ferreira:1979:SMT

[FFCB79] Luiz G. Ferreira, Adalberto Fazzio, Huberto Closs, and Luiz M. Brescansin. Study of the muffin-tin approximation in the multiple-scattering method. International Journal of Quantum Chemistry, 16(5):1021–1031, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Flament:1979:EMM

[FG79] J. P. Flament and H. P. Gervais. Equations-of-motion method: Calculation of the k lowest or highest solutions. International Journal of Quantum Chemistry, 16(6):1347–1356, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Figeys:1977:RII

[FGV77] H. P. Figeys, P. Geerlings, and C. Van Alsenoy. Rotational invariance of INDO theories including d-orbitals into the basis set. International Journal of Quantum Chemistry, 11(5):705– 713, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Froelich:1977:NCT

[FHB77] Piotr Froelich, Michael Hehenberger, and Erkki Brändas. A note on the complex transformation method and the extended virial theorem for treating resonances in the continuous spectrum. International Journal of Quantum Chemistry, 12(S11): 295–299, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Ford:1970:MMS

[FHP70] B. Ford, G. G. Hall, and J. C. Packer. Molecular modelling with spherical Gaussians. International Journal of Quantum Chemistry, 4(5):533–539, September 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fink:1973:PBB

[Fin73] William H. Fink. Projection between basis sets: A means of defining the electronic structure of functional subunits. International Journal of Quantum Chemistry, 7(6):1045–1050, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fratev:1970:WBP

[FJP70] F. Fratev, R. Janoschek, and H. Preuss. Wellenmechanische Behandlung der Protonisierung von F- und CH_{3}- monosubstituierten Äthylenen. (German) [Wave-mechanical treatment of the protonation of F- and CH_{3}-monosubstituted ethylene]. International Journal of Quantum Chemistry, 4(5): 529–532, September 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fraga:1976:HAD

[FKS76] Serafin Fraga, Jacek Karwowski, and K. M. S. Saxena. Handbook of atomic data. Elsevier, Amsterdam, The Netherlands, 1976. ISBN 0-444-41461-4. 551 pp. LCCN QC173.397 .F7.

Frischleder:1979:QCS

[FL79] H. Frischleder and R. Lochmann. Quantum-chemical and semiclassical calculations of intermolecular interactions of phospholipids. International Journal of Quantum Chemistry, 16(2):

203–213, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Flannery:1979:AMC

[Fla79] M. R. Flannery. Atomic and molecular collision processes in rare-gas-halide lasers and rare-gas excimer lasers. International Journal of Quantum Chemistry, 16(S13):501–529, March 11– 17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Flurry:1972:SSM

[Flu72] R. L. Flurry, Jr. Site symmetry in molecular point groups. International Journal of Quantum Chemistry, 6(S6):455–458, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Farberov:1972:TBM

[FM72] D. S. Farberov and A. N. Men. On the translation between MO and HL descriptions of co-ordination compounds. International Journal of Quantum Chemistry, 6(4):613–616, July 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Farantos:1978:CDC

[FM78a] S. C. Farantos and J. N. Murrell. Classical dynamics of the O + ClO → Cl + O_{2} and Cl + O_{3} → ClO + O_{2} reactions. International Journal of Quantum Chemistry, 14(5):659–674, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Frishberg:1978:NDM

[FM78b] C. A. Frishberg and L. J. Massa. Notes on density matrix model for coherent x-ray diffraction. International Journal of Quantum Chemistry, 13(6):801–810, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Findley:1979:GGC

[FM79] G. L. Findley and S. P. McGlynn. A generalized genetic code. International Journal of Quantum Chemistry, 16(S6):313–327, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Freeman:1975:EDI

[FMRG75] A. J. Freeman, H. W. Myron, J. Rath, and R. P. Gupta. Electronically driven instabilities in metals. International Journal of Quantum Chemistry, 9(S9):535–550, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Franchini:1970:EGF

[FMZ70] P. F. Franchini, R. Moccia, and M. Zandomeneghi. Extended group function calculations for H_{2}O, NH_{3} and CH_{4}. International Journal of Quantum Chemistry, 4(5):487–501, September 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fletcher:1973:DTM

[FN73] G. C. Fletcher and B. A. Nudel. Dependence of transition metal energy bands on Hubbard parameters. International Journal of Quantum Chemistry, 7(S7):619–627, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Foglio:1971:CDM

[Fog71] Mario E. Foglio. Concentration dependence of the macroscopic magnetic anisotropy and of the ferrimagnetic line width in the gallium substituted europium iron garnet. International Journal of Quantum Chemistry, 5(S5):653–659, January 18– 23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Fox:1975:MPQ

[Fox75] Sidney W. Fox. The matrix for the protobiological quantum: Cosmic casino or shapes of molecules? International Journal of Quantum Chemistry, 9(S2):307–320, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Firsht:1977:CMS

[FP77] D. Firsht and B. T. Pickup. On convergence in many-shell SCF theory. International Journal of Quantum Chemistry, 12 (4):765–776, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fricke:1972:WCO

[Fri72] B. Fricke. What can one learn with muons about atomic physics? International Journal of Quantum Chemistry, 6(S6): 43–46, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Fritsche:1974:BSC

[FRM74] Lothar Fritsche and Mohammad Rafat-Mehr. Band structure calculation for realistic (non-muffin-tin) potentials based on the rigorous cellular method. International Journal of Quantum Chemistry, 8(S8):457–463, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

FroeseFischer:1970:CIS

[Fro70] Charlotte Froese Fischer. Configuration interaction in the spinpolarized Hartree–Fock approximation. International Journal of Quantum Chemistry, 5(S4):95–102, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

FroeseFischer:1974:MHF

[Fro74] Charlotte Froese Fischer. A multiconfiguration Hartree–Fock approach to atomic structure calculations. International Journal of Quantum Chemistry, 8(S8):5–15, January 20–26, 1974.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

FroeseFischer:1975:CSL

[Fro75] Charlotte Froese Fischer. A correlation study of Li ground state by the MCHF procedure. International Journal of Quantum Chemistry, 9(2):273–285, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fromhold:1979:DES

[Fro79] A. T. Fromhold. Difference equation solutions for hopping transport of ions through membranes. International Journal of Quantum Chemistry, 16(S6):329–342, March 4–10, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Frisman:1979:ISS

[FSV79] E. V. Frisman, S. V. Slonitsky, and A. N. Veselkov. Influence of solvent structure on the conformation of the native DNA molecule. International Journal of Quantum Chemistry, 16(4): 847–855, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fabian:1977:MLC

[FTN77] J. Fabian and G. Tröger-Naake. MO–LCAO-calculations on polymethines. V. PPP-type calculations and configuration analyses of simple prototype dyes in terms of molecular subsystems. International Journal of Quantum Chemistry, 11(2): 259–269, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fimple:1976:FDA

[FU76] W. R. Fimple and M. J. Unwin. A finite-difference approach to the electron correlation problem. International Journal of Quantum Chemistry, 10(4):643–658, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fukui:1970:OST

[Fuk70] Ken’ichi Fukui. Orientation and stereoselection: Theory of orientation and stereoselection, volume 1 of Topics in current chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. 85 pp. LCCN QD1 .F58.

Fukui:1975:TOS

[Fuk75] Ken’ichi Fukui. Theory of orientation and stereoselection, volume 2 of Reactivity and structure. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. ISBN 0-387-07426-0. 134 pp. LCCN QD501 .F92.

Fimple:1975:FDS

[FW75] W. R. Fimple and S. P. White. A finite-difference solution of the Hartree–Fock equations for diatomic molecules. International Journal of Quantum Chemistry, 9(2):301–324, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Findley:1977:MRT

[FWFM77] G. L. Findley, K. Wittel, W. S. Felps, and S. P. McGlynn. Molecular Rydberg transitions. VIII. The geometry of ethylene in the R_{1s} state. International Journal of Quantum Chemistry, 12(S11):229–243, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Freeman:1979:ESC

[FWJ^{+}79] A. J. Freeman, C. S. Wang, T. Jarlborg, M. Weinert, F. Wagner, and C. W. Chu. Electronic structure of CuCl. International Journal of Quantum Chemistry, 16(S13):445–451, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Fukui:1971:NTI

[FY71] Kenichi Fukui and Tokio Yamabe. A note on the theory of interatomic long-range forces. International Journal of Quantum Chemistry, 5(4):478, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fedyanin:1979:EMO

[FY79] V. K. Fedyanin and L. V. Yakushevich. Energy migration in one-dimensional structures with resonance interaction. International Journal of Quantum Chemistry, 16(2):215–219, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gabbay:1976:RAS

[Gab76] Edmond J. Gabbay. A rational approach to the synthesis of template inhibitory antitumor drugs. International Journal of Quantum Chemistry, 10(S3):217–229, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gallup:1972:EPM

[Gal72a] G. A. Gallup. The N-electron problem and matrices representing the symmetric groups. International Journal of Quantum Chemistry, 6(4):761–778, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gallup:1972:VBC

[Gal72b] G. A. Gallup. Valence-bond calculations and matrix elements between two tableau functions of non-orthogonal orbitals. International Journal of Quantum Chemistry, 6(5): 899–910, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gallup:1974:SEW

[Gal74] G. A. Gallup. Symmetries of n-electron wave functions under various spatial groups. International Journal of Quantum Chemistry, 8(2):267–291, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garton:1974:GGE

[Gar74] D. Garton. Gradient of the general energy expression. International Journal of Quantum Chemistry, 8(4):619–621, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garrod:1978:DMM

[Gar78] Claude Garrod. Density matrix methods in orbital optimization for MCSCF calculations. International Journal of Quantum Chemistry, 13(6):769–775, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gaspar:1971:MEP

[Gás71] R. Gáspár. Many-electron problems III scaling of the density distributions and the potential fields. International Journal of Quantum Chemistry, 5(3):311–317, May 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1971:PAP

[GB71] O. Goscinski and E. Brändas. Padé approximants to physical properties via inner projections. International Journal of Quantum Chemistry, 5(2):131–156, March 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gruninger:1972:HTF

[GB72] John Gruninger and Richard Bailo. Hulthén transform functions for the 2s orbitals of atoms. International Journal of Quantum Chemistry, 6(S6):47–50, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Grein:1975:MME

[GB75] F. Grein and A. Banerjee. A multiconfiguration method for excited states of atoms and molecules. International Journal of Quantum Chemistry, 9(S9):147–154, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Garciaa-bach:1977:VLS

[GbK77a] M. A. Garcíaa-bach and D. J. Klein. Variational localized-site cluster expansions. VII. Higher ansätze and the linear Heisenberg model. International Journal of Quantum Chemistry, 12 (2):273–289, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Guse:1977:PEC

[GBK77b] M. P. Guse, R. J. Blint, and A. B. Kunz. Potential energy curves for NiH and NiH_{2}. International Journal of Quantum Chemistry, 11(5):725–732, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1973:PAP

[GBL^{+}73] O. Goscinski, E. Brändas, B. Laskowski, P. Van Leuven, and W. Th. A. M. Van Der Lugt. Padé approximants to physical properties via inner projections. International Journal of Quantum Chemistry, 7(1):133–134, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gupta:1974:CHF

[GBM74] A. Gupta, A. K. Bhattacharya, and P. K. Mukherjee. Coupled Hartree Fock calculation of static dipole polarizabilities and shielding factors of open shell systems. International Journal of Quantum Chemistry, 8(1):97–105, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ghosh:1970:TCS

[GBS70] Kunal Ghosh, Subirnath Bhattacharyya, and Purnendranath Sen. Theoretical calculations on σ-systems, part 1: Ethers and chloro-substituted ethers. International Journal of Quantum Chemistry, 4(4):347–353, July 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

George:1979:CIO

[GBT79] P. George, C. W. Bock, and M. Trachtman. Contributions from inner and outer shell electron energies to reaction heats for C_{1}, C_{2}, and C_{3} hydrocarbons. International Journal of Quantum Chemistry, 15(4):423–443, April 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

George:1978:ADR

[GBTB78] Philip George, Charles W. Bock, Mendel Trachtman, and Alistair M. Brett. Attractive-dominant and repulsive-dominant hydrocarbon reactions. International Journal of Quantum Chemistry, 13(3):271–295, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Godefroid:1977:MTS

[GBV77] M. Godefroid, J. J. Berger, and G. Verhaegen. Multiconfigurational transition state calculations of atomic oscillator strengths. The resonance transition of beryllium. International Journal of Quantum Chemistry, 12(S11):119–123, January 16– 22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Georghiou:1975:NSR

[GC75] S. Georghiou and Jorge E. Churchich. Nanosecond spectroscopy of retinol bound to proteins. International Journal of Quantum Chemistry, 9(S2):331–337, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Goddard:1977:SMO

[GC77] John D. Goddard and Imre G. Czizmadia. A study of molecular one-electron properties in terms of localized molecular orbital components. International Journal of Quantum Chemistry, 12 (1):133–143, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gruninger:1973:EPQ

[GCJ73] John Gruninger, William Clements, and Steven Jaworowicz. Effective principal quantum numbers and integral transform functions. International Journal of Quantum Chemistry, 7(S7): 103–109, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Gresh:1979:IIR

[GCP79] N. Gresh, P. Claverie, and A. Pullman. Intermolecular interactions: Reproduction of the results of ab initio supermolecule computations by an additive procedure. International Journal of Quantum Chemistry, 16(S13):243–253, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Garrod:1976:DMV

[GF76] Claude Garrod and Mario A. Fusco. A density matrix variational calculation for atomic Be. International Journal of Quantum Chemistry, 10(3):495–510, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gaspar:1979:IMC

[GG79a] R. Gáspáar and R. Gáspár, Jr. Ab initio molecular calculations with pseudopotentials: calculations of double-zeta quality on BeH_{2}, BH_{3}, CH_{4}, and C_{2}H_{6}. International Journal of Quantum Chemistry, 15(6):559–566, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gaspar:1979:IMFb

[GG79b] R. Gáspár, Jr. and R. Gáspáar. Ab initio molecular fragment calculations with pseudopotentials. International Journal of Quantum Chemistry, 16(1):57–64, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gaspar:1979:IMFa

[GG79c] R. Gáspár, Jr. and R. Gáspáar. Ab initio molecular fragment calculations with pseudopotentials: Hydrocarbon calculations of double-zeta quality. International Journal of Quantum Chemistry, 15(6):567–578, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garritz:1979:SPM

[GGCK79] Andoni Garritz, José Luis Gázquez, Miguel Castro, and Jaime Keller. Space partitioning in multiple scattering techniques. I. Hydrogen molecular ion and hydrogen molecule. International Journal of Quantum Chemistry, 15(6):731–744, June 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Green:1975:TFL

[GGJ75] A. E. S. Green, R. H. Garvey, and C. H. Jackman. A Thomas– Fermi-like analytic independent particle model for atoms and ions. International Journal of Quantum Chemistry, 9(S9):43– 50, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Green:1977:AYS

[GGJ77] A. E. S. Green, R. H. Garvey, and C. H. Jackman. Analytic yield spectra for electrons on H_{2}. International Journal of Quantum Chemistry, 12(S11):97–103, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Grundke:1972:PDF

[GH72a] E. W. Grundke and D. Henderson. Pair distribution functions for hard-sphere mixtures. International Journal of Quantum Chemistry, 6(S6):383–385, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Guest:1972:LOC

[GH72b] M. F. Guest and I. H. Hillier. Localized orbital calculations of the bonding in SO, SO_{2}F^{2}, ClO_{3}F and SOCL_{2}. International Journal of Quantum Chemistry, 6(5):967–973, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garrod:1978:IFD

[GH78] Claude Garrod and J. Michael Hannon. Inequalities for fermion density matrices. International Journal of Quantum Chemistry, 13(1):125–141, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Govil:1979:OPB

[GH79] Girjesh Govil and Ramakrishna V. Hosur. Organization of phospholipids in biological membranes. International Journal of Quantum Chemistry, 16(1):19–29, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gunnarsson:1977:MCU

[GHJ77] O. Gunnarsson, J. Harris, and R. O. Jones. Molecular calculations using the muffin-tin orbital method. International Journal of Quantum Chemistry, 12(S11):71–74, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Goddard:1979:GDC

[GHS79] John D. Goddard, Nicholas C. Handy, and Henry F. Schaefer III. Generalization of the direct configuration interaction method to the Hartree–Fock interacting space for doublets, quartets, and open-shell singlets. International Journal of Quantum Chemistry, 16(S13):471, March 11–17, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Gianinetti:1973:OMO

[Gia73] Ermanno Gianinetti. Optimization of molecular orbitals: Easy parametrizations of little rotations suitable for functional iteration. International Journal of Quantum Chemistry, 7(5): 869–876, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gianturco:1976:SSP

[Gia76] F. A. Gianturco. On the scattering of slow protons by diatomic targets. International Journal of Quantum Chemistry, 10(1):37–56, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gingerich:1978:VBA

[Gin78] Karl A. Gingerich. Valence bond approach for calculating bond energies of diatomic intermetallic compounds with multiple bonds and comparison with experiment. International Journal of Quantum Chemistry, 14(S12):489–496, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Gunnarsson:1976:SDF

[GJ76] O. Gunnarsson and P. Johansson. The spin-density-functional formalism for quantum mechanical calculations: Test on diatomic molecules with an efficient numerical method. International Journal of Quantum Chemistry, 10(2):307–323, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gunnarsson:1975:CBS

[GJLL75] O. Gunnarsson, P. Johansson, S. Lundqvist, and B. I. Lundqvist. Chemical binding in small molecules by the spindensity-functional formalism. International Journal of Quantum Chemistry, 9(S9):83–93, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Geller:1979:ADA

[GJP79] M. Geller, A. Jaworski, and A. Pohorille. Atomic dipole approximation and energies of interactions between purine and pyrimidine bases. I. Electrostatic interactions of adenine with uracil, thymine, thiouracils, dihydrouracil, and 5-fluorouracil. International Journal of Quantum Chemistry, 15(4):369–377, April 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gubanov:1975:XRE

[GK75] V. A. Gubanov and E. Z. Kurmaev. X-ray emission spectra and electronic structure of TiC, TiN, and TiO. International Journal of Quantum Chemistry, 9(S9):297–310, January 19– 25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Green:1976:SOA

[GK76] A. E. S. Green and G. J. Kutcher. Some observations on atomic independent-particle models. International Journal of Quantum Chemistry, 10(S10):135–140, January 19–24, 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gilardi:1978:FST

[GK78] Richard Gilardi and Jerome Karle. Fibrous structures: Their character and the role of interatomic forces in their analysis.

International Journal of Quantum Chemistry, 14(S5):103–126, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Guerillot:1974:ECS

[GL74] C. R. Guerillot and R. Lissillour. Electronic correlation studies. II. Self-correlated field method. Application to ground state and first^{1}P,^{3}P excited states of two-electron atomic systems. International Journal of Quantum Chemistry, 8(6):825–838, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gianturco:1979:EMS

[GLS79] F. A. Gianturco, U. T. Lamanna, and S. Salvini. Electron– molecule scattering at low and intermediate energies: A Glauber–eikonal approximation for simple targets. International Journal of Quantum Chemistry, 16(S13):579–600, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Golebiewski:1973:MCG

[GM73a] A. Gołebiewski and J. Mrozek. Modified Cartesian Gaussian functions and their use in quantum chemistry. International Journal of Quantum Chemistry, 7(5):1021–1028, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Golebiewski:1973:MGF

[GM73b] A. Gołebiewski and J. Mrozek. Modified Gaussian functions and their use in quantum chemistry. I. Integrals. International Journal of Quantum Chemistry, 7(3):623–634, May 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Graovac:1975:CFT

[GMG75] Ante Graovac, Hendrik J. Monkhorst, and M. L. Glasser. Computation of Fourier transform quantities in Hartree–Fock calculations for simple crystals. International Journal of Quantum Chemistry, 9(2):243–259, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Guenzburger:1977:ESC

[GML77] Diana Guenzburger, Bruno Maffeo, and Sven Larsson. Electronic structure calculation of [Fe(CN)_{6}]^{−3} by the SCF–MS Xα method: Study of optical transitions, X-ray photoelectron spectrum, and Mössbauer isomer shifts. International Journal of Quantum Chemistry, 12(2):383–396, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Graovac:1973:SOM

[GMZ73] Ante Graovac, Hendrik J. Monkhorst, and Tomislav Zivkovic. Slater orbital molecular integrals with numerical Fourier transform methods. I. (coplanar) multicenter exchange integrals over 1s orbitals. International Journal of Quantum Chemistry, 7 (2):233–251, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gardre:1977:EMD

[GN77] S. R. Gardre and P. T. Narasimhan. Electron momentum distributions from valence-bond wave functions. International Journal of Quantum Chemistry, 12(1):105–113, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1968:CEP

[GÖ68] Osvaldo Goscinski and Yngve Öhrn. Coupling of equivalent particles in a field of given symmetry. International Journal of Quantum Chemistry, 2(6):845–856, November 1968. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See errata [GÖJ^{+}73].

Goscinski:1973:ECE

[GÖJ^{+}73] O. Goscinski, Y. Öhrn, A. Norman Jette, E. Lombardi, R. Ritter, and L. Jansen. Errata: Coupling of Equivalent Particles in a Field of Given Symmetry. International Journal of Quantum Chemistry, 7(5):1039–1043, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [GÖ68].

Gazquez:1979:EGM

[GOK79] José Luis Gázquez, Elba Ortíz, and Jaime Keller. Electron gas model for inhomogeneous systems. International Journal of Quantum Chemistry, 16(S13):377–385, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Gold:1970:SEP

[Gol70] Harvey J. Gold. Specific enhancement of proton tunneling: A possible factor in enzyme catalysis. International Journal of Quantum Chemistry, 5(S4):353–354, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Golden:1973:TDQ

[Gol73] Sidney Golden. Time-dependent quantum statistical inequalities. International Journal of Quantum Chemistry, 7(S7):577– 584, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Golebiewski:1979:NIS

[Goł79] A. Gołebiewski. New iterative scheme for a simultaneous calculation of m first eigenstates of a real symmetric matrix. International Journal of Quantum Chemistry, 15(6):693–700, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goodisman:1976:VET

[Goo76] Jerry Goodisman. Valence-electron theory using the Thomas– Fermi–Dirac model: Application to the K atom and the K^{−} ion. International Journal of Quantum Chemistry, 10(2):341– 358, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gopinathan:1976:HTS

[Gop76] M. S. Gopinathan. On the hypervirial theorem and the scaling problem. International Journal of Quantum Chemistry, 10(6): 975–983, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gora:1971:FCF

[Gor71] Edwin K. Gora. On formulas in closed form for Van Vleck expansions. International Journal of Quantum Chemistry, 5(S5): 773, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Gora:1972:FCF

[Gor72] Edwin K. Gora. On formulas in closed form for Van Vleck expansions. International Journal of Quantum Chemistry, 6 (4):681–700, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1975:TOD

[Gos75] O. Goscinski. On transition operators for describing ionization and excitation. International Journal of Quantum Chemistry, 9(S9):221–227, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Goscinski:1978:BRBb

[Gos78a] Osvaldo Goscinski. Book review: Principles of Ultraviolet Photoelectron Spectroscopy. J. W. Rabalais. Wiley-Interscience Monographs in Chemical Physics, New York, 1977. Price: $22.50. International Journal of Quantum Chemistry, 13(2): 263, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Rab77].

Goscinski:1978:BRBa

[Gos78b] Osvaldo Goscinski. Book review: Quanta: A handbook of concepts. P. W. Atkins. Oxford chemistry series, 1974. price: £3.50 (paperback), £6.50 (hardcover). International Journal of Quantum Chemistry, 13(1):159, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Atk74, Atk91].

Gouyet:1973:CAN

[Gou73] J. F. Gouyet. Le champ autocoherent non hermitien. (French) [The non-Hermitian self consistent field]. International Journal of Quantum Chemistry, 7(2):139–153, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gilles:1972:ISG

[GP72] J.-M. F. Gilles and J. Philippot. Internal symmetry groups of non-rigid molecules. International Journal of Quantum Chemistry, 6(2):225–261, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

George:1974:QMD

[GP74] C. George and I. Prigogine. Quantum mechanics of dissipative systems and noncanonical formalism. International Journal of Quantum Chemistry, 8(S8):335–346, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Gilles:1978:SNM

[GP78] J. M. F. Gilles and J. Philippot. Symmetry of nonrigid molecules and isomerization processes. International Journal of Quantum Chemistry, 14(3):299–311, September 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1979:END

[GP79] O. Goscinski and A. Palma. Electron and nuclear density matrices and the separation of electronic and nuclear motion. International Journal of Quantum Chemistry, 15(2):197–205, February 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gubanov:1971:EAS

[GPC^{+}71] V. A. Gubanov, L. A. Pereliaeva, A. K. Chirkov, G. N. Yastchenko, and R. O. Matevosian. Electronic absorption spectra for free hydrazyl radicals, their ionic forms and initial hydrazines. International Journal of Quantum Chemistry, 5(5): 513–524, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Giessner-Prettre:1973:MOC

[GPP73] C. Giessner-Prettre and B. Pullman. Molecular orbital calculations of the long-range proton-proton coupling constants in purine and pyrimidine nucleosides. International Journal of Quantum Chemistry, 7(S7):295–305, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Goldblum:1979:UOM

[GPP79] A. Goldblum, D. Perahia, and A. Pullman. Use of the overlap multipole expansion for approximating molecular electrostatic potentials. International Journal of Quantum Chemistry, 15 (1):121–129, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grau:1977:ABL

[Gra77] D. Grau. An application of the Brillouin–Löwdin perturbation expansion: Lower bounds to the energy eigenvalues of the rigid rotator in an electric field. International Journal of Quantum Chemistry, 11(6):931–941, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goodall:1974:ICN

[GRB74] Marcus C. Goodall, William O. Romine, and Ronald J. Bradley. Ion channels and neurotransmitters. International Journal of Quantum Chemistry, 8(S1):233–243, January 17–19, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Grein:1971:ASE

[Gre71] Friedrich Grein. All single excitations in multiconfigurational wave functions. International Journal of Quantum Chemistry, 5(S5):165–167, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Garduno:1977:PPB

[GRE^{+}77] Ramon Garduño, Robert Rein, John T. Egan, Yves Coeckelenbergh, and Robert D. Macelroy. Purine-purine base pairs and the origin of transversion-type mutation. International Journal of Quantum Chemistry, 12(S4):197–204, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gupta:1975:CHF

[GRM75] A. Gupta, H. P. Roy, and P. K. Mukherjee. Coupled Hartree– Fock calculation of static and shielding factors for open shell atomic systems. International Journal of Quantum Chemistry, 9(1):1–8, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gruninger:1970:RAF

[Gru70] John Gruninger. Radial auxiliary functions for many electron integrals. International Journal of Quantum Chemistry, 5(S4): 385–389, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Gombas:1970:TTL

[GS70a] P. Gombás and T. Szondy. The theoretical treatment of the Li_{2}-molecule. International Journal of Quantum Chemistry, 4 (6):603–612, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1970:SPT

[GS70b] O. Goscinski and N. Stepanov. Successive partitioning technique and Feenberg perturbation theory. International Journal of Quantum Chemistry, 4(6):545–559, November 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gelbart:1978:SSA

[GSE78] William M. Gelbart, Paul R. Stannard, and Mark L. Elert. Some symmetry aspects of the local-mode description of vibrational structure. International Journal of Quantum Chemistry, 14(5):703–708, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Green:1975:QMB

[GT75] H. S. Green and T. Triffet. Quantum mechanics and the brain. International Journal of Quantum Chemistry, 9(S2):289–296, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

George:1977:CEC

[GTBB77] Philip George, Mendel Trachtman, Alistair M. Brett, and Charles W. Bock. Correlation energy contributions to reaction heats. International Journal of Quantum Chemistry, 12(1): 61–81, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grabenstetter:1976:GGS

[GTG76] J. E. Grabenstetter, T. J. Tseng, and F. Grein. Generation of genealogical spin eigenfunctions. International Journal of Quantum Chemistry, 10(1):143–149, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Guberman:1979:PCD

[Gub79] Steven L. Guberman. Potential curves for dissociative recombination of O_{2}^{+}. International Journal of Quantum Chemistry, 16(S13):531–540, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Gervy:1977:SMI

[GV77] D. Gervy and G. Verhaegen. Simplified methods for ab initio calculations. The valence states of CH_{2} and CH. International Journal of Quantum Chemistry, 12(1):115–131, July 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goodisman:1975:MPT

[GY75] J. Goodisman and K. Yonei. Minimum principles for the Thomas–Fermi–Dirac density. International Journal of Quantum Chemistry, 9(S9):59–66, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Garrett:1972:TSDa

[GZ72a] Thomas B. Garrett and Daniel Zeroka. Theoretical studies of diamagnetic properties of the hydrogen molecule ion. I. Approximate variation-perturbation calculation. International Journal of Quantum Chemistry, 6(4):651–661, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garrett:1972:TSDb

[GZ72b] Thomas B. Garrett and Daniel Zeroka. Theoretical studies of diamagnetic properties of the hydrogen molecule ion II. Effect of varying internuclear separation. International Journal of Quantum Chemistry, 6(4):663–668, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garrett:1973:E

[GZMM73] Thomas B. Garrett, Daniel Zeroka, Paul McGuire, and David A. Micha. Errata. International Journal of Quantum Chemistry, 7(3):635–636, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hackmeyer:1974:GIM

[Hac74] M. Hackmeyer. A general ab initio molecular multiconfiguration self-consistent field algorithm. International Journal of Quantum Chemistry, 8(5):783–788, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Halevi:1971:PDB

[Hal71] E. A. Halevi. Polarization differences between isotopic molecules. International Journal of Quantum Chemistry, 5 (S5):367–370, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Hall:1973:GIG

[Hal73] G. G. Hall. Gauge invariant Gaussian orbitals and the ab initio calculation of diamagnetic susceptibility for molecules. International Journal of Quantum Chemistry, 7(1):15–25, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hall:1975:CBE

[Hal75a] G. G. Hall. Charges and bond energies in the analysis of quantum chemistry calculations. International Journal of Quantum Chemistry, 9(S9):279–284, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Hall:1975:SMS

[Hal75b] Michael B. Hall. A semiquantitative model of spin-orbit coupling in doublet states and its application to the photoelectron spectra of diatomic halogens. International Journal of Quantum Chemistry, 9(S9):237–243, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Hall:1978:GMO

[Hal78] Michael B. Hall. Generalized-molecular-orbital theory: Simple multiconfiguration self-consistent-field method. International Journal of Quantum Chemistry, 14(5):613–621, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hall:1979:WCQ

[Hal79a] George G. Hall. What can quantum chemistry contribute to biology? International Journal of Quantum Chemistry, 16(1): 51–56, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hall:1979:GMO

[Hal79b] Michael B. Hall. Generalized molecular orbital theory: Ground state and ionization potentials of water and dinitrogen. International Journal of Quantum Chemistry, 16(S13):195–203, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Ham:1971:JTE

[Ham71] Frank S. Ham. The Jahn–Teller effect. International Journal of Quantum Chemistry, 5(S5):191–199, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Handler:1970:MTF

[Han70a] George S. Handler. 2-matrices from Thomas–Fermi theory. International Journal of Quantum Chemistry, 5(S4):359–361, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Hansen:1970:ATD

[Han70b] Aage E. Hansen. Alternative time-dependent Hamiltonians by infinitesimal transformations. International Journal of Quantum Chemistry, 5(S4):473–478, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Handler:1975:HSL

[Han75] George S. Handler. Hermitian symmetry and low-order statistical density matrices. International Journal of Quantum Chemistry, 9(S9):35–42, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Handler:1977:LRC

[Han77a] George S. Handler. Linear response correction to the kinetic energy from Thomas–Fermi theory. A simple model. International Journal of Quantum Chemistry, 12(S11):397–401, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Hanson:1977:WR

[Han77b] Harold P. Hanson. Welcoming remarks. International Journal of Quantum Chemistry, 12(S11):15–16, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Hanson:1979:ALE

[Han79] Louise Karle Hanson. Axial ligand effects on iron and manganese porphyrins: Extended Hückel calculations of cyt P450 analogs and of O_{2} binding to iron and manganese. International Journal of Quantum Chemistry, 16(S6):73–87, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Harcourt:1970:VSI

[Har70a] R. D. Harcourt. Valence structures for N_{2}O_{4}. III. International Journal of Quantum Chemistry, 4(2):173–189, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harriman:1970:DMA

[Har70b] John E. Harriman. A density matrix approach to multiconfiguration calculations. International Journal of Quantum Chemistry, 5(S4):363–370, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Harcourt:1971:IVF

[Har71a] R. D. Harcourt. Increased-valence formulae and the bonding of oxygen to haemoglobin. International Journal of Quantum Chemistry, 5(5):479–495, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harriman:1971:DMA

[Har71b] J. E. Harriman. A density matrix approach to multiconfiguration calculations. International Journal of Quantum Chemistry, 4(??):363–??, ?? 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harrison:1971:CZF

[Har71c] James F. Harrison. The calculation of the zero field splitting parameters in carbenes. International Journal of Quantum Chemistry, 5(S5):285–293, January 18–23, 1971. CODEN IJQCB2.

ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Harmison:1975:BST

[Har75] Charles R. Harmison. Biophysical studies on two conformations of a human plasma lipoprotein. International Journal of Quantum Chemistry, 9(S2):117–128, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Harcourt:1977:VFF

[Har77a] R. D. Harcourt. Valence formulas for the Fe(II)O_{2} linkage of oxyhemoglobin and cytochrome P-450-dependent mono– oxygenases. International Journal of Quantum Chemistry, 12 (S4):143–153, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Harriman:1977:CGM

[Har77b] John E. Harriman. A comparison of geometric methods with other methods for the characterization of density matrices. International Journal of Quantum Chemistry, 12(6):1039–1048, December 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harris:1977:CCM

[Har77c] Frank E. Harris. Coupled-cluster method for excitation energies. International Journal of Quantum Chemistry, 12(S11): 403–411, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Harris:1978:HLM

[Har78] Frank E. Harris. Hartree-like methods in electronic structure theory. International Journal of Quantum Chemistry, 13(2): 189–198, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harriman:1979:GDM

[Har79a] John E. Harriman. Geometry of density matrices. III. Spin components. International Journal of Quantum Chemistry, 15 (6):611–643, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harris:1979:RON

[Har79b] J. Harris. The role of occupation numbers in HKS theory. International Journal of Quantum Chemistry, 16(S13):189–193, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Hatch:1975:SSS

[Hat75] Dorian M. Hatch. Selected stable states in large quantum systems. International Journal of Quantum Chemistry, 9(4):649– 655, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hannesson:1979:FNM

[HB79] Thorsteinn Hannesson and S. M. Blinder. Finite nucleus model for the Fermi contact interaction. International Journal of Quantum Chemistry, 15(1):7–14, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hadinger:1977:MET

[HBB77] G. Hadinger, N. Bessis, and G. Bessis. Matrix elements and their selection rules from ladder operator considerations. International Journal of Quantum Chemistry, 11(3):451–472, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hehenberger:1978:WTS

[HBE78] Michael Hehenberger, Erkki Brändas, and Nils Elander. Weyl’s theory for a system of coupled second-order differential equations. International Journal of Quantum Chemistry, 14(S12): 67–71, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Harrison:1972:AMC

[HC72] R. W. Harrison and L. C. Cusachs. Alternate method of choosing α in Herman–Skillman HFS program. International Journal of Quantum Chemistry, 6(S6):217–219, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Houjer:1978:SAM

[HC78] Germund Höujer and Johnson Chung. Some aspects of the model potential method. International Journal of Quantum Chemistry, 14(5):623–634, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Henkel:1975:PCF

[HCIE75] J. H. Henkel, T. C. Collins, J. L. Iveyt, and R. N. Euwema. Polarization calculations for ferroelectric sodium nitrite. International Journal of Quantum Chemistry, 9(S9):529–533, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Hobza:1979:REF

[HČZ79] P. Hobza, P. Čársky, and R. Zahradník. Role of entropy in formation of van der Waals complexes. International Journal of Quantum Chemistry, 16(2):257–260, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Horing:1971:CLP

[HDG71] Norman J. Morgenstern Horing, R. W. Danz, and M. L. Glasser. Coupled longitudinal phonon-plasmon spectrum in quantizing magnetic field. International Journal of Quantum Chemistry, 5(S5):765–769, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Hehenberger:1977:DRE

[HE77] Michael Hehenberger and Nils Elander. Determination of resonance energies and nonradiative lifetimes of rotationally predissociating diatomic molecules: The B^{2}Σ^{−} state of CH. International Journal of Quantum Chemistry, 12(S11):289–293, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Hearn:1971:FDR

[Hea71] Dwight D. Hearn. Frequency distribution of radiation emitted from excited atomic systems. International Journal of Quantum Chemistry, 5(5):565–574, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hendekovic:1974:MCM

[Hen74] J. Hendeković. Method of complex molecular orbitals. International Journal of Quantum Chemistry, 8(5):799–816, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hendekovic:1976:CMO

[Hen76] Josip Hendeković. Complex molecular orbital method: Openshell theory. International Journal of Quantum Chemistry, 10 (6):1025–1035, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hendekovic:1977:HFS

[Hen77] J. Hendeković. Hermitian form of the secular equation in the complex molecular orbitals method. International Journal of Quantum Chemistry, 12(4):781–783, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Henderson:1978:HCP

[Hen78] George A. Henderson. Highly correlated particle densities and idempotent one-densities. International Journal of Quantum Chemistry, 13(1):143–148, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Henry:1979:LEE

[Hen79] Ronald J. W. Henry. Low-energy-electron–diatomic molecule scattering. International Journal of Quantum Chemistry, 16 (S13):633–644, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Herndon:1974:QTA

[Her74] William C. Herndon. Quantum theory of aromatic hydrocarbon carcinogenesis. International Journal of Quantum Chemistry, 8(S1):123–134, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Heller:1972:EDN

[HF72] Donald F. Heller and Karl F. Freed. Energy dependence of nonradiative decay in polyatomic molecules. International Journal of Quantum Chemistry, 6(S6):267–277, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Heidrich:1975:QUM

[HG75] D. Heidrich and M. Grimmer. Quantenchemische Untersuchungen zum Mechanismus der elektrophilen Substitution. I. Zur Potentialhyperfläche des systems Benzol/H^{+}. (German) [Quantum chemical studies on the mechanism of electrophilic substitution. I. potential hypersurface of the benzene/H^{+} system]. International Journal of Quantum Chemistry, 9(5):923–940, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hedinger:1978:YSP

[HG78] R. A. Hedinger and A. E. S. Green. Yield spectrum for protons impacting on helium. International Journal of Quantum Chemistry, 14(S12):377–387, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Haley:1977:CME

[HH77] Lawrence V. Haley and Hendrik F. Hameka. Calculation of molecular electric polarizabilities and dipole moments. II. The LiH molecule. International Journal of Quantum Chemistry, 11(5):733–741, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hall:1970:MPM

[HHR70] G. G. Hall, J. Hyslop, and D. Rees. A minimum principle for molecular systems allowing the use of discontinuous wave functions. International Journal of Quantum Chemistry, 4(1): 5–20, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hilal:1979:CSS

[Hil79] Rifaat Hilal. Closed-shell SCF–CI MO treatment of all-reactive electrons in complex heteroorganic molecules. International Journal of Quantum Chemistry, 15(1):37–47, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hinchliffe:1977:ESP

[Hin77] Alan Hinchliffe. The electronic structure and properties of pyrazine, its cation, and anion. International Journal of Quantum Chemistry, 11(5):767–773, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hirschfelder:1970:PEP

[Hir70] Joseph O. Hirschfelder. Primitive exchange perturbation theory, IV. International Journal of Quantum Chemistry, 5(S4): 257–269, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Hojer:1974:SNP

[HK74] Germund Höjer and Jaime Keller. A short note on the possible stabilization of the local water structure in the liquid due to an oxygen or a nitrogen molecule. International Journal of Quantum Chemistry, 8(S8):177–178, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Harris:1971:EHF

[HKM71] Frank E. Harris, Lalit Kumar, and Hendrik J. Monkhorst. “exact” Hartree–Fock results for atomic hydrogen crystals. International Journal of Quantum Chemistry, 5(S5):527–531, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Hariharan:1979:EMP

[HKP79] P. C. Hariharan, Joyce J. Kaufman, and Carlo Petrongolo. Electrostatic molecular potential contour maps generated from ab-initio MODPOT/VRDDO/MERGE wave functions of carcinogenic benzo(a)pyrene and its metabolites. International Journal of Quantum Chemistry, 16(S6):223–257, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Handler:1978:ICE

[HL78a] George S. Handler and Everett W. Larson. The information content of an exact eigenfunction of a physical Hamiltonian. International Journal of Quantum Chemistry, 14(S12):405–406, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Hopkinson:1978:IMO

[HL78b] Alan C. Hopkinson and Min H. Lien. An ab initio molecular orbital study of the deprotonation of amines. International Journal of Quantum Chemistry, 13(3):349–366, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Halkjaer:1979:PCN

[HL79] Ole Halkjaer and Jan Linderberg. Potential curves and nonadiabatic coupling matrix elements for the O^{+}–Ne system. International Journal of Quantum Chemistry, 16(S13):475–484, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Harrison:1973:SMO

[HLGC73] R. W. Harrison, H. J. Lader, L. P. Gary, and L. C. Cusachs. Semiempirical molecular orbital calculations: The ionization potentials of HgI_{2}, the orbitals of the heavier elements, and the approximation of electron repulsion integrals. International Journal of Quantum Chemistry, 7(S7):307–310, January 21– 27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Hardisson:1977:IDT

[HLMV77] A. Hardisson, R. Lefebvre, F. Mauricio, and S. Velasco. The interference in the decay of two \frac{1}{2} spins in a molecular medium, studied by the Nakajima–Zwanzig technique. International Journal of Quantum Chemistry, 11(2):301–315, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hopkinson:1977:NEM

[HLYC77] Alan C. Hopkinson, Min H. Lien, Keith Yates, and Imre G. Csizmadia. A non-empirical molecular orbital study of valence tautomers of C_{2}H_{3}N. International Journal of Quantum Chemistry, 12(2):355–368, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Huzinaga:1971:CVS

[HMC71] S. Huzinaga, D. Mcwilliams, and A. A. Cantu. A comment on the V state of ethylene. International Journal of Quantum Chemistry, 5(S5):119–121, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Hardisson:1977:ALS

[HMV77] A. Hardisson, F. Mauricio, and S. Velasco. Absorption line shape of two coupled oscillators decaying in two different mediums. International Journal of Quantum Chemistry, 11(6):979– 989, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hardisson:1978:ALS

[HMV78] A. Hardisson, F. Mauricio, and S. Velasco. Absorption line shape of two coupled oscillators decaying in the same bath. International Journal of Quantum Chemistry, 14(1):65–69, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hatano:1978:CES

[HNT78] Yasuyo Hatano, Tsutomu Nomura, and Kiyoshi Tanaka. Calculation of excited states of H_{2}O and NH_{3} by the one-center expansion approximation. International Journal of Quantum Chemistry, 13(2):207–220, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ha:1973:PPP

[HO73] Tae-Kyu Ha and Chester T. O’Konski. Properties of pyridine and pyrazine from ab initio molecular orbital wave functions. International Journal of Quantum Chemistry, 7(3):609– 622, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hojer:1979:TPEa

[Höj79a] Germund Höjer. Two-parameter exponential-type basis functions for atomic calculations. International Journal of Quantum Chemistry, 15(4):389–401, April 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hojer:1979:TPEb

[Höj79b] Germund Höjer. Two-parameter exponential-type basis functions for atomic calculations. II. International Journal of Quantum Chemistry, 16(S13):77–80, March 11–17, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Holtje:1977:IEC

[Höl77] Hans-Dieter Höltje. Interaction energy calculations on interactions between pharmacon models and ionic receptor site models. International Journal of Quantum Chemistry, 12(S4):245– 252, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Hopper:1978:ESH

[Hop78] Darrel G. Hopper. The electronic structure of HeH_{2}^{+}. International Journal of Quantum Chemistry, 14(S12):305–322, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Horing:1971:NPM

[Hor71] Norman J. Morgenstern Horing. Nondegenerate plasma modes in quantizing magnetic field. International Journal of Quantum Chemistry, 5(S5):763, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Hosoya:1972:SRO

[Hos72] Haruo Hosoya. Studies on Rydberg orbitals. IV. Basic formulas for the one-electron perturbation calculation of molecular Rydberg excited states. International Journal of Quantum Chemistry, 6(5):801–817, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hosur:1978:QMT

[Hos78] Ramakrishna V. Hosur. Quantum-mechanical theory for transport across biological membranes. International Journal of Quantum Chemistry, 13(3):411–428, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harter:1977:BCC

[HP77a] William G. Harter and Chris W. Patterson. Bands, clusters, and crystal field splitting: Understanding SF_{6} rotational levels. International Journal of Quantum Chemistry, 12(S11):479– 492, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Huang:1977:FSG

[HP77b] Hsing Hua Huang and Ali H. Pakiari. Floating spherical Gaussian orbital open-shell calculations on the four-electron H_{4} system. International Journal of Quantum Chemistry, 12(4):593– 608, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Heaton:1972:CIC

[HPK72] Marie M. Heaton, Aaron Pipano, and Joyce J. Kaufman. Configuration interaction calculations on O_{3} and O_{3}^{−}. International Journal of Quantum Chemistry, 6(S6):181–186, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Hariharan:1979:NIM

[HPK79] P. C. Hariharan, Herbert E. Popkie, and Joyce J. Kaufman. Nonempirical ab initio MODPOT, VRDDO, and MODPOT/ VRDDO calculations. X. The attack of the simplest ultimate carcinogen, CH_{3}^{+}, on guanine by a MERGE technique and a possible fundamental difference between methylating versus ethylating ultimate carcinogens. International Journal of Quantum Chemistry, 16(S13):255–265, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Hofmann:1979:TCA

[HPW79] Hans-Jörg Hofmann, Gustav Peinel, and Thilo Weller. Theoretical conformational analysis considering solvent effects. International Journal of Quantum Chemistry, 16(2):379–386, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hyslop:1976:ISE

[HR76] J. Hyslop and D. Rees. An interative scheme for electronic systems using one-electron Green’s functions. International Journal of Quantum Chemistry, 10(4):669–681, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harcourt:1979:VBS

[HR79] Richard D. Harcourt and Walter Roso. Valence-bond studies of four-electron three-center bonding units. II. FNO, HNO, LiNO, LiON, and HF_{2}^{−}. International Journal of Quantum Chemistry, 16(5):1033–1049, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Herman:1976:OCD

[HRBB76] F. Herman, W. E. Rudge, I. P. Batra, and B. I. Bennett. Orbital charge distributions for the TCNQ and TTF molecules. International Journal of Quantum Chemistry, 10(S10):167–181, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Hillier:1970:ICU

[HS70a] I. H. Hillier and V. R. Saunders. Ab initio calculations, using a small Gaussian basis set, of the electronic structure of the sulphate ion. International Journal of Quantum Chemistry, 4 (2):203–215, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hillier:1970:NSP

[HS70b] I. H. Hillier and V. R. Saunders. A new SCF procedure and its applications to ab initio calculations of the states of the fluorosulphate radical. International Journal of Quantum Chemistry, 4(5):503–518, September 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hankey:1971:AFS

[HS71] Alex Hankey and H. Eugene Stanley. An alternate formulation of the static scaling hypothesis. International Journal of Quantum Chemistry, 5(S5):593–604, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Harriss:1975:SLS

[HS75] D. K. Harriss and I. G. Solev. On the solution of the leastsquares local energy variance minimization equations. International Journal of Quantum Chemistry, 9(6):975–980, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hameka:1976:ECF

[HS76] Hendrik F. Hameka and E. Nørby Svendsen. Errors in the calculations of first-order and second-order energy perturbations. International Journal of Quantum Chemistry, 10(2):249–259, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hameka:1977:VMC

[HS77] Hendrik F. Hameka and E. Nørby Svendsen. A variational method for calculating dynamic polarizabilities. International Journal of Quantum Chemistry, 11(1):129–135, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Handler:1979:DFT

[HS79] George S. Handler and Darwin W. Smith. Density functional theory of quantum mechanical systems. II. Accurate potential and energy functionals and their representations. International Journal of Quantum Chemistry, 16(S13):111–114, March 11– 17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Harrison:1978:HIP

[HSBD78] S. W. Harrison, S. Swaminathan, David L. Beveridge, and Robert Ditchfield. Heuristic intermolecular potential function for the methane–water interaction based on ab initio quantummechanical calculations. International Journal of Quantum Chemistry, 14(3):319–327, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Horino:1975:ESE

[HT75] Yasuyo Horino and Hiroshi Tatewaki. The electronic structure of the excited states for B, C^{+}, and C arising from the configuration 1s^{2}2s2p^{n} — the effect of removal of the “equivalence” restriction. International Journal of Quantum Chemistry, 9 (2):287–299, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Howat:1977:DSO

[HTG77a] G. Howat, M. Trsic, and O. Goscinski. Differential scaling of σ and π orbitals, and the geometric approximation for molecular polarizabilities. International Journal of Quantum Chemistry, 12(5):963–965, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Howat:1977:GAM

[HTG77b] G. Howat, M. Trsic, and O. Goscinski. Geometric approximation for molecular polarizabiities. International Journal of Quantum Chemistry, 11(2):283–292, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Henkel:1974:LDN

[HU74] J. H. Henkel and C. A. Uzes. Level-dependent new Tamm– Dancoff calculations. International Journal of Quantum Chemistry, 8(S8):73–79, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Huber:1971:CDS

[Hub71] D. L. Huber. Critical dynamics of the staggered magnetization in uniaxial antiferromagnets. International Journal of Quantum Chemistry, 5(S5):667–671, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Hudson:1972:A

[Hud72] R. L. Hudson. Announcement. International Journal of Quantum Chemistry, 6(3):599, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hunter:1974:CPA

[Hun74] Geoffrey Hunter. Conditional probability amplitude analysis of coupled harmonic oscillators. International Journal of Quantum Chemistry, 8(S8):413–420, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Hunter:1975:CPA

[Hun75a] Geoffrey Hunter. Conditional probability amplitudes in wave mechanics. International Journal of Quantum Chemistry, 9 (2):237–242, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hunter:1975:IPC

[Hun75b] Geoffrey Hunter. Ionization potentials and conditional amplitudes. International Journal of Quantum Chemistry, 9(S9): 311–315, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Hurley:1976:ECS

[Hur76a] Andrew Crowther Hurley. Electron correlation in small molecules, volume 7 of Theoretical chemistry, a series of monographs. Academic Press, New York, USA, 1976. ISBN 0-12-362450-9. viii + 276 pp. LCCN QD461 .H958.

Hurley:1976:IET

[Hur76b] Andrew Crowther Hurley. Introduction to the electron theory of small molecules. Academic Press, New York, USA, 1976. ISBN 0-12-362460-6. x + 329 pp. LCCN QD461 .H96.

Hurley:1977:GIP

[Hur77] A. C. Hurley. Group integration, projected basis functions, and correlation rules for linear molecules. International Journal of Quantum Chemistry, 12(S11):223–228, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Hirschfelder:1978:ECC

[HY78] Joseph O. Hirschfelder and Kuo-Ho Yang. Equations of change in configuration space. International Journal of Quantum Chemistry, 14(5):695–702, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hyman:1978:IAX

[HYL78] Arthur S. Hyman, Stuart I. Yaniger, and Joel F. Liebman. Interrelations among x-ray scattering, electron densities, and ionization potentials. International Journal of Quantum Chemistry, 14(6):757–766, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Imai:1974:DCD

[ICS74] Ren Imai, Tsung-Tai Chen, and Jack Simons. Direct calculation of density matrices: Natural orbitals and occupation numbers of model conjugated molecules. International Journal of Quantum Chemistry, 8(S8):323–333, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

IHaya:1973:MOR

[IM73] Y. J. I’Haya and F. Matsukawa. Magneto-optical rotation in molecules. V. The Verdet constant for the oxygen molecule. International Journal of Quantum Chemistry, 7(S7):181–184, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Ihaya:1979:TPG

[IMN79] Y. J. I’haya, Tetsuo Morikawa, and Junji Nakano. Theoretical prediction of g tensors and hyperfine coupling tensors of triplet molecules. International Journal of Quantum Chemistry, 16(S13):205–214, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

IHaya:1971:SJT

[INI71] Y. J. I’Haya, M. Nakayama, and T. Iwabuchi. Static Jahn– Teller distortion in the tri-phenylene mono-negative ion. International Journal of Quantum Chemistry, 5(S5):227–232, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

IHaya:1971:RQC

[INY^{+}71] Y. J. I’Haya, T. Nakamura, Y. Yagi, T. Sano, and H. Ito. Remark on quantum chemical interpretation of the induced optical activity of DNA-dye complexes. International Journal of Quantum Chemistry, 5(S5):361–365, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Ivanov:1975:PTD

[IS75] L. N. Ivanov and U. I. Safronova. Perturbation theory for degenerate states of atomic and molecular systems. International Journal of Quantum Chemistry, 9(4):711–719, July 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ivey:1974:LDE

[Ive74] Jerry L. Ivey. Local density exchange approximations versus Hartree–Fock for ground-state calculations. International Journal of Quantum Chemistry, 8(S8):117–122, January 20– 26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Ivanov:1979:CPD

[IZZ^{+}79] V. I. Ivanov, V. B. Zhurkin, S. K. Zavriev, Yu. P. Lysov, L. E. Minchenkova, E. E. Minyat, M. D. Frank-Kamenetskii, and A. K. Schyolkina. Conformational possibilities of double-helical nucleic acids: Theory and experiment. International Journal of Quantum Chemistry, 16(1):189–201, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jain:1970:SSP

[Jai70] D. C. Jain. A study of some potential energy functions for diatomic molecules. International Journal of Quantum Chemistry, 4(6):579–586, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jakubowski:1976:BOOa

[Jak76a] P. Jakubowski. Best optimized one-electron wave functions. I. The general procedure of optimization. International Journal of Quantum Chemistry, 10(5):719–731, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jakubowski:1976:BOOb

[Jak76b] P. Jakubowski. Best optimized one-electron wave-functions. II. Isoelectronic series of Li, Be, B, and C. International Journal of Quantum Chemistry, 10(5):733–737, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jakubowski:1976:BOOc

[Jak76c] P. Jakubowski. Best optimized one-electron wave-functions. III. Direct examination of optimization effectiveness. International Journal of Quantum Chemistry, 10(5):739–742, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jakubowski:1976:BOOd

[Jak76d] P. Jakubowski. Best optimized one-electron wave-functions. IV. Ionization energies of atoms. International Journal of Quantum Chemistry, 10(5):743–746, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jamieson:1970:TDH

[Jam70] M. J. Jamieson. Time-Dependent Hartree–Fock theory for atoms. International Journal of Quantum Chemistry, 5(S4): 103–115, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Jankowski:1976:CPS

[Jan76] K. Jankowski. A characterization of pairs of subspaces for quantum chemical applications of the Galerkin–Petrov method. International Journal of Quantum Chemistry, 10(4):683–697, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jansen:1977:PLS

[Jan77] L. Jansen. P.-O. Löwdin’s scientific and other activities. International Journal of Quantum Chemistry, 11(6):897–905, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jucys:1970:TTS

[JBG70] A. P. Jucys, A. A. Bandzaitis, and J. J. Grudzinskas. Theory of Two Shells of Atomic Electrons Using Non-Orthogonal Radial Orbitals. International Journal of Quantum Chemistry, 4(3): 341, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jeziorski:1976:FOP

[JBP76] Bogumił Jeziorski, Marek Bulski, and Lucjan Piela. First-order perturbation treatment of the short-range repulsion in a system of many closed-shell atoms or molecules. International Journal of Quantum Chemistry, 10(2):281–297, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jackels:1974:ERO

[JD74] Charles F. Jackels and Ernest R. Davidson. Equivalencerestricted open-shell SCF theory. International Journal of Quantum Chemistry, 8(5):707–714, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jaworski:1979:CNC

[JE79] Andrzej Jaworski and Irena Ekiel. Calculations of^{1}H NMR coupling constants for conformational studies of isomeric pentofuranosyl nucleosides. International Journal of Quantum Chemistry, 16(3):615–622, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jehle:1973:PCI

[Jeh73] Herbert Jehle. A program concerning intermolecular forces and tissue-specific cell contacts. International Journal of Quantum Chemistry, 7(S7):585–596, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Jette:1973:CME

[Jet73] A. Norman Jette. A convenient method for evaluating two center integrals. International Journal of Quantum Chemistry, 7(1):131–132, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jucys:1972:GSU

[JG72] A. P. Jucys and J. J. Grudzinskas. The general scheme of using non-orthogonal radial orbitals in a complex electronic configuration of the atom. International Journal of Quantum Chemistry, 6(3):455–464, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Johnson:1974:MOS

[JG74] Carl Lynn Johnson and Jack Peter Green. Molecular orbital studies on tryptamines active on the LSD receptor of the rat fundus strip. International Journal of Quantum Chemistry, 8 (S1):159–167, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Julg:1978:VNI

[JJ78] Andre Julg and Pierre Julg. Vers une nouvelle interprétation de la liaison chimique?. (French) [Toward a new interpretation of the chemical bond]. International Journal of Quantum Chemistry, 13(4):483–497, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jorgensen:1970:TDH

[JL70] Poul Jørgensen and Jan Linderberg. Time-dependent Hartree– Fock calculations in the Pariser–Parr–Pople model. Applications to aniline, azulene and pyridine. International Journal of Quantum Chemistry, 4(6):587–602, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jucys:1974:EHF

[JL74] A. P. Jucys and V. M. Lazauskas. Extended Hartree–Fock self-consistent field with some results for berylliumlike atomic systems. International Journal of Quantum Chemistry, 8(6): 839–855, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Josefson:1973:AMS

[JM73] Clarence M. Josefson and Boris Musulin. An accurate method for solution of one-electron perturbation equations. International Journal of Quantum Chemistry, 7(S7):117–122, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Johnson:1976:RAS

[JM76a] K. H. Johnson and R. P. Messmer. Recent applications of the SCF–Xα scattered-wave method to systems of catalytic and biocatalytic importance. International Journal of Quantum Chemistry, 10(S10):147–153, January 19–24, 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Julg:1976:HAM

[JM76b] Andre Julg and Francis Marinelli. Halogénures alcalins: un modèle incluant une délocalisation à courte distance: application à la stabilité relative des divers arrangements cristallins. (French) [Alkali halides: a model with a short delocalization distance: application to the relative stability of various crystalline arrangements]. International Journal of Quantum Chemistry, 10(6):1037–1047, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jankowski:1971:EEC

[JMM71] K. Jankowski, Gulzari Malli, and J. H. Maysilles. Effect of electron correlation on the spin–spin interaction in many-electron atomic systems. International Journal of Quantum Chemistry, 5(S5):643–652, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Julg:1978:DCD

[JMP78] André Julg, Francis Marinelli, and Alain Pellégatti. Determination of the charge distribution in nonmetallic crystals. I. Theory and application to tetrahedrally coordinated solids. International Journal of Quantum Chemistry, 14(2):181–190, August 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jankowski:1979:EPC

[JMP79] K. Jankowski, P. Malinowski, and M. Polasik. Electron pair correlation energies for ZN^{2+}. International Journal of Quantum Chemistry, 16(1):65–70, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jucys:1972:GTE

[JNŽ72] A. P. Jucys, E. P. Našlėnas, and P. S. Žvirblis. The general theory of the extended method of calculation of atomic structures. International Journal of Quantum Chemistry, 6(3):465– 472, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Johnson:1970:GSW

[Joh70] K. H. Johnson. Generalized scattered-wave approach to molecular-orbital theory. International Journal of Quantum Chemistry, 5(S4):153–164, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Johnson:1973:RAS

[Joh73] K. H. Johnson. Recent applications of the SCF–Xα scatteredwave method to complex molecules and solids. International Journal of Quantum Chemistry, 7(S7):347–351, January 21– 27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Johansen:1975:ICE

[Joh75] Helge Johansen. Ab initio calculation of the electronic structure of TCNQ and its ions. International Journal of Quantum Chemistry, 9(3):459–471, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Johnson:1977:SOE

[Joh77] K. H. Johnson. Spin-orbital electronegativity, the Xα method, and reactivity at transition-metal interfaces. International Journal of Quantum Chemistry, 12(S11):39–60, January 16– 22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Jolly:1979:IMH

[Jol79] Pratibha Jolly. Improved minimization for the Hylleraas sixparameter wave function. International Journal of Quantum Chemistry, 16(5):1149–1150, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jordan:1973:UUH

[Jor73] K. D. Jordan. Use of Unrestricted Hartree–Fock wave functions in RPA calculations. International Journal of Quantum Chemistry, 7(S7):491–504, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Jordan:1975:AAC

[Jor75] K. D. Jordan. Applications of analytic continuation in the construction of potential energy curves. International Journal of Quantum Chemistry, 9(S9):325–336, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Jorgensen:1977:OPD

[Jør77] Flemming Jørgensen. Orthonormalization, polar decomposition, and transformation to an effective Hamiltonian. International Journal of Quantum Chemistry, 12(2):397–426, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jorgensen:1978:OEF

[Jør78] Flemming Jørgensen. Orientation of the Eckart frame in a polyatomic molecule by symmetric orthonormalization. International Journal of Quantum Chemistry, 14(1):55–63, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Johnson:1973:GOO

[JP73] J. W. Johnson and R. D. Poshusta. Gaussian orbitals optimized for lower bounds of hydrogenic atoms. International Journal of Quantum Chemistry, 7(5):951–958, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Johnson:1977:LBM

[JP77] J. W. Johnson and R. D. Poshusta. Lower bound molecular orbitals for H, HeH^{++}, and H. International Journal of Quantum Chemistry, 11(5):885–894, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jorgensen:1978:SEM

[JPTM78] N. H. Jørgensen, P. B. Pedersen, E. W. Thulstrup, and Josef Michl. Semiempirical π-electron models for the calculation of MCD B terms for systems with approximate alternant pairing symmetry. MCD of biphenylene. International Journal of Quantum Chemistry, 14(S12):419–431, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Johansen:1974:ICE

[JR74] Helge Johansen and Björn Roos. Ab initio calculation of the electronic structure and geometry of Ni(CH_{3})_{2}. International Journal of Quantum Chemistry, 8(S8):137–143, January 20– 26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Jucys:1970:ILD

[JS70] A. P. Jucys and V. J. Stasiukaitis. Inter-level differences in iron group atomic systems in the extended method of calculation. International Journal of Quantum Chemistry, 4(3):333– 335, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Johnson:1971:SMO

[JS71] K. H. Johnson and F. C. Smith, Jr. SCF molecular-orbital studies of the sulphate ion by the scattered-wave model. International Journal of Quantum Chemistry, 5(S5):429–441, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Jaszunski:1977:ATC

[JS77] Michal Jaszuński and Andrzej J. Sadlej. Analytic techniques for the coupled multiconfiguration SCF perturbation theory. International Journal of Quantum Chemistry, 11(2):233–245, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jeziorski:1978:SFC

[JSC78] Bogumił Jeziorski, Krzysztof Szalewicz, and Grzegorz Chałasiński. Symmetry forcing and convergence properties of perturbation expansions for molecular interaction energies. International Journal of Quantum Chemistry, 14(3):271–287, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jucys:1975:NRO

[JSK75a] A. P. Jucys, K. V. Sabas, and Z. J. Kupliauskis. Nonorthogonal radial orbitals for the configuration l^{n}l. International Journal of Quantum Chemistry, 9(4):721–741, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jucys:1975:SOI

[JSK75b] A. P. Jucys, K. V. Sabas, and Z. J. Kupliauskis. The spin-orbit interaction in the configuration nl^{N}n’l using nonorthogonal radial orbitals. International Journal of Quantum Chemistry, 9 (4):743–754, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Javor:1977:RLE

[JTR77] Frank Javor, Gerald F. Thomas, and Stuart M. Rothstein. Reduced local energy as a criterion for the accuracy of approximate H_{2} wave-functions. International Journal of Quantum Chemistry, 11(1):59–71, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Julg:1978:CGM

[Jul78] André Julg. Crystals as giant molecules, volume 9 of Lecture notes in chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. ISBN 0-387-08946-2. vii + 135 pp. LCCN QD921 .J8.

Junker:1978:CCC

[Jun78] B. R. Junker. Complex coordinate calculations of resonances in N-electron atoms. International Journal of Quantum Chemistry, 14(4):371–382, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Johnson:1978:NTM

[JVM78] K. H. Johnson, D. D. Vvedensky, and R. P. Messmer. New theoretical model for magnetic impurities and the Kondo effect. International Journal of Quantum Chemistry, 14(S12):437–439, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Johnson:1972:DES

[JW72] K. H. Johnson and U. Wahlgren. Determination of the electronic structures of metal complexes by the SCF–Xα scatteredwave method. International Journal of Quantum Chemistry, 6 (S6):243–255, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Jafri:1976:III

[JW76] J. A. Jafri and J. L. Whitten. Iterative introduction of integral approximations with error bounds in SCF calculations. International Journal of Quantum Chemistry, 10(3):383–391, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1978:MFS

[JW78] Herbert W. Jones and C. A. Weatherford. A modified form of Sharma’s formula for STO Löwdin alpha functions with recurrence relations for the coefficient matrix. International Journal of Quantum Chemistry, 14(S12):483–488, March 12–18, 1978.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Jeziorek:1979:ICI

[JŻ79a] Danuta Jeziorek and Bronisław Żurawski. Ab initio calculations including electron correlation for the minimum energy path of the (^{1}A_{1}) CH_{2} + (^{1}Σ)H_{2} → (^{1}A_{1}) CH_{4} insertion reaction. International Journal of Quantum Chemistry, 16(2):277– 292, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jeziorek:1979:GAM

[JŻ79b] Danuta Jeziorek and Bronislaw Żurawski. Geometry of ammonia molecule in the lowest triplet state estimated theoretically. International Journal of Quantum Chemistry, 16(2):261–264, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kleindienst:1976:LFV

[KA76] H. Kleindienst and W. Altmann. I. Lineare Fehlerminimisierung Ein Verfahren zur Eigenwertberechnung bei Schrödinger-Operatoren. (German) [I. Linear error minimization: a method for the eigenvalue calculation of Schrödinger operators]. International Journal of Quantum Chemistry, 10(5):873–885, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaijser:1977:DCP

[Kai77] Per Kaijser. On directional Compton profiles and the J(0) surface. International Journal of Quantum Chemistry, 12(S11): 321–325, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Kang:1974:CES

[Kan74] Sungzong Kang. Conformation and electronic structures of pilocarpine, a muscarinic agonist. International Journal of Quantum Chemistry, 8(S1):109–117, January 17–19, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kapoor:1973:SPA

[Kap73] K. L. Kapoor. The site of protonation in amino-substituted pteridines. International Journal of Quantum Chemistry, 7 (1):27–34, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaplan:1979:MSI

[Kap79] I. G. Kaplan. Modern state of intermolecular interaction theory. International Journal of Quantum Chemistry, 16(3):445–465, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karlsson:1970:HFM

[Kar70a] Göran Karlsson. Hartree–Fock MO–LCAO equations with charge-dependent atomic integrals. International Journal of Quantum Chemistry, 4(3):289–295, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karlsson:1970:SCPa

[Kar70b] Göran Karlsson. Self-consistent pseudo-eigenvalue equations in parametric charge-dependent MO methods. International Journal of Quantum Chemistry, 4(2):127–138, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karlsson:1970:SCPb

[Kar70c] Göran Karlsson. Self-consistent pseudo-eigenvalue equations in parametric charge-dependent MO methods. International Journal of Quantum Chemistry, 4(3):341–342, May 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karle:1978:FAC

[Kar78] Isabella L. Karle. Factors affecting conformations of cyclic polypeptides in the crystalline state. International Journal of Quantum Chemistry, 14(S5):91–101, March 3–11, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kashiwagi:1976:IAB

[Kas76] Hiroshi Kashiwagi. Integral approximations on the basis of semiorthogonalized orbitals. International Journal of Quantum Chemistry, 10(1):135–141, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Katriel:1972:SLH

[Kat72] J. Katriel. The spherical limit of H: a symmetry dilemma. International Journal of Quantum Chemistry, 6(3):541–548, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1970:BMC

[Kau70a] Joyce J. Kaufman. The building up of molecular correlation energies from the molecular correlation energies of their fragments. International Journal of Quantum Chemistry, 5(S4): 139–140, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Kaufman:1970:MPA

[Kau70b] Joyce J. Kaufman. Mulliken population analysis in CNDO and INDO LCAO–MO–SCF methods. International Journal of Quantum Chemistry, 5(S4):205–208, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Kaufman:1973:PUC

[Kau73] Joyce J. Kaufman. Proposed use of charge–density contour maps from ab initio SCF calculations to calibrate sphere sizes for the MS–Xα method. International Journal of Quantum Chemistry, 7(S7):369–370, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Kaufman:1974:DRP

[Kau74a] Joyce J. Kaufman. Discussion remarks on a possible commonality in the etiology of chemical carcinogenesis and allergy. International Journal of Quantum Chemistry, 8(S1):119–122, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1974:SPI

[Kau74b] Joyce J. Kaufman. A suggested procedure to improve the description of lone pairs in the PCILO or more general ab initio perturbative configuration interaction schemes based on localized orbitals. International Journal of Quantum Chemistry, 8 (S1):197–199, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1976:MND

[Kau76] Joyce J. Kaufman. Molecular neurotransmitter disorders — theoretical and chemicophysical approach to their understanding and treatment. International Journal of Quantum Chemistry, 10(S3):187–215, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1977:TAP

[Kau77] Joyce J. Kaufman. Theoretical approaches to pharmacology. International Journal of Quantum Chemistry, 12(S4):375–412, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1979:QCP

[Kau79a] Joyce J. Kaufman. Quantum chemical and physicochemical influences on structure–activity relations and drug design. International Journal of Quantum Chemistry, 16(2):221–241, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1979:SSO

[Kau79b] Joyce J. Kaufman. Spin, symmetry, and orbital filling restrictions indicate the necessity for ab-initio configuration interaction calculations in several fundamental chemical carcinogenesis problems. International Journal of Quantum Chemistry, 16(S6):503–510, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klahn:1977:CLI

[KB77] Bruno Klahn and Werner A. Bingel. Completeness and linear independence of basis sets used in quantum chemistry. International Journal of Quantum Chemistry, 11(6):943–957, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kleindienst:1976:IEB

[KBA76] H. Kleindienst, E. Baumgarten, and W. Altmann. II. Eigenwertberechnung beim Dreikörperproblem. (German) [II. Eigenvalue calculation in the two-body problem]. International Journal of Quantum Chemistry, 10(5):887–899, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Koutecky:1972:DMH

[KBK72] Jaroslav Koutecký and Vlasta Bonačić-Koutecký. Direct minimization of the Hartree–Fock energy for alternant open shell molecules in the PPP model. International Journal of Quantum Chemistry, 6(S6):171–179, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Koutecky:1978:NSP

[KBKČD78] Jaroslav Koutecky, Vlasta Bonačič-Koutecký, Jiří Čížek, and Detlef Döhnert. Nature of the “sudden polarization” effect and its role in photochemistry. International Journal of Quantum Chemistry, 14(S12):357–369, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Kari:1972:GWF

[KC72] R. E. Kari and I. G. Csizmadia. Gaussian wave functions for CH_{3} and NH. International Journal of Quantum Chemistry, 6(3):401–406, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kang:1973:MOA

[KC73] Sungzong Kang and Moon-Hae Cho. Molecular orbital aspects of substituent effects II. International Journal of Quantum Chemistry, 7(S7):319–327, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Klein:1974:SAR

[KC74] D. J. Klein and A. A. Cantu. Symmetry adaptation and the relation between the spin-free and space-spin formulations of electronic structure. International Journal of Quantum Chemistry, 8(2):223–233, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kari:1977:EEG

[KC77] Roy Kari and Imre G. Csizmadia. The energetic effects of p, d, and f Gaussian polarization functions on closed-shell AH_{n} oxygen and sulfur hydrides. International Journal of Quantum Chemistry, 11(3):441–450, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kai:1979:MXM

[KC79] Aldwyn H. Tang Kai and Leon L. Combs. A MS–Xα–MT study of cyclobutane. International Journal of Quantum Chemistry, 16(S13):175–182, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Katriel:1974:SSD

[KD74] J. Katriel and E. Domany. A study of the symmetry dilemma: Second-order transitions. International Journal of Quantum Chemistry, 8(4):559–564, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kothekar:1977:MOC

[KD77] V. Kothekar and S. Dutta. Molecular orbital calculations of metal ion interaction with nucleic acid bases. I. Binding of Cu(II) with adenine, guanine, uracil, and cytosine. International Journal of Quantum Chemistry, 12(3):505–514, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kobeissi:1979:AER

[KD79a] Hafez Kobeissi and Mounzer Dagher. Analytic expression of the rotation-vibration eigenfunction of any electronic potential of a diatomic molecule. International Journal of Quantum Chemistry, 16(S13):169–174, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Kothekar:1979:RCI

[KD79b] V. Kothekar and S. Dutta. Role of Ca^{+2} ion in the abortifacient action of prostaglandins. I. Molecular-orbital and conformation-energy calculations of PGF_{2α}. International Journal of Quantum Chemistry, 15(5):481–489, May 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kibler:1979:DAC

[KE79] M. Kibler and E. Elbaz. Diagram approach to character formulas for finite and compact groups. International Journal of Quantum Chemistry, 16(6):1161–1194, December 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Keller:1975:SCM

[Kel75] Jaime Keller. SC Cellular multiple scattering in molecular electronic structure calculations. International Journal of Quantum Chemistry, 9(4):583–604, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karlsson:1971:SCP

[KFY71] Göran Karlsson, Kenichi Fukui, and Tokio Yamabe. Selfconsistent pseudo-eigenvalue equations in parametric chargedependent MO methods. International Journal of Quantum Chemistry, 5(4):477–478, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kibler:1976:CGC

[KG76] M. R. Kibler and P. A. M. Guichon. Clebsch–Gordan coefficients for chains of groups of interest in quantum chemistry. International Journal of Quantum Chemistry, 10(1):87–111, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kibler:1977:CGC

[KG77] M. R. Kibler and G. Grenet. Clebsch–Gordan coefficients for chains of groups of interest in quantum chemistry. II. The chain SU(2) ⊃ D′_{∞} ⊃ D′_{4} ⊃ D′_{2}. International Journal of Quantum Chemistry, 11(2):359–379, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kunz:1976:AHS

[KGB76] A. Barry Kunz, M. P. Guse, and R. J. Blint. On the adsorption of hydrogen by supported transition-metal catalysts and MgO. International Journal of Quantum Chemistry, 10(S10): 283–288, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klein:1977:VLSb

[KGB77] D. J. Klein and M. A. García-Bach. Variational localized-site cluster expansions. VIII. Projection of spin symmetries. International Journal of Quantum Chemistry, 12(2):291–303, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kosevich:1979:ELM

[KGP79] A. M. Kosevich, V. L. Galkin, and M. L. Polyakov. Effect of low-molecular impurity on the melting of DNA-type macromolecules. International Journal of Quantum Chemistry, 16(3): 623–629, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kanazawa:1979:SMS

[KHI79] Tomoko Kanazawa, Haruo Hosoya, and Suehiro Iwata. Semiempirical MO study on the abnormal bond orders of large networks of highly branched polyenes. International Journal of Quantum Chemistry, 15(2):243–257, February 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Khokhlov:1979:CFL

[Kho79] A. R. Khokhlov. Conditions of formation of liquid crystals from biopolymer solutions. International Journal of Quantum Chemistry, 16(4):857–862, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kibler:1975:APT

[Kib75a] Maurice R. Kibler. An alternative parametrization for the theory of complex spectra. International Journal of Quantum Chemistry, 9(3):421–429, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kibler:1975:CBP

[Kib75b] Maurice R. Kibler. Comparison between the point-charge electrostatic model and the angular overlap model: The general case. International Journal of Quantum Chemistry, 9(3):403– 420, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kijewski:1972:TMC

[Kij72] L. J. Kijewski. Test of the g-matrix condition on the twomatrix satisfying the Pauli restriction. International Journal of Quantum Chemistry, 6(S6):73–78, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Kiss:1975:TIE

[KJ75] A. I. Kiss and F. Joó. Theoretical investigation of the πelectronic structure and spectra of protonated aromatic carbonyl compounds. International Journal of Quantum Chemistry, 9(2):261–271, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kurtz:1978:TSF

[KJ78] H. A. Kurtz and K. D. Jordan. Theoretical study of [F^{−}; e^{+}] and [CN^{−}; e^{+}]. International Journal of Quantum Chemistry, 14(6):747–755, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1972:QCC

[KK72] Joyce J. Kaufman and Ellen Kerman. Quantum chemical calculations on antipsychotic drugs and narcotic agents. International Journal of Quantum Chemistry, 6(S6):319–335, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Kaufman:1974:SPD

[KK74a] Joyce J. Kaufman and Ellen Kerman. The structure of psychotropic drugs (including theoretical prediction of a new class of effective neuroleptics). International Journal of Quantum Chemistry, 8(S1):259–287, January 17–19, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Konig:1974:ITO

[KK74b] E. König and S. Kremer. Irreducible tensor operator methods in intermediate-field coupling. International Journal of Quantum Chemistry, 8(3):347–362, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1975:PQC

[KK75] Joyce J. Kaufman and Walter S. Koski. Physicochemical, quantum chemical, and other theoretical studies of the mechanism of action of CNS agents: Anesthetics, narcotics and narcotic antagonists, and psychotropic drugs. International Journal of Quantum Chemistry, 9(S2):35–57, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1976:QCC

[KK76a] Joyce J. Kaufman and Ellen Kerman. Quantum chemical calculations on thioridazine. International Journal of Quantum Chemistry, 10(4):559–567, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1976:SCP

[KK76b] E. S. Kryachko and Yu. A. Kruglyak. On the solution of the N-completeness problem. International Journal of Quantum Chemistry, 10(6):937–940, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kuprievich:1976:CSO

[KK76c] V. A. Kuprievich and V. E. Klymenko. Computation scheme for optimizing multiconfigurational wave-functions. International Journal of Quantum Chemistry, 10(6):941–949, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1977:CPN

[KK77a] Joyce J. Kaufman and Ellen Kerman. Conformational profile of nalorphine by PCILO calculations. International Journal of Quantum Chemistry, 11(1):181–184, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Konig:1977:CFP

[KK77b] E. König and S. Kremer. The concept of fractional parentage for arbitrary molecular point groups. International Journal of Quantum Chemistry, 12(3):457–469, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Konig:1977:ITO

[KK77c] E. König and S. Kremer. Irreducible tensor operator methods in strong-field coupling. International Journal of Quantum Chemistry, 12(6):1017–1031, December 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kuprievich:1977:MOE

[KK77d] V. A. Kuprievich and V. E. Klymenko. Modified one-electron Hamiltonian method in the MC SCF theory. International Journal of Quantum Chemistry, 12(2):327–334, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kertesz:1978:EBS

[KKA78] M. Kertesz, J. Koller, and A. Ažman. Energy band structure of (SN)_{x} chain: Unrestricted Hartree–Fock and charge density wave solutions. International Journal of Quantum Chemistry, 14(3):239–243, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1970:IPS

[KKK70] Joyce J. Kaufman, Ellen Kerman, and Walter S. Koski. Implications of photoelectron spectroscopic measurements for compounds which produce no parent ion: Contribution to the session on photoelectron spectroscopy. International Journal of Quantum Chemistry, 5(S4):391–394, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Kaufman:1974:QCO

[KKK74] Joyce J. Kaufman, Ellen Kerman, and Walter S. Koski. Quantum chemical, other theoretical and physicochemical studies on narcotics and narcotic antagonists to understand their mechanism of action. International Journal of Quantum Chemistry, 8(S1):289–313, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kuprievich:1979:REC

[KKK79] V. A. Kuprievich, Z. G. Kudritskaya, and V. E. Klymenko. Role of electron correlation in the quantum-mechanical calculations of the Coulomb interaction energy in the DNA base pairs. International Journal of Quantum Chemistry, 16(1):71–77, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kharkyanen:1979:MCS

[KKKP79] V. N. Kharkyanen, L. N. Khristoforov, V. V. Kukhtin, and E. G. Petrov. Mechanisms of charge separation in bacterial photosynthesis. International Journal of Quantum Chemistry, 16(4):877–882, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kuprievich:1970:FCI

[KKM70] V. A. Kuprievich, Yuri A. Kruglyak, and E. V. Mozdor. Full configuration interaction for the benzyl radical. International Journal of Quantum Chemistry, 4(1):73–87, January 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1977:SCT

[KKP77] Joyce J. Kaufman, Walter S. Koski, and David Peat. A systems and control theoretic approach to dynamic neurotransmitter balance: Normal, abnormal, and “catastrophic”. International Journal of Quantum Chemistry, 11(3):485–503, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karo:1973:RAM

[KKW73] Arnold Karo, Morris Krauss, and Arnold C. Wahl. Recent applications of the multiconfiguration self-consistent field method to polarizabilities, excited states, van der Waals forces, and triatomic surfaces. International Journal of Quantum Chemistry, 7(S7):143–159, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Krizan:1973:IBM

[KL73] John E. Krizan and Brian F. Lloyd. Interactions between molecules in the neighborhood of a long linear molecule. International Journal of Quantum Chemistry, 7(5):991–998, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kamel:1975:DME

[KL75] M. A. Kamel and L. N. Labzowsky. Density matrix expansions and their application in the theory of the many-electron systems. International Journal of Quantum Chemistry, 9(1):35– 46, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kleier:1977:MOS

[KL77a] Daniel A. Kleier and William N. Lipscomb. Molecular orbital study of polypeptides. Conformational and electronic structure of polyglycine. International Journal of Quantum Chemistry, 12(S4):73–86, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Knirk:1977:MIP

[KL77b] Dwayne L. Knirk and Richard C. Liedtke. Model interaction potentials for atom-diatom scattering. International Journal of Quantum Chemistry, 12(S11):327–336, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Koutecky:1977:SQS

[KL77c] Jaroslav Koutecký and Alexandre Laforgue. Second quantization for systems with a constant number of particles in the Dirac notation. International Journal of Quantum Chemistry, 11(3):505–523, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kai:1978:CSM

[KL78] Aldwyn Tang Kai and Sven Larsson. A comparative study of multiple scattering calculations on CrF_{6}^{3−}. International Journal of Quantum Chemistry, 13(3):367–374, March 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kwiatkowski:1979:MSM

[KL79] Józef S. Kwiatkowski and Bogdan Lesyng. Modified π–SCF MO CI calculations for anionic forms of pyrimidinic nucleic acid bases. International Journal of Quantum Chemistry, 16 (S6):391–408, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klasinc:1978:APS

[Kla78] L. Klasinc. Application of photoelectron spectroscopy to biologically active molecules and their constituent parts. V. Amino acid methyl esters. International Journal of Quantum Chemistry, 14(S5):373–380, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement:

Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klein:1970:SAP

[Kle70a] Douglas J. Klein. Symmetry adapted perturbation theory for interatomic and intermolecular exchange interactions. International Journal of Quantum Chemistry, 5(S4):271–283, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Klessinger:1970:SCG

[Kle70b] Martin Klessinger. Self-consistent group calculations on polyatomic molecules V. Molecules with a double or triple bond. International Journal of Quantum Chemistry, 4(2):191–202, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Klein:1974:CKS

[Kle74a] D. J. Klein. Comments on kets not strongly symmetry adapted. International Journal of Quantum Chemistry, 8(2):235–245, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kleinpoppen:1974:SPC

[Kle74b] Hans Kleinpoppen. Spin, polarization, and coherence effects in electron atom collision processes. International Journal of Quantum Chemistry, 8(S8):59–72, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Klein:1977:VLSa

[Kle77] D. J. Klein. Variational localized-site cluster expansions. VI. General theory revisited. International Journal of Quantum Chemistry, 12(2):255–271, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Klein:1979:LRO

[Kle79] D. J. Klein. Long-range order for spin pairing in valence bond theory. International Journal of Quantum Chemistry, 16(S13):

293–303, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Kliger:1979:SPS

[Kli79] David S. Kliger. Spectroscopic and photochemical studies of model visual pigment chromophores. International Journal of Quantum Chemistry, 16(4):809–817, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1972:TCS

[KM72] Joyce J. Kaufman and Albert A. Manian. Topological conformational similarities among antipsychotic drugs, narcotics and biogenic amines: a summary. International Journal of Quantum Chemistry, 6(S6):375–381, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Kagawa:1974:ECS

[KM74a] Takashi Kagawa and Tomokazu Murai. Electron correlation study for two-electron atoms by a simple correlated wave function. International Journal of Quantum Chemistry, 8(5):755– 764, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kiss:1974:MMC

[KM74b] A. I. Kiss and A. Martin. Molecules-in-molecules calculations with fixed resonance integrals. International Journal of Quantum Chemistry, 8(3):427–433, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kitaura:1976:NED

[KM76] Kazuo Kitaura and Keiji Morokuma. A new energy decomposition scheme for molecular interactions within the Hartree–Fock approximation. International Journal of Quantum Chemistry, 10(2):325–340, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kutzelnigg:1977:DEG

[KMLB77] W. Kutzelnigg, A. Meunier, B. Lévy, and G. Berthier. On the dependence of the energy given by a CI limited to double substitutions with respect to the number of electrons. International Journal of Quantum Chemistry, 12(4):777–779, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kumar:1977:ESS

[KMO77] Lalit Kumar, Hendrik J. Monkhorst, and Jens Oddershede. Electronic-structure studies of solides. V. Rigorous Hartree– Fock treatment of metallic hydrogen using a plane-wave basis. International Journal of Quantum Chemistry, 12(1):145–160, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Knezevic:1977:SSP

[KMT77] G. Knežević, M. Marinković, and B. Tošić. On some specific properties of systems endowed with dipole–dipole interactions. International Journal of Quantum Chemistry, 12(5):785–794, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kikuchi:1979:TSH

[KN79] Takeshi Kikuchi and Kichisuke Nishimoto. Theoretical studies of hemoproteins. I. Mathematical description of the allosteric effect. International Journal of Quantum Chemistry, 15(4):379–387, April 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kouba:1971:NOI

[KÖ71] Joseph E. Kouba and Yngve Öhrn. Natural orbital iterations for the ground state of nitric oxide. International Journal of Quantum Chemistry, 5(5):539–548, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kobe:1978:IGI

[Kob78] Donald H. Kobe. Implications of gauge invariance for length versus velocity forms of the interaction with electric dipole radiation. International Journal of Quantum Chemistry, 14(S12): 73–86, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Kochanski:1974:DDT

[Koc74] Elise Kochanski. Description of the dispersion term between two molecules and two atoms from an ab initio perturbative procedure: H_{2} + H_{2} and Ne + Ne. International Journal of Quantum Chemistry, 8(S8):219–227, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Koelling:1974:ESR

[Koe74] D. D. Koelling. On the efficient solution of the relativistic APW secular equation for both eigenvalues and eigenvectors. International Journal of Quantum Chemistry, 8(S8):473–481, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Kolos:1974:LIR

[Koł74] W. Kołos. Long- and intermediate-range interaction between two ground state hydrogen atoms. International Journal of Quantum Chemistry, 8(S8):241–252, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Kolos:1975:LIR

[Koł75] W. Kołos. Long- and intermediate-range interaction between hydrogen atoms in the B^{1}Σ state. International Journal of Quantum Chemistry, 9(1):133–142, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kolos:1976:LIR

[Koł76] W. Kołos. Long- and intermediate-range interaction in three lowest sigma states of the HeH^{+} ion. International Journal of Quantum Chemistry, 10(2):217–224, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kothekar:1976:MOI

[Kot76] V. Kothekar. Molecular orbital interpretation of isomer shifts and quadrupole splitting of the Mössbauer γ-line in the compounds of antimony and tin. International Journal of Quantum Chemistry, 10(6):993–1005, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kwiatkowski:1977:CHC

[KOW77] T. Kwiatkowski, S. Olszewski, and A. Wierzbicki. Cubic harmonics in Cartesian coordinates. International Journal of Quantum Chemistry, 11(1):21–47, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kunert:1978:ISNb

[KOW78] A. Kunert, S. Olszewski, and A. Wierzbicki. Incompletely symmetric non-Bloch electron states in perfect cubic crystals. II. Density of states in the FCC lattice. International Journal of Quantum Chemistry, 14(2):153–162, August 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1971:NAL

[KP71a] Joyce J. Kaufman and R. Predney. Non-applicability for the Li^{+}H_{2} ion molecule system of an INDO potential energy surface or of the approximate monopole-transition moment long range force expressions. International Journal of Quantum Chemistry, 5(S5):235–258, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Kijewski:1971:NCC

[KP71b] L. J. Kijewski and J. K. Percus. Numerical calculations for C^{2+} using reduced density matrices. International Journal of Quantum Chemistry, 5(S5):67–69, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Kaufman:1972:EIF

[KP72] Joyce J. Kaufman and Robert Predney. Extension of INDO formalism to d orbitals and parameters for second-row atoms.

International Journal of Quantum Chemistry, 6(S6):231–242, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Kaufman:1976:SPP

[KP76] Joyce J. Kaufman and H. E. Popkie. Similarity in the pattern per atom-pair in benzene of contributions to total energy or total overlap populations. International Journal of Quantum Chemistry, 10(3):543–544, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Krishnan:1978:AFO

[KP78] R. Krishnan and J. A. Pople. Approximate fourth-order perturbation theory of the electron correlation energy. International Journal of Quantum Chemistry, 14(1):91–100, July 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Khutorsky:1979:CAD

[KP79a] V. E. Khutorsky and V. I. Poltev. Conformational analysis of double-helical polynucleotides. International Journal of Quantum Chemistry, 16(1):153–157, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kwiatkowski:1979:ISP

[KP79b] J. S. Kwiatkowski and B. Pullman. Ab initio study of a purine nucleoside: Adenosine. International Journal of Quantum Chemistry, 15(5):499–510, May 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kothekar:1978:IMO

[KPD78] V. Kothekar, A. Pullman, and D. Demoulin. Ab initio molecular-orbital study of the binding of Zn^{II} with SH_{2} and SH^{−}. International Journal of Quantum Chemistry, 14(6): 779–791, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Klein:1970:CSF

[KPJM70] D. J. Klein, R. D. Poshusta, B. R. Junker, and F. A. Matsen. Computation in the spin-free formulation of the polyelectronic problem. International Journal of Quantum Chemistry, 5(S4):

141–146, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Kaufman:1973:CPP

[KPKC73] Joyce J. Kaufman, Harry J. T. Preston, Ellen Kerman, and L. Chopin Cusachs. Comparison for pyrrole and pyrazole of orbital energies and population analyses from ab-initio SCF, CNDO/2, INDO, extended Hückel, and ARCANA calculations. International Journal of Quantum Chemistry, 7(S7):249–260, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Kurtz:1976:CAE

[KPÖ76] Henry Kurtz, George Purvis, and Yngve Öhrn. Calculation of the amplitudes to the electron propagator from a minimal basis CI calculation on N_{2}. International Journal of Quantum Chemistry, 10(S10):331–336, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1977:IAR

[KPP77] Joyce J. Kaufman, Herbert E. Popkie, and Harry J. T. Preston. Ab-initio and approximately rigorous calculations on small, medium, and large systems. International Journal of Quantum Chemistry, 11(6):1005–1015, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1978:INM

[KPP78a] Joyce J. Kaufman, Herbert E. Popkie, and Harry J. T. Preston. Ab initio and nonempirical MODPOT/VRDDO calculations on drugs, carcinogens, suspected teratogens, and biomolecules. International Journal of Quantum Chemistry, 14(S5):201–218, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1978:MCNb

[KPP78b] Joyce J. Kaufman, Herbert E. Popkie, and Harry J. T. Preston. Molecular calculations with the nonempirical ab initio MODPOT, VRDDO, and MODPOT/VRDDO procedures. VIII. charge delocalization in the anions of aromatic carboxylic acids and phenolic compounds. International Journal of Quantum Chemistry, 14(S12):283–291, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Kwiatkowski:1978:AOW

[KPP78c] J. S. Kwiatkowski, D. Perahia, and B. Pullman. Aspects of oxopyridine-water interactions. International Journal of Quantum Chemistry, 14(S12):249–256, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Kaufman:1978:MCNa

[KPPH78] Joyce J. Kaufman, Herbert E. Popkie, Surat Palalikit, and P. C. Hariharan. Molecular calculations with the nonempirical ab initio MODPOT, VRDDO, and MODPOT /VRDDO procedures. IX. Carcinogenic benzo(a)pyrene and its metabolites using a MERGE technique. International Journal of Quantum Chemistry, 14(6):793–800, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaplan:1971:ICE

[KR71] I. G. Kaplan and O. B. Rodimova. Ab initio calculation of electronic states of the ions H_{4}^{+} and H_{5}^{+}. International Journal of Quantum Chemistry, 5(6):669–682, November 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaplan:1973:MEG

[KR73a] I. G. Kaplan and O. B. Rodimova. Matrix elements of general configuration of nonorthogonalized orbitals in state with definite spin. International Journal of Quantum Chemistry, 7(6): 1203–1220, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kumari:1973:SDD

[KR73b] Usha Kumari and N. K. Ray. Spin density distributions in some fluorinated radical cations. International Journal of Quantum Chemistry, 7(2):215–222, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaplan:1976:GTC

[KR76] I. G. Kaplan and O. B. Rodimova. Group theoretical classification of states of a molecular system at definite states of its constituent parts. International Journal of Quantum Chemistry, 10(4):699–714, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kolos:1978:ASM

[KR78] Włodzimierz Kołos and Elżbieta Radzio. Application of the statistical method in the theory of intermolecular interactions. International Journal of Quantum Chemistry, 13(5):627–634, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Klasinc:1978:PSH

[KRKŠ78] L. Klasinc, B. Ruščič, F. Kajfež, and V. Šunjić. Photoelectron spectroscopy of the heterocycles imidazole and methylimidazoles. International Journal of Quantum Chemistry, 14(S5): 367–371, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kryachko:1977:PSR

[Kry77] Eugene S. Kryachko. Permutational symmetry of reduced density matrices. International Journal of Quantum Chemistry, 11(4):577–589, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1978:RDM

[Kry78] Eugene S. Kryachko. Reduced-density-matrix theory and algebraic structures. International Journal of Quantum Chemistry, 13(6):719–730, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kuprievich:1972:MST

[KS72] V. A. Kuprievich and O. V. Shramko. The MC SCF theory: Method of one-electron Hamiltonian. International Journal of Quantum Chemistry, 6(2):327–336, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kari:1973:DME

[KS73a] R. Kari and B. T. Sutcliffe. Direct minimization of the energy functional in LCAO–MO calculations. International Journal of Quantum Chemistry, 7(3):459–473, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kashiwagi:1973:GLO

[KS73b] Hiroshi Kashiwagi and Fukashi Sasaki. A generalization of the Löwdin orthogonalization. International Journal of Quantum Chemistry, 7(S7):515–520, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Kuprievich:1975:ICS

[KS75] V. A. Kuprievich and O. V. Shramko. Improved convergence of self-consistence procedures in the MC SCF theory. International Journal of Quantum Chemistry, 9(6):1009–1020, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kuprievich:1979:TSE

[KS79] V. A. Kuprievich and O. V. Shramko. Theoretical study of electronic spectra of the DNA bases. International Journal of Quantum Chemistry, 16(4):833–837, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kleier:1976:COS

[KSL76] Daniel A. Kleier, Steve Scheiner, and William N. Lipscomb. Comments on orbital steering. International Journal of Quantum Chemistry, 10(S3):161–169, January 12–14, 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1978:GAB

[KSP78] Joyce J. Kaufman, Denis Salahub, and Harry J. T. Preston. General applicability of a big molecule Gaussian SCF/CI program for calculations of excited metastables and of negative ion bound states and resonances. I. Stabilization method: H^{−}. International Journal of Quantum Chemistry, 14(4):515–518, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kapur:1973:SRL

[KSS73] K. Kapur, K. V. Sane, and P. K. Srivastava. Some remarks on Linderberg’s relation. International Journal of Quantum Chemistry, 7(5):1029–1031, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Klimo:1976:SEH

[KT76] Viliam Klimo and Jozef Tiňo. Spin-extended Hartree–Fock calculation for hydrocarbon π-radicals. International Journal of Quantum Chemistry, 10(5):761–769, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kashiwagi:1977:TIA

[KT77] Hiroshi Kashiwagi and Toshikazu Takada. Test of an integral approximation scheme based on semiorthogonalized orbitals in LCAO SCF MO CI calculations of naphthalene. International Journal of Quantum Chemistry, 12(3):449–456, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kashiwagi:1978:IMO

[KTO^{+}78] Hiroshi Kashiwagi, Toshikazu Takada, Shigeru Obara, Eisaku Miyoshi, and Kimio Ohno. Ab initio molecular orbital calculations of the cobalt porphine complex. I. LCAO SCF MO calculation of low-spin, high-spin, and π-ionized states of Coporphine. International Journal of Quantum Chemistry, 14(1): 13–27, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kruglyak:1970:SES

[KU70] Yuri A. Kruglyak and I. I. Ukrainsky. Study of the electronic structure of alternant radicals by the DODS method. International Journal of Quantum Chemistry, 4(1):57–72, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kummer:1977:ARB

[Kum77] H. Kummer. About the relationship between some necessary conditions for N-representability. International Journal of Quantum Chemistry, 12(6):1033–1038, December 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kunert:1978:ISNa

[Kun78a] A. Kunert. Incompletely symmetric non-Bloch electron states in perfect cubic crystals. I. Eigenproblem and its solution. International Journal of Quantum Chemistry, 14(2):127–152, August 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kunert:1978:ISNc

[Kun78b] A. Kunert. Incompletely symmetric non-Bloch electron states in perfect cubic crystals. III. Density of dynamical observables in the FCC lattice. International Journal of Quantum Chemistry, 14(2):163–169, August 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kutzelnigg:1978:ABP

[Kut78] Werner Kutzelnigg. Asymptotic behavior of the primitive function of different “symmetry-adapted” perturbation schemes for the H ground state. International Journal of Quantum Chemistry, 14(1):101–120, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kulkarni:1979:CSP

[KVSG79] Vithal M. Kulkarni, N. Vasanthkumar, Anil Saran, and Girjesh Govil. Conformational structure of propranolol: A β-adrenergic blocking drug studied by NMR and PCILO methods. International Journal of Quantum Chemistry, 16(S6):153–170, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kennard:1976:CAF

[KW76] E. B. Kennard and J. T. Waber. Calculated anisotropies in the Fermi surface of alkali metals. International Journal of Quantum Chemistry, 10(2):261–280, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kunz:1979:ETP

[KWC79] A. B. Kunz, R. S. Weidman, and T. C. Collins. Electronic and transport properties of CuCl. International Journal of Quantum Chemistry, 16(S13):453–465, March 11–17, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Karlsson:1973:DOC

[KZ73] Göran Karlsson and Michael C. Zerner. Determination of onecentre core integrals from the average energies of atomic configurations. International Journal of Quantum Chemistry, 7 (1):35–49, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laughlin:1970:PGE

[LA70] C. Laughlin and A. T. Amos. A perturbed generalized eigenvalue equation for the helium atom. International Journal of Quantum Chemistry, 4(6):613–623, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lo:1976:CAF

[LA76] B. W. N. Lo and Y. A. Ang. Calculation of the G(a) auxiliary function in Gaussian orbit–orbit integrals. International Journal of Quantum Chemistry, 10(6):1087–1088, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ladik:1970:EBD

[Lad70] J. Ladik. Energy bands in DNA. International Journal of Quantum Chemistry, 5(S4):307–317, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Ladik:1974:AVE

[Lad74a] János Ladik. All valence electron band structures of simple periodic protein models. International Journal of Quantum Chemistry, 8(S1):5–11, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ladik:1974:EBS

[Lad74b] János Ladik. The energy band structure and conduction properties of DNA. International Journal of Quantum Chemistry, 8(S1):65–69, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ladik:1975:NBS

[Lad75a] János Ladik. A note on the band structure of the TCNQ-TTF system. International Journal of Quantum Chemistry, 9 (S9):563–567, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Ladik:1975:TES

[Lad75b] János J. Ladik. Toward the electronic structure of real DNA. International Journal of Quantum Chemistry, 9(S2):133–143, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ladik:1976:EBS

[Lad76] János Ladik. Energy-band structure and charge transfer in biopolymers. International Journal of Quantum Chemistry, 10 (S3):237–246, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Laidlaw:1973:SCS

[Lai73] W. G. Laidlaw. Stability conditions for the solutions of the Hartree–Fock equations. VII. Stability of some closed-shell nonalternant systems. International Journal of Quantum Chemistry, 7(1):87–108, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laki:1975:AAN

[Lak75] Koloman Laki. Actin as an ancient nucleotide-binding protein. International Journal of Quantum Chemistry, 9(S2):297– 305, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Laki:1976:TYG

[Lak76] K. Laki. Twenty years a-growing with Albert Szent-Györgyi. International Journal of Quantum Chemistry, 10(S3):3–5, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lyslo:1971:TIS

[LAM71] Gudmund Lyslo, Kjell Aashamar, and John Midtdal. A theoretical investigation of the singly excited two-electron atomic states 1s2s^{3} S and 1s3s^{3} S. International Journal of Quantum Chemistry, 5(5):583–598, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Langhoff:1973:STF

[Lan73] P. W. Langhoff. Separation theorem for first-order paircorrelation equations. International Journal of Quantum Chemistry, 7(S7):443–448, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Langhoff:1974:SIA

[Lan74] P. W. Langhoff. Stieltjes-integral approximations to elementary dispersion relations. International Journal of Quantum Chemistry, 8(S8):347–361, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Langhoff:1977:AST

[Lan77] P. W. Langhoff. Applications of the Stieltjes–Tchebycheff procedure for atomic and molecular photoionization calculations in Hilbert space. International Journal of Quantum Chemistry, 12 (S11):301–310, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Larson:1973:RRD

[Lar73] Everett G. Larson. Representable reference density matrices induced by information theory. International Journal of Quantum Chemistry, 7(5):853–867, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larson:1975:PHO

[Lar75] Everett G. Larson. Particle-hole operators as state generators defects in the random phase approximation (as evidenced by calculations on the frozen-K-shell model of the B^{+} ion). International Journal of Quantum Chemistry, 9(S9):199–212, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Larsson:1976:BRB

[Lar76] Sven Larsson. Book review: Quantum theory of the chemical bond. Author: H. F. Hameka. Published by: Hafner Press, New York and London, 1975. International Journal of Quantum Chemistry, 10(3):553, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larson:1978:CRS

[Lar78a] Everett G. Larson. Comparisons of the RPA, SCRPA, Tamm– Dancoff, and full CI methods by analysis of their transition density matrices, oscillator strengths, and energy moments of oscillator strengths for the electric dipole transitions from the ground state of the B^{+} ion (frozen K-shell model). International Journal of Quantum Chemistry, 13(6):743–767, June 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larson:1978:RRD

[Lar78b] Everett G. Larson. Representable reference density matrices applied to the graphical description of the Coulomb hole in the 1s^{2}(2p)^{2} ^{1}S state of the B^{+} ion. International Journal of Quantum Chemistry, 13(6):791–799, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larsson:1978:BRB

[Lar78c] Sven Larsson. Book review: Applications of MO theory in organic chemistry (Progress in Theoretical Organic Chemistry, Vol. 2). I. G. Csizmadia, Ed. Elsevier Scientific Publishing Company, Amsterdam and New York, 1976. Price: $69.50, 626 pp. International Journal of Quantum Chemistry, 13(4):563, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Csi77].

Larson:1979:RSR

[Lar79] Everett G. Larson. The role of symmetry in representing reduced density operators and reduced transition density operators: General formulation with specific application to atomic systems. International Journal of Quantum Chemistry, 16 (S13):121–140, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Lathouwers:1976:NAC

[Lat76] Luc Lathouwers. A new approach to canonical orthonormalisation. International Journal of Quantum Chemistry, 10(3): 413–418, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lathouwers:1977:PIB

[Lat77] Luc Lathouwers. Parseval’s identity for biorthonormal systems. International Journal of Quantum Chemistry, 11(2):277–281, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laurenzi:1979:PMH

[Lau79] Bernard J. Laurenzi. Power moments of hydrogenic Green’s functions and Green’s functions of the second kind. International Journal of Quantum Chemistry, 15(6):717–729, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Letcher:1972:ULO

[LAV72] J. H. Letcher, I. Absar, and J. R. Van Wazer. Use of localized orbitals to circumvent convergence difficulties in LCAO–MO– SCF calculations. International Journal of Quantum Chemistry, 6(S6):451–452, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Layzer:1977:SMS

[Lay77] David Layzer. The structure of matter and the structure of the astronomical universe. International Journal of Quantum Chemistry, 12(S11):637–645, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Lazzeretti:1978:USF

[Laz78] Paolo Lazzeretti. On the use of symmetry in first-order perturbed Hartree–Fock theory. International Journal of Quantum Chemistry, 13(3):375–390, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lazzeretti:1979:GAN

[Laz79] P. Lazzeretti. Geometric approximation to nuclear spin–spin coupling constants in the water molecule. International Journal of Quantum Chemistry, 15(2):181–196, February 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lindner:1973:NFE

[LB73a] P. Lindner and K-F. Berggren. A nearly free electron model for the dihydrides of the Alkaline-Earth metals. International Journal of Quantum Chemistry, 7(4):667–674, July 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lyast:1973:ALE

[LB73b] I. Ts. Lyast and A. Ya. Biktimirov. Application of Local Energy Method to the Estimation of the Correlation in Geminals. International Journal of Quantum Chemistry, 7(4):641–646, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Levy:1977:RBO

[LB77] B. Lévy and G. Berthier. Relationship between orthogonalization and orbital localization procedures. International Journal of Quantum Chemistry, 12(3):579–590, September 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larsson:1979:BDN

[LB79a] S. Larsson and M. Braga. Back donation in Ni(CO)_{4}: Comments on papers by K. H. Johnson and by K. H. Johnson and U. Wahlgren. International Journal of Quantum Chemistry, 15(1):1–5, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Les:1979:MCC

[LB79b] A. Les and I. Ortega Blake. Magnesium and calcium cationligand interactions within the pseudopotential approach. I. Cation-water interaction. International Journal of Quantum Chemistry, 16(S6):189–201, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lindholm:1978:IES

[LBÅF78] E. Lindholm, G. Bieri, L. Åsbrink, and C. Fridh. Interpretation of electron spectra. III. Spectra of formamide, studied with HAM /3. International Journal of Quantum Chemistry, 14(6): 737–740, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Linz:1977:EEC

[LC77] Peter Linz and Tien Chi Chen. Exact eigenvalues for the Coulomb potential with cut-off. International Journal of Quantum Chemistry, 11(6):991–1000, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leopold:1976:SLC

[LCK76] J. G. Leopold, M. Cohen, and J. Katriel. A study of Löwdin’s criterion for completeness of basis sets. International Journal of Quantum Chemistry, 10(6):1049–1056, November 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lazzeretti:1976:CED

[LCP76] Paolo Lazzeretti, Beniamino Cadioli, and Ugo Pincelli. Calculations of electric dipole polarizabilities of polyatomic molecules. International Journal of Quantum Chemistry, 10(5):771–780, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Langlet:1979:SSE

[LCPP79] J. Langlet, P. Claverie, B. Pullman, and D. Piazzola. Studies of solvent effects. IV. Study of hydration of the dimethyl phosphate anion (DMP^{−}) and of the solvent effect upon its conformation. International Journal of Quantum Chemistry, 16(S6):409–437, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Larson:1971:PHC

[LD71] Everett G. Larson and Orville W. Day. The particle-hole component of the 2-particle Green function — its relation to the canonical particle-hole states and its N-representability. International Journal of Quantum Chemistry, 5(S5):729–744, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Langhoff:1973:ICS

[LD73] Stephen R. Langhoff and Ernest R. Davidson. An Ab Initio calculation of the spin dipole-dipole parameters for methylene. International Journal of Quantum Chemistry, 7(4):759–777, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Langhoff:1974:CIC

[LD74] Stephen R. Langhoff and Ernest R. Davidson. Configuration interaction calculations on the nitrogen molecule. International Journal of Quantum Chemistry, 8(1):61–72, January 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lavery:1979:PTS

[LDP79] Richard Lavery, Marçal De Oliveira, and Bernard Pullman. A preliminary theoretical study of the acid catalyzed hydration of glyoxal, methylglyoxal, and other simple aldehydes. International Journal of Quantum Chemistry, 16(S6):459–466, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lin:1972:SEP

[LE72] Teng Ke Lin and Darrell D. Ebbing. A separated-electron pair study of the electronic structure of the Li_{2}O molecule. International Journal of Quantum Chemistry, 6(2):297–311, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leclerc:1976:CSU

[Lec76a] Jean-Claude Leclerc. Comparative study of unconventional 1s basis functions for the ^{1}Σ ground state of H_{2} and He. International Journal of Quantum Chemistry, 10(3):439–454, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leclerc:1976:HSG

[Lec76b] Jean-Claude Leclerc. Hybrid Slater–Gaussian-type 1s orbitals for the ^{1}Σ ground state of H_{2} and He. International Journal of Quantum Chemistry, 10(5):781–789, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Langhoff:1973:CIS

[LED73] Stephen R. Langhoff, Stephen T. Elbert, and Ernest R. Davidson. A configuration interaction study of the spin dipoledipole parameters for formaldehyde and methylene. International Journal of Quantum Chemistry, 7(5):999–1019, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leibovici:1974:SIF

[Lei74] Claude Leibovici. Structure de l’ion fluoronium FH et chemin de la protonation du fluorure d’hydrogène. (French) [Structure of the fluoronium ion FH and the protonation path of hydrogen fluoride]. International Journal of Quantum Chemistry, 8 (2):193–200, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Letoquart:1974:SKP

[Lét74] D. Létoquart. Simplification of Kronecker products for the representations of point groups having only one n-fold axis with n ≥ 3. International Journal of Quantum Chemistry, 8(4):627– 632, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Levy:1970:MMS

[Lev70] B. Levy. Molecular MC–SCF calculations. International Journal of Quantum Chemistry, 4(3):297–313, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Levin:1978:MBS

[Lev78a] F. S. Levin. Many-body scattering theory methods as a basis for molecular structure calculations. International Journal of Quantum Chemistry, 14(S12):109–130, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Levy:1978:PGV

[Lev78b] Mel Levy. Pointwise and generalized virial theorems. International Journal of Quantum Chemistry, 14(S12):343–344, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Lewis:1978:CPM

[Lew78] T. J. Lewis. Conduction in protein and methylglyoxal-protein complexes. International Journal of Quantum Chemistry, 14 (S5):149–158, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Laurenzi:1977:ECF

[LF77] Bernard J. Laurenzi and Alan Flamberg. Electronic computation of first-order wave functions using Green’s functions. International Journal of Quantum Chemistry, 11(5):869–880, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leite:1979:VCM

[LFF79] José R. Leite, Armando C. Ferraz, and L. G. Ferreira. Variational cellular method in molecular and crystal electronic structure calculations. International Journal of Quantum Chemistry, 16(S13):395–402, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Lindner:1970:GPE

[LG70] P. Lindner and O. Goscinski. Generalized polarizabilities and energy expressions. International Journal of Quantum Chemistry, 5(S4):251–255, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Lowdin:1971:SPT

[LG71] P.-O. Löwdin and O. Goscinski. Studies in perturbation theory. XIV. Treatment of constants of motion, degeneracies and symmetry properties by means of multidimensional partitioning. International Journal of Quantum Chemistry, 5(S5):685– 705, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Lissillour:1975:ECS

[LG75] R. Lissillour and C. R. Guerillot. Electronic correlation studies. III. Self-correlated field method. Application to ^{2}S ground state and ^{2}P excited state of three-electron atomic systems. International Journal of Quantum Chemistry, 9(4):627–634, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Liakus:1977:DSC

[LG77] A. Liakus and G. A. Gallup. A double self-consistent-field method for electron correlation: Applications to LiH, H_{2}O, and cyclopropane. International Journal of Quantum Chemistry, 11(2):213–231, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Langhoff:1979:PMO

[LGAM79] P. W. Langhoff, A. Gerwer, C. Asaro, and B. V. Mckoy. Photoionization in molecular oxygen. International Journal of Quantum Chemistry, 16(S13):645–657, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Lazukova:1977:EAS

[LGM77] N. I. Lazukova, V. A. Gubanov, and V. G. Mokerov. Electronic absorption spectra of V_{2}O_{5}. International Journal of Quantum Chemistry, 12(5):915–923, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lazukova:1975:EFG

[LGP75] N. I. Lazukova, V. A. Gubanov, and R. N. Pletnev. Electric field gradient calculation in V_{2}O_{5}. International Journal of Quantum Chemistry, 9(4):691–700, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Llaguno:1973:MOS

[LGR73] Claro T. Llaguno, Shiv K. Gupta, and Stuart M. Rothstein. Molecular orbital studies on small molecules using H_{2}^{+}-type elliptical basis orbitals. Application to H_{2}^{+}, H_{2}, He_{2}^{++} and H_{3}^{+}. International Journal of Quantum Chemistry, 7(4):819–834, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larson:1974:CUG

[LH74] Everett G. Larson and Andreas F. Hall. Concepts useful in the graphical representation of the Coulomb hole: Application to the description of the ground state of the four-electron boron ion B^{+}. International Journal of Quantum Chemistry, 8(S8): 49–57, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Langhoff:1975:QTR

[LH75] P. W. Langhoff and W. R. Heffner. Quantum theory of radiative transition phenomena. International Journal of Quantum Chemistry, 9(S9):461–470, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Langhoff:1976:BGF

[LH76] P. W. Langhoff and A. J. Hernandez. On the Brueckner and Goldstone forms of the linked-cluster theorem. International Journal of Quantum Chemistry, 10(S10):337–351, January 19– 24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lochmann:1977:PCC

[LH77] R. Lochmann and H.-J. Hofmann. PCILO Calculations on charge transfer complexes. International Journal of Quantum Chemistry, 11(3):427–433, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1979:PCT

[LH79] R. Lochmann and P. Hobza. PCILO calculations on “true” van der Waals complexes. International Journal of Quantum Chemistry, 15(1):73–81, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Loew:1977:MEA

[LHK77] Gilda H. Loew, Leonard M. Hjelmeland, and Robert F. Kirchner. Models for the enzymatically active state of cytochrome p-450. International Journal of Quantum Chemistry, 12(S4): 225–244, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Li:1976:CDM

[Li76] C. H. Li. Calculation of dipole moments for diatomic molecules with multiple-scattering Xα wave functions. International Journal of Quantum Chemistry, 10(S10):193–202, January 19– 24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Liberman:1976:ESM

[Lib76] David A. Liberman. Equation of state of molecular hydrogen at high pressure. International Journal of Quantum Chemistry, 10 (S10):297–303, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lim:1974:SEP

[Lim74a] T. K. Lim. SCF equations for pure spin states with manyparticle interactions. International Journal of Quantum Chemistry, 8(4):539–546, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1974:WHS

[Lim74b] T. K. Lim. Waller–Hartree spin-free method. International Journal of Quantum Chemistry, 8(4):523–537, May 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1975:SDW

[Lim75a] T. K. Lim. Spin densities in the Waller–Hartree spin-free method. International Journal of Quantum Chemistry, 9(3): 489–493, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1975:YOW

[Lim75b] T. K. Lim. Young operators and the Waller–Hartree spinfree method. International Journal of Quantum Chemistry, 9(6): 981–985, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1976:EPS

[Lim76a] T. K. Lim. Exclusion principle and the symmetry adapted wave-functions. International Journal of Quantum Chemistry, 10(2):299–305, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1976:MET

[Lim76b] T. K. Lim. Many-electron theory in the Waller–Hartree spinfree method. International Journal of Quantum Chemistry, 10 (5):791–797, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1977:RDA

[Lim77a] T. K. Lim. The role of double antisymmetrization. International Journal of Quantum Chemistry, 11(2):341–348, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1977:SGW

[Lim77b] T. K. Lim. The symmetric groups and the Waller–Hartree spinfree method. International Journal of Quantum Chemistry, 11 (3):393–398, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lindner:1970:LBE

[Lin70] Peter Lindner. Lower bounds for ϵ_{2} in the 1/z-expansion for heliumlike ions. International Journal of Quantum Chemistry, 4(2):217–218, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lindgren:1971:SEA

[Lin71] Ingvar Lindgren. A statistical exchange approximation for localized electrons. International Journal of Quantum Chemistry, 5(S5):411–420, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Lindner:1973:BRB

[Lin73] Peter Lindner. Book review: The quantum theory of atoms, molecules and photons. Author: John Avery. Published by: McGraw-Hill, London, 1972. Price £5·50. No. of pages: 378. International Journal of Quantum Chemistry, 7(5):1037, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Ave72].

Linderberg:1977:PCG

[Lin77] Jan Linderberg. A program for the calculation of ground state and excited state energy surfaces. International Journal of Quantum Chemistry, 12(S11):353–357, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Lindgren:1978:CCA

[Lin78] Ingvar Lindgren. A coupled-cluster approach to the many-body perturbation theory for open-shell systems. International Journal of Quantum Chemistry, 14(S12):33–58, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Lipari:1972:FPE

[Lip72] Nunzio O. Lipari. First principle energy bands and optical properties of insulating crystals. International Journal of Quantum Chemistry, 6(S6):403–410, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Liu:1971:PCL

[Liu71a] B. Liu. Potential curve for linear symmetric H_{3}. International Journal of Quantum Chemistry, 5(S5):123–131, January 18– 23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Liu:1971:MDR

[Liu71b] N. L. Huang Liu. Magnon dispersion relation in europium iron garnet. International Journal of Quantum Chemistry, 5(S5): 661–665, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Loew:1975:QCC

[LJBN75] Gilda H. Loew, J. Randal Jester, Donald Berkowitz, and Richard C. Newth. Quantum chemical calculations of flexible opiate narcotics: Methadone, meperidine, and prodines. International Journal of Quantum Chemistry, 9(S2):25–34, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Loew:1978:BNC

[LK78] Gilda H. Loew and Robert F. Kirchner. Binding of O_{2}, NO, and CO to model active sites in ferrous heme proteins: Ligand geometry, electronic structure, and quadrupole splittings. International Journal of Quantum Chemistry, 14(S5):403–415, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Liedtke:1977:FPC

[LKH77] Richard C. Liedtke, Dwayne L. Knirk, and Edward F. Hayes. Feshbach projection calculations of compound state resonances in atom-diatom scattering. International Journal of Quantum Chemistry, 12(S11):337–341, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Likhtenstein:1979:SPE

[LKK^{+}79] G. I. Likhtenstein, A. I. Kotelnikov, A. W. Kulikov, L. A. Syrtsova, V. R. Bogatyrenko, A. I. Melnikov, E. N. Frolov, and A. I. Berg. Some peculiarities of the electronic transfer in redox enzymes. International Journal of Quantum Chemistry, 16(3): 419–435, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1970:CLB

[LL70] P. O. Löwdin and T. K. Lim. Calculation of lower bounds to energy eigenvalues by reduced density matrices and the representability problem. International Journal of Quantum Chemistry, 3S(??):697–??, ?? 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Labarre:1972:SEC

[LL72] J.-F. Labarre and C. Leibovici. Structure électronique des complexes acide-base de Lewis. I. Structure électronique et conformation moléculaire des molécules F_{3} P·BH_{3} et F_{2}HP·BH_{3}. (French) [Electronic structure of Lewis acid-base complexes. I. Electronic structure and molecular conformation of the molecules F_{3}P·BH_{3} and F_{2}HP·BH_{3}]. International Journal of Quantum Chemistry, 6(4):625–637, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Labzowsky:1973:CMS

[LL73] L. N. Labzowsky and Yu. E. Lozovik. Conjugated molecules in strong magnetic fields. International Journal of Quantum Chemistry, 7(5):985–989, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laskowski:1975:APH

[LL75] Bernard Laskowski and Sten Lunell. Angularly projected Hartree–Fock calculation of the hyperfine interaction in Li 2^{2}P. International Journal of Quantum Chemistry, 9(S9):175–182, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Laki:1976:NET

[LL76] Koloman Laki and Jànos Ladik. A note on the “electronic theory” of cancer. International Journal of Quantum Chemistry, 10(S3):51–57, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lindner:1972:ESS

[LM72] P. Lindner and O. Mårtensson. Electronic structure of some odorant molecules of the benzaldehyde type. International Journal of Quantum Chemistry, 6(S6):363–366, January 17– 22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Lo:1975:OOM

[Lo75] B. W. N. Lo. Orbit-orbit molecular integrals over Gaussian orbitals. International Journal of Quantum Chemistry, 9(6): 1055–1067, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Linderberg:1977:SVP

[LÖ77] Jan Linderberg and Yngve Öhrn. State vectors and propagators in many-electron theory. A unified approach. International Journal of Quantum Chemistry, 12(1):161–191, July 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Locqueneux:1972:TQP

[Loc72] R. Locqueneux. Théorie quantique des phénomènes de biréfringence électrique et magnétique des molécules. (French) [Quantum theory of the phenomena of molecular electric and magnetic birefrigence]. International Journal of Quantum Chemistry, 6(1):1–21, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1976:CCU

[Loc76] R. Lochmann. Crystal calculation using the PCILO method. International Journal of Quantum Chemistry, 10(5):901–908, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1977:CII

[Loc77a] R. Lochmann. Calculation of intermolecular interactions within the PCILO framework using fixed wave-functions. International Journal of Quantum Chemistry, 11(2):293–300, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1977:CMC

[Loc77b] R. Lochmann. Calculation of molecular crystals within the PCILOCC framework. International Journal of Quantum Chemistry, 12(5):851–858, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1977:PCH

[Loc77c] R. Lochmann. PCILO calculations on hydrogen bonded complexes. International Journal of Quantum Chemistry, 12(5): 841–850, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1977:PCI

[Loc77d] R. Lochmann. Problems of the calculation of intermolecular interactions on complexes with strong delocalized π-bonds within the PCILO framework. International Journal of Quantum Chemistry, 12(5):795–803, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lohr:1976:HOR

[Loh76] Lawrence L. Lohr, Jr. Hybrid orbitals and the Runge–Lenz vector. International Journal of Quantum Chemistry, 10(5): 799–809, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lonneaus:1973:OAE

[Lon73] G. Lonneaus. Opening address at the Edward U. Condon symposium. International Journal of Quantum Chemistry, 7(S7): 1, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Lonaeus:1974:OA

[Lon74] Gunnar Lonaeus. Opening address. International Journal of Quantum Chemistry, 8(S8):1, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Lonaeus:1975:OS

[Lon75] Gunnar Lonaeus. Opening statement. International Journal of Quantum Chemistry, 9(S9):1, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Lindner:1973:SEI

[LÖS73] Peter Lindner, Yngve Öhrn, and John R. Sabin. A semiempirical investigation of the electronic structure and stability of the oxycumulenes. International Journal of Quantum Chemistry, 7(S7):261–267, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Lowdin:1970:SPI

[Löw70] Per-Olov Löwdin. Some properties of inner projections. International Journal of Quantum Chemistry, 5(S4):231–237, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Lowdin:1971:SPI

[Löw71a] P. O. Löwdin. Some properties of inner projections. International Journal of Quantum Chemistry, 4(??):231–237, ?? 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1971:I

[Löw71b] Per-Olov Löwdin. Introduction. International Journal of Quantum Chemistry, 5(S5):xv–xvi, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Lowdin:1973:O

[Löw73] Per-Olov Löwdin. Obituary. International Journal of Quantum Chemistry, 7(1):i, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1977:SCM

[Löw77] Per-Olov Löwdin. Some current models of carcinogenesis. International Journal of Quantum Chemistry, 12(S4):185–196, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1978:I

[Löw78] Per-Olov Löwdin. Introduction. International Journal of Quantum Chemistry, 14(4):iii–ii, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1979:P

[Löw79] Per-Olov Löwdin. Preface. International Journal of Quantum Chemistry, 16(1):1–2, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ledovskaya:1974:SHLa

[LP74a] E. M. Ledovskaya and M. I. Petrashen. Symmetry in the Heitler-London method. I. The connection of the molecular point group with its electron permutation group in the Heitler-London method. International Journal of Quantum Chemistry, 8(2):163–169, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ledovskaya:1974:SHLb

[LP74b] E. M. Ledovskaya and M. I. Petrashen. Symmetry in the Heitler-London method. II. The determination of the allowed molecular multiplets. International Journal of Quantum Chemistry, 8(2):171–177, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Linderberg:1974:DPH

[LP74c] Jan Linderberg and Domingo Prato. Dynamic polarizability of helium: a random phase approximation calculation. International Journal of Quantum Chemistry, 8(6):901–913, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lavery:1979:MQC

[LP79a] Richard Lavery and Bernard Pullman. Model quantum chemical studies of the electronic structure and aspects of reactivity of diol epoxides of aromatic hydrocarbons. International Journal of Quantum Chemistry, 15(3):271–280, March 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lavery:1979:TMSb

[LP79b] Richard Lavery and Bernard Pullman. A theoretical model study of the mechanism of action of the enzyme glyoxylase I. International Journal of Quantum Chemistry, 16(S6):467–474, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lavery:1979:TMSa

[LP79c] Richard Lavery and Bernard Pullman. Theoretical model study of the reactivity of benzo(a)pyrene diol epoxide with the amino groups of the nucleic acid bases. International Journal of Quantum Chemistry, 16(1):175–188, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lavery:1978:MQC

[LPP78] Richard Lavery, Alberte Pullman, and Bernard Pullman. Model quantum-chemical studies on the reaction between the candidate proximate carcinogen benzo(a)pyrene-7,8-dihydrodiol-9,10-epoxide and guanine. International Journal of Quantum Chemistry, 14(S5):21–34, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lombardi:1974:VTC

[LPT^{+}74] E. Lombardi, L. Pirola, G. Tarantini, L. Jansen, and R. Ritter. Validity of the three-center, four-electron model for stability of rare gas halides on the basis of exchange perturbation theory. International Journal of Quantum Chemistry, 8(3):335– 345, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Linder:1977:ABL

[LR77] Bruno Linder and David A. Rabenold. On the attraction between linear chain molecules and the many-body problem. International Journal of Quantum Chemistry, 12(S4):443–449, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lipari:1977:EIV

[LRD^{+}77] N. O. Lipari, M. J. Rice, C. B. Duke, R. Bozio, A. Girlando, and C. Pecile. Electron-intramolecular vibration coupling in TTF– TCNQ systems. International Journal of Quantum Chemistry, 12(S11):583–594, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Lipari:1978:EIV

[LRD^{+}78] N. O. Lipari, M. J. Rice, C. B. Duke, R. Bozio, A. Girlando, and C. Pecile. Electron-intramolecular vibration coupling in TTF-TCNQ systems. International Journal of Quantum Chemistry, 14(S12):545, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Lombardi:1973:EMB

[LRJ73] E. Lombardi, R. Ritter, and L. Jansen. Effect of many-body interactions on the stability of rare-gas halides. International Journal of Quantum Chemistry, 7(2):155–171, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larson:1970:DMA

[LS70] Everett G. Larson and Darwin W. Smith. A density matrix analysis of Slater’s hyper-Hartree–Fock method. International Journal of Quantum Chemistry, 5(S4):87–94, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Larson:1971:DMA

[LS71] E. G. Larson and D. W. Smith. A density matrix analysis of Slater’s hyper-Hartree–Fock method. International Journal of Quantum Chemistry, 4(??):87–??, ?? 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lindner:1972:EDS

[LS72] Peter Lindner and John R. Sabin. Electron distribution in a short A-type hydrogen bond. International Journal of Quantum Chemistry, 6(S6):301–304, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Linderberg:1974:MEA

[LS74a] Jan Linderberg and Lloyd Seamans. Matrix elements in all valence electron models. International Journal of Quantum Chemistry, 8(6):925–940, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Loubriel:1974:TOE

[LS74b] G. M. Loubriel and R. G. Selsby. A technique for orbital exponent optimization in ab initio HF SCF LCAO MO calculations. International Journal of Quantum Chemistry, 8(4):547– 557, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lozes:1979:EES

[LS79] Richard Lozes and John R. Sabin. Excited electronic states of a hydrogen bond: Bifluoride ion. International Journal of Quantum Chemistry, 16(2):273–275, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Loew:1979:AAC

[LSB^{+}79] Gilda Loew, B. S. Sudhindra, S. Burt, G. R. Pack, and R. Macelroy. Aromatic amine carcinogenesis: Activation and interaction with nucleic acid bases. International Journal of Quantum Chemistry, 16(S6):259–281, March 4–10, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ladik:1977:EBS

[LSOC77] János Ladik, Sándor Suhai, Peter Otto, and Thomas C. Collins. Estimation of the band structure of DNA on the basis of an ab initio SCF LCAO band structure of polycytosine. International Journal of Quantum Chemistry, 12(S4):55–63, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ladik:1978:ESB

[LSS78] János Ladik, Sándor Suhai, and Max Seel. Electronic structure of biopolymers and possible mechanisms of chemical carcinogenesis. International Journal of Quantum Chemistry, 14(S5):

35–49, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ley:1974:EAV

[LTR74] Barbara E. Ley, Tony Thorpe, and Stuart M. Rothstein. Error analysis of variationally optimized upper- and lower-bound wave functions for H. International Journal of Quantum Chemistry, 8(6):971–980, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ludena:1971:ECP

[Lud71] E. V. Ludeña. The electronic correlation problem and loge localization. International Journal of Quantum Chemistry, 5 (4):395–409, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ludena:1975:VPD

[Lud75] Eduardo V. Ludeña. Variational principles for discontinuous wave functions and the independent particle model of electronic structure. International Journal of Quantum Chemistry, 9(6): 1069–1085, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ludena:1977:ASL

[Lud77] Eduardo V. Ludeña. Atomic SCF loge localized wave functions. International Journal of Quantum Chemistry, 11(5):715–723, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lundqvist:1975:DET

[Lun75] S. Lundqvist. Dynamical extensions of the Thomas–Fermi theory. International Journal of Quantum Chemistry, 9(S9):23– 30, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Lundqvist:1977:SME

[Lun77] Stig Lundqvist. Strong many-electron effects in photoelectron spectra. International Journal of Quantum Chemistry, 12(S11):

379–388, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Lunell:1979:CUP

[Lun79] Sten Lunell. Comparison of UHF and PHF methods for hyperfine structure calculations. International Journal of Quantum Chemistry, 15(1):97–107, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laskowski:1971:POM

[LV71a] B. Laskowski and P. Van Leuven. Parametrization of the orbital mixing coefficients in the AMO method. International Journal of Quantum Chemistry, 5(S5):707–714, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Laskowski:1971:TFO

[LV71b] B. Laskowski and P. Van Leuven. Theta-function overlap approximation applied to the energy of cyclic polyenes. International Journal of Quantum Chemistry, 5(2):235–238, March 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Liu:1971:ESN

[LV71c] H. P. D. Liu and G. Verhaegen. Electronic states of NH and OH^{+}. International Journal of Quantum Chemistry, 5(S5):103– 118, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Laskowski:1972:TFO

[LV72] B. Laskowski and P. Van Leuven. Theta-function overlap approximation applied to the energy of cyclic polyenes. International Journal of Quantum Chemistry, 6(1):193–196, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lathouwers:1978:MSG

[LV78] L. Lathouwers and P. Van Leuven. Molecular spectra and the generator coordinate method. International Journal of Quantum Chemistry, 14(S12):371–375, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Larsson:1974:ESO

[LVDC74] Sven Larsson, Eeva-Kaarina Viinikka, Manoel L. De Siqueira, and John W. D. Connolly. The electronic structure of octahedral transition metal halides as calculated by the multiple scattering method. International Journal of Quantum Chemistry, 8(S8):145–160, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Lochmann:1976:CII

[LW76] R. Lochmann and T. Weller. Calculation of intermolecular interactions within the PCILO framework. International Journal of Quantum Chemistry, 10(5):909–916, September 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lin:1977:QMG

[LW77] Ying-Wei Lin and George Wolken, Jr. Quantum model for gassolid energy transfer: Role of phonon polarization and propagation direction. International Journal of Quantum Chemistry, 12 (S11):555–571, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Laughlin:1974:BWP

[LWA74] C. Laughlin, M. R. Woodward, and A. T. Amos. Brillouin– Wigner perturbation theory and the generalized eigenvalue equation. International Journal of Quantum Chemistry, 8(4): 491–498, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lyast:1978:APW

[Lya78] I. Ts. Lyast. The antisymmetrized product of weak nonorthogonal geminals. International Journal of Quantum Chemistry, 13(1):83–88, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lazzeretti:1977:CMS

[LZ77] P. Lazzeretti and R. Zanasi. Calculations of the magnetic shielding constants of heavy nuclei in polyatomic molecules. International Journal of Quantum Chemistry, 12(1):93–103, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lazzeretti:1979:USF

[LZ79] Paolo Lazzeretti and Riccardo Zanasi. On the use of symmetry in first-order perturbed HF theory. II. International Journal of Quantum Chemistry, 15(6):645–653, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Maggiora:1979:ARC

[Mag79] Gerald M. Maggiora. Assessment of reaction center special-pair chlorophyll models. International Journal of Quantum Chemistry, 16(2):331–352, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Maksic:1971:CJJ

[Mak71] Z. B. Maksić. Calculations of J_{C13−C13} and J_{C13−H1} coupling constants in hydrocarbons by the maximum overlap method. International Journal of Quantum Chemistry, 5(S5):301–306, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Malrieu:1971:HOCa

[Mal71a] J. P. Malrieu. High orders corrections to the van der Waals– London forces. I. A model problem: Two interacting hydrogen molecules is the minimal basis set. International Journal of Quantum Chemistry, 5(4):435–453, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Malrieu:1971:HOCb

[Mal71b] J. P. Malrieu. High orders corrections to the van der Waals– London forces. II. Interaction of two molecules with isotropic polarizabilities. International Journal of Quantum Chemistry, 5(4):455–467, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Manne:1977:LDE

[Man77] Rolf Manne. The linked-diagram expansion of the ground state of a many-electron system: A time-independent derivation. International Journal of Quantum Chemistry, 12(S11): 175–192, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Marinelli:1973:TOM

[Mar73a] Francis Marinelli. Theorie des orbitals moléculaires: recherche d’une base STO homogène pour les atomes alcalins et leur série isoélectronique. (French) [Molecular orbital theory: search for a homogeneous STO basis for the alkali atoms and their isoelectronic series]. International Journal of Quantum Chemistry, 7 (1):109–125, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Marois:1973:TPB

[Mar73b] M. (Maurice) Marois, editor. From theoretical physics to biology, proceedings, Versailles, June 21–26, 1971. S. Karger, Basel, Switzerland, 1973. ISBN 3-8055-1578-2. LCCN QH505 .I59 1971.

Marston:1976:WR

[Mar76] President Robert Q. Marston. Welcoming remarks. International Journal of Quantum Chemistry, 10(S3):1, January 12– 14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Maartensson:1978:BRB

[Mår78] Olle Mårtensson. Book review: Quantum pharmacology. W. G. Richards. Butterworths, London and Boston, 1977. price:

$24.95, 213 pp. International Journal of Quantum Chemistry, 13(4):563–564, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Ric77].

Marinov:1979:PET

[Mar79] B. S. Marinov. Photoinduced electron transport across proteincontaining membranes. International Journal of Quantum Chemistry, 16(4):739–743, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsuoka:1971:GEM

[Mat71a] Osamu Matsuoka. Gaussian expansion method for molecular integrals of molecular properties. International Journal of Quantum Chemistry, 5(1):1–11, January 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matthias:1971:S

[Mat71b] Bernd T. Matthias. Superconductivity in 1970. International Journal of Quantum Chemistry, 5(S5):771–772, January 18– 23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Matsen:1972:MSF

[Mat72a] F. A. Matsen. Magnetism and spin-free quantum chemistry. International Journal of Quantum Chemistry, 6(S6):411–417, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Matthias:1972:S

[Mat72b] Bernd T. Matthias. Superconductivity in 1971. International Journal of Quantum Chemistry, 6(S6):429, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Matsuoka:1973:MIR

[Mat73a] Osamu Matsuoka. Molecular integrals of relativistic effects with Gaussian-type orbitals. International Journal of Quantum Chemistry, 7(2):365–381, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matthias:1973:S

[Mat73b] Bernd T. Matthias. Superconductivity in 1972. International Journal of Quantum Chemistry, 7(S7):659, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Matsen:1974:UGF

[Mat74a] F. A. Matsen. The unitary group formulation of the N-particle problem. International Journal of Quantum Chemistry, 8(S8): 379–388, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Matthias:1974:S

[Mat74b] Bernd T. Matthias. Superconductivity in 1973. International Journal of Quantum Chemistry, 8(S8):499–500, January 20– 26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Matthias:1975:S

[Mat75] Bernd T. Matthias. Superconductivity in 1974. International Journal of Quantum Chemistry, 9(S9):579–580, January 19– 25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Matsen:1976:SFQa

[Mat76a] F. A. Matsen. Spin-free quantum chemistry. XVII. The Hückel– Hubbard theory and the Woodward–Hoffmann rules. International Journal of Quantum Chemistry, 10(3):511–524, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1976:SFQb

[Mat76b] F. A. Matsen. Spin-free quantum chemistry. XVIII. The unitary group formulation of the many-electron problem. International Journal of Quantum Chemistry, 10(3):525–542, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsuoka:1976:MIR

[Mat76c] Osamu Matsuoka. Molecular integrals of relativistic effects with Gaussian-type orbitals. International Journal of Quantum Chemistry, 10(2):379, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matthias:1976:SSS

[Mat76d] Bernd T. Matthias. Symmetries of superconducting sulfides. International Journal of Quantum Chemistry, 10(S10):435– 436, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Matthias:1977:STB

[Mat77] Bernd T. Matthias. Superconductivity of ternary borides. International Journal of Quantum Chemistry, 12(S11):647–649, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Matsen:1978:TTI

[Mat78] F. A. Matsen. A theory of thermal isomerization rates. International Journal of Quantum Chemistry, 14(S12):529–542, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Matsen:1979:UGF

[Mat79a] F. A. Matsen. The unitary-group formulation of coupledcluster many-electron theory. International Journal of Quantum Chemistry, 16(S13):305–309, March 11–17, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Matthias:1979:SSP

[Mat79b] Bernd T. Matthias. Superconductivity, p-state pairing, and magnetism. International Journal of Quantum Chemistry, 16 (S13):467–469, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Mayer:1974:SPEa

[May74a] I. Mayer. Spin projected EHF method: Calculations for a four-electron model system. International Journal of Quantum Chemistry, 8(3):363–372, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1974:SPEb

[May74b] I. Mayer. Spin-projected extended Hartree–Fock equations. II. Odd-electron systems. International Journal of Quantum Chemistry, 8(6):893–899, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1975:E

[May75] I. Mayer. Erratum. International Journal of Quantum Chemistry, 9(3):555, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1978:SPE

[May78] I. Mayer. Spin-projected EHF method. IV. Comparison of potential curves given by different one-electron methods. International Journal of Quantum Chemistry, 14(1):29–38, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukherjee:1971:SCC

[MBM71] P. K. Mukherjee, A. K. Bhattacharya, and A. Mukherji. Selfconsistent calculation of excited ^{3}P state wave functions of atoms. International Journal of Quantum Chemistry, 5(6): 647–655, November 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mendelsohn:1973:RHF

[MBM73a] Lawrence B. Mendelsohn, Frank Biggs, and Joseph B. Mann. Relativistic Hartree–Fock Compton profiles. International Journal of Quantum Chemistry, 7(S7):395–407, January 21– 27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Mortola:1973:ISP

[MBM73b] Albert P. Mortola, Harold Basch, and Jules W. Moskowitz. An Ab Initio study of the permanganate ion. International Journal of Quantum Chemistry, 7(4):725–737, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mishra:1979:EBE

[MBMM79] R. K. Mishra, K. Bhaumik, S. C. Mathur, and S. Mitra. Excitons and Bose–Einstein condensation in living systems. International Journal of Quantum Chemistry, 16(4):691–706, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mcdowell:1979:IBS

[Mcd79] Keith Mcdowell. The incomplete basis set problem. III. Virtual orbitals and the CIBS expansion of Hartree–Fock energies. International Journal of Quantum Chemistry, 16(S13): 51–58, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Men:1970:GTM

[MCF^{+}70] A. N. Men, V. I. Cherepanov, D. S. Farberov, V. Ja. Mitrofanov, and Gr. I. Chufarov. Group-theoretical method for determining permitted terms of the electronic states of polyatomic molecules taking account of spin-orbit interaction. International Journal of Quantum Chemistry, 4(1):109–119, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Men:1973:GTM

[MCM73] B. A. Men, V. L. Cherepanov, and A. N. Men. Group theoretical methods for determining permitted terms of the electronic states of complexes in crystals. III. The use of plethysm for classification of the permitted terms of impurity complexes in crystal. International Journal of Quantum Chemistry, 7(4): 739–743, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1973:SFQc

[MCP73] F. A. Matsen, J. G. Cosgrove, and J. M. Picone. Spin-free quantum chemistry. XIV. The infinite interaction range model for ferromagnetism. International Journal of Quantum Chemistry, 7(6):1077–1090, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matzke:1972:CSE

[MCST72] P. Matzke, O. Chacon, E. Sanhueza, and M. Trsic. Comparison of several expansions in the calculation of static electric dipole π polarizability of conjugated molecules by perturbation theory. The ground and the first excited singlet states. International Journal of Quantum Chemistry, 6(3):407–412, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moiseyev:1978:CCS

[MCW78] N. Moiseyev, P. R. Certain, and F. Weinhold. Complexcoordinate studies of helium autoionizing resonances. International Journal of Quantum Chemistry, 14(6):727–736, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Momany:1978:CEC

[MDA78] F. A. Momany, L. G. Drake, and J. R. AuBuchon. Conformational energy calculations on the growth hormone inhibitor: Somatostatin. International Journal of Quantum Chemistry, 14(S5):381–391, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mezey:1979:DAI

[MDC79] Paul G. Mezey, Raymond Daudel, and Imre G. Csizmadia. Dependence of approximate ab initio molecular loge sizes on the quality of basis functions. International Journal of Quantum Chemistry, 16(5):1009–1019, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morrison:1975:EKT

[MDS75] Robert C. Morrison, Orville W. Day, and Darwin W. Smith. An extension of Koopmans’ theorem III. Ionization energies of the open-shell atoms Li and B. International Journal of Quantum Chemistry, 9(S9):229–235, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Miles:1978:CEP

[ME78] Douglas L. Miles and Henry Eyring. Conformational effects of purine N3 electronic properties on drug design and cyclic nucleotide metabolism. International Journal of Quantum Chemistry, 14(S5):173–190, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mehler:1973:IPP

[Meh73a] E. L. Mehler. Independent pair-potential correlated wave functions. International Journal of Quantum Chemistry, 7(S7):437– 442, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Mehler:1973:PGF

[Meh73b] E. L. Mehler. Product of group-function expansions for correlated wave functions. International Journal of Quantum Chemistry, 7(2):187–193, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mehler:1978:SCN

[Meh78] Ernest L. Mehler. Self-consistent, nonorthogonal group function approximation: Application to two interacting, rigid HF molecules. International Journal of Quantum Chemistry, 14 (S12):407–417, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Mehler:1979:ICP

[Meh79] E. L. Mehler. Ab-initio calculations on protein secondary structure: Preliminary studies of hydrogen bonding in the α helix and the 3_{10} helix. International Journal of Quantum Chemistry, 16(S6):377–379, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Memory:1975:ESP

[Mem75] J. D. Memory. Electrophilic superdelocalizability and the Pullman theory of carcinogenesis. International Journal of Quantum Chemistry, 9(S2):179–180, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Memory:1977:SCR

[Mem77] J. D. Memory. A suggestion concerning the role of the L-region in carcinogenesis by polycyclic hydrocarbons. International Journal of Quantum Chemistry, 11(1):179, January 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Memory:1979:ESC

[Mem79] J. D. Memory. Electrophilic superdelocalizability and carcinogenesis by polycyclic aromatic hydrocarbons — Pullman theory. International Journal of Quantum Chemistry, 15(4):363– 368, April 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Menne:1971:PCH

[Men71] T. J. Menne. Phonon contribution to the hyperfine coupling constant. International Journal of Quantum Chemistry, 5(S5): 295–299, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Messmer:1973:COX

[Mes73a] R. P. Messmer. A calculation of the optical and X-ray photoemission transitions of K_{2}PtCl_{4}. International Journal of Quantum Chemistry, 7(S7):371–376, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Mestechkin:1973:SDS

[Mes73b] M. M. Mestechkin. Some developments in the spin-extended Hartree–Fock method. International Journal of Quantum Chemistry, 7(3):425–457, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mestechkin:1978:RHF

[Mes78] M. M. Mestechkin. Restricted Hartree–Fock method instability. International Journal of Quantum Chemistry, 13(4):469–481, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mestechkin:1979:ITP

[Mes79] M. Mestechkin. Instability threshold and peculiar solutions of Hartree–Fock equations. International Journal of Quantum Chemistry, 15(6):601–610, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Massa:1970:GMO

[MEW70] L. J. Massa, S. Ehrenson, and M. Wolfsberg. Gaussian molecular orbital calculations of the barrier to internal rotation in the ethyl cation. International Journal of Quantum Chemistry, 4 (6):625–630, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Meyer:1971:IEW

[Mey71] Wilfried Meyer. Ionization energies of water from PNO-CI calculations. International Journal of Quantum Chemistry, 5(S5): 341–348, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Matsen:1973:SFQa

[MF73] F. A. Matsen and A. L. Ford. Spin-free quantum chemistry. XII. Coarse structure magnetic theory. International Journal of Quantum Chemistry, 7(6):1051–1061, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morovic:1976:PNQ

[MFRE76] T. Morović, B. Fricke, A. Rosén, and D. E. Ellis. The problem of noncharacteristic quasimolecular X-rays in heavy ion collision. International Journal of Quantum Chemistry, 10(S10): 227–230, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mrozek:1977:SSE

[MG77] J. Mrozek and A. Gołebiewski. SCF scheme for excited states and analysis of the orbital relaxation effect for H_{2}O and FNO. International Journal of Quantum Chemistry, 12(1):207–213, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matthai:1978:FFB

[MGM78] C. C. Matthai, P. J. Grout, and N. H. March. Force fields in d-band metals. International Journal of Quantum Chemistry, 14(S12):443–459, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Malenkov:1979:III

[MGT79] G. G. Malenkov, A. V. Gagua, and V. P. Timofeev. Influence of intermolecular interactions on the DNA conformation. International Journal of Quantum Chemistry, 16(3):655–668, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Monkhorst:1972:ACF

[MH72] Hendrik J. Monkhorst and Frank E. Harris. Accurate calculation of Fourier transform of two-center Slater orbital products. International Journal of Quantum Chemistry, 6(4):601– 607, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Michels:1970:MOS

[MHA70] H. H. Michels, F. E. Harris, and J. B. Addison. Molecular orbital studies of the stability of CH_{5}^{+} and BH_{5}. International Journal of Quantum Chemistry, 5(S4):149–151, January 19– 24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Murphy:1974:MOS

[MHH74] M. F. Murphy, V. B. Haarstad, and F. B. Hahn. Molecular orbital studies of the mast cell zinc-histamine storage complex. International Journal of Quantum Chemistry, 8(S1):149– 157, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Marron:1970:SHF

[MHPS70] M. T. Marron, N. C. Handy, R. G. Parr, and H. J. Silverstone. Solution of the Hartree–Fock problem by expansion onto nested bases. International Journal of Quantum Chemistry, 4(3):245– 255, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

McIntosh:1976:LDS

[MHRS76] Harold V. McIntosh, Michael Hehenberger, and Rodolfo Reyes-Sanchez. Lattice dynamics with second neighbor interactions. II. Green’s formula. International Journal of Quantum Chemistry, 10(2):191–216, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

McIntosh:1977:LDS

[MHRS77] Harold V. McIntosh, Michael Hehenberger, and Rodolfo Reyes-Sanchez. Lattice dynamics with second-neighbor interactions. III. Green’s matrix. International Journal of Quantum Chemistry, 11(2):189–211, February 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Michels:1978:ESN

[MHW78] H. H. Michels, R. H. Hobbs, and L. A. Wright. Electronic structure of the noble-gas dimer ions. International Journal of Quantum Chemistry, 14(S12):257–269, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Michels:1978:ESE

[MHWC78] H. H. Michels, R. H. Hobbs, L. A. Wright, and J. W. D. Connolly. Electronic structure of excimer molecular lasers. International Journal of Quantum Chemistry, 13(2):169–187, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morikawa:1978:MMO

[MI78] Tetsuo Morikawa and Y. J. I’haya. Maximum and minimum overlap, localized, and hybrid orbitals for atoms and molecules by means of orthonormality-constrained variation. International Journal of Quantum Chemistry, 13(2):199–206, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Micha:1974:RDT

[Mic74] David A. Micha. Recent developments in the theory of reactive molecular collisions. International Journal of Quantum Chemistry, 8(S8):263–269, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Micha:1976:OFR

[Mic76a] David A. Micha. Operator formalisms of reactive molecular scattering. International Journal of Quantum Chemistry, 10 (S10):259–266, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Michl:1976:ESR

[Mic76b] Josef Michl. π-electron systems: Recent developments in electronic spectroscopy and the PPP model. International Journal of Quantum Chemistry, 10(S10):107–117, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Miller:1971:ESA

[Mil71] Kenneth J. Miller. Electron scattering off atoms for excitations to valence and Rydberg states. International Journal of Quantum Chemistry, 5(S5):71–81, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Miller:1974:PMS

[Mil74] Kenneth J. Miller. Partitioning of molecular systems and subsystem replacement by single-centered perturbations: I. A partitioning theorem and the pseudo substituent technique. International Journal of Quantum Chemistry, 8(S8):179–186, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Mintmire:1979:LMX

[Min79] J. W. Mintmire. LCAO methods in the Xα formalism: A comparison of results for the F_{2} molecule. International Journal of Quantum Chemistry, 16(S13):163–168, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Mitra:1972:EMH

[MK72] R. P. Mitra and K. L. Kapoor. Effect of methyl and hydroxyl substituents on the π-electronic spectra and ionization potential of p-benzoquinone. International Journal of Quantum Chemistry, 6(2):387–399, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1975:SPEb

[MK75a] I. Mayer and M. Kertész. Spin-projected EHF method. III. Applications to π-electron systems. International Journal of Quantum Chemistry, 9(3):527–536, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1975:SPEa

[MK75b] I. Mayer and J. Kondász. Spin-projected EHF method. II. The equations for successive optimization of the orbitals in the many-electron case. International Journal of Quantum Chemistry, 9(3):517–526, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1976:CDD

[MK76a] I. Mayer and M. Kertész. A comparison of different DODS methods when the number of electrons increases. International Journal of Quantum Chemistry, 10(6):961–966, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moiseyev:1976:SCS

[MK76b] N. Moiseyev and J. Katriel. Stability, continuity, and symmetry of variational wave-functions. International Journal of Quantum Chemistry, 10(6):1011–1023, November 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mestechkin:1978:SDO

[MK78] M. M. Mestechkin and G. T. Klimko. Spin-dependent operators in the spin-free quantum chemistry. International Journal of Quantum Chemistry, 13(5):579–596, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Makhanek:1977:LTM

[MKF77] A. G. Makhanek, V. S. Korolkov, and A. F. Fedorov. Laplace transform method in two-photon spectroscopy theory. International Journal of Quantum Chemistry, 12(6):971–983, December 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moiseyev:1977:WBQ

[MKT77] N. Moiseyev, J. Katriel, and C. Trindle. A weakly bound, quantum-mechanical Buridan’s ass. International Journal of Quantum Chemistry, 12(4):759–763, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mohraz:1976:COP

[ML76] Manijeh Mohraz and Lawrence L. Lohr, Jr. A comparison of orthogonalized plane wave and augmented plane wave methods for calculating photodetachment cross-sections. International Journal of Quantum Chemistry, 10(5):811–835, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Meunier:1979:DCP

[ML79] A. Meunier and B. Levy. Difficulties in the computation of physical observables due to a truncated CI. International Journal of Quantum Chemistry, 16(5):955–972, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Millie:1972:AEC

[MLB72] Philippe Millié, Bernard Lévy, and Gaston Berthier. All electron calculations of open-shell polyatomic molecules. III. Perturbation treatment of the spin polarization in vinyl and methyl radicals. International Journal of Quantum Chemistry, 6(1): 155–166, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1973:SPE

[MLB73] I. Mayer, J. Ladik, and G. Biczó. Spin projected extended Hartree–Fock equations. International Journal of Quantum Chemistry, 7(3):583–608, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Meunier:1976:PMC

[MLB76] A. Meunier, B. Levy, and G. Berthier. The N^{2} problem in molecular CI calculations. International Journal of Quantum Chemistry, 10(6):1061–1070, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Men:1975:SCS

[MLM75] B. A. Men, M. L. Leschinsky, and A. N. Men. The “seniority” classification of spin Hamiltonian and states of complexes and molecules in approximation of one-centered plethysms. International Journal of Quantum Chemistry, 9(4):669–681, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukherjee:1970:MCS

[MM70] N. G. Mukherjee and R. McWeeny. Multi-configuration SCF calculations on LiH. International Journal of Quantum Chemistry, 4(1):97–107, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mcguire:1972:CCI

[MM72a] Paul Mcguire and David A. Micha. Coupled-channel investigation of rotationally and vibrationally inelastic collisions between He and H_{2}. International Journal of Quantum Chemistry, 6(S6):111–132, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Moitra:1972:SCC

[MM72b] R. K. Moitra and P. K. Mukherjee. Self consistent calculation of dynamic multipole polarizabilities and excited state wave functions of open shell ions: Li sequence. International Journal of Quantum Chemistry, 6(2):211–224, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matcha:1974:UTT

[MM74] Robert L. Matcha and Mac B. Milleur. Unified treatment of two-electron integrals. II. One-center, electron repulsion, spinspin and spin-orbit integrals containing r_{ij}. International Journal of Quantum Chemistry, 8(5):635–642, September 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1977:SM

[MM77a] Joseph Edward Mayer and Maria Goeppert Mayer. Statistical mechanics. John Wiley, New York, NY, USA, 2d edition, 1977. ISBN 0-471-57985-8. xv + 491 pp. LCCN QC174.8 .M37 1977.

Mueller:1977:PGS

[MM77b] F. M. Mueller and H. W. Myron. Point-group symmetry and transition metal superconductivity: Nb Mo Tc alloys. International Journal of Quantum Chemistry, 11(6):1035–1041, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Maeder:1970:AEC

[MMB70] Fabienne Maeder, Philippe Millie, and Gaston Berthier. Allelectron calculations of open-shell polyatomic molecules: II. Construction of quasi-localized molecular orbitals in vinyl and formaldiminoxyl radicals. International Journal of Quantum Chemistry, 5(S4):179–186, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Martin:1978:TMC

[MMD78] Richard L. Martin, Larry E. McMurchie, and Ernest R. Davidson. Transition moment closure tests for ethylene. International Journal of Quantum Chemistry, 13(2):161–168, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Makarov:1979:IIG

[MME79] A. A. Makarov, D. R. Monaselidze, and N. G. Esipova. Intermolecular interactions of globular proteins in the crystal state. International Journal of Quantum Chemistry, 16(3):437–444, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukherjee:1971:SCV

[MMM71] P. K. Mukherjee, R. K. Moitra, and A. Mukherji. A selfconsistent variation-perturbation theory for open shell atomic systems: Calculation of polarizabilities and shielding factors. International Journal of Quantum Chemistry, 5(6):637–646, November 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukhopadhyay:1975:NPH

[MMM75] Atri Mukhopadhyay, Raj Kumar Moitra, and Debashis Mukherjee. On a new partitioning of the Hamiltonian in manybody calculation of pair-correlation energies in closed-shell systems. International Journal of Quantum Chemistry, 9(3):545– 554, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mortola:1974:AMS

[MMR74] Albert P. Mortola, Jules W. Moskowitz, and Notker Rösch. Application of the multiple scattering Xα method to the dirheniumoctachloride anion Re_{2}Cl_{8}^{2−}. International Journal of Quantum Chemistry, 8(S8):161–167, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Moitra:1970:SDQ

[MMS70] R. K. Moitra, P. K. Mukherjee, and S. Sengupta. A study of dynamic quadrupolar and octupolar excitations and calculations on excited d and f states of atom and ions: He-sequence. International Journal of Quantum Chemistry, 4(5):465–486, September 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1979:SFQ

[MN79] F. A. Matsen and C. J. Nelin. Spin-free Quantum Chemistry. XXI. Hartree–Fock theory. International Journal of Quantum Chemistry, 15(6):751–767, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Merrifield:1977:RRP

[MO77] D. L. Merrifield and N. S. Ostlund. A rigid-rotor potential surface for the hydrogen molecule dimer. International Journal of Quantum Chemistry, 12(S11):245–250, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Moccia:1973:TDV

[Moc73] R. Moccia. Time-dependent variational principle. International Journal of Quantum Chemistry, 7(4):779–783, July 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moccia:1974:SDF

[Moc74] R. Moccia. Static and dynamic first- and second-order properties by variational wave functions. International Journal of Quantum Chemistry, 8(2):293–314, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moffat:1979:CII

[Mof79] J. B. Moffat. Cyanide–isocyanide isomerization in the structural isomers of cyanogen isocyanate. International Journal of Quantum Chemistry, 15(5):547–557, May 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mohammad:1977:SVE

[Moh77] S. Noor Mohammad. Selected valence electron SCF MO calculations on alkali metal diatomics. International Journal of Quantum Chemistry, 12(5):813–825, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mohammad:1978:LPE

[Moh78] S. Noor Mohammad. London potential-energy surfaces for symmetrical alkali trimers. International Journal of Quantum Chemistry, 13(3):429–442, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mohammad:1979:IIP

[Moh79] S. Noor Mohammad. Improved interaction potential for alkali halide molecules. International Journal of Quantum Chemistry, 15(6):769–782, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Monkhorst:1977:CPC

[Mon77] Hendrik J. Monkhorst. Calculation of properties with the coupled-cluster method. International Journal of Quantum Chemistry, 12(S11):421–432, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Morikawa:1975:OFF

[Mor75] Tetsuo Morikawa. Operator formalism for forming ket vectors under some constraints. International Journal of Quantum Chemistry, 9(6):969–973, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morikawa:1978:MCV

[Mor78] Tetsuo Morikawa. Metric-constrained variation method for atoms and molecules. International Journal of Quantum Chemistry, 14(3):313–318, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moseley:1973:IQC

[Mos73] W. D. Moseley, Jr. Integrals of quantum chemical interest in bases of Laguerre functions. I. Functions of order 2l + 2. International Journal of Quantum Chemistry, 7(2):285–300, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Misurkin:1975:ESC

[MOV75] I. A. Misurkin, A. A. Ovchinnikov, and G. A. Vinogradov. Electronic structure and collective excitations of long molecules with conjugation in the main chain. International Journal of Quantum Chemistry, 9(4):605–625, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mantione:1971:QMI

[MP71a] M.-J. Mantione and B. Pullman. A quantum mechanical investigation of the bathochromic shift in visual pigments. International Journal of Quantum Chemistry, 5(S5):349–360, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Meath:1971:ADR

[MP71b] William J. Meath and Edwin A. Power. Angular dependence of the retarded dispersion energy. International Journal of Quantum Chemistry, 5(5):549–556, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

McMahon:1975:PRS

[MP75] James E. McMahon and James M. Pipas. Predicting RNA secondary interactions from primary sequence. International Journal of Quantum Chemistry, 9(S2):129–131, January 16– 18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Moskalenko:1979:BCP

[MPMK79] S. A. Moskalenko, E. P. Pokatilov, M. F. Miglei, and E. S. Kiselyova. Bose condensation of phonons in biological systems. International Journal of Quantum Chemistry, 16(4):745–752, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Markovsky:1979:QCI

[MPN79] L. N. Markovsky, O. M. Polumbrik, and A. M. Nesterenko. Quantum-chemical investigation of spatial and electronic structure of verdazyl and its derivatives. International Journal of Quantum Chemistry, 16(4):891–895, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Millie:1970:NTB

[MPS70] Philippe Millie, Lionel Praud, and Josiane Serre. The nature of the “triple bond” of 1,2-benzyne. International Journal of Quantum Chemistry, 5(S4):187–193, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Matsen:1975:SFQa

[MPW75] F. A. Matsen, J. M. Picone, and T. L. Welsher. Spin-free quantum chemistry. XV Spin-only neutron diffraction. International Journal of Quantum Chemistry, 9(1):157–169, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mishra:1972:VES

[MR72] P. C. Mishra and D. K. Rai. Variable electronegativity SCF-MO calculations on the electronic structure and spectra of some substituted benzenes I. Monosubstitutions. International Journal of Quantum Chemistry, 6(1):47–55, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Majorino:1973:AMP

[MR73] G. F. Majorino and C. Rubino. On the accuracy of the manypoints local procedures. International Journal of Quantum Chemistry, 7(4):675–686, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Malli:1971:SSS

[MS71a] Gulzari Malli and K. M. S. Saxena. Spin-spin splitting factors for f^{6} electron configurations. International Journal of Quantum Chemistry, 5(S5):605–612, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Martin:1971:DES

[MS71b] T. W. Martin and L. L. Swift, Jr. Discovery and ESR spectra of matrix stabilized hydronium radicals H_{3}O and D_{3}O. International Journal of Quantum Chemistry, 5(S5):371, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Musher:1971:CRI

[MS71c] Jeremy I. Musher and Jerome M. Schulman. On the calculation of Raman intensities. International Journal of Quantum Chemistry, 5(S5):183–190, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Mathur:1972:MES

[MS72a] S. C. Mathur and Jai Singh. Molecular electronic structure of metal phthalocyanines. International Journal of Quantum Chemistry, 6(1):57–81, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mathur:1972:MOZ

[MS72b] S. C. Mathur and Jai Singh. Molecular orbitals of zinc phthalocyanine. International Journal of Quantum Chemistry, 6 (4):747–759, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mathur:1974:MME

[MS74] S. C. Mathur and Jai Singh. More on molecular electronic structure of metal phthalocyanines. International Journal of Quantum Chemistry, 8(1):79–82, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Murty:1975:MCO

[MS75] J. S. Murty and C. R. Sarma. A method for the construction of orthogonal spin eigenfunctions. International Journal of Quantum Chemistry, 9(6):1097–1107, November 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Messmer:1976:RAS

[MS76a] R. P. Messmer and D. R. Salahub. Recent applications of the SCF–Xα–SW method to some inorganic systems. International Journal of Quantum Chemistry, 10(S10):183–191, January 19– 24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mintmire:1976:IPS

[MS76b] John W. Mintmire and John R. Sabin. Intermolecular potential studies of hydrogen-molecule interactions with rare-gas atoms. International Journal of Quantum Chemistry, 10(S10): 213–221, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mathur:1977:TPT

[MSKM77] S. C. Mathur, D. C. Singh, B. Kumar, and S. Mitra. Twoparameter ω-technique for MO calculations. International Journal of Quantum Chemistry, 11(5):759–766, May 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morrison:1973:PSM

[MSL73] Robert C. Morrison, Darwin W. Smith, and Everett G. Larson. Partitioning schemes for the 2-matrix and the pair density. International Journal of Quantum Chemistry, 7(5):837–852, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukherjee:1970:SCC

[MSM70] P. K. Mukherjee, S. Sengupta, and A. Mukherji. Self-consistent calculation of excited ^{1}P state wave functions of atoms. International Journal of Quantum Chemistry, 4(2):139–147, March 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1973:SFQb

[MSP73] F. A. Matsen, J. E. Suger, and J. M. Picone. Spin-free quantum chemistry. XIII. Spin waves. International Journal of Quantum Chemistry, 7(6):1063–1075, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mineyev:1979:OOI

[MSSF79] A. P. Mineyev, Yu. A. Sharonov, N. A. Sharonova, and V. A. Figlovsky. Orbit–orbit interaction in the NO–Fe(II) hemoproteins studied by the low-temperature magnetic circular dichroism: Dynamic Jahn–Teller effect detected by MCD. International Journal of Quantum Chemistry, 16(4):883–889, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Muftakhova:1976:CES

[Muf76] F. I. Muftakhova. On calculation of energy states of polyatomic molecules and clusters in a solid. International Journal of Quantum Chemistry, 10(6):1077–1079, November 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukherjee:1974:CCE

[Muk74] Debashis Mukherjee. A comparative calculation on excited state energies of some conjugated hydrocarbons. International Journal of Quantum Chemistry, 8(2):247–253, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukherjee:1975:OOS

[Muk75] Debashis Mukherjee. Orbital optimisation in single open-shell configurations: A sequential unconstrained minimisation technique. International Journal of Quantum Chemistry, 9(6): 943–949, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mulliken:1971:GSB

[Mul71a] Robert S. Mulliken. The ground state of BeH. International Journal of Quantum Chemistry, 5(S5):95–101, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Mulliken:1971:LES

[Mul71b] Robert S. Mulliken. The lowest excited states of BH. International Journal of Quantum Chemistry, 5(S5):83–94, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Mulliken:1974:MON

[Mul74] Robert S. Mulliken. Molecular orbitals of nitrogen at small internuclear distances. International Journal of Quantum Chemistry, 8(5):817–821, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Men:1975:MCO

[MVM75] B. A. Men, P. T. Varshavsky, and A. N. Men. Many-centered and one-centered plethysms in the problem of many centers. International Journal of Quantum Chemistry, 9(4):657–668, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mehler:1975:OCE

[MVN75] E. L. Mehler, G. A. Van Der Velde, and W. C. Nieuwpoort. Orbital correlation effects II. Potential curve and ionization potential of boron hydride. International Journal of Quantum Chemistry, 9(S9):245–254, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Maggiora:1974:ESP

[MW74a] Gerald M. Maggiora and Ludwik J. Weimann. Electronic structure of porphyrins. III. All-valence-electron SCF–MO–CI calculations of the excited singlet states of dianion and free base reduced porphins. International Journal of Quantum Chemistry, 8(S1):179–195, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mestechkin:1974:MFG

[MW74b] M. M. Mestechkin and G. E. Whyman. Matrix formulation of the generalized Hartree–Fock methods. International Journal of Quantum Chemistry, 8(1):45–60, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1975:SFQb

[MW75a] F. A. Matsen and T. L. Welsher. Spin-free quantum chemistry. XVI. Spin correlation. International Journal of Quantum Chemistry, 9(1):171–188, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mestechkin:1975:DMO

[MW75b] M. M. Mestechkin and G. E. Whyman. Density matrix in the open shell theory. International Journal of Quantum Chemistry, 9(5):761–773, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Minchin:1975:EOE

[MW75c] P. E. H. Minchin and B. G. Wybourne. Effective operators and the Z-expansion method. International Journal of Quantum Chemistry, 9(1):23–34, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1977:SFQb

[MW77] F. A. Matsen and T. L. Welsher. Spin-free quantum chemistry. XX. The alternancy quantum number. International Journal of Quantum Chemistry, 12(6):1001–1015, December 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1977:SFQa

[MWY77] F. A. Matsen, T. L. Welsher, and B. Yurke. Spin-free quantum chemistry. XIX. Particle-hole and pairing symmetries. International Journal of Quantum Chemistry, 12(6):985–999, December 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1977:UQG

[MYTC77] P. G. Mezey, K. Yates, G. Theodorakopoulos, and I. G. Csizmadia. Uniform quality Gaussian basis sets for organo-silicon compounds. International Journal of Quantum Chemistry, 12 (2):247–254, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nalewajski:1978:SAV

[Nal78] Roman F. Nalewajski. Some applications of the virial theorem to molecular force fields: The zero virial reaction coordinate and diatomic potentials from the normalized kinetic field functions. International Journal of Quantum Chemistry, 14(S12): 87–102, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Nam:1975:NMC

[Nam75] Sang Boo Nam. A new method for calculating the g factor in a solid. International Journal of Quantum Chemistry, 9(S9):551– 553, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Nir:1972:PC

[NAR72] Shlomo Nir, Steve Adams, and Robert Rein. On polarizability calculations. International Journal of Quantum Chemistry, 6(S6):295–300, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Nauckhoff:1977:IR

[Nau77] Carl-Henric Nauckhoff. Introductory remarks. International Journal of Quantum Chemistry, 12(S11):11–14, January 16– 22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Nauckhoff:1978:IR

[Nau78] Carl Henric Nauckhoff. Introductory remarks. International Journal of Quantum Chemistry, 14(S12):1–4, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Nazaroff:1972:TAL

[Naz72] George V. Nazaroff. A theory of autoionization lifetimes. International Journal of Quantum Chemistry, 6(S6):87–102, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Naqvi:1972:NCR

[NB72] K. Razi Naqvi and W. Byers Brown. The non-crossing rule in molecular quantum mechanics. International Journal of Quantum Chemistry, 6(2):271–279, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nicolaides:1978:TDC

[NB78] Cleanthes A. Nicolaides and Donald R. Beck. Time dependence, complex scaling, and the calculation of resonances in manyelectron systems. International Journal of Quantum Chemistry, 14(4):457–513, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Neely:1975:UMO

[Nee75] W. Brock Neely. The use of molecular orbital calculations as an aid in screening aromatic amines as potential carcinogens. International Journal of Quantum Chemistry, 9(S2):171– 177, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Nelin:1979:UGF

[Nel79] C. J. Nelin. The unitary-group formulation of the open-shell random-phase approximation. International Journal of Quantum Chemistry, 16(S13):311–320, March 11–17, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Nerbrant:1975:AMB

[Ner75] Per-Olof Nerbrant. Application of many-body Green’s functions to the calculation of molecular ionization potentials. International Journal of Quantum Chemistry, 9(5):901–916, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nerbrant:1976:UPT

[Ner76] Per-Olof Nerbrant. The use of partitioning technique in the solution of operator equations for ionizations and excitations. International Journal of Quantum Chemistry, 10(2):233–247, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nesbet:1970:EPC

[Nes70] R. K. Nesbet. Electronic pair correlation in atoms and molecules. International Journal of Quantum Chemistry, 5 (S4):117–125, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Nesbet:1971:EPC

[Nes71] R. K. Nesbet. Electronic pair correlation in atoms and molecules. International Journal of Quantum Chemistry, 4(??): 117–??, ?? 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nesbet:1975:AMR

[Nes75] R. K. Nesbet. Atom-molecule reactive collisions: A combined experimental and theoretical research project. International Journal of Quantum Chemistry, 9(S9):397–401, January 19– 25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Nesbet:1977:CEE

[Nes77] R. K. Nesbet. Correlation effects in electron scattering and attachment by open-shell atoms. International Journal of Quantum Chemistry, 12(S11):263–269, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Norbeck:1973:FRH

[NG73] J. M. Norbeck and G. A. Gallup. The first-row hydrides and influences of orbital scaling on formally charged valence bond structures. International Journal of Quantum Chemistry, 7 (S7):161–168, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Nir:1977:DMS

[NGR^{+}77] Shlomo Nir, Ramon Garduño, Robert Rein, Yves Coeckelenbergh, Robert D. Macelroy, and John T. Egan. Display, manipulation, and simulation of macromolecules. International Journal of Quantum Chemistry, 12(S4):135–141, January 9– 12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Nakayama:1970:MCLa

[NI70a] M. Nakayama and Y. J. I’Haya. A multi-configuration LCAO– MO study for complex unsaturated molecules. I. General theory and its application to the benzene anion. International Journal of Quantum Chemistry, 4(1):21–42, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nakayama:1970:MCLb

[NI70b] M. Nakayama and Y. J. I’Haya. A multi-configuration LCAO– MO study for complex unsaturated molecules. II. Application to the benzene cation. International Journal of Quantum Chemistry, 4(1):43–55, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Narita:1974:DBO

[NI74] Susumu Narita and Y. J. I’haya. Dual-basis orthonormalityconstrained variation method. International Journal of Quantum Chemistry, 8(5):765–775, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nir:1975:ACF

[Nir75] Shlomo Nir. Application of continuum field approach to calculation of hydrogen bonding energies in liquid water. International Journal of Quantum Chemistry, 9(S2):263–270, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Nishikawa:1977:ECF

[Nis77] Kiyoshi Nishikawa. Exact calculations of Franck–Condon overlaps and of matrix elements of some potentials by means of the coherent state representation. International Journal of Quantum Chemistry, 12(5):859–873, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nakashima:1977:TRP

[NJF^{+}77] Tadayoshi Nakashima, John R. Jungck, Sidney W. Fox, E. Lederer, and B. C. Das. A test for randomness in peptides isolated from a thermal polyamino acid. International Journal of Quantum Chemistry, 12(S4):65–72, January 9–12, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Nagata:1979:MAC

[NKKY79] Chikayoshi Nagata, Masahiko Kodama, Teruyuki Kimura, and Tamié Yamaguchi. Metabolic activation of chemical carcinogens and binding of metabolites with nucleic acid bases. International Journal of Quantum Chemistry, 16(4):917–930, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nevis:1977:SSR

[NM77] Arnold H. Nevis and Arun K. Majumdar. State space representation of nerve excitation. International Journal of Quantum Chemistry, 12(S4):155–160, January 9–12, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Niyogi:1977:ESE

[NMS77] B. Guha Niyogi, D. C. Mukherjee, and A. B. Sannigrahi. Effect of semi-empirical parameters on the simple random-phase approximation calculations of conjugated hydrocarbons. International Journal of Quantum Chemistry, 12(4):639–654, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nanda:1972:UHF

[NN72] D. N. Nanda and P. T. Narasimhan. Unrestricted Hartree– Fock calculations of spin densities with orthogonalized atomic orbitals: Proton and ^{13}C hyperfine splittings in some pi hydrocarbon radicals. International Journal of Quantum Chemistry, 6(5):819–841, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nanda:1974:AVS

[NN74] D. N. Nanda and P. T. Narasimhan. Approximate valence-shell electron unrestricted Hartree–Fock calculations with orthogonalized atomic orbitals: Proton hyperfine splittings in benzyl and related radicals. International Journal of Quantum Chemistry, 8(3):451–465, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nanda:1977:IRA

[NN77] D. N. Nanda and P. T. Narasimhan. On invariance requirements in approximate SCF MO Theory. International Journal of Quantum Chemistry, 12(1):215–223, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nordling:1975:EAL

[Nor75] J. O. Nordling. Effect of approximate linear dependence on the solution of the eigenvalue problem: A study by means of the super-secular equation. International Journal of Quantum Chemistry, 9(4):561–581, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Novaro:1973:CGT

[Nov73] O. Novaro. Comment on the group theoretical justification of the aufbau scheme. International Journal of Quantum Chemistry, 7(S7):53–56, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement:

Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Nagy:1979:MOS

[NPL^{+}79] L. Turi Nagy, P. Pelikán, M. Liška, M. Haring, M. Čeppan, and M. Breza. Molecular orbital study of molecular nitrogen fixation. International Journal of Quantum Chemistry, 16(3): 485–500, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nerbrant:1979:FOP

[NRS79] P.-O. Nerbrant, B. Roos, and A. J. Sadlej. First-order properties and the Hellmann–Feynman theorem in the case of a limited CI wave function. International Journal of Quantum Chemistry, 15(2):135–145, February 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Naray-Szabo:1972:DHF

[NS72] G. Náray-Szabó. Decoupled Hartree–Fock methods. I. Calculations for atoms with orthogonalized orbitals. International Journal of Quantum Chemistry, 6(4):779–785, July 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Naray-Szabo:1973:DHF

[NS73] G. Náray-Szabó. Decoupled Hartree–Fock methods. II. Calculation of the potential curves of diatomic molecules. International Journal of Quantum Chemistry, 7(3):569–581, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Naray-Szabo:1975:APW

[NS75] G. Náray-Szabó. All-pair wave function and reduced variational equation for electronic systems. International Journal of Quantum Chemistry, 9(1):9–21, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Niyogi:1978:TIS

[NS78] B. Guha Niyogi and A. B. Sannigrahi. On the triplet instability in semiempirical RPA calculations of conjugated hydrocarbons. International Journal of Quantum Chemistry, 14(3):337–340, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Naray-Szabo:1979:EIM

[NS79a] G. Náray-Szabó. Electrostatic isopotential maps for large biomolecules. International Journal of Quantum Chemistry, 16 (2):265–272, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nemukhin:1979:DMS

[NS79b] A. V. Nemukhin and N. F. Stepanov. Diatomics-in-molecules study of LiOH. International Journal of Quantum Chemistry, 15(1):49–56, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nepras:1979:IAB

[NT79] Miloš Nepraš and Miloš Titz. Influence of annelation by benzene ring on electronic absorption spectra and on character of the singlet π–π* transitions at polynuclear p-quinones. International Journal of Quantum Chemistry, 16(3):543–555, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nelson:1979:CMR

[NTN^{+}79] Donald J. Nelson, Anthony D. Theoharides, Andra C. Nieburgs, Richard K. Murray, Federico Gonzalez-Fernandez, and Daeg S. Brenner. Carbon-13 magnetic resonance study of lanthanide-substituted muscle calcium binding parvalbumins. International Journal of Quantum Chemistry, 16(1):159–174, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nuttall:1978:SAC

[Nut78] J. Nuttall. Scattering amplitudes and complex scaling. International Journal of Quantum Chemistry, 14(4):519–522, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Novaro:1975:FBT

[NY75] O. Novaro and F. Yáñez. Four-body terms in multibody expansions for noble gas clusters. International Journal of Quantum Chemistry, 9(S9):359–365, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Novaro:1976:GTB

[NY76] O. Novaro and F. Yáñez. Gaussian-type bases for the calculation of the interaction energy between neon atoms. International Journal of Quantum Chemistry, 10(3):545–550, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ortega-Campos:1971:LDS

[OCRMLR^{+}71] Jesús Ortega-Campos, Isidro Romero-Medina, Evodio Lopez-Rojas, Leonel Torres-Hernández, and Harold V. Mcintosh. Lattice dynamics with second neighbor interactions. International Journal of Quantum Chemistry, 5(S5):201–225, January 18– 23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Odabasi:1973:SEA

[Oda73] Halis Odabasi. Some evidence about the dynamical group SO(4, 2) symmetries of the periodic table of elements. International Journal of Quantum Chemistry, 7(S7):23–33, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Oddershede:1977:DEE

[OJB77] Jens Oddershede, Poul Jørgensen, and Nelson H. F. Beebe. Determination of excitation energies and transition moments in a second order polarization propagator approach. application to the Be atom and the CH^{+} molecule. International Journal of Quantum Chemistry, 12(4):655–670, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Olson:1975:EIT

[OK75] Mark L. Olson and Daniel D. Konowalow. Energy of interaction of two hydrogen atoms using contracted Gaussian type valence-bond wavefunctions. International Journal of Quantum Chemistry, 9(6):1113–1129, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oddershede:1974:HFL

[OKM74] Jens Oddershede, Lalit Kumar, and Hendrik J. Monkhorst. On the Hartree–Fock limit for metallic hydrogen with plane wave basis. International Journal of Quantum Chemistry, 8(S8):447– 450, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Ohrn:1978:CCG

[ÖL78] Y. Öhrn and J. Linderberg. Characteristics of the consistent ground state of the random phase of approximation. International Journal of Quantum Chemistry, ??(??):1–18, ?? 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ohrn:1979:CCG

[ÖL79a] Yngve Öhrn and Jan Linderberg. Characteristics of the consistent ground state of the random phase approximation. International Journal of Quantum Chemistry, 15(3):343–353, March 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ohrn:1979:LOH

[ÖL79b] Yngve Öhrn and Jan Linderberg. Linda and Ole Halkjaer. International Journal of Quantum Chemistry, 16(S13):485, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Oie:1976:ICL

[OMC76] Tetsuro Oie, Gerald M. Maggiora, and Ralph E. Christoffersen. Ab initio calculations on large molecules using molecular fragments: Comparison of charge distribution and molecular electrostatic potential for ethyl chlorophyllide a and related molecules. International Journal of Quantum Chemistry, 10 (S3):119–133, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Onipko:1979:TES

[Oni79] A. I. Onipko. Theory of the excitation spectrum of nondiagonal disordered systems. International Journal of Quantum Chemistry, 16(3):557–571, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Otto:1977:ISS

[OSL77] Peter Otto, Sándor Suhai, and János Ladik. Ab initio supermolecule study of charge transfer in the glyoxal–formamide and in the H_{2} S–formamide systems. International Journal of Quantum Chemistry, 12(S4):451–457, January 9–12, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ostlund:1979:CCM

[Ost79] Neil S. Ostlund. Chemistry, computers, and microelectronics: Present and future prospects. International Journal of Quantum Chemistry, 16(S13):15–38, March 11–17, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Overhauser:1973:NPE

[Ove73] A. W. Overhauser. Nonlocal properties of exchange and correlation potentials in band-structure calculations. International Journal of Quantum Chemistry, 7(S7):605, January 21– 27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Pullman:1974:ISH

[PA74] A. Pullman and A. M. Armbruster. An ab initio study of the hydration and ammoniation of ammonium ions. International Journal of Quantum Chemistry, 8(S8):169–176, January 20– 26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Pullman:1977:E

[PA77] A. Pullman and A-M. Armbruster. Erratum. International Journal of Quantum Chemistry, 11(4):701, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Paldus:1977:AGM

[PAČ77] J. Paldus, B. G. Adams, and J. Čížek. Application of graphical methods of spin algebras to limited CI approaches. I. Closed shell case. International Journal of Quantum Chemistry, 11 (5):813–848, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Painter:1971:ADV

[Pai71] G. S. Painter. Application of the discrete variational method to the electronic structure of LiF. International Journal of Quantum Chemistry, 5(S5):501–513, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Painter:1973:RDT

[Pai73] G. S. Painter. Recent developments for treating non-muffin-tin potentials and charge densities in band structure and cluster calculations. International Journal of Quantum Chemistry, 7 (S7):545–549, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Palting:1973:DDP

[Pal73] Pancracio Palting. Direct determination of pseudonatural orbitals. International Journal of Quantum Chemistry, 7(4):717– 723, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Paldus:1975:PCU

[Pal75] Josef Paldus. A pattern calculus for the unitary group approach to the electronic correlation problem. International Journal of Quantum Chemistry, 9(S9):165–174, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Poshusta:1975:UGS

[PAM75] R. D. Poshusta, V. P. Agrawal, and W. D. Moseley. On the use of Gaussian shell type basis orbitals for single-center expansions. I. Evaluation of integrals. International Journal of Quantum Chemistry, 9(4):635–647, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Partridge:1972:BEM

[Par72] R. H. Partridge. Bond exciton model of alkenes and diene polymers. International Journal of Quantum Chemistry, 6(1): 167–179, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Parr:1973:SCF

[Par73] Robert G. Parr. The scientific contributions of S. F. Boys. International Journal of Quantum Chemistry, 7(S7):123–126, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Parr:1977:EHF

[Par77] Robert G. Parr. Erich Hückel and Friedrich Hund — pioneers in quantum chemistry. International Journal of Quantum Chemistry, 12(S11):29–37, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Pauncz:1977:BDS

[Pau77] Ruben Pauncz. Branching diagram and Serber-type spin functions. Algorithms for their construction and special properties. International Journal of Quantum Chemistry, 12(2):369–382, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Paul:1978:MBM

[Pau78] Carlton H. Paul. Models of biological molecules based on wave function electron densities. International Journal of Quantum Chemistry, 14(S5):441–447, March 3–11, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pullman:1977:ISN

[PB77] Alberte Pullman and Hélène Berthod. An ab initio study of a nucleoside: Uridine. International Journal of Quantum Chemistry, 12(S4):327–336, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pokrant:1974:NTV

[PBC74a] M. A. Pokrant, A. A. Broyles, and R. L. Coldwell. Nonzero temperature variational principle. International Journal of Quantum Chemistry, 8(S8):403–412, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Pullman:1974:ECP

[PBC74b] B. Pullman, H. Berthod, and P. H. Courriere. The exploration of the conformational properties of biological phenethylamines by molecular orbital techniques. International Journal of Quantum Chemistry, 8(S1):93–108, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pullman:1976:CLI

[PBG76] A. Pullman, H. Berthod, and N. Gresh. Cation-ligand interactions: ab initio SCF studies of the binding properties of alkali, alkaline-earth, and ammonium ions. International Journal of Quantum Chemistry, 10(S10):59–76, January 19–24, 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pantelides:1978:GFS

[PBPL78] S. T. Pantelides, J. Bernholc, J. Pollmann, and N. O. Lipari. Green’s function scattering-theoretic methods for point defects, surfaces, and interfaces in solids. International Journal of Quantum Chemistry, 14(S12):507–521, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Pople:1976:TMI

[PBS76] John A. Pople, J. Stephen Binkley, and Rolf Seeger. Theoretical models incorporating electron correlation. International Journal of Quantum Chemistry, 10(S10):1–19, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pullman:1976:QMS

[PC76] B. Pullman and P. H. Courriere. Quantum-mechanical studies on the conformation of sphingomyelins. International Journal of Quantum Chemistry, 10(S3):29–38, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Paldus:1974:CEL

[PČH74] J. Paldus, J. Čižek, and I. HubaČ. Correlation effects in the low-lying excited states of the PPP models of alternant hydrocarbons. II. State correlation diagrams. International Journal of Quantum Chemistry, 8(S8):293–303, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Paldus:1978:RBU

[PČL78] J. Paldus, J. Čížk, and A. Laforgue. The relationship between the unrestricted and projected Hartree–Fock methods in a simple three-electron model system. International Journal of Quantum Chemistry, 13(1):41–65, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Petke:1977:ICL

[PCMS77] J. D. Petke, Ralph E. Christoffersen, Gerald M. Maggiora, and Lester L. Shipman. Ab initio calculations on large molecules using molecular fragments. SCF-MO-CI Studies on low-lying singlet and triplet states of pyrazine. International Journal of Quantum Chemistry, 12(S4):343–355, January 9–12, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Politzer:1977:CAE

[PD77] Peter Politzer and Kenneth C. Daiker. A comparative analysis of the electrostatic potentials of some polycyclic aromatic hydrocarbons. International Journal of Quantum Chemistry, 12(S4):317–325, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Piela:1978:EPE

[PD78a] Lucjan Piela and Joseph Delhalle. An efficient procedure to evaluate long-range Coulombic interactions within the framework of the LCAO-CO method for infinite polymers. International Journal of Quantum Chemistry, 13(5):605–617, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Politzer:1978:SPE

[PD78b] Peter Politzer and Kenneth C. Daiker. Some potential-energy relationships for isoelectronic atomic series. International Journal of Quantum Chemistry, 14(3):245–251, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pullman:1979:MSA

[PD79] A. Pullman and D. Demoulin. Model studies on the active site of carbonic anhydrase: Ligand properties and CO_{2} binding. International Journal of Quantum Chemistry, 16(3):641–653, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Politzer:1979:RHB

[PDE79] Peter Politzer, Kenneth C. Daiker, and Vernon M. Estes. The role of hydrogen bonding in some diol epoxides. International Journal of Quantum Chemistry, 16(S6):47–53, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Politzer:1978:ENI

[PDEB78] Peter Politzer, Kenneth C. Daiker, Vernon M. Estes, and Michael Baughman. Epoxide–nucleophile interactions: Acidcatalyzed reaction of ethylene oxide with water. International Journal of Quantum Chemistry, 14(S5):291–299, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Peat:1974:ESO

[Pea74a] F. David Peat. The emergence of structure and organization from physical systems. International Journal of Quantum Chemistry, 8(S1):213–219, January 17–19, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Peat:1974:SAD

[Pea74b] F. David Peat. Suggested applications of the density matrix. International Journal of Quantum Chemistry, 8(S8):313–321, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Pellegatti:1977:SES

[Pel77] Alain Pellégatti. Simple extended STO basis sets. Helium. International Journal of Quantum Chemistry, 12(3):545–547, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Percus:1978:RMS

[Per78] J. K. Percus. The role of model systems in the few-body reduction of the N-fermion problem. International Journal of Quantum Chemistry, 13(1):89–124, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Petersson:1977:PCG

[Pet77] G. A. Petersson. The pairwise correlated generalized valence bond model of electronic structure. II. A simple physical model for electron correlation. International Journal of Quantum Chemistry, 11(5):775–812, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pethig:1978:EPP

[Pet78] Ronald Pethig. Electronic properties of protein-methylglyoxal complexes: Strong evidence for energy-band conduction. International Journal of Quantum Chemistry, 14(S5):159–171, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Petrov:1979:MET

[Pet79] E. G. Petrov. Mechanisms of electron transfer through proteins. International Journal of Quantum Chemistry, 16(1):133–152, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pohl:1974:IHT

[PF74] Herbert A. Pohl and Donald R. Fowler. An improved Hellmanntype pseudopotential for atoms and molecules. International Journal of Quantum Chemistry, 8(3):435–450, May 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ptitsyn:1979:MPF

[PF79] O. B. Ptitsyn and A. V. Finkelstein. Mechanism of protein folding. International Journal of Quantum Chemistry, 16(3): 407–418, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pullman:1977:UPS

[PGDM77] A. Pullman, N. Gresh, J. P. Daudey, and J. W. Moskowitz. The use of pseudopotentials for the study of cation-ligand and intermolecular interactions. International Journal of Quantum Chemistry, 12(S11):501–507, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Prado:1979:IQM

[PGPP79] F. Ribas Prado, C. Giessner-Prettre, and B. Pullman. Ab initio quantum-mechanical calculations of the variation of the magnetic shielding constant of hydrogen and carbon 13 nuclei of the dimethylphosphate anion as a function of molecular conformation: A model study for nucleic acid constituents. International Journal of Quantum Chemistry, 16(S6):491–501, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Perlin:1977:PTP

[PGQ77] Y. U. E. Perlin, S. H. N. Gifeisman, and Pham Van Quyet. Perturbation theory in the path-integral representation and its application to some simple quantum systems. International Journal of Quantum Chemistry, 12(2):233–246, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Patterson:1977:UBM

[PH77] Chris W. Patterson and William G. Harter. Unitary bases for molecular electronic configurations. International Journal of Quantum Chemistry, 12(S11):445–457, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Pendergast:1976:MSQ

[PHL^{+}76] Phil Pendergast, Edward F. Hayes, Richard C. Liedtke, Maurice E. Schwartz, Stephen Rothenberg, and Peter A. Kollman. MOLE: A system for quantum chemistry. II. recent developments. International Journal of Quantum Chemistry, 10(S10): 77–83, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Piela:1971:PCE

[Pie71] Lucjan Piela. Perturbational calculation of the exchange forces in the two lowest states of the hydrogen molecule. International Journal of Quantum Chemistry, 5(1):85–99, January 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pipano:1971:ICP

[PK71] A. Pipano and Joyce J. Kaufman. Ab initio calculation of potential energy curves for the ion-molecule reaction O^{+} + N_{2} → NO^{+} + N. International Journal of Quantum Chemistry, 5 (S5):233–234, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Petelin:1972:REF

[PK72] A. N. Petelin and A. A. Kiselev. The Renner effect in fouratomic molecules. International Journal of Quantum Chemistry, 6(4):701–716, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Preston:1973:ISC

[PK73] Harry J. T. Preston and Joyce J. Kaufman. Ab-initio SCF calculations on pyrrole and pyrazole. International Journal of Quantum Chemistry, 7(S7):207–215, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Popkie:1975:TCC

[PK75] Herbert E. Popkie and Joyce J. Kaufman. Test of chargeconserving integral approximations for a variable retention of diatomic differential overlap (VRDDO) procedure for semi-ab initio molecular orbital calculations on large molecules. International Journal of Quantum Chemistry, 9(S2):279–288, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Popkie:1976:ILM

[PK76a] Herbert E. Popkie and Joyce J. Kaufman. Ab initio LCAO– MO–SCF calculation of chlorpromazine and promazine. International Journal of Quantum Chemistry, 10(4):569–580, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Popkie:1976:MCM

[PK76b] Herbert E. Popkie and Joyce J. Kaufman. Molecular calculations with the MODPOT, VRDO, and MODPOT/VRDO procedures I. HF, F_{2}, HCl, Cl_{2}, formamide, pyrrole, pyridine, and nitrobenzene. International Journal of Quantum Chemistry, 10(S10):47–57, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Popkie:1977:MCMa

[PK77a] Herbert E. Popkie and Joyce J. Kaufman. Molecular calculations with the MODPOT, VRDDO, and MODPOT /VRDDO procedures. IV. Boron hydrides and carboranes. International Journal of Quantum Chemistry, 12(5):937–961, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Popkie:1977:MCMb

[PK77b] Herbert E. Popkie and Joyce J. Kaufman. Molecular calculations with the MODPOT, VRDDO, and MODPOT/VRDDO procedures. V. Ab initio and MODPOT LCAO-MO-SCF calculations on the chlorofluoromethanes. International Journal of Quantum Chemistry, 12(S11):433–443, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Preston:1977:MXC

[PK77c] Harry J. T. Preston and Joyce J. Kaufman. MS–Xα calculations of the electron affinities and ionization potentials of some heavy metals and their hydrides and fluorides. International Journal of Quantum Chemistry, 12(3):471–484, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Preston:1975:MXC

[PKK75] Harry J. T. Preston, Joyce J. Kaufman, and Walter S. Koski. MS–Xα calculations of polyhedral boron compounds. International Journal of Quantum Chemistry, 9(S9):137–145, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Pople:1978:ECT

[PKSB78] J. A. Pople, R. Krishnan, H. B. Schlegel, and J. S. Binkley. Electron correlation theories and their application to the study of simple reaction potential surfaces. International Journal of Quantum Chemistry, 14(5):545–560, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pople:1979:DSH

[PKSB79] J. A. Pople, R. Krishnan, H. B. Schlegel, and J. S. Binkley. Derivative studies in Hartree–Fock and Møller–Plesset theories. International Journal of Quantum Chemistry, 16(S13): 225–241, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Pack:1977:OSS

[PL77] George R. Pack and Gilda H. Loew. The origins of sequence specificity of ethidium nucleic acid intercalation. International Journal of Quantum Chemistry, 12(S4):87–96, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pack:1979:SSM

[PL79] George R. Pack and Gilda H. Loew. Semiempirical studies of the mechanism of models for the N-hydroxylation of amines by cytochrome P450. International Journal of Quantum Chemistry, 16(S6):381–390, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Praud:1973:ISO

[PLMB73] Lionel Praud, Bernard Levy, Philippe Millie, and Gaston Berthier. Ab initio second-order perturbation calculations for excited and ionized states of benzene isomers. International Journal of Quantum Chemistry, 7(S7):185–191, January 21– 27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Plotnikov:1979:RPR

[Plo79] V. G. Plotnikov. Regularities of the processes of radiationless conversion in polyatomic molecules. International Journal of Quantum Chemistry, 16(3):527–541, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Peng:1978:UPS

[PLSL78] S. Peng, J. Lin, M. Shahbaz, and P. R. Lebreton. Ultraviolet photoelectron studies of 5-trifluoromethyluracil: Electronic susceptibility of substituted uracils to nucleophilic attack. International Journal of Quantum Chemistry, 14(S5):301–309, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pack:1978:CSS

[PLYM78] George R. Pack, Gilda H. Loew, Shinichi Yamabe, and Keiji Morokuma. Comparative study of semiempirical methods for calculating interactions between large molecules with an application to the actinomycin–guanine complex. International Journal of Quantum Chemistry, 14(S5):417–432, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Papaconstantopoulos:1971:BSF

[PN71] Dimitrios A. Papaconstantopoulos and David J. Nagel. Band structure and Fermi surfaces of ordered intermetallic compounds TiFe, TiCo and TiNi. International Journal of Quantum Chemistry, 5(S5):515–526, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Pilkington:1976:MEA

[PN76] Phillip W. Pilkington and Stanley C. Neely. Matrix elements for an approximate ab initio π-electron theory. International Journal of Quantum Chemistry, 10(3):473–493, May 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Petersson:1978:PCG

[PNC^{+}78] G. A. Petersson, M. R. Nyden, J. T. Chupka, P. B. Ryan, and H. D. Todd. The pairwise-correlated generalized valence bond model of electronic structure. V. A multiconfiguration overlap approximation for GVB pair energies. International Journal of Quantum Chemistry, 14(S12):201–218, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Papaconstantopoulos:1978:CXR

[PNJB78] Dimitrios A. Papaconstantopoulos, David J. Nagel, and Christine Jones-Bjorklund. Calculations of X-Ray band spectra: Application to vanadium. International Journal of Quantum Chemistry, 14(S12):497–506, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Purvis:1977:AMI

[PÖ77] George D. Purvis and Yngve Öhrn. Accurate molecular ionization potentials from the electron propagator. International Journal of Quantum Chemistry, 12(S11):359–363, January 16– 22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Polak:1970:OHO

[Pol70] Rudolf Polák. Optimum hybrid orbitals in localized orbitals. International Journal of Quantum Chemistry, 4(3):271–287, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poltev:1979:SII

[Pol79] V. I. Poltev. Simulation of intermolecular and intramolecular interactions of nucleic acid subunits by means of atom–atom potential functions. International Journal of Quantum Chemistry, 16(4):863–868, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pople:1971:ESW

[Pop71] J. A. Pople. Electronic states and wave functions associated with orbital energy crossing. International Journal of Quantum Chemistry, 5(S5):175–182, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Popov:1979:QAC

[Pop79] E. M. Popov. Quantitative approach to conformations of proteins. International Journal of Quantum Chemistry, 16(4):707– 737, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Porta:1977:MHC

[Por77] Jordi Porta. A model for the helix–coil transition of DNA. International Journal of Quantum Chemistry, 11(1):49–57, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Porta:1978:TDH

[Por78] Jordi Porta. Thermodynamics of the DNA helix-coil transition. International Journal of Quantum Chemistry, 13(4):499–507, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poshusta:1970:CDS

[Pos70] R. D. Poshusta. Comment: Discussion on small molecules, session Al. International Journal of Quantum Chemistry, 5(S4): 147, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Poshusta:1977:TGL

[Pos77] R. D. Poshusta. Tetrahedral Gaussian Lobe Functions. International Journal of Quantum Chemistry, 12(1):225–226, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poshusta:1978:SPE

[Pos78] R. D. Poshusta. Singer polymals. I. Evaluation of matrix elements. International Journal of Quantum Chemistry, 13(1):27– 39, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poshusta:1979:SPI

[Pos79] R. D. Poshusta. Singer polymals. II. Tempering methods. International Journal of Quantum Chemistry, 16(S13):59–66, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Prohofsky:1971:SQM

[PP71] E. W. Prohofsky and R. C. Purdom. Simple quantum mechanics and complicated ultrasonic attenuation. International Journal of Quantum Chemistry, 5(S5):563–567, January 18– 23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Port:1974:QMS

[PP74] G. N. J. Port and A. Pullman. Quantum-mechanical studies of environmental effects on biomolecules. III. Ab initio model studies of the hydration of peptides and proteins. International Journal of Quantum Chemistry, 8(S1):21–32, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Paine:1979:DTD

[PP79] Douglas A. Paine and William L. Pensinger. A dynamical theory describing superconductant DNA. International Journal of Quantum Chemistry, 15(3):333–341, March 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pullman:1978:SIS

[PPB78] Bernard Pullman, Alberte Pullman, and Helene Berthod. SCF ab initio study of the “through-water” versus “direct” binding of the Na^{+} and Mg^{2+} cations to the phosphate anion. International Journal of Quantum Chemistry, 14(S5):79–90, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Perahia:1972:MOC

[PPC72] D. Perahia, B. Pullman, and P. Claverie. Molecular orbital calculations on the conformation of polypeptides and proteins X. The conformational energy maps of the cysteinyl and methionyl residues. International Journal of Quantum Chemistry, 6(S6):337–351, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Pyykko:1973:HLR

[PPI73] P. Pyykkö, E. Pajanne, and Mitio Inokuti. Hydrogen-like relativistic corrections for electric and magnetic hyperfine integrals. International Journal of Quantum Chemistry, 7(4):785– 806, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Petrongolo:1978:ILM

[PPK78] Carlo Petrongolo, Harry J. T. Preston, and Joyce J. Kaufman. Ab initio LCAO–MO–SCF calculation of the electrostatic molecular potential of chlorpromazine and promazine. International Journal of Quantum Chemistry, 13(4):457–468, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Perahia:1979:MEP

[PPP79] D. Perahia, A. Pullman, and B. Pullman. Molecular electrostatic potential of the B-DNA helix. V. Poly(dG·dC) and poly(dA·dT). International Journal of Quantum Chemistry, 16(S6):353–363, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Prigogine:1975:DSD

[Pri75] I. Prigogine. Dissipative structures, dynamics and entropy. International Journal of Quantum Chemistry, 9(S9):443–456, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Phillips:1970:UHF

[PS70] Donald H. Phillips and John C. Schug. Unrestricted Hartree– Fock SCF calculations on some hypothetical molecules with triplet ground states. International Journal of Quantum Chemistry, 5(S4):221–229, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Palting:1973:STU

[PS73] Pancracio Palting and William A. Sanders. Simple test of the utility of modified Hartree–Fock orbitals: The He atom. International Journal of Quantum Chemistry, 7(2):301–307, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Primas:1975:DFR

[PS75a] H. Primas and M. Schleicher. A density functional representation of quantum chemistry. II. Local quantum field theories of molecular matter in terms of the charge density operator do not work. International Journal of Quantum Chemistry, 9(5): 855–870, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pullman:1975:QMS

[PS75b] Bernard Pullman and Anil Saran. Quantum-mechanical studies on the conformation of phospholipids. International Journal of Quantum Chemistry, 9(S2):71–97, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Philippot:1979:OAN

[PS79a] J. Philippot and I. Sengers. Optical activity of nonrigid molecules. International Journal of Quantum Chemistry, 15 (6):713–715, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Prasad:1979:PAG

[PS79b] C. V. Prasad and K. Sundaram. On the polarity of the amide group and its impact on dipeptide conformation. International Journal of Quantum Chemistry, 15(6):783–792, June 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pople:1977:VCI

[PSK77] J. A. Pople, R. Seeger, and R. Krishnan. Variational configuration interaction methods and comparison with perturbation theory. International Journal of Quantum Chemistry, 12(S11): 149–163, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Petrongolo:1975:UEM

[PT75] Carlo Petrongolo and Jacopo Tomasi. The use of the electrostatic molecular potential in quantum pharmacology. I. Ab initio results. International Journal of Quantum Chemistry, 9 (S2):181–190, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pamuk:1978:SEC

[PT78] H. Önder Pamuk and Carl Trindle. Semiempirical estimation of correlation energy corrections to ionization potentials and dissociation energies for open-shell systems. International Journal of Quantum Chemistry, 14(S12):271–282, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Petrasen:1969:AGQ

[PTUK69] Marija Ivanovna Petraseń, Evgenij D. Trifonov, Armin Uhlmann, and Adolf Kühnel. Anwendung der Gruppentheorie in der Quantenmechanik. (German) [Application of group theory in quantum mechanics]. Akademie-Verlag, Berlin, Germany, 1969. x + 240 pp. LCCN ???? URL http://www.gbv.

de/dms/ilmenau/toc/021694478.PD.

Pullman:1970:MOC

[Pul70] B. Pullman. Molecular orbital calculations on the conformation of biomolecules. International Journal of Quantum Chemistry, 5(S4):319–340, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Pullman:1974:ISB

[Pul74] Alberte Pullman. Ab initio studies on the binding of Na^{+} and K^{+} to the fundamental components of depsipeptides. International Journal of Quantum Chemistry, 8(S1):33–42, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pullman:1976:QMM

[Pul76] Bernard Pullman, editor. Quantum mechanics of molecular conformations. Perspectives in quantum chemistry and biochemistry. John Wiley, New York, NY, USA, 1976. ISBN 0-471-01489-3. ix + 412 pp. LCCN QD462.5 .Q36.

Pullman:1977:CQM

[Pul77] B. Pullman. Contribution of quantum-mechanical computations to the elucidation of the solution structure of tRNA by NMR techniques. International Journal of Quantum Chemistry, 12(S4):3–10, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pullman:1978:IID

[Pul78a] Bernard Pullman, editor. Intermolecular interactions, from diatomics to biopolymers. Perspectives in quantum chemistry and biochemistry; v. 2. John Wiley, New York, NY, USA, 1978. ISBN 0-471-99507-X. ix + 447 pp. LCCN QD462.5 .I57.

Pullman:1978:NMR

[Pul78b] Bernard Pullman, editor. Nuclear magnetic resonance spectroscopy in molecular biology: proceedings of the Eleventh Jerusalem Symposium on Quantum Chemistry and Biochemistry, held in Jerusalem, Israel, April 3–7, 1978, volume 11 of The Jerusalem symposia on quantum chemistry and biochemistry. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1978. ISBN 90-277-0932-7. LCCN QH324.9.N8 J47 1978.

Pullman:1979:RDM

[Pul79a] B. Pullman. Recent developments on the mechanism of chemical carcinogenesis by aromatic hydrocarbons. International Journal of Quantum Chemistry, 16(3):669–689, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pullman:1979:MIA

[Pul79b] Bernard Pullman. 14th Meeting of the international academy of quantum molecular science. International Journal of Quantum Chemistry, 15(1):131, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pullman:1979:R

[Pul79c] Bernard Pullman. Reminiscences. International Journal of Quantum Chemistry, 16(S6):33–45, March 4–10, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Purcell:1975:QSA

[Pur75] William Paul Purcell. Quantitative structure–activity relationships of psychotropic agents. International Journal of Quantum Chemistry, 9(S2):191–196, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pulfer:1974:HIEb

[PW74a] J. D. Pulfer and M. A. Whitehead. The Hammerstein integral equation: a general technique for constructing a rapidly convergent Padé-type approximation to the logarithmic derivative. International Journal of Quantum Chemistry, 8(5):693–706, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pulfer:1974:HIEa

[PW74b] J. D. Pulfer and M. A. Whitehead. Hammerstein integral equivalent of Riccati’s equation. International Journal of Quantum Chemistry, 8(5):677–692, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Paldus:1979:CIM

[PW79] J. Paldus and P. E. S. Wormer. Configuration interaction matrix elements. II. Graphical approach to the relationship between unitary group generators and permutations. International Journal of Quantum Chemistry, 16(6):1321–1335, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pullman:1979:MEP

[PZP79] Alberte Pullman, Chrystyna Zakrzewska, and David Perahia. Molecular electrostatic potential of the B-DNA helix. I. Region of the guanine–cytosine base pair. International Journal of Quantum Chemistry, 16(2):395–403, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ra:1976:API

[Ra76a] Øystein Ra. Asymptotic properties of inverses of cyclic overlap matrices in the large separation limit. International Journal of Quantum Chemistry, 10(1):57–69, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ra:1976:OOO

[Ra76b] Øystein Ra. Orthogonalization of overlapping orbitals by optimal polynomials. I. Polynomials constructed from spectral radii. International Journal of Quantum Chemistry, 10(1):5– 20, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rabalais:1977:PUP

[Rab77] J. Wayne Rabalais. Principles of ultraviolet photoelectron spectroscopy. Wiley-Interscience monographs in chemical physics.

John Wiley, New York, NY, USA, 1977. ISBN 0-471-70285-4. xv + 454 pp. LCCN QC454.P48 R3.

Raffenetti:1975:OET

[Raf75] Richard C. Raffenetti. Optimal even-tempered Gaussian atomic orbital bases: First-row atoms. International Journal of Quantum Chemistry, 9(S9):289–295, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Ramachandran:1978:CPT

[Ram78] G. N. Ramachandran. Confirmation of Pauling’s theory that vitamin C improves immunity to infection. International Journal of Quantum Chemistry, 14(S5):15–20, March 3–11, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Randic:1974:HMO

[Ran74] Milan Randić. Hybridization by the maximum overlap method. International Journal of Quantum Chemistry, 8(5):643–676, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1978:GTA

[Ran78] Milan Randić. Graph-theoretical analysis of structure-property and structure-activity correlations. International Journal of Quantum Chemistry, 14(S5):245–255, March 3–11, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Randic:1979:SPG

[Ran79] Milan Randić. Symmetry properties of graphs of interest in chemistry. II. Desargues–Levi graph. International Journal of Quantum Chemistry, 15(6):663–682, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ratner:1971:SPH

[Rat71] Mark Ratner. Single particle-hole approximation to twoparticle Green function for ethylene. International Journal of Quantum Chemistry, 5(S5):161–163, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Ratner:1978:TAI

[Rat78] Mark A. Ratner. Theoretical approaches to intramolecular electron transfer processes. International Journal of Quantum Chemistry, 14(5):675–694, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Reid:1975:IHO

[RB75a] Charles E. Reid and Erkki J. Brändas. Integrals of harmonic oscillator eigenfunctions over powers of x. International Journal of Quantum Chemistry, 9(1):189–191, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roy:1975:RCH

[RB75b] H. P. Roy and A. K. Bhattacharya. Rearrangement correction to higher polarizabilities of two-electron systems. International Journal of Quantum Chemistry, 9(3):495–504, May 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Redmon:1977:SSC

[RB77] L. T. Redmon and J. C. Browne. Sum-over-states calculations of the polarizabilities of the ground and first excited states of the lithium atom. International Journal of Quantum Chemistry, 12(S11):311–319, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Rohmer:1977:EVS

[RBDV77] Marie-Madeleine Rohmer, Michael Barry, Alain Dedieu, and Alain Veillard. End-on versus side-on coordination of dioxygen: An ab initio calculation for peroxotitaniumporphyrin. International Journal of Quantum Chemistry, 12(S4):337–342, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Robb:1970:GSE

[RC70] M. A. Robb and I. G. Csizmadia. The generalized separated electron pair model. 1. An application to NH_{3}. International Journal of Quantum Chemistry, 4(4):365–387, July 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Robb:1971:GSE

[RC71] M. A. Robb and I. G. Csizmadia. The generalized separated electron pair model. II. An application to NH, NH_{3}, NH, NH^{2−} and N^{3−}. International Journal of Quantum Chemistry, 5(6): 605–635, November 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Robb:1972:GSE

[RC72] M. A. Robb and I. G. Csizmadia. The generalized separated electron pair model. III. An application to three localization schemes for CO. International Journal of Quantum Chemistry, 6(2):367–382, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Romano:1978:MCS

[RC78] Silvano Romano and Enrico Clementi. Monte Carlo simulation of water solvent with biomolecules. Glycine and the corresponding zwitterion. International Journal of Quantum Chemistry, 14(6):839–850, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rein:1975:EPB

[RCE75] Robert Rein, Yves Coeckelenbergh, and John T. Egan. Elaboration of the principle of base complementarity and the elements of a theory of point mutations. International Journal of Quantum Chemistry, 9(S2):145–153, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ruedenberg:1979:MOT

[RCE79] K. Ruedenberg, L. M. Cheung, and S. T. Elbert. MCSCF optimization through combined use of natural orbitals and the Brillouin–Levy–Berthier theorem. International Journal of Quantum Chemistry, 16(5):1069–1101, November 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roustan:1979:RAS

[RD79] J. Roustan and M. Dugay. Refined AMOS for some organic semiconductors. International Journal of Quantum Chemistry, 15(1):15–35, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roothaan:1979:IMM

[RDH79] Clemens C. J. Roothaan, John Detrich, and Darrel G. Hopper. An improved MCSCF method. International Journal of Quantum Chemistry, 16(S13):93–101, March 11–17, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Robinson:1970:EBD

[RE70] Peter D. Robinson and Saul T. Epstein. On energy bounds derived from the conjugate eigenvalue problem. International Journal of Quantum Chemistry, 4(5):453–463, September 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Redmon:1979:RRT

[Red79] Michael J. Redmon. Recent results from three-dimensional quantum reactive scattering theory. International Journal of Quantum Chemistry, 16(S13):559–568, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Rein:1970:REP

[Rei70] Robert Rein. Remarks on elementary processes in quantum biology. International Journal of Quantum Chemistry, 5(S4): 341–352, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Reid:1972:IHL

[Rei72a] C. E. Reid. Intermediate Hamiltonians for the lithium atom. International Journal of Quantum Chemistry, 6(4):793–795, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Reid:1972:SPT

[Rei72b] Charles E. Reid. A simplified proof of a theorem about group representations. International Journal of Quantum Chemistry, 6(2):383–386, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Reinhardt:1976:MCC

[Rei76] William P. Reinhardt. Method of complex coordinates: application to the Stark effect in hydrogen. International Journal of Quantum Chemistry, 10(S10):359–367, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Resta:1978:SCH

[Res78] Raffaele Resta. Study of correlation holes. I. Number–sum rules and infinite system. International Journal of Quantum Chemistry, 14(2):171–179, August 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roy:1975:FDP

[RGM75] H. P. Roy, A. Gupta, and P. K. Mukherjee. Frequencydependent polarizability of open-shell atomic systems. International Journal of Quantum Chemistry, 9(1):75–81, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roy:1977:CHF

[RGM77] H. P. Roy, A. Gupta, and P. K. Mukherjee. Coupled Hartree– Fock calculation of static dipole polarizations of open-shell ions. International Journal of Quantum Chemistry, 12(1):1–9, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Richards:1977:QP

[Ric77] W. G. (William Graham) Richards. Quantum pharmacology. Butterworths, London, UK, 1977. ISBN 0-408-70919-7. xiii + 213 pp. LCCN RM301.7 .R5.

Ritchie:1976:EEH

[Rit76] A. G. Ritchie. Electronic energies of the hydrogen molecule D′Π_{u} and B′′Σ states. International Journal of Quantum Chemistry, 10(6):1071–1073, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rosengren:1973:DEG

[RJ73] A. Rosengren and B. Johansson. DODS and the electron gas. International Journal of Quantum Chemistry, 7(5):1033–1035, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rudzikas:1976:GSF

[RK76] Z. B. Rudzikas and J. M. Kaniauskas. Generalized spherical functions in the theory of many-electron atoms. International Journal of Quantum Chemistry, 10(5):837–852, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rosenkrantz:1977:MXM

[RK77] Marcy E. Rosenkrantz and Daniel D. Konowalow. The MT Xα_{R} method. The determination and use of “molecular” α values in molecular multiple scattering Xα calculations. International Journal of Quantum Chemistry, 12(4):707–719, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rosenkrantz:1979:CPW

[RK79] Marcy E. Rosenkrantz and Daniel D. Konowalow. Composition of the partial wave expansion as a parameter in the MSXα scheme. International Journal of Quantum Chemistry, 16(6): 1301–1305, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rein:1979:TIE

[RKE79] Robert Rein and Thomas Kieber-Emmons. Theory of ionic effects on intermolecular interactions in polyions: Application to ethidium binding to DNA. International Journal of Quantum Chemistry, 16(S6):481–489, March 4–10, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Renugopalakrishnan:1977:NMR

[RKOU77] V. Renugopalakrishnan, A. Khaled, K. Okamoto, and D. W. Urry. Nuclear magnetic resonance and conformational energy calculations of repeat peptides of tropoelastin: Correlation of 1J(^{15}N–^{1}H) with nonplanarity of peptide moiety. International Journal of Quantum Chemistry, 12(S4):97–110, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Rai:1977:BTT

[RL77] D. K. Rai and J. Ladik. Breit-type three-electron equation in the Pauli approximation. International Journal of Quantum Chemistry, 12(5):925–935, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rohmer:1979:ESP

[RL79] Marie-Madeleine Rohmer and Gilda H. Loew. Electronic structure and properties of model oxy and carboxy ferrous cytochrome P450: Comparison of semiempirical and ab-initio calculations. International Journal of Quantum Chemistry, 16 (S6):93–104, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Riera:1972:GRC

[RM72] A. Riera and W. J. Meath. Gaussian representations of charge overlap effects in intermolecular forces. International Journal of Quantum Chemistry, 6(3):501–508, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Riera:1973:NSO

[RM73] A. Riera and William J. Meath. Nondiagonal second order intermolecular forces for interactions involving molecules. International Journal of Quantum Chemistry, 7(5):959–970, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Redmon:1974:IPD

[RM74] Michael J. Redmon and David A. Micha. Interaction potentials and dynamics for Li + F collisions. International Journal of Quantum Chemistry, 8(S8):253–262, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Relyea:1979:MTP

[RM79] Lawrence R. Relyea and David A. Micha. Multichannel treatment of Penning ionization in He* (1s2s,^{3} S) + Ar with discretization of the electronic continuum. International Journal of Quantum Chemistry, 16(S13):569–577, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Ray:1978:FSG

[RMB78] N. K. Ray, S. P. Mehandru, and Shobha Bhargava. Floating spherical Gaussian orbital model study of some organometallic systems. I. LiCH_{3}, HBeCH_{3}, and CH_{3}BeCH_{3}. International Journal of Quantum Chemistry, 13(4):529–536, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Robinson:1971:IFW

[RMP71] G. Robinson, N. H. March, and R. C. Perrin. Interatomic forces over a wide range of nuclear separations. International Journal of Quantum Chemistry, 5(S5):261–276, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Roby:1971:TEC

[Rob71] Keith R. Roby. On the theory of electron correlation in atoms and molecules: Relation between cluster expansion theory and the correlated wave functions method. International Journal of Quantum Chemistry, 5(2):119–130, March 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Robin:1972:ESR

[Rob72a] M. B. Robin. Experimental study of the Rydberg orbitals of large molecules. International Journal of Quantum Chemistry, 6(S6):257–266, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Roby:1972:TEC

[Rob72b] Keith R. Roby. On the theory of electron correlation in atoms and molecules. II. General cluster expansion theory and the general correlated wave functions method. International Journal of Quantum Chemistry, 6(1):101–123, January 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roeggen:1973:APA

[Røe73] I. Røeggen. Applications of Padé approximants of type II in partitioning technique. International Journal of Quantum Chemistry, 7(6):1193–1201, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roeggen:1975:RTP

[Røe75] I. Røeggen. Resolvent technique and Padé approximants in configuration interaction calculations. International Journal of Quantum Chemistry, 9(3):431–443, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rosen:1974:RQT

[Ros74] Robert Rosen. The role of quantum theory in biology. International Journal of Quantum Chemistry, 8(S1):229–232, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Rosen:1975:UC

[Ros75] Philip Rosen. Ultraviolet carcinogenesis. International Journal of Quantum Chemistry, 9(3):473–478, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rosen:1978:RSO

[Ros78a] Arne Rosén. Relativistic symmetry orbitals for the double groups C_{2v}, C_{∞v}, D_{∞h}, and O_{h}. International Journal of Quantum Chemistry, 13(4):509–528, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rosina:1978:ATB

[Ros78b] M. Rosina. Application of the two-body density matrix of the ground state for calculations of some excited states. International Journal of Quantum Chemistry, 13(6):737–742, June 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rouse:1973:CAS

[Rou73] Robert A. Rouse. Conformational analysis of saturated heterocycles substituted final ozonides. International Journal of Quantum Chemistry, 7(S7):289–294, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Rouse:1974:CAS

[Rou74] Robert A. Rouse. Conformational analysis of saturated heterocycles ab initio studies of ethylene ozonides. International Journal of Quantum Chemistry, 8(S8):201–207, January 20– 26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Rouzo:1978:DMS

[Rou78] H. Le Rouzo. On the defects of molecular symmetry-lobe orbitals. International Journal of Quantum Chemistry, 14(1): 121–126, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rouzo:1979:USS

[Rou79] H. Le Rouzo. On the use of spatial symmetry in atomic-integral calculations: An efficient permutational approach. International Journal of Quantum Chemistry, 15(1):57–71, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ryback:1978:AMC

[RPK78] W. Ryback, R. Poirier, and R. Kari. An application of the method of conjugate gradients to the calculation of minimal basis set localized orbitals. International Journal of Quantum Chemistry, 13(1):1–16, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rabinowitz:1972:CMQ

[RR72] James R. Rabinowitz and Robert Rein. Calculation of molecular quadrupole moments and a demonstration of the importance of overlap densities in the theory of polyatomic molecules. International Journal of Quantum Chemistry, 6(4):669–680, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rastogi:1977:FSG

[RR77] R. C. Rastogi and N. K. Ray. The floating spherical Gaussian orbital model and shape of B_{2}H_{5} ion. International Journal of Quantum Chemistry, 11(3):435–440, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rein:1975:TSC

[RRNS75] Robert Rein, V. Renugopalakrishnan, S. Nir, and T. J. Swissler. Theoretical studies on the conformation of peptides in solution I. Conformation of N-acetyl glycine N-methyl amide in solution. International Journal of Quantum Chemistry, 9(S2):99– 107, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Renugopalakrishnan:1975:CAG

[RRS75] V. Renugopalakrishnan, M. Renugopalakrishnan, and Bibudhendra Sarkar. Conformational analysis of β-glycine, l-alanine, and bisglycinato-Cu(II) complex. International Journal of Quantum Chemistry, 9(S2):109–116, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Raimondi:1976:OPV

[RS76] Mario Raimondi and Massimo Simonetta. The orthogonality problem in valence bond calculations. International Journal of Quantum Chemistry, 10(6):1057–1059, November 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rouzo:1978:ADGa

[RS78a] H. Le Rouzo and B. Silvi. Angular dependence of Gaussian-Lobe orbitals. I. Analysis of standard p- and d-orbitals. International Journal of Quantum Chemistry, 13(3):297–310, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rouzo:1978:ADGb

[RS78b] H. Le Rouzo and B. Silvi. Angular dependence of Gaussian-Lobe orbitals. II. Set of axial Gaussian-Lobe orbitals. International Journal of Quantum Chemistry, 13(3):311–324, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rouzo:1978:ADGc

[RS78c] H. Le Rouzo and B. Silvi. Angular dependence of Gaussian-Lobe orbitals. III. Polyhedric lobe edifices. International Journal of Quantum Chemistry, 13(3):325–330, March 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roothaan:1979:ASN

[RS79] C. J. Clemens Roothaan and George A. Soukup. Accurate and stable numerical Hartree–Fock calculations for atoms. I. The 1 s^{2} ground state of H^{−}, He, Li^{+}, and Be^{++}. International Journal of Quantum Chemistry, 15(5):449–462, May 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Renugopalakrishnan:1978:CSC

[RSK^{+}78] V. Renugopalakrishnan, H. Sugano, M. A. Khaled, R. S. Rapaka, and D. W. Urry. Conformational studies of cyclohexapeptide analogs of elastin sequences: Cyclo(Ala-Pro-Gly-Ala-Pro-Gly). International Journal of Quantum Chemistry, 14(S5): 69–78, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Rabinowitz:1972:TIT

[RSR72] James R. Rabinowitz, T. J. Swissler, and Robert Rein. Towards an interaction theory for large molecules. International Journal of Quantum Chemistry, 6(S6):353–361, January 17–22, 1972.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Raina:1975:SWM

[RSS75] V. Raina, P. K. Srivastava, and K. V. Sane. Square well model and the functional form for the resonance integral. International Journal of Quantum Chemistry, 9(2):379–381, March 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roy:1977:GEC

[RT77] C. L. Roy and S. K. Tripathy. Generalized existence conditions for surface states of a semi-infinite Kronig–Penney crystal. International Journal of Quantum Chemistry, 11(1):17–20, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ratner:1977:PCI

[RTM77] Mark A. Ratner, Sid Topiol, and J. W. Moskowitz. Pseudopotential calculations. IV. Some results for zinc difluoride. International Journal of Quantum Chemistry, 12(S11):493–499, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Raimondi:1974:IVB

[RTS74] M. Raimondi, G. F. Tantardini, and M. Simonetta. Ab initio valence bond treatment of simple molecules. International Journal of Quantum Chemistry, 8(S8):285–291, January 20– 26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Renugopalakrishnan:1976:CEC

[RU76] V. Renugopalakrishnan and D. W. Urry. Conformational energy calculations demonstrating the β turn and other secondary structural features in the repeat pentapeptide of elastin. International Journal of Quantum Chemistry, 10(S3):13–19, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ruelle:1979:MBP

[Rue79] Jean-Louis Ruelle. Modified bielectronic potential method for correlation energy. International Journal of Quantum Chemistry, 15(3):293–309, March 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rosen:1974:RCF

[RW74] A. Rosén and J. T. Waber. Relativistic crystal-field radial integrals for rare-earth ions. International Journal of Quantum Chemistry, 8(S8):127–136, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Redmon:1975:TDQ

[RW75] Michael J. Redmon and Robert E. Wyatt. Three-dimensional quantum mechanical studies of the H + H_{2} and F + H_{2} reactions. International Journal of Quantum Chemistry, 9(S9): 403–414, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Redmon:1977:CMR

[RW77] Michael J. Redmon and Robert E. Wyatt. Computational methods for reactive scattering. International Journal of Quantum Chemistry, 12(S11):343–351, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Randic:1979:GTS

[RW79] Milan Randić and Charles L. Wilkins. Graph theoretical study of structural similarity in benzomorphans. International Journal of Quantum Chemistry, 16(S6):55–71, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Richards:1971:BIM

[RWFS71] W. G. (William Graham) Richards, T. E. H. Walker, L. Farnell, and P. R. Scott, editors. A Bibliography of ab initio molecular wave functions. Supplement for 1970–1973. Oxford science research papers. Clarendon Press, New York, NY, USA, 1971. ???? pp. LCCN Z5524.P6 R52 Suppl.

Samanta:1977:PTH

[SA77] S. R. Samanta and M. A. Ali. Perturbation treatment of the Hartree–Fock equations of 1s2p3s ^{4}P^{θ} state of the lithium isoelectronic sequence. International Journal of Quantum Chemistry, 12(3):561–568, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sabin:1971:CCS

[Sab71] John R. Sabin. A comment concerning S − H − S type hydrogen bonds. International Journal of Quantum Chemistry, 5(S5):133–136, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sabatier:1973:ACP

[Sab73] Pierre C. Sabatier. Ambiguities in the construction of potentials from phase shifts. International Journal of Quantum Chemistry, 7(S7):421–425, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sabin:1975:BRB

[Sab75a] John R. Sabin. Book review: Bibliography of ab initio molecular wave functions. Supplement for 1970–1973. Authors: W. G. Richards, T. E. H. Walker, L. Farnell, and P. R. Scott. Published by: Oxford University Press, London, 1974. Price: $29.95 (£9.50) No. of Pages: 358. International Journal of Quantum Chemistry, 9(4):757–758, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [RWFS71].

Sabin:1975:EN

[Sab75b] John R. Sabin. Editorial note. International Journal of Quantum Chemistry, 9(S2):345, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sack:1972:AFM

[Sac72] R. A. Sack. An alternative formalism for the method of intermediate Hamiltonians. International Journal of Quantum Chemistry, 6(5):989–999, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Salhany:1972:IPS

[Sal72] J. M. Salhany. The influence of protons and salts on the reactivity of hemoglobin with oxygen. International Journal of Quantum Chemistry, 6(S6):367–374, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Salmon:1973:CCS

[Sal73] Lydia S. Salmon. Convergence of the completely separated bipolar expansion of r_{12}^{−1}. International Journal of Quantum Chemistry, 7(3):411–424, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Salem:1979:MCR

[Sal79] L. Salem. The mechanism of the chemical reaction, with recent developments pertaining to heterogeneous catalysis. International Journal of Quantum Chemistry, 16(S13):321–330, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Sambe:1975:ASM

[Sam75] Hideo Sambe. An approximate SCF MO theory of molecules based on Slater’s Xα method. International Journal of Quantum Chemistry, 9(S9):95–104, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Sarkisov:1979:SNG

[SAM79] G. N. Sarkisov, Yu. V. Agrafonov, and G. A. Martynov. Solvation of nonpolar groups of biomacromolecules. International Journal of Quantum Chemistry, 16(4):869–875, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sannigrahi:1971:ESE

[San71] A. B. Sannigrahi. Effect of semi-empirical parameters on the triplet energy levels and triplet–triplet transition in benzene. International Journal of Quantum Chemistry, 5(3):265–272, May 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sannigrahi:1972:SRC

[San72] A. B. Sannigrahi. Some remarks on the core resonance integral as used in the semi-empirical π-electron calculations. International Journal of Quantum Chemistry, 6(2):281–287, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sarma:1975:SPE

[Sar75] C. R. Sarma. Spin projections for even electron systems. International Journal of Quantum Chemistry, 9(6):1109–1112, November 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sasaki:1974:MEC

[Sas74] Fukashi Sasaki. Matrix elements in configuration interaction calculations. International Journal of Quantum Chemistry, 8 (4):605–617, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sasaki:1977:ECI

[Sas77] Fukashi Sasaki. Effectiveness of configuration interaction calculations for large molecules. International Journal of Quantum Chemistry, 12(S11):125–130, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Schieve:1975:HTD

[SB75] W. C. Schieve and Adi R. Bulsara. The higher time derivatives of the generalized Liapunov function. International Journal of Quantum Chemistry, 9(S9):471–475, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Smeyers:1978:HPH

[SB78] Yves G. Smeyers and Angel M. BruceñNta. Half-projected Hartree–Fock model for computing potential-energy surfaces. International Journal of Quantum Chemistry, 14(5):641–648, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shillady:1979:DOG

[SBB79] Donald D. Shillady and Sheryl Baldwin-Boisclair. Dipoleoptimized Gaussian orbitals for rapid computation of electrostatic molecular potential contour maps. International Journal of Quantum Chemistry, 16(S6):105–114, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sutton:1970:MCM

[SBD^{+}70] P. Sutton, P. Bertoncini, G. Das, T. L. Gilbert, Arnold C. Wahl, and O. Sinanoğlu. Methods for correlating molecules and some optimized valence configuration results on the diatomic molecules Li_{2}, Be_{2}, B_{2}C_{2}, N_{2}, F_{2}; BN, BeO, LiF, HeNe; CO and BF. International Journal of Quantum Chemistry, 4(6): 633–635, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stoddart:1970:OBP

[SBM70] J. C. Stoddart, A. M. Beattie, and N. H. March. One-body potentials including electron interactions, and electron densities in crystals. International Journal of Quantum Chemistry, 5(S4):35–53, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Singh:1972:LRI

[SBM72] T. R. Singh, J. F. Bukta, and W. J. Meath. Long range interaction energies using Gaussian basis sets and a one center method. International Journal of Quantum Chemistry, 6 (2):201–209, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sperber:1973:SAMa

[SC73] Gunnar Sperber and Jean-Louis Calais. A study of the AMO method as applied to the lithium metal. II. The first-order density matrix and the shape of the Fermi surface. International Journal of Quantum Chemistry, 7(3):521–536, May 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schwarz:1976:MWF

[SC76a] W. H. E. Schwarz and T. C. Chang. Multiconfiguration wave functions for highly excited states by the Generalized Brillouin Theorem method. International Journal of Quantum Chemistry, 10(S10):91–97, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Slater:1976:ROS

[SC76b] J. C. Slater and J. W. D. Connolly. Remarks on the overlapping-sphere method for molecular orbitals. International Journal of Quantum Chemistry, 10(S10):141–146, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sokalski:1978:AEP

[SC78a] W. Andrzej Sokalski and Henryk Chojnacki. Approximate exchange perturbation study of intermolecular interactions in molecular complexes. International Journal of Quantum Chemistry, 13(5):679–692, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Spangler:1978:DBS

[SC78b] Dale Spangler and Ralph E. Christoffersen. Development of basis sets for molecular calculations. Comparison of tetrahedral Gaussian lobe functions and Cartesian Gaussian basis sets.

International Journal of Quantum Chemistry, 14(S5):127–135, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Singhal:1979:PDA

[SC79] S. P. Singhal and J. Callaway. Point defects in aluminum: Single vacancy. International Journal of Quantum Chemistry, 16 (S13):363–366, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Schaffer:1971:EOI

[Sch71a] Claus E. Schäffer. Energy operators as idempotents in the angular overlap model of the ligand field. International Journal of Quantum Chemistry, 5(S5):379–390, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Schurr:1971:TDQa

[Sch71b] J. M. Schurr. Time-dependent quantum theory I. An absorber in an intense radiation field. International Journal of Quantum Chemistry, 5(1):13–34, January 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schurr:1971:TDQb

[Sch71c] J. M. Schurr. Time-dependent quantum theory II. Absorption of light by dimers: Quantum theory and classical analogy. International Journal of Quantum Chemistry, 5(1):35–65, January 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schurr:1971:TDQc

[Sch71d] J. Michael Schurr. Time-dependent quantum theory III. Model for vibrational relaxation in crystals. International Journal of Quantum Chemistry, 5(3):239–263, May 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schurr:1971:TDQd

[Sch71e] J. Michael Schurr. Time-dependent quantum theory. IV. Effect of lattice relaxation on the optical spectra. International Journal of Quantum Chemistry, 5(2):221–234, March 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scherr:1972:SNE

[Sch72] Charles W. Scherr. A simple non-exponential function for use in atomic and molecular wave functions for use in atomic and molecular wave functions. International Journal of Quantum Chemistry, 6(S6):51–58, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Schluter:1973:PCD

[Sch73a] Michael Schlüter. Pseudopotential charge densities of trigonal Se and Te. International Journal of Quantum Chemistry, 7 (S7):527–532, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Schmidtke:1973:TOS

[Sch73b] Hans-Herbert Schmidtke. Über die Topologie und Orbitalstruktur Symmetrischer Moleküle. (German) [On the topology and orbital structure of symmetric molecules]. International Journal of Quantum Chemistry, 7(1):127–130, January 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schieve:1976:GEP

[Sch76a] W. C. Schieve. Generalized entropy: Its properties. International Journal of Quantum Chemistry, 10(S10):375–382, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Schurr:1976:LAH

[Sch76b] J. Michael Schurr. Light absorption at high intensities. Comparison of quantum theory and semi-classical results. International Journal of Quantum Chemistry, 10(2):359–377, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scharfenberg:1977:NAS

[Sch77a] Peter Scharfenberg. A new algorithm for the symmetric (Löwdin) orthonormalization. International Journal of Quantum Chemistry, 12(5):827–839, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schmelzer:1977:ESH

[Sch77b] A. Schmelzer. Equivalent spherical harmonics. International Journal of Quantum Chemistry, 11(4):561–576, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schonhammer:1977:THC

[Sch77c] K. Schönhammer. Theory of hydrogen chemisorption on transition metals. International Journal of Quantum Chemistry, 12 (S11):517–527, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Scharfenberg:1978:SRT

[Sch78] Peter Scharfenberg. Some remarks on a transformation satisfying the constraint of maximum similarity of concerned functions. International Journal of Quantum Chemistry, 13(1):155– 158, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scofield:1971:ASE

[Sco71] D. F. Scofield. Aluminum SCOPW energy band dependence upon the local exchange approximation. International Journal of Quantum Chemistry, 5(S5):489–495, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Scofield:1973:NLO

[Sco73] D. F. Scofield. A note on Löwdin orthogonalization and the square root of a positive self-adjoint matrix. International Journal of Quantum Chemistry, 7(3):561–568, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scofield:1974:ASC

[Sco74] D. F. Scofield. Approximate self-consistent Green’s functions for solids. International Journal of Quantum Chemistry, 8(S8): 421–436, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Song:1975:ESP

[SCYB75] Pill-Soon Song, Chen-An Chin, Iwao Yamazaki, and Hiroaki Baba. Electronic structure and photoreactivity of the triplet states of skin-sensitizing psoralens. International Journal of Quantum Chemistry, 9(S2):1–8, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Song:1977:ESP

[SCYB77] Pill-Soon Song, Chen-An Chin, Iwao Yamazaki, and Hiroaki Baba. Excited states of photobiological receptors. II. Chlorophylls, phytochrome, and stentorin. International Journal of Quantum Chemistry, 12(S4):305–315, January 9–12, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Siu:1970:SGS

[SD70] Albert K. Q. Siu and Ernest R. Davidson. A study of the ground state wave function of carbon monoxide. International Journal of Quantum Chemistry, 4(3):223–243, May 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smith:1974:NTO

[SD74] Darwin W. Smith and Orville W. Day. Natural transition orbitals and Rowe’s equations of motion. International Journal of Quantum Chemistry, 8(S8):511–513, January 20–26, 1974.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Sarma:1977:CPN

[SD77] Ramaswamy H. Sarma and Steven S. Danyluk. Conformational properties of nucleic acids in solution. International Journal of Quantum Chemistry, 12(S4):269–276, January 9–12, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sannigrahi:1978:SVE

[SD78] A. B. Sannigrahi and B. R. De. Selected valence-electron splitshell MO calculations on ArF^{+}, KrF^{+}, and XeF^{+}. International Journal of Quantum Chemistry, 14(3):329–331, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sarma:1979:SFC

[SD79a] C. R. Sarma and K. V. Dinesha. Spin-free configurationinteraction study of molecules using single-parameter alternate molecular orbitals. International Journal of Quantum Chemistry, 15(6):579–588, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sarma:1979:SCS

[SD79b] C. R. Sarma and K. V. Dinesha. Subduction coefficients for < 2^{N/2−S}, 1^{2S} >↓< 2^{N_{1}/2−S_{1}}, 1^{2S_{1}} > ⊗ < 2^{N_{2}/2} − S_{2} >, 1^{2S_{2}} > of U(n) ↓ U(n_{1}) ⊗ U(n_{2}). International Journal of Quantum Chemistry, 16(6):1195–1207, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smeyers:1974:HPP

[SDB74] Yves G. Smeyers and G. Delgado-Barrio. Half-projected and projected Hartree–Fock calculations for singlet ground states. II. Lithium hydride. International Journal of Quantum Chemistry, 8(5):733–743, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smeyers:1976:AHP

[SDB76] Yves. G. Smeyers and G. Delgado-Barrio. Analysis of the halfprojected Hartree–Fock function: Density matrix, natural orbitals, and configuration interaction equivalence. International Journal of Quantum Chemistry, 10(3):461–472, May 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scofield:1972:CEA

[SDD72] D. F. Scofield, N. C. Dutta, and C. M. Dutta. The correlation energy of atomic fluorine. International Journal of Quantum Chemistry, 6(S6):9–17, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sanchez:1977:UGL

[SDD^{+}77] M. Sanchez, R. Daudel, P. D. Dacre, R. McWeeny, S. Kwun, and C. Valdemoro. Using group (or Loge) functions to explore the transferability of chemical bonds. International Journal of Quantum Chemistry, 11(3):415–425, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shim:1979:IHF

[SDJ79] Irene Shim, Jens Peder Dahl, and Helge Johansen. Ab initio Hartree–Fock and configuration-interaction treatment of the interaction between two nickel atoms. International Journal of Quantum Chemistry, 15(3):311–331, March 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sherman:1974:PMS

[SDM74] Douglas Franklyn Sherman, Daniel Dilella, and Kenneth J. Miller. Partitioning of molecular systems and subsystem replacement by single-centered perturbations: II. Applications of the pseudo substituent technique within Hückel molecular orbital theory. International Journal of Quantum Chemistry, 8 (S8):187–199, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Smeyers:1973:HPP

[SDS73] Yves G. Smeyers and L. Doreste-Suarez. Half-projected and Projected Hartree–Fock Calculations for Singlet Ground States. i. four-Electron Atomic Systems. International Journal of Quantum Chemistry, 7(4):687–698, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shukla:1979:DIB

[SE79] G. C. Shukla and S. I. Easa. Dispersion interaction between helium pair from hydrodynamic analogy to quantum mechanics. International Journal of Quantum Chemistry, 16(6):1239–1247, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Selsby:1975:VCS

[Sel75] R. G. Selsby. Variable ζ calculation for self-consistent screening parameters in ab initio molecular theory. International Journal of Quantum Chemistry, 9(1):83–102, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Surratt:1973:HFL

[SEW73] G. T. Surratt, R. N. Euwema, and D. L. Wilhite. Hartree– Fock lattice constant and bulk modulus of diamond. International Journal of Quantum Chemistry, 7(S7):607–611, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Seybold:1976:WTF

[Sey76] Paul G. Seybold. Why are there four bases in DNA? International Journal of Quantum Chemistry, 10(S3):39–43, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Steinborn:1975:LTB

[SF75] E. O. Steinborn and E. Filter. Laplace-type, bipolar, and polypolar expansions of r^{N}Y_{L}^{M}(θ, φ) and Helmholtz harmonics j_{L}(kr)Y_{L}^{M}(θ, φ) for molecular calculations. International Journal of Quantum Chemistry, 9(S9):435–442, January 19– 25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Sambe:1976:CBX

[SF76] Hideo Sambe and Ronald H. Felton. Connection between the Xα method and ligand field theory. International Journal of Quantum Chemistry, 10(S10):155–158, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Szent-Gyorgyi:1976:ETC

[SG76a] Albert Szent-Györgyi. The electronic theory of cancer. International Journal of Quantum Chemistry, 10(S3):45–50, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Szent-Gyorgyi:1976:SRM

[SG76b] Albert Szent-Györgyi. Some reminiscences of my life as a scientist. International Journal of Quantum Chemistry, 10(S3): 7–12, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Selsby:1977:SEMb

[SG77a] R. G. Selsby and Alec Grimison. A semi-empirical MO theory for ionization potentials and electron affinities. II. Vertical and adiabatic values, benzenoid and nonbenzenoid aromatic hydrocarbons, and conjugated molecules with heteroatoms. International Journal of Quantum Chemistry, 12(3):527–544, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Szent-Gyorgyi:1977:PRR

[SG77b] Albert Szent-Györgyi. Protein radicals, regulations, and cancer. International Journal of Quantum Chemistry, 12(S4):179– 184, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Szent-Gyorgyi:1977:WR

[SG77c] Albert Szent-Györgyi. Welcoming remarks. International Journal of Quantum Chemistry, 12(S4):1, January 9–12, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Szent-Gyorgyi:1978:LS

[SGM78] Albert Szent-Györgyi and J. A. McLaughlin. The living state. International Journal of Quantum Chemistry, 14(S5):137–141, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Saunders:1973:LSM

[SH73] V. R. Saunders and I. H. Hillier. A “Level–Shifting” method for converging closed shell Hartree–Fock wave functions. International Journal of Quantum Chemistry, 7(4):699–705, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Svendsen:1974:CMO

[SH74] E. Nørby Svendsen and Hendrik F. Hameka. Calculation of the magneto–optical activity of the hydrogen molecule. International Journal of Quantum Chemistry, 8(5):789–797, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sims:1975:CCH

[SH75a] James S. Sims and Stanley A. Hagstrom. Combined CI–HY studies of atomic states. II. Compact wave functions for the be ground state. International Journal of Quantum Chemistry, 9 (1):149–156, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Svendsen:1975:ERM

[SH75b] E. Nørby Svendsen and Hendrik F. Hameka. Effect of the rotational motion of molecules on their optical and magneto-optical activity. International Journal of Quantum Chemistry, 9(5): 797–815, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shankland:1971:FTS

[Sha71] D. G. Shankland. Fourier transformation by smooth interpolation. International Journal of Quantum Chemistry, 5(S5):497– 500, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sharma:1976:ETC

[Sha76] R. R. Sharma. Evaluations of two-center integrals using αfunctions. International Journal of Quantum Chemistry, 10(6): 1075–1076, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sharma:1977:AMC

[Sha77a] K. K. Sharma. Additivity model for calculations of UHF spin densities and charge densities in some methyl-substituted radical anions. International Journal of Quantum Chemistry, 11 (5):753–758, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shavitt:1977:GTC

[Sha77b] Isaiah Shavitt. Graph theoretical concepts for the unitary group approach to the many-electron correlation problem. International Journal of Quantum Chemistry, 12(S11):131–148, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Sharma:1978:AMC

[Sha78a] K. K. Sharma. Additivity model for calculations of UHF spin densities in some aza-aromatic radical anions. International Journal of Quantum Chemistry, 13(5):635–639, May 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shavitt:1978:MEE

[Sha78b] Isaiah Shavitt. Matrix element evaluation in the unitary group approach to the electron correlation problem. International Journal of Quantum Chemistry, 14(S12):5–32, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Sakamoto:1973:CIE

[SHI73] Kazuyoshi Sakamoto, Taisen Hayashi, and Y. J. I’Haya. Core integrals of 2pπ-electron evaluated from Heisenberg’s equation of motion. International Journal of Quantum Chemistry, 7 (S7):337–345, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sasaki:1971:ERE

[SHO71] Fukashi Sasaki, Yasuyo Horino, and Kimio Ohno. The effect of removal of “equivalence” restrictions on the electronic energies of atoms. International Journal of Quantum Chemistry, 5(S5): 21–33, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sims:1976:CCH

[SHR76] James S. Sims, Stanley A. Hagstrom, and John R. Rumble, Jr. Combined CI–HY studies of atomic states. IV. The four lowest 1S and four lowest ^{1}P states of He and the lowest ^{1}S and ^{1}P states of H^{−1}. International Journal of Quantum Chemistry, 10(5):853–866, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shull:1977:BRBb

[Shu77a] Harrison Shull. Book review: Quantum mechanics of molecular conformations. Edited by Bernard Pullman. John Wiley & Sons, London, 1976. 412 pages. Price: £16.00 ($30.00). International Journal of Quantum Chemistry, 11(6):1045, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Pul76].

Shull:1977:BRBa

[Shu77b] Harrison Shull. Book review: Quantum science. Methods and structure. Edited by J.-L. Calais, O. Goscinski, J. Linderberg, and Y. Öhrn. Plenum Press, New York, 1976. 595 pages. Price: $49.50. International Journal of Quantum Chemistry, 11(6):

1043, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [CGLÖ76].

Shultz:1977:CTM

[Shu77c] Mary Jane Shultz. A canonical transformation method for the treatment of a strong coupling problem. International Journal of Quantum Chemistry, 12(S11):251–257, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Silver:1971:PEM

[Sil71] David M. Silver. Poisson equation for molecular exchange, hybrid and Coulomb electron repulsion integrals. International Journal of Quantum Chemistry, 5(5):505–512, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simons:1972:UEC

[Sim72] Jack Simons. The use of explicitly correlated, partially antisymmetric wave functions in atomic and molecular calculations. International Journal of Quantum Chemistry, 6(3):439–448, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shibuya:1975:DSE

[SIM75] Tai-Ichi Shibuya, Yasumasa J. I’haya, and Vincent McKoy. Dynamical screening by Σ-electrons in π-electron transitions: Treatment in the equations-of-motion method. International Journal of Quantum Chemistry, 9(3):505–516, May 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simons:1977:NDB

[Sim77a] Jack Simons. A Note on Differences Between Operator-Level and Function-Level Equations of Motion. International Journal of Quantum Chemistry, 12(1):227–229, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simons:1977:SST

[Sim77b] Jack Simons. A survey of some theoretical studies of negative ions. International Journal of Quantum Chemistry, 11(6):971– 978, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simon:1978:RCS

[Sim78a] Barry Simon. Resonances and complex scaling: A rigorous overview. International Journal of Quantum Chemistry, 14(4): 529–542, October 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simons:1978:DMC

[Sim78b] Jack Simons. Dynamics of molecules in contact with an external medium at equilibrium. International Journal of Quantum Chemistry, 13(4):553–562, April 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simons:1978:NAS

[Sim78c] Jack Simons. Nature of the autodetaching sub ^{2}P_{1/2} threshold states of the alkali anions. International Journal of Quantum Chemistry, 14(3):333–336, September 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simonetta:1979:CSS

[Sim79] Massimo Simonetta. Colloquium on the structure small molecules adsorbed on surfaces. International Journal of Quantum Chemistry, 15(3):359, March 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sinanoglu:1973:ABD

[Sin73a] Oktay Sinanoğlu. On the agreement between dipole length and dipole velocity calculated oscillator strengths. International Journal of Quantum Chemistry, 7(S7):65–67, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sinanoglu:1973:RCM

[Sin73b] Oktay Sinanoğlu. Recent CI methods for the calculation of pair correlations and more-electron clusters in ground-state atoms. International Journal of Quantum Chemistry, 7(S7):57– 63, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sinanoglu:1973:RDN

[Sin73c] Oktay Sinanoğlu. Remarks on dynamical and noncompact groups in physics and chemistry. International Journal of Quantum Chemistry, 7(S7):45–52, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sinanoglu:1975:CDT

[Sin75] Oktay Sinanoğlu. Charge densities and transition densities from the theory of non-closed-shell states and their experimental tests. International Journal of Quantum Chemistry, 9(S9): 155–164, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Sisler:1974:RW

[Sis74] Harry H. Sisler. Remarks of welcome. International Journal of Quantum Chemistry, 8(S1):3–4, January 17–19, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sklenar:1979:MSB

[SJ79] Heinz Sklenar and Joachim Jäger. Molecular structure– biological activity relationships on the basis of quantumchemical calculations. International Journal of Quantum Chemistry, 16(3):467–484, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smith:1979:DFT

[SJH79] Darwin W. Smith, Salem Jagannathan, and George S. Handler. Density functional theory of atomic structure. I. Exchange and correlation potentials for two-electron atoms. International Journal of Quantum Chemistry, 16(S13):103–110, March 11– 17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Sahni:1971:CEV

[SK71] V. Sahni and J. B. Krieger. Calculation of expectation values correct to second order applied to the ground state of the helium isoelectronic sequence. International Journal of Quantum Chemistry, 5(S5):47–58, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sahni:1972:RAC

[SK72] V. Sahni and J. B. Krieger. Results of an analytic calculation for the coherent atomic scattering factor for the ground state of the helium isoelectronic sequence. International Journal of Quantum Chemistry, 6(S6):103–109, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Seitz:1973:CHS

[SK73] W. A. Seitz and D. J. Klein. Computations on Heisenberg Spin Models. International Journal of Quantum Chemistry, 7(4):647–665, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Steslicka:1979:SSE

[SK79] Maria Steślicka and Krzysztof Kempa. Surface structure and energy spectra of localized states. International Journal of Quantum Chemistry, 15(3):281–291, March 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shilov:1979:MSR

[SKL79] V. V. Shilov, N. E. Kruglyak, and Yu. S. Lipatov. Modeling of the short-range order for polymers by the ideal peak method. International Journal of Quantum Chemistry, 16(4):799–808, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sinkkonen:1977:IBC

[SKS77] J. Sinkkonen, K. Kaski, and T. Stubb. Interactions between conduction and magnetic electrons in europium chalcogenides. International Journal of Quantum Chemistry, 12(S11):595– 606, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Shaw:1972:PSI

[SKW72] Jerry C. Shaw, J. B. Ketterson, and L. R. Windmiller. Phase shift inversion of Fermi surface data for some of the transition metals. International Journal of Quantum Chemistry, 6 (S6):395–402, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Stanley:1970:DRT

[SL70] H. Eugene Stanley and M. Howard Lee. Diagrammatic representation of the two-spin correlation function for the generalized Heisenberg model. International Journal of Quantum Chemistry, 5(S4):407–418, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Suhai:1973:CME

[SL73] S. Suhai and J. Ladik. CNDO/2 and MINDO/2 energy band structures of some homopolynucleotides. International Journal of Quantum Chemistry, 7(3):547–560, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saute:1979:GSQ

[SL79] M. Saute and A. Laforgue. Generalization of the second quantization operators: Electron attachment and detachment operators. International Journal of Quantum Chemistry, 16(3): 501–507, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slater:1970:SCF

[Sla70a] J. C. Slater. The self-consistent field for crystals. International Journal of Quantum Chemistry, 3S(??):727–?? (or 158–??), ?? 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slater:1970:NSP

[Sla70b] John C. Slater. Note on the space part of anti-symmetric wave functions in the many-electron problem. International Journal of Quantum Chemistry, 4(6):561–570, November 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slater:1971:TEA

[Sla71a] J. C. Slater. Treatment of exchange in atomic, molecular, and solid-state theory. International Journal of Quantum Chemistry, 5(S5):403–409, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Slater:1971:IPJ

[Sla71b] John C. Slater. Introduction of Professor J. H. Van Vleck, Sanibel Symposium, Monday, January 18, 1971. International Journal of Quantum Chemistry, 5(S5):1–2, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Slater:1973:FPX

[Sla73a] J. C. Slater. Future prospects for the X_{α} method. International Journal of Quantum Chemistry, 7(S7):533–544, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Slater:1973:IUC

[Sla73b] J. C. Slater. Introduction of E. U. Condon, Sanibel Symposium January 22, 1973. International Journal of Quantum Chemistry, 7(S7):3–5, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Slater:1974:RIM

[Sla74] J. C. Slater. Recent improvements in the muffin-tin method. International Journal of Quantum Chemistry, 8(S8):81–88, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Slater:1975:CTX

[Sla75] J. C. Slater. Comparison of TFD and Xα methods for molecules and solids. International Journal of Quantum Chemistry, 9 (S9):7–21, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Slanina:1979:MCT

[Sla79] Z. Slanina. Multimolecular clusters: Their isomerism and effective characteristics evaluated by quantum chemistry. International Journal of Quantum Chemistry, 16(1):79–86, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schug:1977:CME

[SLB77] John C. Schug, Byron H. Lengsfield, and Dana A. Brewer. Calculation of molecular electronic spectra by projectedunrestricted Hartree–Fock theory. International Journal of Quantum Chemistry, 11(4):591–604, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simpson:1978:ILM

[SLC78] R. W. Simpson, N. F. Lane, and R. C. Chaney. Ab initio LCAO-MO cluster-type calculation of the self-consistent electronic screening charge density around a single hydrogen impurity in a nickel crystal. International Journal of Quantum Chemistry, 14(S12):433–436, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Seiders:1978:BEM

[SLJ78] Barbara A. B. Seiders, William L. Luken, and Kenneth D. Jordan. Binding of an electron to a molecular dipole: BeF. International Journal of Quantum Chemistry, 14(6):741–746, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smith:1970:IAS

[SLM70] D. W. Smith, E. G. Larson, and R. C. Morrison. On the interpretative aspects of second-order reduced density matrices. International Journal of Quantum Chemistry, 3S(??):689–??, ?? 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Strelko:1979:QCI

[SLOKP79] V. V. Strelko, K. D. Lavrinenko-Ometzinskaja, T. A. Karaseva, and L. V. Pisarzhevsky. Quantum-chemical investigation of cation motion mechanism in modes of biological channels with phosphate groups. International Journal of Quantum Chemistry, 16(4):753–760, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saxena:1971:SOOa

[SM71a] K. M. S. Saxena and Gulzari Malli. Spin-other-orbit and spinorbit splitting factors for f^{6} electron configurations. International Journal of Quantum Chemistry, 5(S5):619–630, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Saxena:1971:SOOb

[SM71b] K. M. S. Saxena and Gulzari Malli. Spin-other-orbit splitting factors for f^{7} electron configurations. International Journal of Quantum Chemistry, 5(S5):631–642, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Saxena:1971:SSS

[SM71c] K. M. S. Saxena and Gulzari Malli. Spin-spin splitting factors for f^{7} electron configurations. International Journal of Quantum Chemistry, 5(S5):613–618, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sannigrahi:1973:SVE

[SM73a] A. B. Sannigrahi and S. Noor Mohammad. Selected valence electron split-shell molecular orbital calculations on the diatomic interhalogen molecules. International Journal of Quantum Chemistry, 7(6):1183–1192, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stoney:1973:WDO

[SM73b] P. Stoney and N. H. March. Wavenumber dependent orbital diamagnetism of nearly free electrons. International Journal of Quantum Chemistry, 7(S7):569–576, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sannigrahi:1976:ASVb

[SM76a] A. B. Sannigrahi and S. Noor Mohammad. On the applicability of the selected valence electron split-shell method to ionic molecules. International Journal of Quantum Chemistry, 10 (4):715, July 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sannigrahi:1976:ASVa

[SM76b] A. B. Sannigrahi and S. Noor Mohammad. On the applicability of the selected valence electron split-shell model to ionic molecules. International Journal of Quantum Chemistry, 10 (1):181–184, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schfeve:1977:MAD

[SM77] W. C. Schfeve and J. W. Middleton. On the Markoffian approximation in the description of relaxation in open systems. International Journal of Quantum Chemistry, 12(S11):625–635, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Srivastava:1979:ESS

[SM79] S. K. Srivastava and P. C. Mishra. Electronic structure, spectra, and mechanism of photodimerization of pyrimidine bases. International Journal of Quantum Chemistry, 16(5):1051–1068, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Savin:1979:ESP

[SMB^{+}79] F. A. Savin, Yu. V. Morozov, A. V. Borodavkin, V. O. Chekhov, E. I. Budowsky, and N. A. Simukova. Electronic structure of the pyrimidine and purine components of nucleic acids in their ground and lower excited singlet and triplet states. International Journal of Quantum Chemistry, 16(4):825–831, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Selsby:1977:SEMa

[SMH77] R. G. Selsby, C. Machin, and M. L. Hernandez. A semiempirical MO theory for ionization potentials and electron affinities. International Journal of Quantum Chemistry, 11(1): 149–161, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smith:1979:SV

[Smi79] V. H. Smith. Sagamore VI. International Journal of Quantum Chemistry, 15(4):445, April 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sheen:1973:PIP

[SMMA73] S. Sheen, R. Marchessault, C. R. Mueller, and T. Anderson. The practical inversion problem in molecular beam scattering. International Journal of Quantum Chemistry, 7(S7):409– 419, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Saran:1977:MOS

[SMP77] Anil Saran, Chanchal K. Mitra, and Bernard Pullman. Molecular orbital studies on nucleoside Antibiotics. I. Conformation of C-C linked antibiotics, formycin and showdomycin. International Journal of Quantum Chemistry, 12(S4):43–54, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Stoddart:1971:LME

[SMW71] J. C. Stoddart, N. H. March, and D. Wiid. Local moments, electron correlation and density functional theory. International Journal of Quantum Chemistry, 5(S5):745–762, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sannigrahi:1979:TIS

[SN79] A. B. Sannigrahi and B. Guha Niyogi. Triplet instability in semiempirical RPA calculations of conjugated systems containing heteroatoms. International Journal of Quantum Chemistry, 16(5):1151–1155, November 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Szabo:1977:IEB

[SO77] Attila Szabo and Neil S. Ostlund. Interaction energies between closed-shell systems: The correlation energy in the random phase approximation. International Journal of Quantum Chemistry, 12(S11):389–395, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Solomon:1971:OLC

[Sol71] C. E. Solomon. Optimized linear combinations of simple exponentials for atomic systems. International Journal of Quantum Chemistry, 5(3):319–333, May 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Soliverez:1973:UIO

[Sol73] Carlos E. Soliverez. The use of irreducible operators for determining the complete set of linearly independent crystal field parameters. International Journal of Quantum Chemistry, 7 (6):1139–1144, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sundaram:1971:TSC

[SP71] K. Sundaram and William P. Purcell. Theoretical study of a 1: 1 complex between quinone and hydroquinone. International Journal of Quantum Chemistry, 5(1):101–110, January 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scheire:1974:ESS

[SP74] L. Scheire and P. Phariseau. On the electronic structure of spheroidal molecules. International Journal of Quantum Chemistry, 8(S8):109–116, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Scheire:1975:CES

[SP75a] L. Scheire and P. Phariseau. Calculation of the energy states in a general molecular potential model a new MSW approach. International Journal of Quantum Chemistry, 9(S9):105–115, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Scheire:1975:ESM

[SP75b] L. Scheire and P. Phariseau. The electronic structure of molecules having arbitrary boundaries. International Journal of Quantum Chemistry, 9(5):887–900, September 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schleicher:1975:DFR

[SP75c] M. Schleicher and H. Primas. A density functional representation of quantum chemistry. III. Rigorous realization of the program in lattice space. International Journal of Quantum Chemistry, 9(5):871–886, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stoll:1975:HFC

[SP75d] H. Stoll and H. Preuss. Hartree–Fock calculation of cohesive energies and equilibrium lattice constants for solid Li and Be. International Journal of Quantum Chemistry, 9(5):775–787, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stepanov:1977:EMI

[SP77] Nikolaj F. Stepanov and Vladimir I. Pupyshev. Errors in molecular integrals: An influence on RHF energy values. International Journal of Quantum Chemistry, 11(1):1–15, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stolarczyk:1979:IPM

[SP79] Leszek Z. Stolarczyk and Lucjan Piela. Invariance properties of the multipole expansion with respect to the choice of the coordinate system. International Journal of Quantum Chemistry, 15(6):701–711, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sparks:1978:ICM

[Spa78] Robert A. Sparks. Ab initio calculations on a minicomputer with 128K bytes of memory. International Journal of Quantum Chemistry, 14(S12):191–199, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Sengupta:1972:GLS

[SPČ72] S. Sengupta, J. Paldus, and J. Čižek. Geminal localization in the Separated-Pair Model II. Excited States of the Φ-electronic model of benzene. International Journal of Quantum Chemistry, 6(S6):153–169, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Saute:1979:CPA

[SPČ79] M. Saute, J. Paldus, and J. Čížek. Correlation problems in atomic and molecular systems. VII. Application of the openshell coupled-cluster approach to simple π-electron model systems. International Journal of Quantum Chemistry, 15(5):463– 479, May 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sperber:1970:ECS

[Spe70] Gunnar Sperber. Electron correlation in S states of the helium atom. International Journal of Quantum Chemistry, 5 (S4):127–130, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Sperber:1971:ARDa

[Spe71a] Gunnar Sperber. Analysis of reduced density matrices in the coordinate representation. I. Definitions and basic formulas. International Journal of Quantum Chemistry, 5(2):177–187, March 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sperber:1971:ARDb

[Spe71b] Gunnar Sperber. Analysis of reduced density matrices in the coordinate representation. II. The structure of closed-shell atoms in the restricted Hartree–Fock approximation. International Journal of Quantum Chemistry, 5(2):189–214, March 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sperber:1972:ARD

[Spe72a] Gunnar Sperber. Analysis of reduced density matrices in the coordinate representation. III. electron density and correlation in the ground states of H_{2} and the H_{6} ring system within some approximations of the simple LCAO type. International Journal of Quantum Chemistry, 6(5):881–898, September 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sperber:1972:CDM

[Spe72b] Gunnar R. Sperber. The computation of density matrix values in configuration space. International Journal of Quantum Chemistry, 6(4):795–797, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sperber:1973:SAMb

[Spe73] Gunnar Sperber. A study of the AMO method as applied to the lithium metal. III. A method for calculating the total electronic energy of crystals in the LCAO approximation. International Journal of Quantum Chemistry, 7(3):537–546, May 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Salmon:1972:QER

[SR72a] Lydia S. Salmon and Klaus Ruedenberg. A quadrupolar expansion for r_{12}^{−1}. International Journal of Quantum Chemistry, 6(2):347–352, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Salmons:1972:EFC

[SR72b] Lydia S. Salmons and Klaus Ruedenberg. An expansion for four-center integrals over Slater-type orbitals. International Journal of Quantum Chemistry, 6(2):353–366, March 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Steinborn:1972:MOE

[SR72c] E. Otto Steinborn and Klaus Ruedenberg. Molecular oneelectron integrals over Slater-type atomic orbitals and irregular solid spherical harmonics. International Journal of Quantum Chemistry, 6(3):413–438, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

So:1977:TSV

[SR77] S. P. So and W. G. Richards. A theoretical study of the vibrational structure in the transition of PH_{2}. International Journal of Quantum Chemistry, 11(1):73–79, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slater:1978:CPH

[SR78] T. F. Slater and P. A. Riley. Carcinogenicity of polycyclic hydrocarbons and their interaction with DNA. International Journal of Quantum Chemistry, 14(S5):143–148, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Srebrenik:1975:RPS

[Sre75] S. Srebrenik. Rayleigh principle for a subspace of a quantum system. International Journal of Quantum Chemistry, 9(S9): 375–383, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Sheina:1979:CTI

[SRE^{+}79] G. G. Sheina, E. D. Radchenko, S. A. Egupov, Yu. P. Blagoy, and V. M. Orlov. Charge transfer interactions between nucleic acid bases and strong acceptors. International Journal of Quantum Chemistry, 16(2):387–394, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Srinivasan:1977:CEM

[Sri77] K. A. R. Sundaram Srinivasan. Conformational energy map for a pair of peptide units using the bond polarisability method. International Journal of Quantum Chemistry, 12(4):671–681, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shavitt:1976:CCI

[SRP76] Isaiah Shavitt, Bruce J. Rosenberg, and Surat Palalikit. Comparison of configuration interaction expansions based on different orbital transformations. International Journal of Quantum Chemistry, 10(S10):33–46, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See erratum [SRP77]. Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Shavitt:1977:ECC

[SRP77] Isaiah Shavitt, Bruce J. Rosenberg, and Surat Palalikit. Erratum: Comparison of configuration interaction expansions based on different orbital transformations. International Journal of Quantum Chemistry, 12(S11):651, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [SRP76]. Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Singh:1970:CTP

[SS70] S. R. Singh and V. H. Smith, Jr. Comparison of two perturbation methods. International Journal of Quantum Chemistry, 4 (5):519–527, September 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Singh:1971:PTH

[SS71a] S. R. Singh and V. H. Smith, Jr. Perturbation treatment of the Hartree–Fock–Slater (Xα) equations for the three-electron ions. International Journal of Quantum Chemistry, 5(S5):421– 428, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Stettler:1971:BSD

[SS71b] John D. Stettler and Romas A. Shatas. Bond stretch in diatomic vibrotors with Fues potential. International Journal of Quantum Chemistry, 5(S5):277–284, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Shibuya:1973:SFWa

[SS73a] Tai-Ichi Shibuya and Oktay Sinanoğlu. Spin-free wave functions in many-electron perturbation theory. I. Closed-shell systems. International Journal of Quantum Chemistry, 7(6):1145– 1158, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shibuya:1973:SFWb

[SS73b] Tai-Ichi Shibuya and Oktay Sinanoğlu. Spin-free wave functions in many-electron perturbation theory. II. Systems with one nonclosed shell. International Journal of Quantum Chemistry, 7(6):1159–1174, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stigter:1973:ERM

[SS73c] Dirk Stigter and John A. Schellman. Electrostatic representations of molecular groups. International Journal of Quantum Chemistry, 7(1):1–13, January 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shinagawa:1974:HMS

[SS74] Yoshiya Shinagawa and Yasuko Shinagawa. Hückel MO studies on diuretics and carbonic anhydrase inhibitors. International Journal of Quantum Chemistry, 8(S1):169–178, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Steslicka:1977:DLS

[SS77a] Maria Steślicka and E. R. Smith. Density of localized states — effect of the next-nearest neighbour interaction. International Journal of Quantum Chemistry, 12(S11):573–581, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Steslicka:1977:LSS

[SS77b] Maria Steślicka and B. Stankiewicz. Localized states in a simple cubic crystal via the next-nearest neighbor approximation. International Journal of Quantum Chemistry, 12(3):433–441, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shinagawa:1978:MOS

[SS78a] Yoshiya Shinagawa and Yasuko Shinagawa. Molecular orbital studies on m- and p-methylation of substituted catechols by catechol O-methyltransferase. International Journal of Quantum Chemistry, 14(S5):269–279, March 3–11, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Smith:1978:MCB

[SS78b] Iden A. Smith and Paul G. Seybold. Methylbenz[a]anthracenes: Correlations between theoretical reactivity indices and carcinogenicity. International Journal of Quantum Chemistry, 14(S5): 311–320, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sawaryn:1979:TSS

[SS79] Andrzej Sawaryn and W. Andrzej Sokalski. Theoretical studies on substrate binding to the active site of carbonic anhydrase. International Journal of Quantum Chemistry, 16(2):293–298, August 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Svendsen:1978:CCP

[SSH78] E. Nørby Svendsen and T. Stroyer-Hansen. CNDO calculation of polarizabilities including polarization functions in the basis set. International Journal of Quantum Chemistry, 13(2):235– 238, February 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sheen:1975:EOR

[SSM75] S. H. Sheen, J. G. Skofronick, and C. R. Mueller. Effect of orbiting resonance on the elastic differential cross sections. International Journal of Quantum Chemistry, 9(5):817–824, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Silverstone:1970:AET

[ST70] Harris J. Silverstone and H. David Todd. Analytical evaluation of three-center one-electron integrals of r^{N}Y_{L}^{M}(θ, φ) with Slater-type atomic orbitals. International Journal of Quantum Chemistry, 5(S4):371–383, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Shillady:1972:STH

[ST72] Donald D. Shillady and Carl Trindle. The shapes of three hypervalent systems of first-row atoms FH_{3}, H_{4}O, and F_{3}H. International Journal of Quantum Chemistry, 6(S6):187–200, January 17–22, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Shillady:1973:ECE

[ST73] Donald D. Shillady and Carl Trindle. Effects of chemical environment on the carbonyl group. International Journal of Quantum Chemistry, 7(S7):269–278, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sundaram:1978:NIM

[ST78] K. Sundaram and R. S. Tyagi. Nonbonded interactions in membrane active cyclic biopolymers. II. Ion binding by valinomycin. International Journal of Quantum Chemistry, 13(1):17– 26, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sakamoto:1979:EUD

[ST79a] Kazuyoshi Sakamoto and Toshitaka Terasaka. Effective utilization of off-diagonal hypervirial relations considering diagonal hypervirial relation: Harmonic oscillator case. International Journal of Quantum Chemistry, 16(6):1357–1368, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sundaram:1979:NIM

[ST79b] K. Sundaram and R. S. Tyagi. Nonbonded interactions in membrane-active cyclic biopolymers. III. Analogs of valinomycin. International Journal of Quantum Chemistry, 15(5): 491–497, May 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stenger:1974:IPE

[Ste74] William Stenger. Intermediate problems for eigenvalues. International Journal of Quantum Chemistry, 8(4):623–625, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stepanov:1977:KSG

[Ste77] N. F. Stepanov. On the Kashiwagi–Sasaki generalization of the Löwdin orthogonalization and the inverse vibrational problem. International Journal of Quantum Chemistry, 12(2):317–326, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stillinger:1978:PMR

[Sti78] Frank H. Stillinger. Polarization model representation of hydrogen fluoride for use in gas- and condensed-phase studies. International Journal of Quantum Chemistry, 14(5):649–657, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stilbs:1979:BRB

[Sti79] Peter Stilbs. Book review: Nuclear magnetic resonance spectroscopy in molecular biology. Proceedings of the Eleventh Jerusalem Symposium on Quantum Chemistry and Biochemistry held in Jerusalem. Israel, April 3–7. 1978. B. Pullman, Ed. D. Riedel Publishing Company, Dordrecht, Holland. 1978. 532 pp. Price: $56.00. International Journal of Quantum Chemistry, 15(6):793, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Pul78b].

Stocks:1971:ESS

[Sto71] G. M. Stocks. Electronic states in some disordered noble metal–transition metal alloys. International Journal of Quantum Chemistry, 5(S5):533–541, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Sundaralingam:1974:ECP

[Sun74a] M. Sundaralingam. Evolution of conformational principles in nucleic acids. International Journal of Quantum Chemistry, 8(S1):81–91, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sundaram:1974:NIM

[Sun74b] K. Sundaram. Nonbonded interactions in membrane active cyclic biopolymers. I. Valinomycin–potassium ion complex. International Journal of Quantum Chemistry, 8(4):565–583, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sundaram:1975:AOT

[Sun75a] K. Sundaram. Applications of an optimization technique in submolecular biology. International Journal of Quantum Chemistry, 9(3):393–402, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sundaram:1975:NIM

[Sun75b] K. Sundaram. Nonbonded interactions in membrane active cyclic biopolymers. I. valinomycin-potassium ion complex. International Journal of Quantum Chemistry, 9(4):759, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sundaralingam:1977:NWC

[Sun77] M. Sundaralingam. Non-Watson–Crick base pairs in ribonucleic acids. International Journal of Quantum Chemistry, 12 (S4):11–23, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sureau:1971:OIP

[Sur71] A. Sureau. Origin of the irrational part contained in the angular numerical factors of matrix elements of the Coulomb operator. International Journal of Quantum Chemistry, 5(6): 599–603, November 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Suzuki:1977:GJP

[Suz77] Noboru Suzuki. Generalized Jansen’s perturbation theory for exchange interactions between aroms or molecules. International Journal of Quantum Chemistry, 12(1):19–28, July 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Svendsen:1979:CRI

[Sve79] E. Nørby Svendsen. Calculation of Raman intensities. International Journal of Quantum Chemistry, 16(S13):553–557, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Slater:1970:SET

[SW70] John C. Slater and John H. Wood. Statistical exchange and the total energy of a crystal. International Journal of Quantum Chemistry, 5(S4):3–34, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Slater:1971:SET

[SW71] J. C. Slater and J. H. Wood. Statistical exchange and the total energy of a crystal. International Journal of Quantum Chemistry, 4(??):3–??, ?? 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Steinborn:1977:ARB

[SW77] E. O. Steinborn and E. J. Weniger. Advantages of reduced Bessel functions as atomic orbitals. An application to H_{2}^{+}. International Journal of Quantum Chemistry, 12(S11):509–516, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Steinborn:1978:RBF

[SW78] E. O. Steinborn and E. J. Weniger. Reduced Bessel functions as atomic orbitals: Some mathematical aspects and an LCAO-MO treatment of HeH^{++}. International Journal of Quantum Chemistry, 14(S12):103–108, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Sundaralingam:1979:RNC

[SW79] M. Sundaralingam and E. Westhof. The “rigid” nucleotide concept in perspective. International Journal of Quantum Chemistry, 16(S6):115–130, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Syren:1979:IGT

[SWF79] Robert M. Syren, Charles R. Windsor, and Sidney W. Fox. Inhibition of glyoxalase I by thermal copolyamino acids. International Journal of Quantum Chemistry, 16(S6):283–288, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Switendick:1971:MHS

[Swi71] A. C. Switendick. Metal hydrides — structure and band structure. International Journal of Quantum Chemistry, 5(S5):459– 470, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Silver:1978:UBS

[SWN78] David M. Silver, Stephen Wilson, and W. C. Nieuwpoort. Universal basis sets and transferability of integrals. International Journal of Quantum Chemistry, 14(5):635–639, November 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stuebing:1977:DAM

[SWP77] Edward W. Stuebing, John H. Weare, and Robert G. Parr. Discontinuous approximate molecular electronic wave-functions. International Journal of Quantum Chemistry, 11(1):81–102, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sundaralingam:1977:PPL

[SY77] M. Sundaralingam and N. Yathindra. Probing possible leftand right-handed polynucleotide helical conformations from nh plots. Glycosyl and backbone torsional variation on handedness of helix. International Journal of Quantum Chemistry, 12 (S4):285–303, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Stwalley:1976:RPL

[SZ76] William C. Stwalley and Warren T. Zemke. The radiative properties of long-range molecules. International Journal of Quantum Chemistry, 10(S10):223–225, January 19–24, 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Talekar:1977:TDA

[Tal77] Subhash V. Talekar. Temperature dependence of activation energies for self-diffusion of water and of alkali ions in aqueous electrolyte solutions. A model for ion selective behavior of biological cells. International Journal of Quantum Chemistry, 12(S4):459–469, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Tanaka:1974:BDW

[Tan74] Kiyoshi Tanaka. Basis dependence of wave functions of the ground state and low-lying excited states of planar formaldehyde. International Journal of Quantum Chemistry, 8(6):981– 993, November 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tatewaki:1973:EEI

[Tat73] Hiroshi Tatewaki. Effect of the exchange integral K_{2s2p} on the 2p-orbitals of the^{1}P and^{3}P states of the beryllium isoelectronic sequence arising from the configuration (1s)^{2}(2s)(2p). International Journal of Quantum Chemistry, 7(5):971–984, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tatum:1976:ACE

[Tat76] J. Patrick Tatum. Angular correlation of electrons in the ground state of helium. International Journal of Quantum Chemistry, 10(6):967–974, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trindle:1973:NSS

[TB73] Carl Trindle and Thomas D. Bouman. Nuclear spin statistics in fluxional molecules. International Journal of Quantum Chemistry, 7(S7):329–335, January 21–27, 1973. CODEN IJQCB2.

ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Thomson:1974:ISE

[TB74] Colin Thomson and Douglas A. Brotchie. An ab initio study of the electronic structure and isotropic hyperfine coupling constants of HCO, FCO, and HBF using different Gaussian basis sets. International Journal of Quantum Chemistry, 8(S8):277– 284, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Trindle:1975:GTR

[TB75] Carl Trindle and Thomas D. Bouman. Group theory and reaction mechanisms: An extension of the McIver–Stanton rules. International Journal of Quantum Chemistry, 9(S9):255–264, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Tal:1978:SED

[TB78] Y. Tal and R. F. W. Bader. Studies of the energy density functional approach. I. Kinetic energy. International Journal of Quantum Chemistry, 14(S12):153–168, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Tapia:1970:IMC

[TBB70] O. Tapia, G. Bessis, and Et. S. Bratož. Interactions moleculaires à courte portée dans le système H_{2}⋯H_{2}: Leur dépendance de l’orientation et de la géométrie interne des deux partenaires. (French) [Short-distance molecular interactions in the H_{2}⋯H_{2} system: their orientatiion dependence and the internal geometry of the two partners]. International Journal of Quantum Chemistry, 5(S4):289–297, January 19–24, 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Trindle:1970:EBF

[TC70] Carl Trindle and Francis S. Collins. Energy-based formalism for the mapping analysis of concerted reactions. International Journal of Quantum Chemistry, 5(S4):195–204, January 19– 24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Trindle:1977:NME

[TDB77] C. Trindle, S. N. Datta, and T. D. Bouman. Nonrigid molecule effects on the energy levels of XeF_{6}. International Journal of Quantum Chemistry, 11(4):627–664, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trindle:1976:IHF

[TG76] Carl Trindle and John K. George. Integral Hellmann–Feynman computations on H_{3} ABH_{n} − H_{2}ABH_{n+1} rearrangements. International Journal of Quantum Chemistry, 10(1):21–35, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Talman:1979:CSA

[TGG79] J. D. Talman, P. S. Ganas, and A. E. S. Green. Comparative studies of atomic independent-particle potentials. International Journal of Quantum Chemistry, 16(S13):67–76, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Tallet:1970:CVC

[TGSL70] A. Tallet, M. Garcia-Sucre, and Et. R. Lefebvre. Le couplage vibronique dans une chaine linéaire de molecules etudié par une méthode auto-cohérente. (French) [Vibronic coupling in a linear chain of molecules studied by a self-consistent field method]. International Journal of Quantum Chemistry, 4(4):397–411, July 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trindle:1978:SFR

[TH78] Carl Trindle and E. A. Halevi. Spin-forbidden reaction pathways in the interaction of singlet and triplet molecular oxygen with acetylene. International Journal of Quantum Chemistry, 14(S5):281–290, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Thomas:1975:NLS

[Tho75a] L. H. Thomas. A new look at some old ideas about quantum theory. International Journal of Quantum Chemistry, 9(S9): 5–6, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Thomas:1975:R

[Tho75b] L. H. Thomas. Reminiscences. International Journal of Quantum Chemistry, 9(S9):3–4, January 19–25, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Thomson:1976:ESU

[Tho76] Colin Thomson. The electronic structure of unstable intermediates. VI. the electronic structure of AlCN, AlNC, PCN, and PNC and comparison with their 1st-row analogues. International Journal of Quantum Chemistry, 10(S10):85–90, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Thulstrup:1975:SIF

[Thu75] Poul W. Thulstrup. Smooth interpolation, Fourier transformation, and two-center overlap integrals for numerical atomic orbitals. International Journal of Quantum Chemistry, 9(5): 789–795, September 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tichy:1979:PCB

[Tic79] Miloň Tichy. Possible connection between lipophilicity and steric substituent constants. International Journal of Quantum Chemistry, 16(3):509–515, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tinoco:1979:CDL

[Tin79] Ignacio Tinoco, Jr. Circular dichroism of large molecules. International Journal of Quantum Chemistry, 16(1):111–117, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tal:1978:NEM

[TK78] Yoram Tal and Jacob Katriel. On the non-existence of maxima in variational computations containing non-linear parameters. International Journal of Quantum Chemistry, 13(1):149–153, January 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trindle:1971:FSG

[TL71] Carl Trindle and Peter W. Lert. Floating spherical Gaussian calculations on systems of protons and electrons. International Journal of Quantum Chemistry, 5(S5):329–340, January 18– 23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Taylor:1972:VHJ

[TL72] W. J. Taylor and R. D. Levine. On the validity of Hamilton– Jacobi mechanics in the mean in quantum mechanics. International Journal of Quantum Chemistry, 6(3):449–453, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thulstrup:1979:BCI

[TL79] Poul W. Thulstrup and Jan Linderberg. Bounds to Coulomb interaction integrals. International Journal of Quantum Chemistry, 16(S13):39–50, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Terasaka:1973:ESP

[TN73] Toshitaka Terasaka and Youko Nedachi. Equation for the screened potential. International Journal of Quantum Chemistry, 7(2):173–186, March 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Theodorakopoulos:1979:OEB

[TNB79] G. Theodorakopoulos, C. A. Nicolaides, and D. R. Beck. Oneelectron binding and Auger energies of sulfur in atomic and molecular states. International Journal of Quantum Chemistry, 16(S13):671–681, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Truhlar:1979:EPA

[TOED79] Donald G. Truhlar, Kunizo Onda, Robert A. Eades, and David A. Dixon. Effective potential approach to electron– molecule scattering theory. International Journal of Quantum Chemistry, 16(S13):601–632, March 11–17, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Tapia:1977:EEH

[TP77a] O. Tapia and E. Poulain. Environmental effects on hydrogen bonded systems: A quantum chemical study of proton relay model systems. International Journal of Quantum Chemistry, 11(3):473–484, March 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Toivonen:1977:RMO

[TP77b] H. T. Toivonen and P. Pyykkö. Relativistic molecular orbitals for the double group D_{3h}. International Journal of Quantum Chemistry, 11(4):695–700, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Topiol:1978:ECP

[TP78] Sid Topiol and John A. Pople. Effective core potential calculations for some hydrocarbons. International Journal of Quantum Chemistry, 14(S12):323–328, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Thomson:1977:QMI

[TPC77] Colin Thomson, David Provan, and Susan Clark. Quantummechanical investigations of unstable intermediates relevant to the mechanism of chemical carcinogenesis by Nalkylnitrosamines. International Journal of Quantum Chemistry, 12(S4):205–215, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Thompson:1970:MSS

[TPM70] J. Thompson, M. Povich, and B. Musulin. Molecular screening and spectroscopic constants. International Journal of Quantum Chemistry, 3S(??):513–??, ?? 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trinajstic:1977:NDH

[Tri77] Nenad Trinajstić. New developments in Hückel theory. International Journal of Quantum Chemistry, 12(S11):469–477, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Thunemann:1977:SCI

[TRPB77] K. H. Thunemann, J. Römelt, S. D. Peyerimhoff, and R. J. Buenker. A study of the convergence in iterative natural orbital procedures. International Journal of Quantum Chemistry, 11 (5):743–752, May 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tantardini:1973:IVB

[TRS73] Gian Franco Tantardini, Mario Raimondi, and Massimo Simonetta. Ab initio valence bond calculations. I. Methylene. International Journal of Quantum Chemistry, 7(5):893–903, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Truhlar:1972:VME

[Tru72] Donald G. Truhlar. Vibrational matrix elements of the quadrupole moment functions of H_{2}, N_{2} and CO. International Journal of Quantum Chemistry, 6(5):975–988, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Truhlar:1973:IHF

[Tru73a] Donald G. Truhlar. Ab initio Hartree–Fock calculations of electronic wave functions for the c^{3}Π_{u} state of H_{2}. International Journal of Quantum Chemistry, 7(6):1175–1182, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Truhlar:1973:ACI

[Tru73b] Donald G. Truhlar. Application of the configuration-interaction method and the random phase approximation to the Ab Initio calculation of electronic excitation energies of H_{2}O. International Journal of Quantum Chemistry, 7(4):807–817, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Truhlar:1976:QPF

[Tru76] Donald G. Truhlar. Quasiclassical predictions of final vibrational state distributions in reactive and nonreactive collisions. International Journal of Quantum Chemistry, 10(S10):239– 250, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Tanaka:1971:CIS

[TS71] Kiyoshi Tanaka and Fukashi Sasaki. Configuration interaction study of X-ray and fast electron scattering factors for light atomic systems. International Journal of Quantum Chemistry, 5(2):157–175, March 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tung:1972:SPP

[TS72] Eleanor W. Tung and William A. Sanders. Simple perturbation and perturbation-variation treatments of the 1sσ_{g} and 2pσ_{u} states of H. International Journal of Quantum Chemistry, 6 (4):717–723, July 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tantardini:1977:IVB

[TS77] G. F. Tantardini and M. Simonetta. Ab initio valence bond calculations. VII. HF, HF^{+}, and H_{2}F^{+}. International Journal of Quantum Chemistry, 12(3):515–525, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tantardini:1978:IVB

[TS78] Gian Franco Tantardini and Massimo Simonetta. Ab initio valence bond calculations. IX. Ionization potentials of ethylene, allyl radical, trans-, and cis-butadiene. International Journal of Quantum Chemistry, 14(S12):131–141, March 12–18, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Tantardini:1979:IVB

[TS79] Gian Franco Tantardini and Massimo Simonetta. Ab initio valence bond calculations. X. Vertical valence ionization potentials of allene and butatriene. International Journal of Quantum Chemistry, 15(6):655–662, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trsic:1976:STE

[TSE76] M. Trsic, J. E. Sanhueza, and Ll. Espinoza. Singlet and triplet excited state electric dipole polarizabilities of conjugated hydrocarbons. International Journal of Quantum Chemistry, 10 (3):429–437, May 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Toyota:1976:IHF

[TTN76] A. Toyota, T. Tanaka, and T. Nakajima. The instabilities of the Hartree–Fock solutions and the lattice instabilities for conjugated hydrocarbons: The bond-order and bond-length alternations. International Journal of Quantum Chemistry, 10(5): 917–935, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tatewaki:1971:EDB

[TTS71] Hiroshi Tatewaki, Hirosiji Taketa, and Fukashi Sasaki. On the energy difference between the ^{1}P and ^{3}P states of the beryllium isoelectronic sequence. International Journal of Quantum Chemistry, 5(3):335–357, May 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tatewaki:1979:CIS

[TTS^{+}79] Hiroshi Tatewaki, Kiyoshi Tanaka, Fukashi Sasaki, Shigeru Obara, Kimio Ohno, and Megumu Yoshimine. Configurationinteraction study of lower excited states of O_{2}: Valence and Rydberg characters of the two lowest ^{3}Σ_{u}^{−} states. International Journal of Quantum Chemistry, 15(5):533–545, May 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Turner:1974:MMO

[Tur74] Almon G. Turner. Methods in molecular orbital theory. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1974. ISBN 0-13-579383-1. xiii + 225 pp. LCCN QD461 .T85.

Trickey:1977:EPR

[TW77] S. B. Trickey and J. P. Worth. Electrons and phonons in rare gas crystals: Numerical studies of simple local density models. International Journal of Quantum Chemistry, 12(S11):529– 538, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solidstate Theory, Collision Phenomena, and Computational Methods.

Tel:1973:CQC

[TWC73] Luis M. Tel, Saul Wolfe, and Imre G. Csizmadia. A comparative quantum chemical study of ethylcarbonium ion and hydroxymethylcarbonium ion. International Journal of Quantum Chemistry, 7(3):475–490, May 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tyutyulkov:1975:GFE

[Tyu75] N. N. Tyutyulkov. A generalized formula for the energies of alternant molecular orbitals. I. Homonuclear molecules. International Journal of Quantum Chemistry, 9(4):683–689, July 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tyutyulkov:1978:GFE

[Tyu78] N. N. Tyutyulkov. A generalized formula for the energies of alternant molecular orbitals. II. Heteronuclear molecules. International Journal of Quantum Chemistry, 13(3):443–454, March 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ukrainsky:1972:ESL

[Ukr72] I. I. Ukrainsky. Electronic structure of long cumulene chains. International Journal of Quantum Chemistry, 6(3):473–489, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Urry:1976:NPC

[ULM76] D. W. Urry, M. M. Long, and L. W. Mitchell. New polypeptide conformations and their roles in molecular diseases of the vascular wall. International Journal of Quantum Chemistry, 10 (S3):107–118, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Umans:1971:MBP

[UP71] Robert S. Umans and Bernard Pullman. The mechanism of binding of polycyclic aromatic hydrocarbons to nucleic acids: A theoretical investigation. International Journal of Quantum Chemistry, 5(5):575–582, September 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Urry:1975:MMI

[Urr75] Dan W. Urry. Molecular mechanisms of ion permeation of membranes. International Journal of Quantum Chemistry, 9 (S2):221–235, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Urry:1977:PMM

[Urr77] D. W. Urry. Polypeptide models for membrane permeation by cations and their relevance to biomembrane depolarization. International Journal of Quantum Chemistry, 12(S4):25–41, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Urry:1978:NHS

[Urr78] D. W. Urry. Nonclassical helical states and diverse biological functions of sequential polypeptides. International Journal of Quantum Chemistry, 14(S5):51–67, March 3–11, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Urry:1979:TCT

[USK^{+}79] D. W. Urry, A. Spisni, M. A. Khaled, M. M. Long, and L. Masotti. Transmembrane channels and their characterization in phospholipid structures. International Journal of Quantum Chemistry, 16(S6):289–303, March 4–10, 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

VanLeuven:1970:GOA

[Van70] P. Van Leuven. Gaussian overlap approximation in the projected Hartree–Fock method. International Journal of Quantum Chemistry, 4(4):355–363, July 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

VanVleck:1971:RFD

[Van71] John H. Van Vleck. Reminiscences of the first decade of quantum mechanics. International Journal of Quantum Chemistry, 5(S5):3–20, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

VanDerLugt:1972:MOC

[Van72] W. Th. A. M. Van Der Lugt. Molecular-orbital calculations on transition metal complexes, charge-transfer spectra and the sequence of metal and ligand orbitals. International Journal of Quantum Chemistry, 6(5):859–880, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

VanDuijnen:1974:MIG

[Van74] P. Th. Van Duijnen. Molecular integrals over generalized Gaussian basis sets. International Journal of Quantum Chemistry, 8(2):179–191, March 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vanderkooi:1975:TMS

[Van75] Garret Vanderkooi. Thermodynamics of membrane structure: Implications for pharmacology. International Journal of Quantum Chemistry, 9(S2):209–219, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Vetchinkin:1972:GFS

[VB72] S. I. Vetchinkin and V. L. Bachrach. The Green’s functions of the Schrödinger equation for the simplest systems. International Journal of Quantum Chemistry, 6(1):143–153, January 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vasilescu:1978:KHN

[VBC78] D. Vasilescu, H. Broch, and D. Cabrol. Kinked helices in nucleic acids — a molecular orbital investigation. International Journal of Quantum Chemistry, 14(S5):345–354, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

vanDijk:1979:CCA

[vDR79] W. van Dijk and M. Razavy. Collinear collision of an atom with a homonuclear diatomic molecule. International Journal of Quantum Chemistry, 16(6):1249–1263, December 1979. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vojtik:1976:TRS

[VF76] J. Vojtík and J. Fišer. Time reversal and spin-adapted wavefunctions. International Journal of Quantum Chemistry, 10(6): 1083–1085, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vasudevan:1978:TSV

[VG78] K. Vasudevan and F. Grein. Theoretical study on the vertical electronic spectrum of carbonyl fluoride, F_{2}CO. International Journal of Quantum Chemistry, 14(6):717–726, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Volkenstein:1979:ECI

[VGS79a] M. V. Volkenstein, I. B. Golovanov, and V. M. Sobolev. Electron-conformational interactions and functioning of enzyme molecules. International Journal of Quantum Chemistry, 16(4):777–789, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Volkenstein:1979:IIP

[VGS79b] M. V. Volkenstein, I. B. Golovanov, and V. M. Sobolev. Interfragment interactions in peptide molecules. International Journal of Quantum Chemistry, 16(4):791–797, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vilallonga:1975:EIA

[VH75] Federico A. Vilallonga and Johanna S. Hunt. The effect of the injection of alkanols and alkanediols beneath dipalmitoyl phosphatidylethanol amine monolayers. International Journal of Quantum Chemistry, 9(S2):325–329, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

VonBoehm:1974:SCS

[VK74] Juhani Von Boehm and Peter Krusius. The self-consistent symmetrized OPW method with an application to crystalline selenium. International Journal of Quantum Chemistry, 8(3):395– 422, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vladimiroff:1970:SSA

[Vla70] T. Vladimiroff. Small, simultaneous adjustments of orbital exponents in LCAO–MO–SCF calculations using self-consistent perturbation theory. International Journal of Quantum Chemistry, 4(1):89–95, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vogler:1979:IUH

[Vog79] H. Vogler. Improved uncoupled Hartree–Fock (IUCHF) perturbation methods and bounds for the second-order energy in coupled Hartree–Fock perturbation theory. International Journal of Quantum Chemistry, 15(6):745–750, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Volkov:1979:SAD

[Vol79] S. N. Volkov. Some aspects of the DNA hypochromic effect theory. International Journal of Quantum Chemistry, 16(1): 119–132, July 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Venkatachalam:1979:CEH

[VU79] C. Mahalingam Venkatachalam and Dan W. Urry. Conformational energetics of β-helical transmembrane channels and voltage-dependent conductance of poly-(Ala-Ala-GIy). International Journal of Quantum Chemistry, 16(S6):343–351, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Weiner:1971:PMN

[WA71] J. H. Weiner and A. Askar. Particle method for the numerical solution of the time-dependent Schrödinger equation. International Journal of Quantum Chemistry, 5(S5):259, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Woodward:1971:PGE

[WAL71] M. R. Woodward, A. T. Amos, and C. Laughlin. A perturbed generalized eigenvalue equation for the helium atom. II. International Journal of Quantum Chemistry, 5(4):387–394, July 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Walker:1977:QMT

[Wal77] Evan Harris Walker. Quantum mechanical tunneling in synaptic and ephaptic transmission. International Journal of Quantum Chemistry, 11(1):103–127, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wannier:1979:ESB

[Wan79] Gregory H. Wannier. Energy spectrum for Bloch electrons in a magnetic field. International Journal of Quantum Chemistry, 16(S13):413–421, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Wu:1973:TDH

[WB73] Fun-Min Wu and Charles L. Beckel. Theoretical determination of H_{2} ground-state spectroscopic properties. International Journal of Quantum Chemistry, 7(S7):135–141, January 21– 27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Whitton:1976:RBR

[WB76] W. N. Whitton and W. Byers Brown. The relationship between the Rayleigh–Schrödinger and asymptotic perturbation theories of intermolecular forces. International Journal of Quantum Chemistry, 10(1):71–86, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wahl:1970:BNI

[WBKL70] A. C. Wahl, P. Bertoncini, K. Kaiser, and R. Land. BISON: a new instrument for the experimentalist. International Journal of Quantum Chemistry, 3S(??):499–??, ?? 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Waber:1975:RET

[WC75a] J. T. Waber and J. M. Canfield. Relativistic effects in the Thomas–Fermi atom. International Journal of Quantum Chemistry, 9(S9):51–58, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Weber:1975:MXC

[WC75b] J. Weber and J. W. D. Connolly. MS–Xα calculation of the barrier to internal rotation in ethylene. International Journal of Quantum Chemistry, 9(S9):129–136, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Weinstein:1976:RCL

[WCK^{+}76] Harel Weinstein, David Chou, Sungzong Kang, Carl Lynn Johnson, and Jack Peter Green. Reactivity characteristics of large molecules and their biological activity: Indolealkylamines on the LSD/serotonin receptor. International Journal of Quantum Chemistry, 10(S3):135–150, January 12–14, 1976. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Webb:1974:NVR

[Web74] Sydney J. Webb. Nutrition and in vivo rotational motion: A microwave study. International Journal of Quantum Chemistry, 8(S1):245–251, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Walter:1974:LMC

[WEH74] C. F. Walter, H. M. Eberspaecher, and J. P. Hughes. An online method for the collection of continuous data for enzymecatalyzed reactions. International Journal of Quantum Chemistry, 8(S1):253–258, January 17–19, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Weinhold:1971:RCU

[Wei71] Frank Weinhold. Remarks on the calculation of upper and lower limits to quantum-mechanical properties. International Journal of Quantum Chemistry, 5(S5):721–728, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Weinstein:1974:PUA

[Wei74] Harel Weinstein. On the possible use of analytic potentials in the scattered-wave method. International Journal of Quantum Chemistry, 8(S8):123–125, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Weinstein:1975:SNQ

[Wei75] Harel Weinstein. Some new quantum chemical procedures for the analysis of drug–receptor interactions. International Journal of Quantum Chemistry, 9(S2):59–69, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Weislinger:1976:NTD

[Wei76] Edmond Weislinger. Nonadditive triple-dipole interaction in D_{3h} configuration by variational method: Results of a hardsphere model and a confined model. International Journal of Quantum Chemistry, 10(S10):203–212, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Weintraub:1977:CEC

[Wei77] H. J. R. Weintraub. Conformational energy calculations on anticonvulsant drugs. International Journal of Quantum Chemistry, 12(S4):111–125, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Weltin:1970:BES

[Wel70] E. Weltin. Brackets to the eigenvalues of the Schrödinger equation, part 1. Tridiagonal matrices. International Journal of Quantum Chemistry, 4(3):257–270, May 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weltin:1971:BES

[Wel71] E. Weltin. Brackets to the eigenvalues of the Schrödinger equation, Part 2. Partial tridiagonalization of bandmatrices. International Journal of Quantum Chemistry, 5(S5):715–719, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Weltin:1975:DON

[Wel75] E. E. Weltin. Direct optimization of nonlinear parameters. International Journal of Quantum Chemistry, 9(S9):337–341, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Weltin:1976:DME

[Wel76] E. E. Weltin. Direct minimization of the energy by simultaneous variation of parameters in nonorthogonal basis functions. I. Method. International Journal of Quantum Chemistry, 10 (1):163–174, January 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weller:1977:PCH

[Wel77a] Th. Weller. PCILO calculations on hydrogen bonded complexes. Dimers involving second row atoms. International Journal of Quantum Chemistry, 12(5):805–811, November 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weltin:1977:DME

[Wel77b] E. E. Weltin. Direct minimization of the energy by simultaneous variation of parameters in nonorthogonal basis functions. II. Real STOS for first row atoms. International Journal of Quantum Chemistry, 11(4):605–611, April 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wendin:1979:CER

[Wen79] Göran Wendin. Collective effects, relaxation, and localization of hole levels in atoms, molecules, solids, and adsorbates. International Journal of Quantum Chemistry, 16(S13):659–670, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Westhaus:1973:CEC

[Wes73] Paul Westhaus. Cluster expansions of canonically transformed Hamiltonians and effective interactions. International Journal of Quantum Chemistry, 7(S7):463–477, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Westhaus:1979:EHQ

[Wes79] Paul A. Westhaus. Effective Hamiltonians and the quasidegeneracy problem: Calculations on ethylene. International Journal of Quantum Chemistry, 16(S13):215–223, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Wepfer:1973:EMD

[WESW73] G. G. Wepfer, R. N. Euwema, G. T. Surratt, and D. L. Wilhite. Electron momentum distribution in diamond. International Journal of Quantum Chemistry, 7(S7):613–618, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Wolf:1977:VME

[WFS77] G. V. Wolf, D. S. Farberov, and V. P. Shirokovskii. Vanadium monoxide energy-spectrum calculation using the augmentedplane-wave and Green’s function methods. International Journal of Quantum Chemistry, 12(3):485–494, September 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Witschel:1974:OHM

[WG74] W. Witschel and B. Grosswendt. Oscillator and hydrogenic matrix elements by operator algebra. International Journal of Quantum Chemistry, 8(4):585–604, May 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weintraub:1975:CSS

[WH75] H. J. R. Weintraub and A. J. Hopfinger. The CAMSEQ software system in drug design calculations. International Journal of Quantum Chemistry, 9(S2):203–208, January 16–18, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wijkman:1978:IR

[Wij78] Gunnar Wijkman. Introductory remarks. International Journal of Quantum Chemistry, 14(S5):9–10, March 3–11, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Williams:1974:MST

[Wil74] Arthur R. Williams. Multiple scattering theory beyond the muffin-tin approximation. International Journal of Quantum Chemistry, 8(S8):89–108, January 20–26, 1974. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Wilson:1977:IHL

[Wil77a] E. Bright Wilson. Impact of the Heitler-London hydrogen molecule paper on chemistry. International Journal of Quantum Chemistry, 12(S11):17–28, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Wilson:1977:DPT

[Wil77b] Stephen Wilson. Diagrammatic perturbation theory: Potential curves for the ground state of the carbon monoxide molecule.

International Journal of Quantum Chemistry, 12(4):609–622, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wilson:1979:EQM

[Wil79a] E. Bright Wilson. Einstein and quantum mechanics. International Journal of Quantum Chemistry, 16(S13):1–4, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Wilson:1979:DMS

[Wil79b] E. Bright Wilson. On the definition of molecular structure in quantum mechanics. International Journal of Quantum Chemistry, 16(S13):5–14, March 11–17, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Quantum Statistics, and Computational Methods.

Witriol:1972:CTM

[Wit72] Norman M. Witriol. Canonical transformations and molecular structure calculations. International Journal of Quantum Chemistry, 6(S6):145–152, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Witschel:1976:GFO

[Wit76] W. Witschel. Generating functions for oscillator matrix elements. International Journal of Quantum Chemistry, 10(6): 951–954, November 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wahlgren:1979:VPS

[WJ79] Ulf Wahlgren and Helge Johansen. Valence photoelectron spectrum of CoO. Ab initio calculations of energies and relative intensities within a limited CI framework. International Journal of Quantum Chemistry, 15(4):403–410, April 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wheelock:1970:SEM

[WJC70] Kenneth S. Wheelock, Hans B. Jonassen, and L. C. Cusachs. Semi-empirical molecular orbital calculations: The neighbor atom potential term in the effective Hamiltonian. International Journal of Quantum Chemistry, 5(S4):209–219, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Walker:1972:CED

[WK72] T. E. Walker and Hugh P. Kelly. The correlation energy difference between CH and the carbon atom. International Journal of Quantum Chemistry, 6(S6):19–24, January 17–22, 1972. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Weinberger:1973:SGS

[WK73] Peter Weinberger and Daniel D. Konowalow. A study of the ground states of N_{2}, O_{2}, and F_{2} and their ESCA spectra by the multiple scattering Xα method. International Journal of Quantum Chemistry, 7(S7):353–367, January 21–27, 1973. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Webb:1977:PVI

[WLS77] S. J. Webb, R. Lee, and M. E. Stoneham. Possible viral involvement in human mammary carcinoma: A microwave and laser-Raman study. International Journal of Quantum Chemistry, 12(S4):277–284, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wu:1979:AEL

[WM79] Irene C. Wu and Kenneth J. Miller. Atomic energy levels and oscillator strengths calculated with a screened potential. International Journal of Quantum Chemistry, 15(1):83–95, January 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weimann:1975:ESA

[WMB75] Ludwik J. Weimann, Gerald M. Maggiora, and Paul E. Blatz. Excited states of all-trans and 11-cis retinal: All valenceelectron SCF-MO-CI calculations. International Journal of Quantum Chemistry, 9(S2):9–24, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wormer:1977:QTC

[WMV77] Paul E. S. Wormer, Fred Mulder, and Ad Van Der Avoird. Quantum theoretical calculations of van der Waals interactions between molecules. Anisotropic long range interactions. International Journal of Quantum Chemistry, 11(6):959–970, June 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wachters:1971:CFS

[WN71] A. J. H. Wachters and W. C. Nieuwpoort. Crystal field splitting and Born repulsion in KNiF_{3}. Contribution to the panel discussion on ligand field theory. International Journal of Quantum Chemistry, 5(S5):391–396, January 18–23, 1971. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Weintraub:1978:CEC

[WN78] H. J. R. Weintraub and D. E. Nichols. Conformational energy calculations of some amphetamine analogs using a new solvation model. International Journal of Quantum Chemistry, 14(S5):321–343, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Weislinger:1974:CQM

[WO74] Edmond Weislinger and Gabriel Olivier. The classical and quantum mechanical virial theorem. International Journal of Quantum Chemistry, 8(S8):389–401, January 20–26, 1974.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Weislinger:1975:VTB

[WO75] Edmond Weislinger and Gabriel Olivier. The virial theorem with boundary conditions applications to the harmonic oscillator and to sine-shaped potentials. International Journal of Quantum Chemistry, 9(S9):425–433, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Weislinger:1976:AQV

[WO76] Edmond Weislinger and Gabriel Olivier. Applications of a quantum virial theorem to Kronig and Penney’s model and to a diatomic molecule in static approximation. International Journal of Quantum Chemistry, 10(2):225–231, March 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weinstein:1977:MMM

[WO77] Harel Weinstein and Roman Osman. Models for molecular mechanisms in drug–receptor interactions. Serotonin and 5hydroxyindole complexes with imidazolium cation. International Journal of Quantum Chemistry, 12(S4):253–268, January 9–12, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Weinstein:1978:TMM

[WOEG78] Harel Weinstein, Roman Osman, W. Daniel Edwards, and Jack P. Green. Theoretical models for molecular mechanisms in biological systems: Tryptamine congeners acting on an LSD– Serotonin receptor. International Journal of Quantum Chemistry, 14(S5):449–461, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wojcik:1976:TIS

[Wój76] Marek J. Wójcik. Theory of the infrared spectra of the hydrogen bond in molecular crystals. International Journal of Quantum Chemistry, 10(5):747–760, September 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wojcik:1978:TII

[Woj78] Marek J. Wojcik. Theoretical interpretation of the infrared spectra of the hydrogen bond. International Journal of Quantum Chemistry, 14(S12):219–231, March 12–18, 1978. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Wolf:1971:UPC

[Wol71] W. P. Wolf. On the use of parameters in crystal field theory and related calculations. International Journal of Quantum Chemistry, 5(S5):397–401, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Woodward:1972:PGE

[Woo72] M. R. Woodward. A perturbed generalized eigenvalue equation for the hydrogen molecular ion. International Journal of Quantum Chemistry, 6(5):911–917, September 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wormer:1979:CIM

[WP79] P. E. S. Wormer and J. Paldus. Configuration interaction matrix elements. I. Algebraic approach to the relationship between unitary group generators and permutations. International Journal of Quantum Chemistry, 16(6):1307–1319, December 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wahlgren:1973:MSS

[WRF73] Ulf Wahlgren, Björn Roos, and Erik Forslind. A MO–SCF study of hydroxyl binding and hydroxyl diffusion in pure and fluorine-containing hydroxyapatite. International Journal of Quantum Chemistry, 7(5):905–925, September 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

White:1970:CPT

[WRS70] Ronald J. White, D. E. Ramaker, and David M. Schrader. Comment on a perturbation treatment of the ground state of twoelectron atoms using the coordinates r_{<} and r_{>}. International Journal of Quantum Chemistry, 4(1):1–3, January 1970. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Webb:1975:DVE

[WS75] S. J. Webb and M. E. Stoneham. The display of in vivo energy states by laser-Raman spectroscopy. International Journal of Quantum Chemistry, 9(S2):339–343, January 16–18, 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wilson:1979:FOT

[WS79] Stephen Wilson and David M. Silver. Fourth-order terms in the diagrammatic perturbation expansion for the electronic energy of atoms and molecules. International Journal of Quantum Chemistry, 15(6):683–692, June 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weintraub:1976:CAP

[WTB^{+}76] H. J. R. Weintraub, M.-D. Tsai, S. R. Byrn, C.-J. Chang, and H. G. Floss. Conformational analysis of pyridoxal amino acid Schiff’s bases. International Journal of Quantum Chemistry, 10(S3):99–105, January 12–14, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wormer:1974:TPM

[WV74] P. E. S. Wormer and A. Van der Avoird. Transformation properties of many-electron wave functions with special attention to the relation between pair-correlated DODS and configuration interaction. International Journal of Quantum Chemistry, 8(5): 715–732, September 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weinhold:1971:FBR

[WW71] F. Weinhold and P. S. C. Wang. On a formula of Braun and Rebane for variational bounds to overlap. International Journal of Quantum Chemistry, 5(2):215–220, March 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wybourne:1973:LAQ

[Wyb73a] B. G. Wybourne. Lie algebras in quantum chemistry: Symmetrized orbitals. International Journal of Quantum Chemistry, 7(6):1117–1137, November 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wybourne:1973:GPY

[Wyb73b] Brian G. Wybourne. The “gruppen pest” yesterday, today, and tomorrow. International Journal of Quantum Chemistry, 7(S7):35–43, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Witkowski:1970:DCV

[WZ70] Andrzej Witkowski and Marek Zgierski. Decoupling conditions for the vibronic equation in dimers. International Journal of Quantum Chemistry, 4(4):427–429, July 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Young:1975:NCP

[YD75] Ralph H. Young and Walter J. Deal. Nonadiabatic core polarization and penetration correction in alkali-like atoms: Model computations on excited states of helium. International Journal of Quantum Chemistry, 9(5):835–853, September 1975. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yonei:1977:ATF

[YG77] K. Yonei and J. Goodisman. Accurate Thomas–Fermi–Dirac calculations for diatomic systems. International Journal of Quantum Chemistry, 11(1):163–178, January 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yang:1976:TAS

[YJ76] Cary Y. Yang and K. H. Johnson. Truncated-atomic-sphere model and overlapping-sphere corrections in the SCF–Xα scattered-wave method. International Journal of Quantum Chemistry, 10(S10):159–165, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Yuan:1974:DCE

[YL74] Jian-Min Yuan and John C. Light. Direct calculation of the electron density via the Milne equation. International Journal of Quantum Chemistry, 8(S8):305–311, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Young:1972:NPM

[You72] Ralph H. Young. New proof of the minimum principle for excited states. International Journal of Quantum Chemistry, 6 (3):596–597, May 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Young:1975:DRE

[You75] Ralph H. Young. Divergence of the R^{−1} expansion for the second-order H–H interaction. International Journal of Quantum Chemistry, 9(1):47–50, January 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yang:1976:HEL

[YR76] Cary Y. Yang and Sohrab Rabii. Hydrogen effect in lead selenide. International Journal of Quantum Chemistry, 10(S10): 313–318, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Yafet:1973:CIS

[YW73] Y. Yafet and R. E. Watson. Configuration interaction and screening in the s^{2}p^{4}d and sp^{6} configurations. International Journal of Quantum Chemistry, 7(S7):93–101, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Zielinski:1978:OGRb

[ZBH^{+}78] Theresa Julia Zielinski, Donna L. Breen, Karen Haydock, Robert Rein, and Robert D. MacElroy. Optimum geometries and relative energies for guanine, the imino-enol tautomer of guanine, the enol tautomer of guanine, adenine, and the imino tautomer of adenine as found by the MINDO/2 SCF MO method. International Journal of Quantum Chemistry, 14(S5):355–365, March 3–11, 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Zeroka:1974:VPH

[Zer74] Daniel Zeroka. Variation of the polarizability of the hydrogen molecule ion and the hydrogen molecule with internuclear separation. International Journal of Quantum Chemistry, 8 (1):91–95, January 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zunger:1976:CFT

[ZF76] A. Zunger and A. J. Freeman. Combined Fourier transform and discrete variational method approach to the self-consistent solution of the electronic band structure problem within the local density formalism. International Journal of Quantum Chemistry, 10(S10):383–403, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Zunger:1977:LDF

[ZF77] Alex Zunger and A. J. Freeman. Local density formalism approach to cohesive properties of solids: Diamond, BN, and LiF. International Journal of Quantum Chemistry, 12(S11): 539–546, January 16–22, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Zhorov:1979:RBS

[ZG79] B. S. Zhorov and V. A. Govyrin. Relationship between spatial structure and pharmacological activity of a series of βadrenomimetics. International Journal of Quantum Chemistry, 16(3):517–525, September 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zivkovic:1977:ERS

[Živ77] Tomislav P. Živković. Existence and reality of solutions of the coupled-cluster equations. International Journal of Quantum Chemistry, 12(S11):413–420, January 16–22, 1977. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-state Theory, Collision Phenomena, and Computational Methods.

Zuvia:1978:CDA

[ZL78] Marilena Zuvia and Eduardo V. Ludeña. Charge distribution analysis in terms of Berlin’s binding and antibinding regions for Li_{2} and F_{2}. International Journal of Quantum Chemistry, 14 (1):1–11, July 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zielinski:1978:OGRa

[ZR78] Theresa Julia Zielinski and Robert Rein. Optimum geometries and relative energies for cytosine, thymine, uracil, the imino tautomer of cytosine, the enol tautomer of thymine, and the enol tautomer of uracil by the MINDO /2 SCF MO method. International Journal of Quantum Chemistry, 14(6):851–860, December 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zielinski:1979:ISE

[ZSR79] Theresa Julia Zielinski, Masayuki Shibata, and Robert Rein. Ab-initio STO-3G energy of tautomerism for uracil. International Journal of Quantum Chemistry, 16(S6):475–480, March 4–10, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Zasukha:1979:TET

[ZV79] V. A. Zasukha and S. V. Volkov. Theory of electron transfer in donor–acceptor pairs. International Journal of Quantum Chemistry, 16(4):761–767, October 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).